A Complete Bibliography of The Journal of the American Statistical Association: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

07 October 2024 Version 1.09

Title word cross-reference

#4059 [CMS87].

(k + 1) [BT78]. −2 log R [Cha76c]. 2^{m} [SA70]. 2 × 2 [BM77a, CA77, Con74b, Con74c, Man74c, Mie74, SGS74]. [p(1 − p)]^{−1/6} [Geb71]. α [RS72]. B(x; 2/3, 2/3) [Geb71]. b_{2} [SB77]. B_{k} [Roh74]. B_{n} [GL71]. X̄ [Dun71]. β [Far76]. C_{α} [Joh75b]. χ^{2} [BN74, Cha76c, GGM70, GGM71, GC78, Jen70, Lin78c, McC78c, RA75, TH73, Yar70, ZR71, ZR74]. χ_{1}^{2} [BN74]. χ_{r}^{2} [Ger75]. D [Tsa76]. F [Bul71, GP74, HF70, Jam77, Lin78c, MTC76, Mor71b, Ols73, TB72, Tik71, Tik72]. g [Gam77]. G_{1} [DM74]. H [Her71, KT77, Ken76, Par76]. K [Che73b, Gee72, HS79, Kle73a, Sax71, BM78, BT78, Mee74b, Rea75, Wac77, Wei78d, WN73, WN74b]. Λ [Sol72a, Sol72b]. L × J × K [MTC76]. M [Tha79b, BM78, WM77]. N [Bor77, Rel72]. n_{k} [KTS75]. π [Rao71b, Rao73]. Q [BG73, BS77a, Pat75b, TB70]. R [Baw72, Hen79, Hen80]. R^{2} [Pie79]. r × c [GM77a, Joh75b]. S [Ols77, Rob72, Sch77a, Sch77b]. S_{U} [SB75]. σ^{2} [JG72b]. \sqrt{b_{1}} [BS73b, SB77]. T [GH78a, GH79, Goo73, Hoc76, SS73a, Bai77, BDH74, DS73, DD75, Gam77, Gho75a, Her71, HL73, Joh78b, Kri72, RR71, Som76, Ton70, WS78, Zel76a, Zel76b]. T^{2} [CB71, EE70, Eve79, GP74]. V^{(s)} [PJ70]. W [Che71a]. X [Chi73b, Cox71, Cox72, How74a]. X + Y [How74a]. Y [Cox71, Cox72, How74a]. Y < X [EG71].

-Charts [Chi73b, Dun71]. -Class [Kle73a]. -Decision [BT78]. -Dimensional [WM77, Che73b]. -Distribution [BDH74, Kri72, Som76, Her71]. -Distributions [HF70]. -Estimation [Tha79b]. -Factor [Baw72]. -Independence [Jam77]. -Intervals [DD75]. -Method [GH78a, Hoc76, Ols77, Sch77a, Sch77b, SS73a]. -Minimax [DM74]. -On- [Cox72, Cox71]. -Optimal [Tsa76]. -Out-of- [BM78]. -Predictions [Cox72, Cox71]. -Quantile [RS72]. -Ratio [TB72]. -Sample [WN73, WN74b, HS79]. -Series [Geo74]. -Statistic [RR71]. -Test [Ger75, Ken76, Par76, BG73, GP74, Ols73, Tik71, Tik72]. -Tests [Her71]. -Transformation [GL71]. -Variables [Gho75a, WS78].

/x [BE75].

11 [CT76]. 129th [Ano70m]. 1969 [Ano70m].

2nd [Kle77a]. 2SLS [MM79].

32-bit-word [Che71b].

71 [BH08, CMS87]. 73 [FH81]. 74 [Krz81]. 75 [JTB81].

80k [FH81]. 81a [Krz81]. 81f [JTB81]. 829 [BH08].

= [BR70, GH79].

Aberrant [Hil75]. Ability [FT70, Yea73]. ABLE [Che75]. Absolute [AF76, BK78, Her71, Sch73b]. Absorption [WM77]. Accelerated [Reg73, SZB79, Sin71c]. Acceptability [BRB76, Esp77]. Acceptance [DE73, Sub72]. Acceptance-Sampling [DE73]. Access [Lea75]. Access-Biased [Lea75]. Accident [Web71]. Accommodate [Dav70]. According [Ceb74, GGE74]. Accounting [FNL77, NL77, YFS71]. Accrued [Bha73b]. Accumulating [HW74]. Accuracy [BRB76, BRB77, CH74a, Cra70, DC77a, DC77b, Esp77, Lin78c, Lue71, NMC71, RA75, RW74, Tuk79b, Wam70, Yar70, CH74b]. Accurate [Zac70]. Actual [Gri72]. Acute [Sho70, Sho72a, Sho73]. Adaptive [Efr76, Fra76b, Fra76c, Gas74, Har74a, HURD72, Hog74a, Hog74b, HFR75, Hog76, Hub74a, Jon79, Owe75, PH76, Pre78a, Pre78b, RS79, Str78, ZF73].

Addendum [KGR79]. Additional [GP74, Hah70a]. Additive [O’B79, Qua79b]. Additivity [McD72, Rob75, Sim73]. Adjusted [Bur79, ZF73]. Adjustment [Den71, Plo79, Rub79, SBY79, SF72b, Wal74]. Admissibility [LaM77, MA79]. Admissible [DY78, FV73]. Admissions [NS74]. Advances [Hoa74]. Advertising [Joh73a, McG77]. Advocate [War75]. Affecting [Man74d]. After [KS71a, Lan74, Lan76, You72a, FH78, FH81]. Against [Ham76, Lee77b, Rob78a, SM79, Sho72b]. Age [Ceb74, CM72, Coh76c, Gor73b, Gup76, Key71, MB74b, Pre70, Rei76, Gup77]. Age-Incidence [Pre70]. Age-Parity-Nuptiality-Specific [Gup76, Gup77]. Age-Specific [Key71]. Aggregate [Cra70, DWD76, Hol72b, RZ71, RV77]. Aggregated [Kwe75]. Aggregation [AW72, Iji71, Lau72a, Ven78]. Aging [BT72]. Agreement [AB78]. Ahead [Gra77]. Air [Bro78d, Hsu79, Koh72, LS73, Smi75a, Wyz78, You72b]. Aitken [Bar76b]. Alarm [CR74, Haf78]. Algorithm [Lad75, SL79, Som77, CC74a, Rel72]. Algorithms [AF76, EF76, Has72, SH73, Lin74]. All-Bias [DGL77]. Allocation [DS73, Lev77, McL75, ST75, SY72]. Allometry [Mos70a]. Almon [Fro75, GP75, SW73, Sch74a, SM77a]. Alternate [FGHA73, RA75]. Alternative [CGD76, Dav71, HM70, Kin72b, Koc73, LMSW70, MM73, MJ79, Pro71, RZ71, Roh76, Sim74a, Thu79, Wac77, Sim75a]. Alternatives [AGA79, All77, BL77a, Bly72b, Bly72c, DSW77a, DSW77b, EM77, GJ78, Goo72b, HH77, HH78, Hoe77, Lee77b, Lin72, Mie72, PH72, PH74b, Pra72, SM79, Sho72b, SW78b, Smi76, Smi77a, Thi77b, Win72b]. Amenorrhea [SBY79]. Amer [CMS87, FH81, JTB81, Krz81]. American [Ano70m, BH08, Ban71, BW72, GS74a, GS75, KZ72]. Among [Bly72b, Bly72c, Goo72b, JK72, Lin72, Pra72, SS79, Wei77c, Win72b, Goo70, Lee70b]. Analys [MM72]. Analyses [Sim75a, Sla74, You72b, Sim74a]. Analysis [AK71, Add70, Ame73, AM75, BH75, BH79, Bea77, BK79, BL79, BT75, BS77b, BG74, BRH76, Ceb74, CD72a, CD74, CGD76, Chr79, Cog73, Cor76, CA79, Cra76c, Cra76d, CH77, Dag71, Dav78, DF72, DO72, DM76a, DH72, Dun75b, EM75, Efr75, Efr76, Fer75, Fie75b, FV73, Fra73a, Fra76b, Fra76c, FKB73, Gil72, Goo70, Goo71b, Goo79d, GO72, Ham76, HD78, HM79b, HD77, Hem74, Hem76, Hig70, HS75, Hog76, Hor78, Hsu79, JG72a, KK73, KH71, KL72, KJT72, KVK71, Lau70, LS73, Lee75, LM71, Lin73c, LK76, Lit78, LB72, LS79, Mal70, ML74, MJ79, McK75b, McK76, McL75, MS73, Par79b, Pau78, PW78, Qua79a, Qua79b, RBRH78a, Rao71a, RO76, Rob79, Rot79, Sch73a, Sch78c, SS74, Sco78, Sen72b, SF72a, Shu70, SS78a, Smi75b, SHH78]. Analysis [Ste79, SM73, SM75, Tar79b, TBH74, WNIO77, WWF74, Web71, WC70, WG72, Wil78, Zel71, Zel75, Zel76a, Zel76b, Zel79a, Zel79b, Zer79, Ame74, Ols74, PH74a, Wor74]. Analysis-1 [CD72a]. Analytic [Wil78]. Analytical [Ber71b, Har75, PS76a]. Analyzing [GS75, Hoc77, Joh70]. Andrews [Hoa74]. Angular [JW78]. Angular-Linear [JW78]. Annual [Ano70m, Cra70, Den71]. Annually [SJL76]. ANOVA [Bur79, Efr78, Has72, Lem77, O’B79, Sri72b]. Antibiotic [TY74].

Anticipations [FT70]. Anticipatory [RW74]. Any [Tho75]. Apology [Cor75]. Application [Ang79, BH79, CR75b, FH79, GM78, GH78b, Goo75a, GKAH71, Joh73a, KL72, LP75, Lor71, Max73, MS74, PW74, Sim77, SF72b, Tar79b, Vin76b, Web71, Wei78a, Yak76, ZR71, ZR74]. Applications [AM75, BT75, BS77b, BS74b, BDH74, EHB78, Gas74, Gei75, GG71, Gok73, Har71, HS74, Har74a, Hog74a, Hog74b, Hsu79, Hub74a, HB73b, Ifr70, Jam72, Koc73, MM73, Sam76a, SL79, Sti76, Sti77c, Sto78, TS75, Tha79b, Tho73b, VG78, WL77, Wit74, JTB80, JTB81, Ano79e]. Applied [CR74, DF74, HD77, Shu70, TY74]. Appraisal [Fie72]. Appraisers [Mee79]. Approach [Ano79h, BH76, Bor77, CB70, Che75, Dav77a, Den71, DG75, DG76, Dun74a, EM73, FH76, FM74, FM77, GE79, GW78, Hal73, Hau76, HST79, IJK76, KJT72, Koc73, Koo73, KM74, Laa79, Leo77, Mar76, PD74, PY74, Qua72, Roy76, SZB79, SD73a, SD73b, Ven78, Wac77, Win72a, WF79]. Approaches [Car77, CGD76, Har77c, Har77d, Rao77a]. Appropriateness [NM70]. Appropriations [Kim70]. Approximate [Dav79b, Dic76a, Dic77b, Eas72, Gho79, GK70, How74a, Hub74b, LR72, SF72a, Sim72b, Sob71, TB77a, Wal70]. Approximately [AKC76, Man74b, Sin71b]. Approximating [GG71, Gok73, Woo71, WC76]. Approximation [BE75, BR77, CH75, CMH73, Geb71, GL71, JS72, JS73, Kra73, Kra75, Lor71, Obr79, SM77a, TW77, Woo76]. Approximations [Bur73, Bur74, CH74a, CH74b, Ebb73b, Hoa77, Lin78c, MG74, SS78b, SK73, Ste78b, Yar70]. Arbitrary [BT74, SM77a]. Area [DeG73, Dic77b, DWD76, GH78b, RO76, Som76, Sto71]. Areas [Eri74, Liu70, SZ70, Sub72, Sud70]. ARIMA [DS78, Sab77]. Arithmetic [Gla76c]. arm [Fox74]. ARMA [McL79, Vin76c]. Armed [Ber78d, Woo79]. Arrangements [NB71, NB72, NBH73]. Arrival [SW78a]. Article [Hay71b]. Articles [Ano71e, Ano71f, Ano72f, Ano72g, Ano72h, Ano72i, Ano73e, Ano74c, Ano74d]. Ascribed [Ste71, Ste73, Wil73]. Aspect [CD74]. Aspects [BT72, Lau72a, OS70]. Assessment [CH78a, Edw75, Hog75a, Hog75b, NL79, RA75, Win75d]. Assessments [PF78]. Assignable [Dun71]. Assigning [Har73]. Assignment [Gav78, Kie78a, Mal78, SM78a, SM78b, Sim78b]. Assoc [CMS87, FH81, JTB81, Krz81]. Associate [Cle70b]. Associated [Cor76, JS75]. Association [Agr76b, Agr77, Ano70m, Ban71, BH08, GK70, GK72, Goo79d, LS73, Sim78a, Sri70]. Assuming [Atw75]. Assumptions [BK76, Lis72, Mur72]. Asymmetric [Car79, FS72, SM79]. asymmetrical [Joh78b]. Asymptotic [AGA79, And71, And74a, And76, Bai79, BK78, Baw72, CT73, CD74, Cha71c, Cro74, DP72, GK72, HM73, Jen70, KvE72, Lee77a, LF71, LF73, Mie72, Nat72, RS73b, RS74a, SW76, Som76, Wal79a, Wei71b, Wei73, Wei75, SW77]. Asymptotically [Bea74, BC79, BM70a, Gro72, Loy76, McL75, Tak71b, CC74a]. Attempt [Mil76]. attenuation [FH78, FH81]. Attitude [Dun75b]. Attributes [Lev77, WL77]. Augmentation [HB73a]. Augmented [Sto78]. August [Ano70m]. Australian [BL79]. Author [Ano79a, McN73]. Autocorrelated [GG76b, Hur72, Spi79b]. Autocorrelation [Smi76]. Autocorrelations [BP70, Chi76, Chi74]. Automatic [FH76]. Automobile [Hol70c, Hol71b, Wil71]. Autoregressions [Jon75]. Autoregressive [AW72, And78a, Bai79, BP70, DM79, DF79b, Fis72, GS73c, HT79, Hsi79, KG70, McC78c, Sab77, Chi74]. Autoregressive-Integrated [BP70]. Auxiliary [Han73, Sin71b, Sri71b]. Available [Ano72r, Ano72s, Ano72t, Ano73n, Ano74m, Ano74n, Ano76n, Ano76o]. Average [BP70, HT79, Rei76, Sab77, TG71a]. Axes [DGL77]. Axis [Jol73].

B [Goo78d]. Back [Ano70a, Ano70b, Ano70c, Ano70d, Ano71a, Ano71b, Ano71c, Ano71d, Ano72a, Ano72b, Ano72c, Ano72d, Ano73a, Ano73b, Ano73c, Ano73d, Ano74a, Ano74b, Ano75a, Ano75b, Ano75c, Ano75d, Ano76a, Ano76b, Ano76c, Ano77a, Ano77b, Ano78a, Ano78b, Ano78c, Ano78d, Ano79b, Ano79c, Ano79d, Ran73]. Balanced [BB76b, Ell77, Hem74, Hem79, KF70, Lee72, Lee73a, Lem77, McC76b, SA73, Sah76]. BAN [Sam78]. Bandit [Ber78d, Woo79]. bandits [Fox74]. Bands [Bow70, Kab76, KM79, SB71, TB77b, TB79]. Banerjee [BH08]. Bartlett [Gla76a, CG78, Gla76b, THP78]. Base [Bro76a]. Based [BS79, Can74, Cha70a, CCM71, CA74, Den71, Fin77b, GG78, Gli72, HM79a, HST79, KvE72, KM74, LK76, MJ79, Meh75, Mie72, MW70, OPRS77, Pat76, PAA70, Qua79a, RBRH78a, SL74a, SL74b, Sly72, SC75b, SC77, Van76, Wei78d]. Basic [WB71]. Bayes [WD72, BP73, Bor77, CR74, CZ77, Dic77b, DD75, Dix76, Dun74a, EM71, EM72, EM73, GC74, Rom72, SS71, Sha73a, SW78a]. Bayes-Fiducial [BP73]. Bayes/Non [GC74]. Bayes/Non-Bayes [GC74]. Bayesian [Jay79, Zel76b, AM73, Bac79, Bas73, Dav78, Dav79b, DD77, DE73, DG75, DG76, DM76a, Fer75, Hal73, HM79b, Hoa70, Hor78, JP74, KK73, Leo77, LB72, MS73, Owe75, PY74, PS76b, RG72b, Sam76b, Ste79, SV76, SM73, SM75, WNIO77, WF79, Zel71, Zel75, Zel76a]. Be [AS77, Ste71, GF73, Ste73, Str74, Str75, Vea75, Wil73]. Behavior [BB77, Dun75b, GC70, Hil75, HMW74, Kim70, Lau72b, LP75, Mie72, NL77, The74, TDL75, Fox74]. Behrens [Gho75b, Sch70a, Wan71]. Belgium [Gup73]. Bell [Vin76b]. Belt [BL79]. Benoît [Goo78d]. Bernoulli [Ant72, BS73a, HB73b]. Bessel [HS74, Kau72]. Best [Baw72, CCM71, CL76, Eck76, Gee72, KN75, OL71, Rot78b, Sho72b, Sob71, Wac77, ZM71, CC74a, Gov71, Gov76]. Beta [Geb71, Jam72, Jam75, JK72, MG74]. Beta-Integral [Geb71]. Beta-Stacy [JK72]. Better [BS73a, KW71, KW74c, Nor76, PS76a, SW71]. Between [AB78, BNW70, Blo77b, Dur70, Fre77, Gab78a, Geo74, Gol75a, Gol73, Goo75b, Krz79, Krz81, LS73, Low74, Mis72, Mor72, NNO76, Pie77, PW78, RC76, Ram78a, Ram78b, Sam70, Sri70, Sto71, Wal74, PS74]. Between-Groups [Krz79, Krz81]. Bhattacharyya [BH08]. Bias [Bai75, BN73, CM74b, CM74c, CMH73, DS74, DGL77, EM78, Gol75a, Hol70c, Hol71b, Isa74, Kar70, Koo73, LP70, McK75c, Mik72a, MB74b, Rub79, SGO71, Sco78, SBY79, SD73a, SD73b, Tho79, Wes75, Wil71]. Biased [GM77b, Lea75, Obe75]. Biases [WM70]. Bickel [Hoa74]. Bidding [Pel71]. Bidirectional [Sha74]. Binary [Agn74, Gol73, Krz75, Moo73b, Mor72, OK76]. Binomial [Bil77, Bre70, Bur75, DM74, Eck76, Geb71, Gho79, HG70, HG71b, Hwa74, Hwa75, KW71, KW74c, KS71b, Lad75, McC72, Mee74a, Mee74b, Nor76, PE78, Pol73b, PS76a, SW71, Ten70, TG71b, Win74b, Wis72, Fis79, Rel72]. Bioassay [Tsu72]. Biopharmaceutical [WL77]. Bird [Sen76a]. Birnbaum [Dur70, Jos76]. Birth [Gup73, Sab77, SMRL70]. Birthday [Nau74, Sap72]. Births [BH79, Gup76, Gup77, Lee74, Lee75]. Bisquare [Gro77]. bit [Che71b]. Bivariate [Ame74, AKC76, Bas71, BBH72, BJ73, BB74, BDH74, CT73, Fry70, HAB71, JS75, Jol73, Kri72, MS76a, MY78, MS73, Sha77a, SB77, SZ73, Ste70a, TS75, Tih72, dO74]. Block [AS78, BB76b, KS75, MS76d, PH72, PH74b, Por73, Por74, Qua79b, Rob75, Sha70a, SW78b]. Blocks [BM77b, Fen79, Qua79b]. BLUE [LMH73]. BLUEs [Low74, Low77]. BLUS [GS72, RS76]. BMDP [And78b, BF78a, BF78b, Dix78, McC78b, Mul78a, Mul78b, Nel78, NFJN78, Rya78, San78, Sch78b]. Bode [Pra73a, Bly71, Bro71a, Dan71, Efr71a, Efr71b, Goo71a, Pra71, Stu71]. Bonferroni [Bai77, DS73]. Book [A.73, A.74a, A.74b, A.74c, AY76, Abo78, Abr79, Ada76, Ade78, Agr76a, Agy78, Aig71, Aig72, Aig75a, Aig75b, Aig75c, Ait72, Alb76, Alb79, Ame70, AZ79, Anb76, And70b, And77, And79, And74b, Ang78, Ano75f, Ano76h, Ano76i, Ano77g, Ano77h, Ano78g, Ano78h, Ano78i, Ano79k, Ano79l, Ano79m, Ano79n, Ant73a, Ant79, Ant73b, Arc73, Arn75, Arn76, Arn78, Arr70, Atk76, Atk79, Atw76, BJ72, Bal74, Bal75, Ban78, Bar79, Bar70a, Bar70b, Bar74, Bas78, Bau71, Bea76, Bec76, Bec73a, Bec73b, Bec74, Beh73, Ben75, Ben72, Ber71a, Ber73a, Ber78a, Ber73b, Ber70, Ber76, Ber78c, Ber78b, Bha71a, Bha79b, Bha79a, Bic73, Bin78b, Bir78a, Bir78b, Bla71, Blo77a, Blo71, Boa73, Boe71, Boe72, Boh75, Boh71, Bon79, Boo77, Boo79, Bor70a, Bou77]. Book [Boz73, Bra70, Bri71, Bri72, Bro78a, Bro78b, Bro71c, Bro71b, Bro71d, Bro72a, Bro72b, Bro72c, Bro79, Bro73, Bru73, Bru78, Bue76, Bue79, Bul72, But71, But79, But76, But77, C.71, C.72a, C.72b, C.72c, C.74a, C.74b, Cam79, Cas78, Cha76a, Cha79, Cha73, Cha75a, Cha74a, Cha78, Cha76b, Cha75b, Cha70b, Cha71d, Che73a, Chi71, Cho76, Cla70a, Cla70b, Cla71, Cle70a, Cle71, Coc70c, Coc70a, Coh71, Coh74b, Coh78, Col75, Col72, Con76a, Con76b, Coo74, Coo71b, Cox73, Cro79a, Cro71, Cru75, Cza70, D.72a, D.72b, D.73a, D.73b, D.74a, D.74b, D.74c, D.74d, D.74e, D.74f, D.74g, D.74h, D.75a, D.75b, D.75c, D.75d, Dal71, Dal77, Dal78, Dal79, Dav72a, Dav72b, Dav73b, Dav76, Dav77b, Dav79a, Dav79c, Daw76, Day79, DeG70, DeG71a, DeG71b, DeG75, DeG76a, DeG76b, DeL79]. Book [DeW71, Dea79, Dem73a, Den70, Den73, Des70, Dev78a, Dev78b, Dev79, Dic70, Dic73, Dic77a, Dok71, Dok72, Dou71, Dra73, Dra71, DG72, Dud78, Dun75a, Dun75b, Dun76a, Dun77, Dun79d, Dun79a, Dun79b, Dun79c, Dun74b, Dwy73, E.74, Eas77, Eat73, ET74, Eis73, EJ75, Ela75, Ela77, Els77, Eng72, Eni71, F.74, F.75, Fab70, Fai77, Far70, Far73, Far72, Fed70, Fei75, Fel79, Fie75a, Fie76a, Fie76b, Fie76c, Fie77a, Fie77b, Fie77c, Fis77, Fis70, FL76, Fox76, Fra70, FF78, Fra78, Fra79, Fra71, Fre78, Fri70, Fri78, Fry71, Fry72a, Fry72b, G.71, Gal73, Gan79, Gan71, Gar78, Gav70, Gav73a, Gav73b, Ger72, Ger77b, Gib73, Gil74a, GE75, Gil76, Gil78, Gle73, Gle74, Gle75, Gli78, God74, God70, Goe74, Goe77, Goe78, Goe79, Gok79, Gol70]. Book [Gol76, Gol77, Gol72, Gol78, Goo72a, Goo78a, Goo78b, Goo78c, Goo79a, Goo79b, Goo79c, Gor73a, Gor76, Gov70, Gre73, Gre75b, Gre74, Gui78, Gul75, Gun79, H.71, H.72a, H72b, H.75a, H.75b, Hab77, Hag78, Hai70, Haj74, Hal74, Ham75, Han75, Har70, Has76, Hay78, Hay76, Hea79, Hen72a, Hes71, Hey71, Hey78, Hil70a, Hil74a, Hil72, Hil73, Hil76, Hil78, Hin77a, Hin76a, Hin76b, Hin77b, Hix76, Hoe76a, Hoe76b, Hog79, Hog72a, Hog78, Hol75, Hol74, Hoo74, Hor76, How74b, Hsi70, Hud76, Hud75, Hum75, Hun72, Hun75, Hun70, Hur73, Hur74, Ign74, Iji70, Imr76, Imr78, Imr79, Ing77, Ism76, Ize78, Ize79, Jay79, Jed73, Jen76a, Jen76b, Joh72, Joh73b, Joh73c, Joh78a, Joi73, Jon73, K.71, K.72a, K.72b, K.72c, K.73, K.74, K.75, Kad73, Kad75a, Kad75b, Kai72]. Book [Kal78b, Kap70, Kas74, Kat70, Kau73, Kei72, Kem70, Kem72, Kem73, Kha77, Kim77, Kin72a, Kin76, Kin77, Kin78, Kin79, Kle77a, Kle77b, Kle70, Kle74, Kle73b, Kme74, Kme77a, Kme77b, Kme77c, Kme77d, Kme79a, Kme79b, Koo77, Koo72a, Koo72b, Koo76, Kot70, Kot72, Kot73, Kou72, Kro71, Kup76, Kup78, Kup79, Kur72, LaC79, LaV75, Lac74, Lac76, Lac77, Lah75, LS71, Lan75a, Lan79, Lap75, Lar74, Lar75, Lav79, Lay75, Lea78a, Leb74, Leh77, Lev75a, Li72, Lil75, Lin78a, Lin73b, Lin77b, Lin78b, Lin79b, Lin79c, Lin79d, Lin76b, Lov76, Lov78, Luc72, Lyt72, M.72, M.74, M.75c, M.75a, M.75b, Mad73, Mag72, Man72, Man74a, Man70, Man71, MYH76, Man75, Mar73, Mar78, Mar70a, Mar70b, Mas79, Mas72, McC74a, McC75a, McC76a, McC73b, McC74c, McC78a, McC74d, McC75b]. Book [McC79a, McC79b, McK75a, McK77, Mee78, Met71a, Mey72, Mey75a, Mey75b, Mic76, Mie71, Mil70a, Mil75, Mil79, MW79, Min71, Mir73, Mit71, Moo72b, Moo72a, Moo72c, Moo73a, Moo74b, Moo74a, Moo75b, Moo75c, Moo75a, Moo76b, Moo76a, Moo76c, Moo77b, Moo78a, Moo78b, Moo78c, Moo79, Mor77, Mor71a, Mor73, Mor70, Muh73, Mun77, Nag79, Nas79, New76, New79, Noe70, Nol73, Nol74, Nor71, Van78, O.70b, O.70c, O.70d, O.70e, O.70a, O.71, Oak74, Ode71, Ode79, Ore79, Ork72, Ork75, Owe71, P.72, P.75a, P.75b, P.75c, Pac70, Pag70, Pai77, PR75a, Par79a, Par73, Pas72, Pas73, Pas76, Pat73, Pec71, Pet77b, Pet78b, Pet79, PA76, Pir73, Plo71, Poi79, Pol73a, Pol75, Pov79, Pra74, Pre76, Pre74, Pre78d, Pro77, Put78, Put79, Q.74a, Q.74b, Qua77a, Qua78, Qua77b, Rag71, Rag75]. Book [Rag78, Ram74, Rao77b, Rav78, Rel76a, Rel76b, Rho75, Rho77, Ric76, Ric72, Rob74, Rob78b, Rod76, Rog78, Roh77, Roh78, Rol72a, Rol72b, Rol76, Ros74a, Ros74b, Ros73, Rot78a, Rot71, Rub74a, Rym71, S.73, S.74a, S.74b, S.74c, S.75, Saa71, San74c, San77b, Sär72, Sär79, Sch79a, Sch71, Sch73c, Sch78a, Sch79b, Sch70b, Sch75a, Sch75b, Scl73, Scu70, Sed78, Sel75, Sen71, Sen72a, Sen75, Sen76b, Sen76c, Sen77, Sen78a, Sen78b, Sen78c, Sen79a, Sen79b, Sen73, Sen79c, Sha76a, Sha75, Sha78a, Sha77b, Sha78b, She79, She72, She75, Sho76, Sho72c, Sho75, Sie72, Sim79a, Sim79b, Sim76, Sim72a, Smi71a, Smi71b, Smi77b, Smi77c, Smi79a, Smi79b, Smi79c, Sok72, Sri72a, Sta78, Sta79, Ste75, Ste70b, Sti78a, Sti73a, Sti75, Sti77a, Sti78b, Sti78c, Sto70, Str72a, Str76, Str71]. Book [Suc74, Sun79, Sut78, Swi73, Swi76, Swi78, Syk74, Tae76, Tae77, Tak71a, Tam72, Tan78, Tat71, Tay74a, Tay74b, Tha79a, Thi77a, Tho73a, Tia75, Tro75, Tru70, Tru77, Tru78a, Tru78b, Tuc78, Tuf70, Tuf79, Urq75, Van72, Van73, Vin76a, Wac79, Wah71, Wal79b, War74, War79, War77, War78b, Was79, Wat79, Wat75, Wat77, Wat78, Web76, Web78a, Weg74, Wei74, Wei76a, Wei76c, Wei76b, Wei77b, Wei78b, Wei71a, Wel73, Wet78, Whe73, Whi76, Whi79a, Wic79a, Wic79b, Wig70, Wig77, Wil76, Wil74, Win74a, Win75a, Win75b, Win75c, Win76a, Win77, Win78a, Win78b, Win79, Win78c, Win76b, Wio75, Wip71, Wol79a, Wol74b, Wol74c, Yac74, YS72, You70, You73, Z.72, Zab76, Zab78, Zah77, Zar70, Zar72b, Zys73, Zys75]. Borges [Geb71]. Both [CA77, Dol76, HG71a, Krz75, War72, Tuk79b]. Bound [EP70, Jam77]. Bounded [Jam77]. Bounded-Sum [Jam77]. Bounding [GK75, Kri79]. Bounds [Gle72, Haw71, HQ75, Kwe75, Man74b, Meh75, Mor72, PR75b, Pat75a, SC75b, WM77, Yu75]. Box [Spi79a, Poi78, Spi78]. Bradley [Bea77, Dav70]. Branching [CR75c, DF74]. Broken [Bof76]. Budget [Lee70a].

Cable [McG77]. Calculating [Hah70a, Hah70b]. Calculation [OPRS77]. Calibration [And70c]. Can [Ste71, Ste73, Wil73]. Canadian [Wil71, BM73b, Hol70c, Hol71b, Hsi79, OS70]. Canner [PS76a]. Canonical [Con74a, Gle76, HAB71, KH71, Krz76, Sri70, Wil78]. Capita [TF73]. Capital [Bha73b, Bha76, Kim70, TG71a, Wei77c]. Capture [Wit74]. Capture-Recapture [Wit74]. Carlo [Ano77c, JTB81, Spi79a, Bha73a, CS73, CM74b, CB71, Eve79, Gil72, Ham76, JTB80, LS75, MG75, Mik75, RG72a, Spi78, SZ73]. Case [Bak74, BBDS71, Bas73, BM78, BL79, BR77, Bro75, BL77b, BL77c, Dud71, EM71, EM72, Eri76, Gal77, Gil74b, Lee77b, MS72a, MM79, McD74, MR70, NS74, RR71, RS73b, Zac73, BBDS72, RS74a]. Cases [GM75, Hil75, SS73a, Wei71b, Wei73, Wei75]. Casinos [VG78]. Categorical [Agr77, Che79, Con72a, Hoc77, LM71, ML74, SK76]. Categories [Goo79d, WG72]. Categorized [CG70, CG71, Cor71]. Category [Agr76b]. Cauchy [BBS74, Can74, Cha70a, Fer78]. Causes [Dun71, JG75]. Cell [FH73, Hei75, ZR71, ZR74]. Cells [HG73]. Censored [CD72b, DL76, Efr77, Eng75, HW74, HM73, Max73, Sel74, TG75b, Tur74, WEB78b, WE79, CC74a]. Censoring [Gov71, Gov76]. Census [CT76, FH79, GOST75, Mil76, Reg73, Roc75]. Centenarians [SP76]. Center [Noe73]. Central [Bul71, Ste78b, WD76]. Centuries [Sti77b]. Certain [Ber78d, BS75, Lit74a, RH71, Sam78, Sob71]. Certainty [Har75].

Chain [Dev73a, Dev73b, GY70, Hen71, TB77a]. Chains [DL72, Yak76]. Challenges [KK79]. Chance [Goo78d, Kis78]. Change [Blo77b, MY78, Sri75, YFS71]. Changes [Ano77e, BJ77, BD73, DL77, Epp76, Eri74, FS72, Kob73, Tei71, Ano77e]. Changing [KS71a, Rei76]. Characteristic [PAA70]. Characteristics [OPRS77, Scl79]. Characterization [DR71, Lau74, Sam76a, Ski70, Wol74a]. Characterizations [Cho77, Dri78, Dri79, Mos70a, Sha70b, Sri74]. Characterizing [Rub74b]. Charged [Hol70c, Hol71b, Wil71]. Charts [Chi73b, Dun71]. Cheaters [Ang74]. Checking [Hau76, Pie72, Pie73]. Chernoff [Cra75]. Chi [Ame76, BE75, BR77, Bur73, Bur74, Cha72, Coh76b, DG73b, Dav77a, DS73, Fle71, Goo71b, GY70, HT72, Hoa77, KP78, Lar78, Lis72, MG74, ML74, Moo77a, Odo70, RB71, SS77b, Ame74]. Chi-Square [Ame76, BE75, BR77, Bur73, Bur74, Cha72, DG73b, Goo71b, GY70, Lis72, MG74, Odo70, RB71, SS77b, Ame74]. Chi-Squared [Coh76b, Hoa77, KP78, Lar78, Moo77a]. Chi-Squares [Dav77a]. Chile [Bet71]. Choice [Agr76b, Bly72b, Bly72c, Dun74a, Goo72b, Lan75b, Lin72, Obr79, Pra72, Wei70, Win72b, Zin72]. Choices [HB72]. Choosing [Che75, KW71, KW74c, Mee74a, Ofo74, PW78, Som70, Som74, Som75]. Cigarette [War78a]. Circumstantial [Fin77b, SC75b, SC77]. Cities [BMB78]. Citrus [Pra70]. Civil [GS73b]. Claim [Ste71, Ste73, Wil73]. Class [Bre79, DC73, DG73a, Fli79, Gle72, GJ78, KS73, Kle73a, Sob71, TL70, Wei77a, WSB73, Wol77a]. Classes [DG73b]. Classical [Mur72]. Classification [BM77a, CA74, CR75a, DG78, Gli72, Gol75b, GR75, GR77, JG72b, Krz75, Krz76, McC72, OK76, Isa74]. Classifications [BS78, GK70, Goo70, GK72, Goo79d]. Classified [Han78]. Classifying [Rea75]. Clinical [Ber78d, Klo76, Wei77a, Wei78a]. Close [Bur73, Rea75, Bur74]. Closed [Sau74, SMRL70]. Cluster [BH75, HD78, Lin73c, LK76, Rao71a, Sud78]. Clustered [Coh76b]. Clustering [BH76, CR75b, Har72, Hub74b, Ran71, Roh74, WL77]. Clusters [Lee79, WN74a]. Cochran [BG73, Pat75b, TB70]. Coefficient [And74a, And76, Cha71a, DH77, FH70, Fro73, GH76a, Gri72, KW79, Kra73, Kra74, Kra76, MM79, May73, Mur70, NM70, Ott73, PS75, SM75, WT76, Zar72a, Kra75]. Coefficients [BBDS71, BBDS72, Cho71, DC71, Gar77, Gri72, Hen79, LM78, MSW74, MY76, Mis72, Pyn79, Raj75, RS73b, RS74a, SU74b, SM77b, Vin76c, ZP79, Hen80]. Cognitive [Edw75, Hog75a, Hog75b, Win75d]. Cohort [CM72]. Coleman [ZR74, ZR71]. Collection [Ven78]. College [Har77b]. Collegiate [Mos70b, Pol73b]. Coloring [HD78]. Colton [MI78]. Columns [Bur79]. Combination [DM74, Fle71, MS76c]. Combinations [BS78, CM73, Dav77a, HM70, Tho76, WS78]. Combined [KS75, Sha70a, Sud78]. Combining [BC79, Coh76a, KP78, LF71, LF72, LF73, Str74, Zec71, Str75]. Comes [DeG73]. Comment [All77, And78b, Bar76a, Bar77, BRB77, BK79, Bin78a, BL77a, Bjö78, Bly71, BT76, Bro71a, Bro77, Bro78d, Bry78, Bue77, Chr79, CH78c, CH79, Cro79b, Dan71, DS79, Dem77, Dic76b, Dix78, DWW79, Edd79, Edw75, Efr76, EM77, Ell79, Fie75b, Flu79, Fow78, Fra76d, Gab78a, Gab79, Gas74, Gav78, Goo71a, Goo72b, Har74a, Har78, Hay71b, Hoe77, Hog76, Hos78, Hub74a, Joh78c, Kem77, Ken76, Kie78b, Kie78a, Koo73, Kru79, KNP79, Lin72, Lin76a, Lin77a, Lin79a, Mal78, Man74c, McC78b, Mie79, Mie74, Mos79, Nel78, Ney79, NFJN78, Ols77, Par76, Pot72, Pra71, Pra72, Pra76, Rao77a, Rob79, Rot79, Rya78, Sam79, San78, Sch78b, Sim78b, Sim79c, Smi75a, SC77, Smi77a, SGS74, Ste73, Stu71, Tar79a, Thi77b, Vil76, Wel79, Wil71, Wil73, Win72b, Win75d, Wol77b, Pra73a]. Comments [Chi73b, Dem73b, Fra73b, Fuc79, Ofo74, Pra73b, Sav70, Som74, Tuk79a]. Commitment [Gor73b]. Common [Car70, Zac73]. Commonly [RB71]. Community [Ban71]. Comovements [Epp79]. Comparative [MM73, Mik75, Ols74, Ram77, Thu79]. Compared [CM76, DuM73, Efr75]. Comparing [BT78, Gar72, Gar73, Joh75a, Mee79, MI78, Pet78a, Pot72, Str72b, Str72c, Wil78, ZG72, Str73]. Comparison [Adi73, Ano77c, AF76, BNW70, Bea74, BA77, BG73, Bha73a, BH72, CS73, Car77, Cha76c, CS74, CA79, Dav70, DC71, Fen79, Fra73a, Ger77a, GG72, Gho79, Gol75b, Har75, Har77a, HH75, IJK76, KTS75, KR78b, KGR79, Krz79, Lev75b, Lin74, Lis72, LF72, Lit74a, LSW78, Nea72, Odo70, Ols77, PB76, Pyn79, Rea75, Sah76, Sam70, Sch77a, Sch77b, SB79, SS73a, SZ73, Sto78, SM78c, Tam79, TB72, Wel77, Krz81, Tin70]. Comparisons [BJN70, DQ72, DD75, EG75, EHB78, Gab78a, Gab78b, GH78a, GH79, Geo74, Hoc76, JK72, KR78b, KGR79, Kno76, Lar78, LS77a, ML74, Mil77, MS76d, PAA70, Ram78a, Ram78b, Ste74, Tam79, TC72, WD72]. Compartment [Met71b, Met71c]. Compatibility [MS74]. Competing [Gil72]. Competitions [Har73]. Competitor [Tho75]. Competitors [EM73, ML74]. Complete [BH76, BDG78, DL76, Eng75, HD78, Hub74b, Qua79b]. Complete-Link [BH76, HD78, Hub74b]. Completely [Zer79]. Completeness [LaM77]. Complex [KL72, SS79]. Complexity [MA79]. Complicated [Woo71, WC76]. Component [CG74, Ebb73c, Ebb73a, GW79, Gre75a, Har77c, Har77d, LS77b, Rao77a]. Components [Ber74, Bro78c, BS78, CG70, Cor71, EP70, FHJ78, Haf78, pH78, HS73, HG73, Haw74, Hol71a, Hol72a, HM70, Krz79, LaM73, LK78, MM73, Rao72, Sah76, SH79, SM73, Krz81, Pri74, CG71]. Composite [Cha74b, TP79, Wol79b]. composition [Hil70b]. Compound [HAB71, Pau78, SSM75]. Compromise [GC74]. Computation [LS77b]. Computational [BRB76, BRB77, DC77a, DC77b, Esp77, NBH73]. Computations [Ber77]. Computer [FS77, KR76, KR90, MM71b, Wam70, Fis79]. Computerized [WC76]. computers [Che71b]. Computing [Her71, KvE72, Sha73a, Lin74]. Concentration [BB71]. Concept [Goo75a]. Concepts [Bir72, Sha77a, Kle77a]. Concerning [Dag71, Mur72, You76a]. Concomitant [Cor76, DF72, Woo79]. Concordance [Kra76]. Condensation [Bof75]. Condition [Ber77, Cra76a, Hol71a, Hol72a].

Conditional [And71, Bar77, Bro77, Bue77, Dem77, Jam75, Kem77, Kie77a, Kie77b, Law73, Lin77a, Ols73, PR75b, Pot72, Str72b, Str72c, Str73, Wol77b]. Conditionality [Bir70, Dur70, Sav70]. Conditions [EE70, HF70, Mal72, MTC76, MS76d]. Confidence [Agn74, AST73, Ano78e, Bar77, Bau72, Bow70, BB76b, Bro77, Bue77, BS78, Bur73, Bur74, Bur75, Car77, Car78, Coh72, Dem77, Dud72, Eas72, Far76, Gho79, GW79, Gro76, Gro77, Har71, Her71, HQ75, How74a, HB73b, Jol73, Kab76, Kem77, KM79, Kie77a, Kie77b, Kre76, Lan73, Lan74, Law73, LR71, Lin77a, Man74b, MP71, Noe73, RR76, SA73, SW76, SW77, SB71, TG75b, TG77, TB77b, TB79, WG71, WT71, Win74b, Wol77b, Joh78b, Lan76]. Confidentiality [Fel72a, Fel72b]. Configuration [Bof76]. Congenital [Che78]. Conjugate [Coz74]. Connecting [HB77]. Connor [Jam72]. Consensus [DeG74, ZR71, ZR74]. Consequences [BYJ73]. Conservative [Het73, MP71]. Consideration [Bro78d, Wyz78]. Considerations [Agr77, KM74]. Consistent [BN75, Fel74a, Fou77, KW74b, KW74a, Lit78]. Constant [BW76, Ger77a]. Constant-Dollar [BW76]. Constrained [BN75, Har74b, Har79, Lie76]. Constraints [Dav78, Jes70, LP70, NB71, NB72]. Constructing [Bau72, Chi72, Chi73a, GJ75, McL75]. Construction [AS78, Gup73, Lea74, Moo77a, Tsa76, WG71, YFS71]. Consumer [CG73, Pau78, SF71]. Consumption [War78a]. Contagious [HG70, HG71b]. Contain [Hah72]. Containing [MR79, McL77, Rot78b]. Contaminated [Che71a]. Context [FHJ78, Owe75, Rev74a, Rev74b, Spi79b]. Contingency [BM76b, BB77, CA77, Coh76b, Con74b, Con74c, DY78, FG70, Fie70, GM77a, GJ78, Goo71b, Goo73, Goo75b, Hom78, IJK76, Joh75b, KJT72, KVK71, Lee77a, Man74c, Mie74, Nat72, Odo70, Pat74, Sim75a, Sim78a, SGS74, Wag70, WG72, Sim74a]. Continuity [Con74b, Con74c, Man74c, Mie74, SGS74]. Continuous [BB74, CA74, Krz75, PW74, Rom72, Sch73d, Sch74b, Poo74]. Contrasts [EHB78, Gam77, KR78b, KGR79]. Control [Gam77, Koh72, Rub79, SK76, WM71, ZF73]. Controlled [Pre78c, Wei78a]. Controlling [NR71]. Conventional [Her76]. Convergence [MM71a, Mic72, Raj75]. Convoluted [Sam76a]. Corporate [Bha73b]. Corrected [Sly72]. correcting [FH78, FH81]. Correction [Bur75, CMS87, Con74b, Con74c, Cow75, FH81, JTB81, KR90, Krz81, Man74c, Mie74, Sch78c, SGS74]. Correlated [BM77b, BG73, Cha71a, Ebb73a, Fel74b, Fel74c, Haf78, HAB71, Ham71, KW74d, Lee74, Vit79, Wei78c]. Correlation [Bof70, BB77, Cho71, Coc70b, DC71, Hau76, Hen79, Hen80, Jen70, KH71, KW79, Kra73, Kra74, Kra75, May73, McL79, NM70, Ott73, PS75, RS73b, RS74a, Sim75b, Sto71, Wes75, WT76, Zar72a, PH74a]. Correlations [Gol73, Mor72, Pie72, Pie73, Sam78]. Corrigenda [AG72b, And76, Ano72e, Ano77c, Ano77e, Ano77d, Ano78f, Ano78e, Ano79e, Ano79f, Ano79h, Ano79g, Arn73, BBDS72, BW70, Bro76b, BF74a, Bur74, BL77b, Chi73a, CH74b, CG71, Cox72, Cra76c, Dri79, Fel74b, FEJM75, Gar73, GH79, Gla76a, GR77, GGM71, GC78, GK70, Gov71, Gov76, Gre71, Gup77, Hah70b, HN79, HM73, Hen72b, Hen80, HBA72, HH78, HG71b, Hol72a, JS73, KW74b, Kat78, Kra75, KV75, Lan76, LL72, Mar72, McC74b, McK76, Mil72, NB72, Pie73, PH74b, Poi74, Por74, Pra73a, Pre78a, Ran73, Rao73, Rev74a, RS74a, SC76, Sch74b, SW77, Sho72a, Sho73, Sim75a, SL74a, Spi79a, Sri73, Str73, Str75, TG79, Tho70, Tin70, WD72, WN74b, ZR74, Zel76b]. Cost [Koh72, MA79]. Cost-Effectiveness [Koh72]. Costs [Mer71]. Countable [Coh76c]. Counterexamples [JB73]. Counting [Bin76, GM77a, Nat76]. Country [DO72]. Counts [CR75b, KT77, Par79b]. Court [Gil74b]. Courts [Gil74b]. Covariance [Che70, CH77, DF72, Ham76, Hau76, Hem76, KS73, KN77, MD74, McE76, PAA70, Rao72, RY78, SW79, WSB73, You76a]. Covariance-Stationary [Hau76]. Covariates [Ger75]. Cox [Spi79a, Efr77, Poi78, Spi78]. Credit [CB70]. Crime [BMB78]. Criminal [GS73b]. Criteria [Cha76c, Ran71]. Criterion [Fuc79, Kar70, Lue71, NMC71, PJ70, Sah76, ZR71, ZR74]. Critical [GG76a, Gla76a, Gla76b, Rah74, San74a, SZ70, WN73, WN74b]. Critique [Bha71c, BW72, Gil74b, KZ72]. Cronholm [BJ70]. Cross [GK70, GK72, Goo79d, Hau76, McC76b, McL79, NS74, PD74, SM77b, You72b]. Cross-Classifications [Goo79d]. Cross-Correlation [Hau76, McL79]. Cross-Section [NS74, You72b]. Cross-Sectionally [SM77b]. Cross-Validation [McC76b, PD74]. Cube [The70]. Cubic [BL77c, Poi73, Poi74, BL77b]. Cumulative [Ano77f, Tih72]. Current [BW76, Chi72, Chi73a]. Current- [BW76]. Curtailed [Sam70]. Curve [AF76, GG78, Ham74, Kri79, Meh75, Sch73b]. Curves [Ang79, BH72, SS78b, Sub72, SV76, Zer79]. Cycles [Lee75].

D [Hoa74]. Daily [LMSW70]. Daniell [Sti73b]. Data [Agr77, Ang79, Bak74, Ber71b, Bha76, BM76a, CVM71, Che79, CR75a, Con72a, Cro79b, CH77, DD77, Dem73b, DO72, Edd79, EM75, Efr77, Efr78, FH79, Fra73b, FKB73, GK75, GS74b, Gla79, GOST75, Goo70, GKAH71, HW74, Har72, HG73, Has72, Haw74, Hem79, Hoc77, HO71, Hog75c, Hsi79, Kno76, KL72, Koc73, Kri79, Lan74, Lan76, Lea74, Lee70a, LM71, Lin71, Lin73a, LS75, Lin79a, Lit78, LB72, MSD73, Man74b, MB74a, ML74, MR79, Max73, McC72, McG73a, McG73b, McG77, Mil76, Min72, Mor71c, Mur72, NJM72, NS74, OW72, OK76, Par79c, Par79d, Pau78, PW74, PB76, Pra73b, Pro71, RW74, Roc75, Rub74b, SA73, San73, Sch73a, Sch78c, SS79, SMRL70, Sim77, SHH78, SZ73, SF72b, Tar79a, Ten70, TG75b, Tur74, TBH74, Ven78, WNIO77, Wei72]. Data [Wei78c, Wel79, Win74b, Wit74, WE79, CC74a, Tuk79a]. David [Fry70]. Death [Pre70]. Debate [SW73]. Debugging [FS77]. Decision [BT78, Boh79, FM74, GS75, HW74, San74a, Ven78, Win72a, Jay79]. Decision-Making [GS75]. Decision-Theoretic [FM74, Ven78, Win72a]. Decisions [HP70, Pre78c]. Decomposition [CT76, TS75]. Decrement [SL79]. Defective [GH74, Hwa72, Hwa74, KS71b]. Defense [Hig70].

Defining [FM74]. Definite [JS72, JS73]. Definitional [OS70]. Degree [Fro75, McD72]. Degrees [Sly72]. Delivered [Ano70m]. Demand [Gil74b, Lau70, Lee70a, Wes79, You72b]. Demands [Pra70]. Demographic [Raj77]. Densities [TC72]. Density [BBDS71, BBDS72, Coo71a, Coo73, Cra73, Gli72, Gol75b, KL76, Kri79, SS77a, SS78b]. Dependence [Ano77e, BBC79, BT72, DY78, Epp76, GR71, Sha77a]. Dependent [Ame73, Ano77e, DC71, Epp76, Kwe75, Lad75, NM70, Poi78]. Derived [Dix76, Gli72]. Design [Add70, Add72, BM73a, Chi73b, Dun71, Fie75b, GGE74, Kar70, KM74, KMS75, LK78, Mal70, McK75b, McK76, MTC76, OS70, PH72, Rob75, SW78b, SZ70, Sti71, Tsu72, Wei78a, Wyn77, Zac73, Zer79, PH74b]. Designed [Gav78, Kie78a, Mal78, SM78a, SM78b, Sim78b]. Designs [AS78, Ban72, BM77b, Bre79, Cle70b, Con74a, Cor76, CG74, DGL77, EHB78, EM78, GH78a, HB73a, HF70, KS75, Lee72, Lee73a, LW76c, MS76d, Mur71, MS72b, Rao71b, Rao73, SS79, Sud78, Tsa76, Wei77a, GH79]. Detect [Sch76]. Detecting [Ang74, BJ77, Hsu79, Hwa72, Mie72, Smi75b]. Detection [Haw74, MY78]. Determinants [Liu70]. Determination [Hor78, PD74]. Determining [UB70, CC74a]. Development [Ang74, OPRS77]. Developments [Add72, Car70, Mil77, Sen76a]. Deviation [AF76, DC73, Her71, WT71]. Deviations [Hah72, Sch73b]. Device [LCM75]. Diagnosis [She78]. Diagnostic [Pie72, Pie73]. Diagonal [Wag70]. Dialing [Wak78]. Dialogues [Bac79]. Diary [SF71]. Dichotomous [Can70, CA74, KK73, Kno76, San74b, Isa74]. Difference [Fre77, Gho75a, Gol75a, Lin71, Lin73a, RC76]. Differences [Ano78e, Gab78a, Ram78a, Ram78b, TG77]. Different [Nat76]. Differential [Dav77a, Roc75, WM70]. Digit [Wak78]. Dimension [Goo78d]. Dimensional [AS78, Che73b, Pat74, Scl79, WM77]. Dimensions [EM78]. Direct [Har72, Hoa77, Hol70b]. Directed [GM75]. Direction [AS77]. Directional [MR79]. Directly [May73]. Dirichlet [DR71, Jam72]. Disability [Ste71, Ste73, Wil73]. Disaggregated [Sla74]. Discarding [McC72]. Discharges [NS74]. Discontinuous [Con72b]. Discrete [And70a, CM73, Hil75, KG71, LC76, PW74, Som77, SK73, Wil78, Zac70]. Discriminant [BRH76, CD72a, CD74, CGD76, CA79, Efr75, FV73, Han78, Hig70, Hor78, Lit78, MD74, McL75, McL77, PW78, RBRH78a, RBRH78b, SU74a]. Discrimination [Dye73, GG72, Krz75, Krz76, Moo73b, O’N78, SB71]. Discriminatory [Hol70c, Hol71b, Wil71]. Discussion [GOST75]. Disease [BBC79, KJT72]. Dispersion [Lem78, SH79, SZ73]. Disproportionate [HG73]. Disrete [Gok73]. Dissolved [GCZH73]. Distant [Rea75]. Distinguishing [SS79]. Distorted [Sob71]. Distortion [LSB76]. Distributed [Arn79, Bin76, Bro78d, CM74a, Car75, DF75, DuM75, FKB73, HB77, LR72, Mee74b, Mer73, NS74, Pes72, Pes76, Sch74a, SM77a, Sim72b, Wyz78]. Distribution [AE72, And71, And74a, And76, Arn70b, BBS74, BD73, BT74, BBH72, BS77a, BJ73, BP73, BS73b, BP70, BRH76, Bul71, BDH74, Bur73, Bur74, CT73, Can74, Cha70a, CCM71, Cha71c, CG78, CF74, Chi74, Chi76, CH75, CH74a, CH74b, CM72, Coh72, Coh76b, DC73, DR71, DE73, DF79b, DQ72, Ell73, Ell77, Eng75, Fel76, Fer78, Fle71, FK76, Fli79, GG78, Gav78, Geb71, Gho75a, GG71, Gla76c, Gle72, Gle76, GH76a, GGM70, GC74, HB70, Hah70a, Hah72, Hal75, HAB71, HM73, Hen80, Her71, HURD72, HFR75, Hol70a, Jam72, JS72, JS73, JK72, Jol73, Jon79, Kap77, Kie78a, KT77, Kno70, Kra73, Kra74, Kra75, Kre76, Kri72, KV73, KV75, Lad75, Lai78, Law73, Lee70b, LW76b, Lee77b, LL71, LR72, LS79, Mal78, Mar76, MS72a, MM79]. Distribution [MJ79, McL79, MS73, MS74, Mor71b, Mor71c, MW70, Noe73, Off72, Pat75b, PJ70, PH72, PH74b, Pol73b, PS76c, RR76, RR71, RS72, Rob77, Rob72, Roh76, SM78a, SM78b, Sel74, SB77, Sic75, Sim78b, SW78b, SS77b, Som76, Sri74, Ste72, Ste78a, Ste76b, Ste78b, TW77, Tei71, Tha79b, TG71b, Tih72, Upt76, Van76, WS78, Wei75, Whi79b, Wil78, WEB78b, WE79, Yar70, You72a, ZF73, Zel71, Fis79, Poo74, Arn73, GGM71, GC78, Hah70b, LL72, Tho70]. Distribution-Free [BRH76, FK76, Fli79, HFR75, Hol70a, Jon79, Kre76, LW76b, Noe73, PH72, PH74b, RR76, RS72, SW78b]. Distributional [Lau70]. Distributions [AR71, Bin78a, Bof75, BS75, Bry78, Che71a, Cho77, CH78c, Con72b, Coz74, Cra76b, Dav79b, Dic76a, DQ72, Dri78, DuM73, DuM75, Edw75, FR71, FS72, Fow78, Fow79, Gal75a, Gok73, GGM70, GGM71, GC78, Gre76, Gro76, HS74, Har78, Hen79, HG70, Hog75a, Hog75b, Hos78, HF70, Ifr70, JS75, Jam75, Joh78c, JW78, JR71, Kau72, Kie78b, KNP79, Lau74, Lit74b, Mar76, Meh72, Mis72, Mos70a, Nau73, Pha71, Pre72, QR78a, QR78b, QR79, Rob77, Sam76a, San73, Sch73d, SS79, Sha77a, Sha70b, Ski70, Smi75b, TL70, WT71, Win75d, Zac70, ZM71, Zac73, Hil70b, JTB80, JTB81, Dri79, HG71b, Sch74b]. District [Gil74b]. Disturbance [Mis72, Pie72, Pie73, Sri72c, Sri73]. Disturbance-Variance [Sri72c, Sri73]. Disturbances [AK71, HB74, Kad70, KG70, Mor78, Sri72c, Sri73]. Dividend [Lau72a]. Dividend-Earnings [Lau72a]. Divisible [Cle70b]. Do [CH78b]. Does [Bly71, Bro71a, Dan71, Efr71a, Efr71b, Goo71a, Pra71, Stu71, Pra73a]. Doing [DeG73]. Dollar [BW76]. Domain [CR72, ST76a]. Domains [TT72]. Dorfman [PE78]. Dorfman-Type [PE78]. Dosage [EZ73]. Double [Ano77d, Che79, CWR78, Gov71, Gov76, Hal75, Han73, HBA72, Hoc77, Kap77, Ram71, SH75, Ten70, ZG72]. Doubly [Bul71, HM73, Tur74]. Down [Lit74a, Lit74b, Lit75]. Drug [GG75, TY74]. Dual [BB72]. Due [WM70]. Duels [FK70]. Duration [Klo76]. Durbin [Bir70, Ken76, Par75, Par76]. Dyad [LW76a]. Dynamic [DH72, HB77, Lee70a, Pie72, Pie73]. Dynamic- [Pie73]. Dynamic-Disturbance [Pie72]. Dynamics [DF74, McC71, McC74b].

Early [HW74]. Earnings [Bor70b, HP74, Lau72a, VG78]. Earnings-Schooling [HP74]. Easier [Mar76]. Econometric [BK79, Chr79, DO72, Kel71, MM73, Mik72b, Mik75, RO76, Rob79, Rot79, ZP79, Zel79a, Zel79b]. Econometrics [Kle71]. Economic [AG72a, BT75, Chi73b, CL76, Dor74, Dun71, Pie77, SF72b, UB70, WG71]. Economic-Linear [AG72a]. Economics [Mar71, Mar72]. Economies [Vin76b]. Economy [BW76]. ED50 [Ham79]. EDF [GH76b, Ste74]. Edgewood [CH74b]. Edgeworth [CH74a]. Edit [FH76]. Editing [NJM72]. edition [Kle77a]. Education [BN73]. Educational [Ban71]. Effect [Agr76b, AW72, Bro78d, CM74b, CR75a, Fab77, GR71, GS77, GP74, Joh73a, McK75c, MMS71, Rub77, SMRL70, Sti76, Tho77, Wyz78]. Effectiveness [BL79, Koh72, McG77]. Effects [Bai75, BT74, Coc70b, DO72, Gla79, Gol75a, Lau70, Lau72b, Lem77, Mal70, McC72, Qua79b, Sah76, Sla73, SA70, SB73, Vin76c]. Efficacies [Sim78a]. Efficiencies [AGA79, KvE72, LMH73, MS76b]. Efficiency [Bal70, Baw72, DS75, Dud71, Efr75, Efr77, Hay71a, Her76, KP79, LP70, Mik71, TT71]. Efficiency-Robust [Mik71]. Efficient [Ham78, Jon79, Kat77, Kat78, KvE72, Tak71b]. EFTA [Car70]. Eight [BMB78, Sti77b]. Elasticities [Wei77c]. Elections [Qua74]. Elementary [Kle77a, Kle77a]. Elicitation [HB72, Sav71]. Elimination [Bas77]. Emanuel [Tuk79a]. Emergence [Sha76a]. Empirical [AB78, BD73, BR70, BJN70, Bor77, CR74, Dun74a, EM71, EM72, EM73, FS77, GY70, HP74, KL76, Lev75b, Mur72, Nat76, Pet77a, Tur70, WG71]. Employment [Hig70]. Endogenous [Kel71, MS72a, McD74, RR71]. Enforcement [WD76]. Entry [DL72, Har73]. Enumeration [KT77]. Environmental [BT75, WD76]. Epidemic [Kel73]. Epidemiologic [KMS75]. Epidemiological [Wit74]. Equality [BS77a, BF74a, BF74b, DC71, Jen70, LG75, LS75, PL76, PAA70, Vit79]. Equation [And74a, And76, Atk78, Dav77a, Hur72, Kin72b]. Equations [Ano72e, Bar76b, DM76a, Fou77, Gre75c, Rev74b, Sri72c, Sri73, TG75a, TG79, Rev74a]. Equipment [FT70]. Equiprobable [GGM70, GGM71, GC78]. Equitable [Sub72]. Equivalence [Bra73, CFY76, CC73, Har75, PS74]. Equivalent [Bur79]. Ergodicity [Coh76c]. Erlang [CG73]. Error [Bai79, BK78, Bor70b, BM76b, CR75a, DS74, DS73, DF72, FNL77, FB73, GR74, GM77b, HG71a, HN77, HN79, Koc73, KM74, Lit74a, McE76, Min72, MM71b, Pra70, RZ71, RG72a, RS76, Reg73, Rob77, Sah76, Sub72, SM73, TS77, Thu79, War72, You72a, Zel75, Zel76a, Zel76b]. Error-Areas [Sub72]. Errors [Arn79, Atw75, Bai79, Bak74, BW70, BHF77, Car79, Cha71a, Coc70b, Fel74a, GG76b, GOST75, GS73c, Haw74, Hoc77, Hur72, KW74b, KW74a, KF70, McC72, Mur72, Nat76, NJM72, Pot72, RW70, RW77, Spi79b, Str72b, Str72c, Str73, TF73, Vin76c, Wan71, WWF74, ZM71]. Errors-In-Variables [WWF74]. Establishing [BK76]. Estate [Bha71c]. Estimate [And74a, And76, Arn70b, Arn73, BN74, Bra73, Gla79, Gol75b, KQ72, Koo73, Maz72, SD73a, SD73b, Wag70, WK71, Woo71, WC76, CC74a]. Estimated [Cha72, GF73, Smi76]. Estimates [Bai75, BM78, Bic75, BM76a, CM74a, CCM71, CFY76, CD72b, CC73, Coh76a, Dix76, FR71, FH79, Fer78, GH78b, Gov71, Gov76, Gro77, Ham79, Hay71a, Hoa74, Hog75c, KS75, KvE72, Lit74a, LS77b, Lon73, MSW74, Mor71c, Nat76, Ram77, RBRH78b, SS71, SP76, Sly72, Str74, Str75, SL74c, TF73]. Estimating [Atw71, Atw75, Ber74, Bin78a, Bin76, BM75, BDG78, Bor77, Bry78, Car79, CR74, CH78c, DG73a, Eri74, Fel74c, FNL77, Fis72, Fow78, GH76a, Han78, Har78, HG73, Hin79, HHD75, Hos78, Joh78c, KL76, KK75, Kie78b, KNP79, KMS75, Lev77, MJ78, May73, MSW74, McC78c, McL77, McN73, Meh72, Mic72, MS72b, NS74, PS76c, Qua72, QR78a, QR78b, QR79, Raj77, RV77, RC74, Sag79, San77a, San73, SL79, SW78a, SJL76, SW78c, TC72, Ten70, TT71, Tho70, Vij75, Wri78]. Estimation [AG70, Alv78, Ang79, BBS74, Bar76b, Bar73, BBH72, Bha73b, BJ74, Bha77, BD78, Bli74, BM70b, Bro76c, BS74b, Bro75, BM76b, Bro78c, BS79, BG70, Can74, Car77, CVM71, CM73, Cha70a, CD76, Che70, CF74, CL76, CZ75, Coo71a, Coo73, Cow75, Cox71, Cox72, Cra73, Cra76a, Cra76b, CR75c, DL76, Dah74, DF72, DM79, DH77, DH72, EM73, Eng75, EG71, ES74, EM78, Fas71, Fie70, FH73, FB73, GH78b, Goo71b, Gor73b, Gri72, GS73c, GM77b, GJ75, Haf78, HG71a, Ham78, Ham74, HN77, HN79, HM73, HJ73, Har77c, Har77d, Har79, HO71, HURD72, HST79, Hur72, Ifr70, Joh74, JG72b, Jon76, JG75, Kar70, KG70, KV73, KM74, LaM73, Laa79, Lai78, Lan74, Lan76, Lee70a, Lee72, Lee73a, Lee73b, LP75, Leo77, Lev73, LK78, Lie76, Lin71, Mal72, MR79]. Estimation [MY76, McC73a, MS70, MS71, Muk76, OW72, OL71, Odo70, Ord75, Par71, Poo74, Pre72, Rao70, Rao72, Rao77a, RS72, Rob77, RW77, RAK74, Ros72, RH73a, RH73b, RC78, Rub74b, Sam76a, SB78a, Sax71, Sax76, SGO71, SBH78, Sed77, Sel74, Sha70a, SH73, Sim72b, SU74b, SL74a, SL74b, ST76b, Sol72a, Sol72b, Sri72c, Sri73, Sti73b, Sti76, SV76, SM77b, Tar79b, Tha79b, TG75b, TT72, Tur74, Vea75, Wei78d, Win72a, Wol79b, YS77, Zac73, DL74, FH78, FH81, Pri74, Sen74, AG72b, Ano79f, Bro76b, KV75]. Estimato [Rao73]. Estimator [AE72, Ame76, CM74b, Cha71a, CC74b, DS74, DM74, DP72, EM75, Far76, Fel74a, KW74b, KW74a, MS72a, MA79, MS76b, Mik72a, Mor78, Pet77a, PS75, Rao71b, Rao75, Sri71b, Tak71b, Wes75, YBJ72]. Estimators [AK71, Atk78, Bar77, BBDS71, BBDS72, BR70, BW70, BM70a, BLM71, Blu74, BYJ73, Bro77, Bue77, Car79, CZ77, DC73, DL77, Dem77, DF79b, DQ72, DM76b, EM71, EM72, Ekb73, Fro73, GC70, Gli72, pH78, Har77a, Hoa75, HH75, JRH73, JP74, JK72, Kem77, Kie77a, Kie77b, Kle73a, Koc73, LaM77, LMH73, Lin77a, MM73, McD74, MG75, McK75c, Mer73, Mik71, Mik72b, Mik75, Mis72, MW70, NL77, RR77, RW70, Sah76, Sam78, SS77a, ST76a, Spi79b, SZ73, Sta75, Str78, TG71b, TL70, Van76, Wei71b, Wei73, Wol77b, ZM71, ZZ72, ZP79, Ano72e, Tin70]. Etimating [Fel74b]. Evaluating [Sly72]. Evaluation [CS73, GM77b, GY70, Hub74b, KTS75, MB74a, MG75, Mil76, Moo73b, Ran71, WNIO77, Yea73]. Events [Bly73, Raj77]. Everyday [Box79]. Evidence [Bir72, Fin77b, HP74, MS76b, SC75b, SC77, Tur70, Dia78]. Evolution [Dob78]. Exact [Ano72e, Ano78e, BBDS71, BS78, CG78, Gla76b, GGM70, Hen79, Hen80, HF70, KS75, KT77, KN77, LM78, Lar78, MS72a, McD74, MS74, Pat75b, PJ70, Sim75b, TG77, ZR71, BBDS72, Gla76a, GGM71, GC78, ZR74]. Exactly [BBDS71, BBDS72, MM79]. Exactness [KM79]. Examination [Mur72]. Examining [Gar77, MM71b]. Examples [JB73]. Excess [Ste71, Ste73, Wil73]. Exchangeably [Arn79]. Existence [JB73, McD74]. Exit [DL72]. Expansion [And74a, HAB71, JS75, Som76, Wal79a]. Expansions [Cra73, Cra76b, GG76c]. Expectancy [Lon73]. Expectation [Ame73, Maz72, Yar70]. Expectations [Car75, Cho77, Hom78, Mor71c, Pes76, Sav71, Tur70, ZR71, ZR74]. Expected [Blu74, Cha76c, Goo71b, Hil74b, TH73, Vea75, ZP79]. Expenditure [Adi73, Ham78, Par71]. Expenditures [Lee70b, SF71]. Experiment [Baw72, CD72a, Jol73, Tsu72]. Experimental [Add70, BNW70, BM73a, EM78, GGE74, KM74, Mal70, NB71, NB72, NBH73, Pet78a, Sti71, Win71]. Experimentation [Bra79a, Bra79b, CH79, DS79, Ell79, Flu79, Gab79, Kru79, Mie79, Mos79, Ney79, Sim79c]. Experiments [Add70, Add72, Baw72, Bea74, BBC79, Coc73, Dav70, Gav78, Ger77a, IJK76, Kie78a, Mal78, MS70, PH73, SM78a, SM78b, Sim78b, SF71]. Explaining [Lea75]. Explicit [Gor73b]. Exploration [UB70]. Exponential [And70a, Ano77d, Ano79f, Arn70b, Arn73, BM78, BBH72, BJ73, BB74, BM75, Bra73, BH72, CFY76, Cho77, CWR78, Cra76a, FM74, GG78, Gov71, Gov76, Kap77, LR71, MG74, Man74b, PS76c, Ram71, Rob77, Sha70b, Sim73, Tha79b, Tho76, WEB78b, YS77]. Exponentiality [Sho72b]. Expressing [Pet77a]. Expressions [Man74b]. Extended [DS73, Zer79]. Extending [Dav70]. Extension [BB74, Ger75, LL71, LL72, McD72, MI78, SS73a]. Extensions [GH78a, GH79, MG70]. Extent [KJT72]. Extinction [CR75c]. Extraneous [Kle73c]. Extreme [HM73, Lit74b, Wes75, dO74].

F [Ano79t, Hoa74]. Face [Sud78]. Face-to-Face [Sud78]. Faces [Che73b, CR75a]. Factor [Baw72, Dic77b, Goo75b, GW79, HP70, NB71, NB72, NBH73, She78, SA70, Wil78]. Factor-Analytic [Wil78]. Factorial [Add72, MS70, MTC76, PH73, SA70, TA78]. Factors [Hah70a, Lem77, Man74d, MTC76, SB73, ZZ72, Hah70b]. Failure [AM73, Bas71, BM75, CS74, KR78a, PR75b, Sri75]. Fair [KK79, Tho75]. Fairly [RR77]. False [Lea74]. Families [And70a, Dye73, Dye74, KN75, PR75b, Pat75a, Zac70]. Family [Bha71b, Bra73, CFY76, CWR78, Hua73, Lee70a, Meh72, Ram71, Sau74, Sha77a, Sim73, JTB80, JTB81, Ano77d]. Fares [Jag73]. Farm [Bha71c]. Faro [Tho73b]. Father [BN73]. Faulkenberry [Dun76b]. Favorable [Bly73, Bof76]. Features [CR75a]. Federal [Gil74b, Lee70b]. Feedback [Pre78c]. Female [CM72]. Fertility [Lee74, Lee75]. Fiducial [BP73]. Field [Bra79a, Bra79b, CH79, DS79, Ell79, Flu79, Gab79, Kru79, Mie79, Mos79, Ney79, Sim79c]. Finding [AR71, Hil79, Hwa74, Koo73, KS71b, PS76a, SD73a, SD73b]. Fingerprint [OPRS77, Scl79]. Finite [Ano72e, BBDS71, BBDS72, Bur75, Car78, Fel74a, FM74, Fox74, Hil79, Kle73a, Laa79, MS72a, MS74, Raj75, RG72b, Rao75, Rev74a, Rev74b, Rev76, RH73a, RH73b, RC78, SBH78, Sha76b, SD75, SZ70, Sri71b, Wil78, Wol79b, YBJ72, You76b]. Finite-Sample [MS72a]. Fire [CR74, Haf78]. First [CM72, DM79, HM73, Smi76, SL74c]. First-Order [DM79]. Fiscal [SW73]. Fisher [Gho75b, LF71, LF73, SC75a, SC76, Sch70a, Som77, Wan71]. Fit [BH76, BR77, Con72b, GG78, GH76b, HT72, HG70, HG71b, KP79, Lar78, MSD73, Moo77a, Pyn79, RA75, SM79, Sch73d, Sch74b, Ste74, TH73, TP79, Tho79, Yar70]. Fitting [AF76, GF73, GK76, Jon75, Sch73b, War72]. Five [Moo73b]. Fix [Fry70]. Fixed [BNW70, Ber74, Car77, CD76, CM76, GM77a, NL79, PS74, SW76, SW77, Zac73, Wor74]. Fixed-Sample-Size [CM76]. Fixed-State [NL79]. Fixed-Width [Car77, CD76, SW76, SW77]. Flatcar [DS78]. Flow [Hsu79]. Flowing [GCZH73]. Follow [Bly71, Bro71a, Dan71, Efr71a, Efr71b, Goo71a, Pra71, Pra73a, SBY79, Stu71]. Follow-up [SBY79]. Food [BNW70]. Football [Har77b, Mos70b, Pol73b]. Forecasting [DS78, DWD76, Lee74, LMSW70, Lon70b, Plo79, RW74, Sab77, SS73b]. Forecasts [BW76, Nel70, WG71]. Form [Con74a, Goo78d, HP74, Hil74c, JS72, KL76, JS73]. Formalizing [Rub77]. Formation [LW76a, Tur70]. Forming [Pre78c]. Forms [GG76c, Gle72, HW75, KTS75, RZ71, Wal79a]. Formula [OPRS77, SC75a, SC76]. Formulation [DF74]. Forthcoming [Ano71e, Ano71f, Ano72f, Ano72g, Ano72h, Ano72i, Ano73e, Ano74c, Ano74d]. Forward [BA77, CA79]. Four [AS78, SZ73]. Four-Dimensional [AS78]. Fourfold [Gol75a]. Fractals [Goo78d]. Fractional [SA70, Sti77c]. FRB [Bro76a, Yea73]. Free [BRH76, FK76, Fli79, GG78, GCZH73, HFR75, Hol70a, Jon79, Kre76, LW76b, Noe73, PH72, PH74b, RR76, RS72, SW78b]. Free-Flowing [GCZH73]. Freedom [McD72]. Frequencies [Cha76c, Goo71b, Sic75]. Frequency [CR72, CM72, Hil74c]. Frequenies [TH73]. Friedman [Ger75, Hol70b]. Front [Ano70e, Ano70f, Ano70g, Ano70h, Ano71g, Ano71h, Ano71i, Ano71j, Ano72j, Ano72k, Ano72l, Ano72m, Ano73f, Ano73g, Ano73h, Ano73i, Ano74e, Ano74f, Ano74g, Ano74h, Ano75e, Ano76d, Ano76e, Ano76f, Ano76g, Ano79i]. Full [DM76a, HS75]. Function [AM73, Bro76b, Bro76c, Can75, CF74, CZ77, DQ72, Efr77, Han78, Hil75, HT79, Hol72b, Joh73a, Kau72, Lev73, MJ79, McL77, Mic72, Muk76, Pet77a, RZ71, San74a, Tik71, Tur74, Poo74]. Functional [HP74, KL76, RZ71]. Functions [Ano77d, BBDS71, BBDS72, Bro75, BS75, CM73, Coc73, CWR78, Cra70, DGL77, Ger77a, GG71, GK76, Hob76, HB73b, Kri79, KV73, KV75, Lan73, MM79, MD74, Mor71b, NL79, OK76, Pat75a, Pet77a, Ram71, RBRH78b, Rob77, SS78b, Sti76, Wil78, Wri78, ZP79, SU74a]. Fundamental [Iji71, LS77a]. Further [Dag71, MS76b, RS76]. Future [FW79, Gas74, Hah72, Har74a, Hog74a, Hog74b, Hub74a, Key72]. Futures [CR72].

Gains [AKC76, Bha73b, Bha76]. Game [Ano79e, Sen76a, SY72, Tho73b]. Gamma [GG71, Gla76c, HB70, Hsu79, Joh75a, MJ79, Mos70a, Pha71, Sho72b, WE79]. Gaussian [BH08, JS73, BB76a, CF74, FM74, FM77, JS72, TS75]. GCL [BBDS72, BBDS71]. Geisser [Sch78c]. General [Ame76, BM76a, CMH73, Ell73, KN77, LaM77, MG70, Nea75, O’B79, SL79, Sho70, Sho72a, Sho73, SF72b, TB77b, Wol73]. General-Acute [Sho70, Sho72a, Sho73]. Generalization [Sim77, WT76]. Generalizations [EM75]. Generalized [Ban72, CFY76, Cra75, HB70, HS74, Hwa75, Jam72, Joh73a, Jon76, Moo77a, Mos70a, Nau74, Pau78, RBRH78b, Sap72, Sha76b, Sri71b, Ste78b, Str78]. Generated [SL79]. generating [Rel72]. Generation [KR76, KR90]. generator [Che71b]. Generators [Nea72]. Geographic [Lon70a]. Geographical [DWW79, Tob79a, Tob79b]. Geography [Ber71b]. Geometric [Cho77, Gla76c, Klo76, Sha70b, Sri74]. Geometry [FG70]. Gilbert [Kle77a]. Gini [Kri79, Meh75]. Given [Hal75, Tho75]. Goddness [GG78]. Goddness-of-Fit [GG78]. Goodness [BH76, BR77, Con72b, GH76b, HT72, KP79, Lar78, MSD73, Pyn79, RA75, SM79, Sch73d, Sch74b, Ste74, TH73, TP79, Tho79, Yar70]. Goodness-of-Fit [BH76, BR77, Con72b, GH76b, HT72, KP79, Lar78, MSD73, Pyn79, SM79, Sch73d, Sch74b, TP79, Tho79]. Goods [BM73b]. Gradient [KL76]. Graph [BH76, HD78]. Graph-Theoretic [BH76]. Graphical [Par79b]. Graphically [Che73b]. Graphs [LK76, Roh74]. Greenhouse [Sch78c]. Group [Bai75, Cle70b, GH74, GR75, GR77, GR74, Gro72, Han78, Hwa72, Hwa74, Hwa75, KS71b, PE78, Pre78c]. Group-Testing [GR74, KS71b]. Grouped [AG70, AG72b, Bof70, CVM71, GK75, Kri79, MSD73, OW72, San73, Sri75, WB71]. Grouping [Bak74, CM74b, Con72a, RW70, ST76b]. Groups [BM77b, Krz79, Krz81]. Growth [Gre75c, Klo73, Zer79]. Guessing [PY74]. Guided [Goo73]. Guilt [SC75b]. Gupta [RA76].

H [Hoa74, Par75]. Haar [Cra76a]. Hacking [Sha76a]. Half [Lee72, Lee73a, McC76b]. Half-Sample [McC76b]. Hampel [Hoa74]. Handling [Con73]. Having [DL72, GK76, Som75, WG72, Goo79d]. Headship [Kob73]. Heart [CH77, MB74a, TBH74]. Heart-Transplant [MB74a]. Heavy [Smi75b]. Heavy-Tailed [Smi75b]. Heroin [BS74b]. Heterogeneity [Sti76]. Heteroscedastic [HH75, HHD75, Rao70, BW70]. Heteroscedasticity [HM79a, Lev73]. Heteroskedastic [Mal72]. Heywood [Eck70]. Hierarchical [Bak74, BH75, BH76, Hub74b]. High [Har77b]. Hildreth [Gri72]. Hill [RW74]. Historical [BH79, Bro76a]. History [Sti73b]. Hoel [PS74]. Hog [LMSW70]. Holdings [Bha73b]. Home [RW77]. Homogeneity [CG78, Gla76a, Gla76b, Lay73, Mee74b, Sam76b, THP78, Wis72].

Homoscedasticity [HR70]. horizon [Fox74]. Horvitz [Rao73, Rao71b]. Hospit [Sho73]. Hospital [NS74]. Hospitals [Sho70, Sho72a]. Hotelling [EE70, Eve79, GP74]. Houck [Gri72]. Household [CH78a, Kob73, Sir70, SW78c]. Houses [Bha71b]. Housing [GH78b, Hua73, KQ70, KQ72]. Huber [Hoa74, Bic75, Tha79b]. Huber-Sense [Tha79b]. Human [FGHA73, Sud72]. Hybrid [KMS75]. Hydrologic [Yak76]. Hypergeometric [Ski70]. Hypotheses [Bur79, Cha74b, Dye73, Dye74, HD77, Hem79, PF78, Pro71, She78, SH76]. Hypothesis [Arn70a, BBH72, Bet71, BR77, Ceb74, Coh74a, Dye74, Har71, LP70, MG70, RY78, TP79, Vit79, Yak76]. Hypothesized [AB78].

Ian [Sha76a]. Ideal [Wei70]. Identifiable [BBDS71, BBDS72]. Identification [CR75b, HB77]. Identified [MM79, Vea75]. Identifying [NJM72]. IFR [BS75, Pat76]. Ignorance [Vil77b]. II [EM72, LF73, ML74, PS76b, RH73b, Sri72b, Wei73]. III [GK70]. Illustration [MB74a]. Imaginary [Jol73]. Impact [BMB78]. Imperfect [Enn75]. Implementation [TS75]. Implied [Roh74]. Imposed [RB71]. Improved [BS73a, Coh72, Fel74b, Fel74c, Gam77, GH78a, GH79, Kra73, Nor76, Vea75, Kra75]. Improvement [Hem74, SS77a]. Improving [Mor78]. Imputation [FH76, Roc75]. In- [Sto71]. Inaccuracy [TH73]. Inadmissibility [Arn70b, Arn73]. Incentives [CH78a]. Incidence [BB76a, BH08, Pre70]. Include [SS73a]. Included [MS72a, McD74, RR71]. Income [Adi73, Bet71, Bha76, Cow75, Fas71, FH79, Hol70b, Hsi79, McC71, McC74b, MJ79, Mur70, SW78c, TF73]. Income-Expenditure [Adi73]. Incomes [TF73]. Incomplete [AS78, BDG78, BB76b, Coz74, Dag71, Fie70, Geb71, HS74, HO71, IJK76, KS75, Lin73a, LS75, Lit78, LB72, Mee74a, MS73, Rub74b, Sch78c, SZ73, SV76, Tar79b]. Incomplete-Data [Rub74b]. Incompletely [Bro75]. Inconsistency [Bar76a, BT76, Dic76b, Fra76d, Lin76a, Pra76, Sto76a, Sto76b, Vil76]. Incorporates [MA79]. Incorporating [Arn70a]. Incorrect [Cle70b]. Increases [War78a]. Increasing [Wri78]. Increment [SL79]. Increment-Decrement [SL79]. Independence [BOLK75, BJ73, BJN70, CA77, DY78, Dri78, Dri79, Fie70, GJ78, Goo75a, Hau76, Hol71a, Hol72a, Jam77, Nat72, Wal70]. Independent [BM78, BW70, CZ75, Dav77a, Fle71, HS73, HR70, How74a, Jam72, Kle73c, KP78, LR71, LF71, LF72, LF73, Mee74b, MS76c, Mor71b, RH71]. Independently [Cha72]. Index [AB78, Ano77f, Ano79a, Ano79s, Ano79t, Ano79u, Ano79v, Bha71b, Bro76a, Kri79, Meh75, SS73b, Yea73, YFS71]. Indexes [BNW70]. Indication [SZ70]. Indicators [JG75]. Indices [Ang74]. Individual [Bha73b, OPRS77, PAA70]. Individuals [Bin76]. Industries [BM73b, Hig70, VG78]. Industry [AG72a, Wil71, Hol70c, Hol71b]. Inequalities [BS75, KS73, TF73, Ton70]. Inequality [Bha76, Dav78, KW74d, Lie76, LP70, MJ79, Mur70]. Inference [And70c, Ano78f, BP73, BG74, Chr73, DD77, DuM73, DuM75, DL72, Efr76, Fra76b, Fra76c, Hin71, HR77, Hog76, TL70, Vil77a, You76a, Jay79]. Inferences [Wei72, Wes77, WEB78b]. Inferential [HB70, WE79]. Infinite [GO72]. Influence [Ham74, Whi79b]. Influential [Coo79]. Information [And74a, And76, Ano70r, Ano71o, Ano72u, Ano73o, Ano74o, Ano75k, Ano76p, Ano77l, Ano78n, Ano79w, BM77b, Bor70b, DS78, DM76a, DuM75, Enn75, Han73, KR78a, MS72a, Mar71, Mar72, Meh72, Mor78, Por73, Sha70a, Sim73, Sri71b, Ste76a, Tar79b, War75, Zec71, Por74]. Initial [Blo77b, McL77]. Inliars [McC72]. Inner [Vil77a]. Innovation [Jon76]. Input [BW76, WM71]. Input-Output [BW76, WM71]. Inspection [Min72, Sti77b]. Instability [Hua73]. instrumental [DL74]. Insurance [Hol70c, Hol71b, Wil71]. Integer [GK76, Sta75]. Integer-Valued [Sta75]. Integral [Geb71]. Integrated [BP70, Sab77]. Integration [MR70]. Intensity [CZ77]. Inter [CG73, KS75, Sha70a]. Inter- [KS75, Sha70a]. Inter-Purchase [CG73]. Interaction [BN74, Goo75b, JG72a, JG72b, Odo70, Ord75, PH73, Pat74]. Interactions [BG74, Goo70, MS70, SA70]. Interarea [Mur70]. Interblock [BM77b]. Intercept [Ram77]. Intercepts [HQ75]. Interdependent [EP70]. Interest [Nel70, Pie77]. Intermediate [HR77, Ano78f]. International [Adi73]. Interpolation [Cla70c, DWW79, Tob79a, Tob79b]. Interpolatory [GG71]. Interpretation [BM77a]. Interpreting [Bro76a, DeG73]. Interpurchase [BB76a, BH08]. Interrelated [Ant72]. Intersection [SW79]. Interstate [Ceb74]. Interval [AST73, Bas73, Car77, CD76, Cox71, Dix76, FM77, Gro76, Lan74, Lan76, MP71, NR71, RS72, Sax71, SW76, SW77, SMRL70, SL74c, TG75b, Win72a, Zac73, Cox72]. Intervals [BB76b, BS78, Car78, Coh72, DD75, Dud72, Dun76b, Far76, Fau73, FW79, Gho79, GW79, Gro77, Hah70a, Hah72, Har71, HB73b, Kap77, Kre76, Law73, LR71, MG74, Noe73, RR76, WN74a, WT71, Joh78b]. Intervention [BL79, BT75]. Interview [Lau72b, LSB76]. Interviewer [BMB78]. Interviewing [Sud78]. Intial [MSW74]. Intra [KS75, Por73, Por74, Sha70a]. Intra-Block [KS75, Por73, Por74, Sha70a]. Intraclass [Haf78]. Intrinsic [BM76a, Key71]. Introduction [Ano79j]. Invariant [BOLK75]. Inventory [ZF73]. Inverse [Ban72, BB76a, BH08, CF74, Hoa70, Hoe72, PS74, RY78, SW71]. Inverses [Moo77a]. Investigation [CB71, Eve79, Her76, KL76, LSB76, Roc75, RB71, Spi78, Spi79a]. Investment [Bha71c, HP70, Kim70]. Involving [HM70, Kau72, Koc73, MB74a, RS74b]. Isolating [GH74]. Isopleth [Sag79]. Isotonic [BB72, CR75c]. Items [Wei70, Zin72]. Iterate [DP72]. Iterated [Bar76b]. Iterative [Bjö78, DL74, Fle75, Hem74, Hem76, MM71a, McL75, RC74, Sag79, Sch73b]. IV [GK72, Woo76].

J [CMS87, FH81, Hoa74, JTB81, Krz81]. Jackknife [Gla79, Sha76b]. Jamaica [BH79]. James [FH79]. Jardine [Roh74]. Jeopardy [LW76c]. Job [OS70]. Johnson [Kle77a, SB75]. Join [GF73]. Joined [DGL77]. Joint [Ell73, Ell77]. Journal [BH08]. Judgments [Pre78c]. Jury [BW72, Fab77, GS73b, GS74a, GS75, GS77, KK79, KZ72].

Kaplan [Pet77a]. Karber [CC73]. Karhunen [BH79]. Kendall [Kra76, Rob72, Sim77, THP78, Wes75]. Kernel [SS77a]. Keynesian [Hol72b]. Knockout [CH78b]. Knowledge [Mee74a]. Known [DE73, Dol76, Fer75, GH76a, Pot72, Sam78, SJL76, Str72b, Str72c, Str73, Str74, Str75, Uth70, War72, Zac73]. Kolmogorov [WN74b, Can75, Con72b, GG76a, Kno70, Wei78c, WN73]. Koop [Koo73, SD73a, SD73b]. Kruskal [KT77]. Kurtosis [Ali74, DL77, Hog72b, SB75].

L.A.M.S.T. [BN74]. Labor [Hol72b, Man74d]. Lack [Pie77]. Lag [Bro78d, CM74a, Fro75, GP75, HB77, Hin79, KK75, LR72, Mer73, Nea72, NS74, Pes72, Pes76, SW73, Sch74a, SM77a, Sim72b, Wyz78]. Lagrange [JS75]. Lags [KK75]. Lambda [JR71]. Languages [Dob78]. Large [Bre79, Ebb73b, EHB78, Het73, Jam77, Lay73, Liu70, Som75, UB70, DL74, Rel72]. Large-Sample [Ebb73b, EHB78, Lay73]. Large-Scale [Bre79]. Larger [Wei78d]. Largest [CD76, Dud71, Ofo74, RS72, RS73b, RS74a, Sax71, Sax76, Som70, Som74, WN74a, Wal70, Wei78d]. Latent [Dun75b, JG75, Mut79, SB78a]. Latest [KR78a]. Lattice [Cor76, Jes73]. Law [AG70, AG72b, Bly71, Bro71a, Dan71, Efr71a, Efr71b, Goo71a, Hil74c, HW75, LP75, Pra71, Pra73a, Sic75, Stu71, The70, Hil70b]. Laws [CZ75, Sim73]. Layout [Ell77, Hoc76, KT77]. Leading [BBDS71, BBDS72, SS73b]. Learning [Est72, Ste78a]. Least [All77, Ame76, AS77, Ano79h, Bal70, BK78, Bea77, BL77a, Bjö78, Bof76, Bra73, BL77b, BL77c, Cha71a, CFY76, DSW77a, DSW77b, DP72, EM77, Ell73, Fle75, FHJ78, GK76, HM79a, Has72, Her76, HH77, HH78, Hoe77, Kel71, Lie76, McE76, Mik72a, Mis72, MM71b, Obe75, RW70, Roy76, Sam79, Smi77a, Spi79b, TT71, Thi77b, Wam70]. Least-Squares [Ame76, Ano79h, Bal70, Bea77, GK76, Her76, HH77, HH78, Lie76, MM71b, Roy76]. Lee [Dun75b]. Left [AE72]. Left-Truncated [AE72]. Legal [GS74a]. Legislation [BL79]. Lehmann [Dav71]. Length [Lau72b]. Level [GR71, Ols73, Sub72, TA78]. Levels [Lar78, ZF73]. Lieberman [Cra75]. Lies [Gra77]. Life [Ano79f, BS74a, Box79, Chi72, Chi73a, Maz72, SL79, SZB79, Sin71c, Sri75, Vea75, WC70, YS77]. Lifetime [Cow75, Fas71]. Light [Hog72b]. Likelihood [Ame76, And71, And74a, And76, Bar76b, Bra73, CVM71, CM73, CM74b, CFY76, CC73, Cra76a, DD77, DeG73, DH77, Dur70, DM76b, Efr77, Fer78, Fie70, Fou77, Gal75b, GGM70, GGM71, GC78, Har77c, Har77d, Hil75, HT79, HO71, Jos76, Kal78a, Lai78, Lev73, Lis72, MM71a, MS72a, MP71, Mor78, Mor71c, Odo70, Par71, Rao77a, RAK74, Sel74, Spi79b, TC72, TG75a, TG79, Tar79b, Tho75, Wag70, Wei71b, Wei73, Wei75, Wor79, Zac70, ZM71, HM73].

Likelihood-Ratio [GGM70, GGM71, GC78]. Limit [WD76]. Limited [And74a, And76, DS78, MS72a, Mor78, Poi78]. Limited-Information [MS72a]. Limiting [EM71, EM72]. Limits [Ano78e, Bur73, Bur75, Eas72, Her71, How74a, Jol73, Lan73, Lem77, TG77, WG71, Win74b, Zac70, Bur74]. LIML [MM79, McD74]. Line [HG71a]. Linear [AG72a, Ano79h, Arn79, BNW70, Bic75, BLM71, Bow70, BM70b, BS78, Can74, CM73, Cha70a, CCM71, Che79, Che70, CL76, Chr73, Cla70c, CZ77, Coo79, Cox71, Cox72, DC73, DL76, Dav77a, Dav78, DM74, DF74, Dol76, DH72, Ebb73b, Ela72, Ell73, Eng75, EF76, Far76, FH70, Fle71, FB73, GC70, Gov71, Gov76, Ham78, Han78, HJ73, Har77b, Hay71a, Hei75, Hoa70, HS75, HH75, HHD75, HM70, JW78, KN75, Kel71, KM79, Kin72b, KJT72, KN77, KV73, KV75, LaM77, LM78, Lan73, LMH73, Mal72, MSW74, McE76, McL77, Mik71, MG70, Mor71b, Muk76, OL71, Par71, Pat75a, Raj75, Ram77, RBRH78b, Rao70, Rao72, RY78, Roy76, SU74a, SH73, Sie78, Sob71, SHH78, SF72b, Sti76, SM77b, Tho76, TS77, TB77b, TB79, WS78, War71]. linear [CC74a]. Linear-Model [Har77b]. Linearity [BK76]. Lines [Cox71, Cox72, Hog75c, LSW78, TB77b, TB79, War72]. Link [BH76, HD78, Hub74b]. Linkages [Key71]. List [Ano75f, Ano76h, Ano76i, Ano77g, Ano77h, Ano78g, Ano78h, Ano78i, Ano79k, Ano79l, Ano79m, Ano79n]. Litter [Gla79]. Livebirth [SL74c]. Local [DWD76, Eri74]. Locally [Bea74, Cle79, MS76c]. Location [Ano77d, BM70a, BLM71, Bro75, Cha70a, CC74b, Coh76a, CWR78, DL77, Dem73b, Dud71, Fra73b, Fry70, Gal75b, GC70, Gro72, Haw77, Het73, Hoa74, Hoa75, Joh74, Jon79, KN75, Kno70, Krz76, KV73, KV75, MS76a, McG73a, McG73b, Mie72, Mik71, PH76, Pra73b, Ram71, RR77, Sol72a, SW78c, Tak71b, TL70, Wor79]. Loeve [BH79]. Log [BE75, Che79, Ham71, Mic72, Zel71]. Log-Linear [Che79]. Log-Normal [Ham71, Mic72, Zel71]. Logarithm [Ham71]. Logistic [CCM71, DL76, Efr75, Eng75, Joh73a, PW78, San74b, SB78a, CC74a]. Logit [HD77, Odo70]. Loglinear [Lee77a]. Lognormal [BM70b, Ebb73b, ES74, Fow79, Meh73, Mos70a, MW70, OW72, Tho70]. Long [Blu74, DuM73, Gro76]. Long-Run [Blu74]. Long-Tailed [DuM73, Gro76]. Longitudinal [GGS73]. Look [BRB76, Bly71, Bro71a, Dan71, Efr71a, Efr71b, Esp77, Goo71a, Hoa70, Pra71, Pra73a, Stu71]. Lorenz [GG78, Kri79, Meh75]. Loss [Bro75, MA79, Sha70a, ZP79]. Losses [Hay71a]. Lots [FM74, FM77]. Lower [SC75b, Wei72, Wes77, Win74b]. Lumpability [TB77a]. Lyman [Kle77a].

M [Dun75b]. Macroeconomic [CF72]. Main [SA70]. Maintained [EP70]. Major [Jol73]. Making [GS75, Pre78c, San74a]. Malaria [New77]. Males [BHF77]. Malformations [Che78]. Managerial [WWF74]. Mandelbrot [Goo78d]. Manuals [And78b, BF78a, BF78b, Dix78, McC78b, Mul78a, Mul78b, Nel78, NFJN78, Rya78, San78, Sch78b]. Manufacturing [TG71a, Wei77c]. Many [DG73b, Jag73]. Marginal [Goo71b, Jes70, Lev73].

Margins [CA77, GM77a]. Marital [SL79]. Marital-Status [SL79]. Market [Car70, DS78, HMW74, Nel70]. Markov [DY78, Dev73a, Dev73b, DL72, GY70, Hen71, MMS71, SL79, TB77a, Yak76]. Markov-Generated [SL79]. Markovian [McC74b, Coh76c, McC71]. Marriage [CM72]. Marriages [Lee75]. Matched [BS77a, Cla70c, HL73, Jen73, McK75c, Rah74, Rub79]. Mathematical [Dia78, OPRS77, Rao71a]. Mathematics [Ano79e]. Matrices [Jen70, MD74, PAA70, SH79, Wor74]. Matrix [Che70, Har72, KN77, McE76, RY78, SW79, SZ73, You76a]. Matter [Ano70a, Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano70g, Ano70h, Ano71a, Ano71b, Ano71c, Ano71d, Ano71g, Ano71h, Ano71i, Ano71j, Ano72a, Ano72b, Ano72c, Ano72d, Ano72j, Ano72k, Ano72l, Ano72m, Ano73a, Ano73b, Ano73c, Ano73d, Ano73f, Ano73g, Ano73h, Ano73i, Ano74a, Ano74b, Ano74e, Ano74f, Ano74g, Ano74h, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano77a, Ano77b, Ano78a, Ano78b, Ano78c, Ano78d, Ano79b, Ano79c, Ano79d, Ano79i, Ran73]. Max [McC78c]. Maxima [CT73]. Maximin [Lee77b]. Maximizing [Bof70]. Maximum [Ame76, And74a, And76, Bar76b, Bra73, CVM71, CM73, CM74b, CFY76, CC73, Cra76a, DH77, DM76b, Enn75, Fer78, Fie70, Har77c, Har77d, HO71, Lai78, MM71a, MS72a, Mor78, Mor71c, Odo70, Par71, Rao77a, RAK74, Sel74, Spi79b, Ste70a, TC72, Tar79b, Tho75, TB72, Wag70, Wei71b, Wei73, ZM71, HM73]. Maximum-Likelihood [Wag70, HM73]. may [Vea75]. McCornack [Cla70c]. McGraw [RW74]. Mean [Bai79, BOLK75, Bor77, BM76b, Car78, CD76, CC73, Dah74, DS74, DE73, Eri70, Fis72, Gla76c, GH76a, GM77b, Haf78, HG73, Her71, How74a, HF70, KM74, Lan73, Lit74a, MR79, MY78, Meh73, MS71, NR71, Ofo74, Sah76, Sax76, Sin71a, SW79, Som70, Som74, SH75, Sri71b, Sri72b, SM78c, Tho70, WK71, You76a, ZM71, Zac73, CC74a]. Means [BT78, Bur79, CZ75, Dud72, Gab78b, GL71, Han78, Hei75, Joh75a, Koc73, Lan74, Lan76, LG75, Lin71, Lin73a, LS75, Muk76, Pat76, PL76, Pot72, Pre78b, Rea75, Str72b, Str72c, Str73, Tam79, War72, Lin74, Pre78a]. Measure [Ali74, BB71, HP70, Mur70, ZR71, ZR74]. Measured [DF72]. Measurement [Cha71a, Coc70b, DG75, DG76, Ebb73c, Pot72, Pra70, RW77, Str72b, Str72c, Str73]. Measurements [And78a, HF70, Sri70]. Measures [Agr76b, Efr78, GK72, MJ79, Pie79, RC76, Sim78a, GK70]. Measuring [Agr77, BH75, Ebb73a, Gil74b, KQ70, Lon70a, McG77]. Mechanism [BM73b]. Median [Ekb73, FW79, Lit74a, MJ78]. Medical [BS77b, MI78, Pet78a, WD78]. Meeting [Ano70m]. Meier [Pet77a]. MELO [ZP79]. Members [Hwa72]. Memory [Lea75, SD75, SB73]. Mendenhall [Kle77a]. Mental [Owe75]. Merits [Vit79]. Method [AR71, Bha73a, CMS76, Dun76b, Eck76, Eri74, Fau73, Fel74b, Fel74c, Fen79, Gab78b, Gei75, GH78a, GH79, GG71, HG73, HT75, Hoc76, Hwa72, Joh75b, Lee72, Lee73a, LF71, LF73, Lit74a, Lit74b, Lit75, LSB76, McC78c, Mic72, Moo77a, MM71b, O’B79, Ols77, PS76a, RW70, Sag79, Sch77a, Sch77b, SS73a, TC72, WC70, WG71, Wil78, Woo71, WC76, CMS87, DL74, Jay79]. Methodology [Har77b, Tuk79b]. Methods [Ano77c, BC79, Ber71b, Bha73a, CS73, CR74, Con73, EG75, Fow79, Gar77, Gar72, Gar73, GS73a, GGS73, Goo73, GLT79, Kal78a, KR78b, KGR79, KvE72, LF72, Lit78, Ord75, Ran71, Roh74, Sch76, SS74, SHH78, SF71, Tho75, Vin76b, Wak78, Ame74]. Metric [BK76, WT75]. Metropolitan [Liu70, Sto71]. Migration [Ceb74, Gre75c, Lon73, Sly72, Sto71]. Migratory [Sen76a]. Mineral [Pel71]. Minimax [DM74, Mur71, Ofo75, PS76a, Sol72a, Sol72b, Str78, Wyn77]. Minimization [Den71]. Minimize [Rom72]. Minimized [AS77]. Minimum [AE72, Ame76, BN74, CMH73, DD75, EM78, Far76, Gre75a, Kar70, Odo70, Rob77, Ste70a, Yar70, ZP79, Ame74]. MINQU [Mal72]. MINQU-Estimation [Mal72]. MINQUE [LS77b, Pri74]. Misclassification [Che79, CA77, Gol75a, Hoc77]. Misclassifications [Ten70]. Mises [Ste72, Upt76]. Misinformation [Dun74a]. Missing [CD72a, CD74, CGD76, CL76, Dor74, HG73, Lin71, Mor71c, PS76b]. Misspecified [Mer73]. Mixed [Gil72, Har79, Joh70, WL77, YBJ72]. Mixed-Model [Har79]. Mixing [Lai78]. Mixture [BS73b, CG70, CG71, Cor71, Fow79, TC72, TS75]. Mixtures [Bin78a, BG77, Bry78, CH78c, Fow78, Har78, Hos78, Ifr70, Joh78c, Kie78b, KNP79, QR78a, QR78b, QR79]. ML [DQ72]. Mobility [KH71, Lon70a]. Modal [MM72]. Mode [Fro73, RC74, Sag79]. Model [AG72a, AW72, Ame76, AKC76, Ano77e, BJ77, Bai79, BB76a, Bas73, BH08, Bic75, BS79, CG73, Chr73, CT76, CM74c, CMH73, Dav70, Dav78, Dav79b, DF74, DO72, Dun74a, Ela72, Epp76, FH70, Fen79, Fer75, FGHA73, Gal75c, GE79, Gil72, Goo75a, GCZH73, Gre75c, Gri72, Gup76, Gup77, HM79b, Har77b, Har79, Hen71, Hin79, HS75, HST79, JG72a, JG72b, JG75, KW74b, KW74a, KM79, Koh72, KN77, Krz76, LaM77, LM78, Lea74, Lee70b, Lem77, LR72, LC76, Max73, MSW74, McC71, McC74b, MI78, Met71b, Met71c, MG70, Moo71, Mut79, NJM72, Nel70, NS74, Pau78, PE78, Pot72, Qua74, RW70, RAK74, Sah76, San74b, SB78a, SB77, Sho70, Sho72a, Sho73, Sri72b, SF72b, Str72b, Str72c, Str73, SL74c, TS77, TB77b, TB79, UB70]. Model [War71, Wei78a, WF79, Wit74, ZZ72, Zel76a, Zel76b, BJ74, Hen72b]. Model-Based [HST79]. Modeling [Cro79b, Edd79, GS74a, GS74b, Hsi79, Lin79a, Par79c, Par79d, PW74, Tar79a, Wel79, Tuk79a]. Models [Agn74, Arn79, Bal70, Bea77, BK79, BHF77, BD78, Bow70, BP70, BM70b, BM76a, Bro78d, BN75, Che79, Che70, Chr79, CF72, Cra76a, DH77, DS78, DS75, Dun75b, DWD76, Ebb73b, ES74, FB73, GF73, Gal75b, GS73b, Goo73, Goo79d, GW79, HJ73, HB77, Hei75, HT79, HH75, HHD75, JW78, Klo76, KJT72, Koc73, LW76a, Lea74, Lee77a, LW76c, Lis72, MM73, Mal72, Maz72, McC78c, McE76, McL79, O’B79, Ord75, PD74, Pes72, Pie72, Poi78, PB76, RC76, Rao70, Rao72, Rob79, Rot79, Sab77, SA73, Sam78, SH79, SH73, SHH78, SM73, SM75, SM77b, WSB73, Wyz78, ZP79, Zel79a, Zel79b, dO74, Sen74, Pie73]. Modification [AM73, Bra79a, Bra79b, Bre70, CH79, DS79, Ell79, Flu79, Gab79, Hoc76, Kru79, Mie79, Mos79, Ney79, Sch74a, Sim79c, Wei78c]. Modified [Ber78d, Bir70, CC74b, GH76b, HS74, HL75, Joh78b, KR78b, KGR79, PE78, SM77a]. Modified-EDF [GH76b]. Modified-Tukey [KGR79, KR78b]. Modulus [Ste70a]. Moment [BE75]. Moments [CS72, GK75, Hil79, Kri79, McD74, PR75b, Pat75a, RS73a, Sin76, TC72, Tho76, You72a]. Monetary [CH78a, CF72, SW73]. Money [Hsi79, Pie77, Sed77]. Monotone [PR75b, Zac70]. Monotonic [GK76]. Monotonicity [Tho79]. Monte [Ano77c, JTB81, Spi79a, Bha73a, CS73, CM74b, CB71, Eve79, Gil72, Ham76, JTB80, LS75, MG75, Mik75, RG72a, Spi78, SZ73]. Monthly [Den71]. Morbidity [Ste71, Ste73, Wil73]. Morgan [Lue71, NMC71]. Mortality [Bro78d, LS73, New77, Smi75a, Wyz78]. Mosimann [Jam72]. Most [Bea74, Kwe75, Zac70]. Movements [Bro76a]. Moving [BP70, HT79, Sab77]. MR [CMS87, FH81, JTB81, Krz81]. MSE [Fel74a]. Multi [DO72, Har73, NB71, NB72, NBH73, Sri71b]. Multi-Auxiliary [Sri71b]. Multi-Country [DO72]. Multi-Entry [Har73]. Multi-Factor [NB71, NB72, NBH73]. Multiauxiliary [Ste76a]. Multicollinearity [Atk78, GM75]. multicomponent [BJ74]. Multidimensional [Goo71b, Goo73, KJT72, KVK71, Lee77a]. Multinomial [Cha76c, DG78, FH73, GR75, GR77, GGM70, GGM71, GC74, GC78, HO71, Lee77b, Leo77, RA75, Rob78a, Wei72, Wit74]. Multiparameter [MM71a, Ste79]. Multiphase [Sri71a]. Multiple [BT74, Boh79, CS73, Chr73, Coc70b, DD75, Dun74a, EG75, EF76, Fen79, FEJM73, FEJM75, Gab78a, Gab78b, GH78a, GH79, Goo70, HT79, Hoc76, JG75, KTS75, KR78b, KGR79, Kno76, LS77a, Lev75b, LP70, Mil77, MS76d, Ols77, Ram78a, Ram78b, Rea75, RS73b, RS74a, Sch77a, Sch77b, SB79, SS73a, Sto78, Sud70, Tam79, TA78, TB72, WT75, WD72, Web71, Wei70, Wel77, Zin72, Chi74]. Multiple-Choice [Dun74a, Wei70, Zin72]. Multiple-Response [TA78]. Multiplicative [LW76a, ZZ72]. Multiplicity [Dun71, Nat76, Sir70, Sir72, SL74b, SL74a]. multiply [CC74a]. Multipurpose [Ano79g, KA78]. Multisample [Ste72, TDL75, Upt76]. Multistage [TT72]. Multistep [Bai79]. Multitype [DF74]. Multivariable [GS74b]. Multivariate [BOLK75, BS77b, BT72, CR75a, Chi76, Gal75a, Ger75, GG72, GS74b, Goo70, Haf78, Haw74, Jam75, Jon76, KA78, KL72, Koc73, Lee79, MA73, McD72, MM72, Mor71c, OK76, Pre72, PS76c, RY78, Rub79, Sag79, Sam78, Sim77, Sol72a, Ste76b, Ton70, WT75, Zel76a, Zel76b, Ols74, Ano79g]. Mutual [Dri78, Dri79]. Mutually [Bly73]. MVU [DQ72]. MVUE [Meh73].

N [Ano79t, BR70]. National [BMB78, CH78a, SB78b, Sud70]. Nations [Klo73]. Natural [Lee75]. Naturally [DeG73]. Naya [Lue71]. Near [Dev73b]. Near-Neighbor [Dev73b]. Necessary [Efr76, Fra76b, Fra76c, Hog76, Mal72, MTC76]. Negative [CS72, HG70, HG71b, LaM73, Mee74b, Pol73b]. Neighbor [Dev73b]. Nested [BS78, CG74, FB73, GW79, Sah76]. Nested-Error [FB73]. Net [Bha71c, Sly72]. Network [Lev77]. Newcomb [Sti73b]. Neyman [TP79].

No [BN74, JG72a, CMS87, FH81, JTB81, Krz81]. Noise [Chi76, Sob71]. Noise-Distorted [Sob71]. Noisy [Dev73a, Dev73b, FK70]. Nominal [MM72, Sri70]. Non [BJN70, Bul71, Kra73, Kra74, Kra75, LaM73, McD72, NR71, Smi76, SM75, Tik71, Zel71, Zel76a, Zel76b]. Non-Additivity [McD72]. Non-Bayes [GC74]. Non-Bayesian [Zel76b, SM75, Zel71, Zel76a]. Non-Central [Bul71]. Non-First- [Smi76]. Non-Negative [LaM73]. Non-Normal [NR71, Tik71]. Non-Null [Kra73, Kra74, Kra75]. Non-Parametric [BJN70]. Nonadditive [HM79b]. Nonadditivity [HJ76]. Nonblack [BHF77]. Noncentral [Gle76, GP74, Hen79, Hen80, Kri72, Ste78b]. Nondefense [Hig70]. Nondifferentiable [dO74]. Nonfarm [Bha71b]. Nonhomogeneous [MMS71]. Nonlinear [Ame76, Bar76b, BD78, BM76a, Coc73, Ekb73, Gal75b, Gal75c, GG76b, Gal77, Hin79, Kel71, Spi79b]. Nonnegative [pH78, Jam77]. Nonnormal [Sin76, TW77]. Nonnull [Kra76]. Nonorthogonal [Cra76c, Cra76d, Hem74, Hem76]. Nonparametric [Ala74, Ano77d, Bha77, BS77b, CB70, CWR78, Cro79b, Dud71, Edd79, FW79, Fry70, Gee72, GGS73, Ham76, Joh74, Kal78a, Lai78, Lin79a, LS79, MS70, Par79c, Par79d, PH73, Ram71, SZB79, SB79, SV76, Tar79a, TDL75, Tuk79a, Tur74, Wel79, Cro74]. Nonrandom [Tho73b]. Nonregular [Gal77]. Nonrespondents [Rub77]. Nonresponse [KK73, SS78a, Sri71a, WM70]. Nonstandard [MS76d, Wei71b, Wei73, Wei75]. Nonstationary [CZ77, Sob71]. Norma [Kle77a]. Normal [AKC76, Atw75, Bac79, BOLK75, Baw72, BT78, Bet71, Bin78a, BG77, BS73b, Bry78, Che71a, CD76, Chr73, CH78c, CH74a, CH74b, Coh72, DC73, DE73, Efr75, Fow78, Fow79, GG76c, Gle72, GH76a, Gre70, Gre71, Haf78, Hah70a, Hah70b, Hah72, Hal75, Ham71, Har78, Hos78, Joh75a, Joh78c, Jol73, Kab76, Kie78b, KR76, KR90, Kno70, KW79, Kra76, KNP79, Lan73, LG75, LS79, MS73, Mic72, Mor71c, Muk76, NR71, O’N78, Ofo75, PL76, PH72, PH74b, QR78a, QR78b, QR79, Rot78b, Sam78, Sax71, ST76b, SH75, SZ73, Ste76b, SSM75, TC72, Tih72, Tik71, Ton70, Wal79a, Whi70, WT71, Wor79, You76b, ZM71, Zac73, Zel71, Che71b]. Normality [Che71a, HB74, Lan74, MA73, SM79, SF72a, Cro74, Lan76]. Normalization [Kle73a]. Normally [Car75]. Note [AM73, BG73, BN74, CD74, CWR78, Cra73, Cra75, Dol76, Dun76b, Ebb73b, Ekb73, Esp77, ES74, HG71a, Ham78, HBA72, Jos76, KQ72, Koo73, LL72, MM79, MJ78, Ols77, Ott73, RA76, Sch77a, Sch77b, SM77a, Sch78c, Sim75b, SK73, SD73a, SD73b, TG75a, TG79, TB70, Tei71, Whi79b, Zin72]. Notes [Con74a]. Notions [Rub77]. Nuisance [Bas77]. Null [Kra73, Kra74, Kra75, Yar70]. Number [Atw75, Bin76, Fer75, GM77a, Gav78, Hil79, Kie78a, Mal78, McC73a, McL77, Raj77, SM78a, SM78b, SP76, Sim78b, Che71b]. Numbers [Bly71, Bro71a, Dan71, Efr71a, Efr71b, Goo71a, KK79, Pra71, Pra73a, Stu71]. Numerically [GLT79]. Nuptiality [Gup76, Gup77].

O [Ano79u, Law73, LMH73]. Objective [Ran71]. Objects [Geo74].

Observation [Vea75]. Observational [Fie75b, McK75b, McK76, Rub79]. Observations [AG70, Bof70, CL76, Coo79, Coz74, Dag71, Dor74, Ebb73a, Fis72, Han78, Haw77, Lau70, McL77, MS73, O’N78, PW74, Sri75, SV76, Wei78d, AG72b]. Observed [Bly71, Bro71a, Dan71, Efr71a, Efr71b, Goo71a, May73, Meh75, Pra71, Pra73a, Stu71]. Obtained [GGE74]. Obtaining [Dun76b, Fau73, FM77, GKAH71, Man74b, SF71]. OC [Sub72]. Occupancy [Sho70, Sho72a, Sho73]. Occupancy-Related [Sho70, Sho72a]. Occupancy-Related-Random [Sho73]. Occupation [Sto71]. Occupational [KH71]. Occupations [Wei77c]. Occurrence [Scl79]. Odds [Ano78e, Gol75a, TG77]. Offices [Man74d, Mer71]. Oil [Pel71]. OLS [RS76]. OLUMV [TL70]. One [Baw72, Bha71b, Bic75, BS77b, Can70, Can75, CH75, Dan73, Efr78, Eve79, Fro75, GG76a, GR71, HQ75, Hoc76, KT77, Kno70, Lem77, LS75, LS79, MR79, McD72, PH76, Rob72, Vea75, WB71, Woo79, Zac73, ZR71, ZR74, May73]. One- [Can75, Eve79]. One-Armed [Woo79]. One-at-a-Time [Dan73]. One-Directional [MR79]. One-Factor [Baw72]. One-Family [Bha71b]. One-Sample [CH75, GR71, PH76, WB71]. One-Sided [BS77b, GG76a, HQ75, Kno70, Lem77]. One-Step [Bic75]. One-Way [Hoc76, KT77, LS79]. One-Way-ANOVA [Lem77]. Open [Met71b, Met71c, SMRL70]. Optimal [AST73, BB71, BBS74, Bas73, BC79, Ber74, Bof75, EZ73, EHB78, EM78, FHJ78, Gro72, Har74b, Jen73, LW76c, Loy76, McL75, MS76c, MS72b, RS79, Sch76, SZ70, SA70, Ste76a, Ste78a, Sti71, Tsa76]. Optimality [GS72, LF71, LF73, McE76, Obe75]. Optimally [BLM71]. Optimization [Moo71, RG72b, VG78]. Optimum [AKC76, Bil77, Che75, Hal75, JB73, Kle73a, Lev77, Man74b, NR71, Ofo74, Sin71b, Som70, Som74, Som75, Sud78, THP78, Uth70]. Order [AB78, Bof75, BS75, CCM71, Dav73a, DM79, Gal75a, GY70, Haw71, Hay71a, KN75, KV73, KV75, McC78c, MW70, Pat75a, Rob78a, Smi76, Sti76, Sti77c, Van76, Wal70, Yak76]. Ordered [Goo79d, Lee77b, PH72, PH74b, Pro71, Sim75a, SW78b, Wei72, WG72, Sim74a]. Ordering [Ali74, LW76b]. Orderings [Tho75]. Orders [BM73b]. Orders-Shipments [BM73b]. Ordinal [Agr76b, Agr77, Sim78a]. Ordinary [All77, Bal70, BL77a, DSW77a, DSW77b, EM77, HM79a, Hoe77, Obe75, Smi77a, Thi77b]. organ [Cro74]. Organization [Ber71b]. Origin [Cox71, Cox72]. Orthogonal [Cra73, Cra76b, GJ75, OK76, SA70]. Oscillatory [Klo73]. Other [DuM73, Hen79, Hen80, HG70, HG71b, ML74, PR75b, Pat75a]. Ott [Kle77a]. Out-Migration [Sto71]. Outcomes [Gol73, Har73, Mor72]. Outlier [Gre76]. Outlier-Prone [Gre76]. Outlier-Resistant [Gre76]. Outliers [Ela72, MR79, Sch76]. Output [BW76, WM71]. Overhaul [Bas73]. Overidentifying [Kad70]. Overlapping [Roh74, Zec71]. Overstatement [FNL77]. Owner [KQ72]. Oxygen [GCZH73].

P [Hoa74]. Pair [Cla70c, McK75c, Rah74]. Pair-Matched [McK75c]. Paired [Bea74, Dav70, EHB78, Geo74, IJK76]. Paired-Comparison [IJK76]. Pairs [HL73, Rea75, Wes75]. Pairwise [CS73, Dri78, Dri79, Gab78a, KR78b, KGR79, Ram78a, Ram78b]. Panel [Bai75, McG77, Pau78, WM70, Dun75b]. Paper [Cra75, Ofo74, Som74]. Papers [Ano70m]. Paradox [Bly72a, Bly73]. Paradoxes [Bly72b, Bly72c, Dav73a, Goo72b, Lin72, Pra72, Sha73b, Win72b]. Parallel [HS73]. Parallelism [Hol70a, LSW78]. Parameter [AE72, AST73, Alv78, Ano79f, BJ77, BBS74, Bil77, Bli74, CM74b, CC74b, DM79, DQ72, Dud71, FR71, Gho79, Hsu79, JK72, Meh72, MW70, OW72, Sol72a, Sol72b, Sta75, Str74, Str75, Tak71b, Tha79b, TL70, WEB78b, WE79, YS77]. Parameters [Atw75, Bas77, BBH72, Bha77, BB76b, Bro75, BG70, Can74, Cha70a, CCM71, Cha72, CD72b, DL76, DM76b, Eng75, Fer78, Gal75c, Gov71, Gov76, Har71, HM73, Het73, HB73b, Kel71, KN77, KV73, KV75, Law73, LP75, LR72, MG74, Mic72, Mor71c, NS74, PF78, PS76c, Rob78a, Ros72, Rub74b, SL74c, TC72, TG75a, Wei78d, ZP79, TG79]. Parametric [BJN70, Boh79, GW78, Ham76]. Paretian [HMW74]. Pareto [AG72b, KV75, AG70, KV73, Van76]. Parity [Gup76, Gup77]. Part [EM71, EM72]. Partial [AB78, Agr77, Bil77, BRH76, Gas74, Han73, Har74a, Hog74a, Hog74b, Hub74a, Rea71]. Partially [Kwe75, Lee72, Lee73a, Wol77a]. Particular [Can75]. Partition [Gut74]. Partitioning [Goo71b]. Party [Qua74]. Parzen [Tuk79a]. Path [Sen72b]. Pattern [Mic72, RY78, TT71]. Patterns [Car70, Goo75a, SL79]. Pay [Jag73]. PBIB [Cle70b]. Pearson [DG73b, SS78b, Woo76]. Pearsonian [Mis72]. People [Fra76a, Jag73]. Percentage [BJN70, Bur73, Bur74, Her71, Hoa77]. Percentages [DF79a]. Percentile [Ang79, GW78, Hog75c, ST76b, TB77b, TB79, Whi70]. Percentiles [MG74, SB71]. Percy [Sti73b]. Peremptory [KK79]. Performance [DG78, TG71b, WWF74]. Performing [Her71]. Periodic [Sri75]. Permanent [Ebb73c, Ebb73a, Hol70b, Hol71a, Hol72a]. Permutation [Som77, TY74]. Permutations [CR75a]. Personal [Bha76, DF75, Sav71]. Pertaining [Goo75b, Wes77]. Petroleum [AG72a, UB70]. Pharmacokinetics [Hay71b, Met71b, Met71c]. Phase [Hin71, Rao75]. philosophy [Jay79]. Piecewise [EF76, Hay71a, MC70, Poi73, Poi74]. Pillai [PJ70]. Pitman [Hoa75]. Places [FH79]. Plan [ZG72]. Plans [CM76, Dan73, FM74, FM77, Her76, Min72, Sam70, SA70, Sub72, TA78]. Plant [FT70, WC70]. Plant-Record [WC70]. Plantation [BH79]. Play [Hoe72, HL75, KW74c, Nor76, PS74, SW71, WD78]. Play-the-Winner [Hoe72, HL75, KW74c, Nor76, PS74, SW71, WD78]. Players [CH78b]. Plot [LK78, Sch78c]. Point [Bar73, JP74, MS71, Obr79, TF73, Wor74]. Points [BJN70, Bro76b, Bro76c, Bur73, Bur74, Che73b, GF73, Her71, Hoa77, Lon70b, Meh75, Whi70]. Poisson [Bur74, Gre71, AE72, BDG78, Bur73, CVM71, CM74b, CZ75, CZ77, DG73a, DM76b, FKB73, GS73b, Gre70, HAB71, Har71, Pau78, Pyn79, Rom72, Sam76a, Sel74, SW78a, Sta75, Ste76b, Web71, Wes77, ZZ72]. Poisson-Distributed [FKB73]. policies [Fox74]. Policy [Jen73, SW73]. Political [Hut72]. Poll [Fie72]. Pollution [Bro78d, Koh72, LS73, Smi75a, Wyz78]. Polynomial [Atw71, Fro75, GF73, Hal73, MS72b, Sti71]. Pool [Coh74a]. Pooling [Sin71a, Sri72b]. Population [Baw72, BM75, Bur75, Car78, DL77, Dah74, DF74, DF75, Eri74, Eri76, Eri70, Gup76, Gup77, Hil79, HST79, Hwa72, Kab76, Key72, Lee75, Mil76, Ofo74, Ofo75, PL76, Rot78b, RC78, SBH78, Sen72b, Som70, Som74, Som75, SH75, Sri71b, SZ73, SSM75, Wol74a, Hil70b, Isa74, Gov71, Gov76]. Population-Distributed [DF75]. Populations [BOLK75, BS73a, BG77, Coh76c, Eck76, FNL77, Gee72, KW71, KW74c, Kra76, Laa79, Lee74, LG75, Lev77, MMS71, Muk76, NL77, Nor76, Pat76, PS76a, RG72b, Rao75, RS74b, RH73a, RH73b, Sax71, Sin76, SW71, SZ70, Sud72, Wac77, Wol79b, Wor79, You76b, Joh78b]. Positive [CS72, HM70, Str74, Str75]. Positively [KW74d]. Possible [War78a]. Post [Dem73b, Fra73b, Lea74, Lee74, Man74d, McG73a, McG73b, Mer71, Pra73b]. Post-Data [Dem73b, Fra73b, Lea74, McG73a, McG73b, Pra73b]. Post-Transition [Lee74]. Posterior [DeG73, Dic76a, Eri70, Hil79, Whi79b]. Postpartum [SBY79]. Potential [Web71]. Power [Ano77d, BH75, BJN70, BR77, Che71a, CWR78, Gab78a, Gal75b, GP74, HJ76, JS75, Ken76, Kno70, LG75, Lis72, Mar76, Nat72, Nau73, Par75, Par76, PF78, PAA70, Ram71, Ram78a, Ram78b, San74a, SM79, Smi75b, Smi76, Ste70a, Tik71, Tik72, Wis72, Yu75]. Powerful [GH76b]. Powers [EG75, Mie72]. Practical [Sch70a]. Pratt [Rah74]. Pre [Ven78]. Pre-Data-Collection [Ven78]. Precision [MS76d]. Predicted [Rob77]. Predictibe [MM72]. Prediction [AW72, Ano79h, Bai79, Bha73a, Dor74, Dun76b, Fau73, FM77, FW79, Hah70a, Hah72, HN77, HN79, Her76, KN75, KR78a, MG74, Roy76, Sob71, WT75, Win71, WM71, Ano77c]. Predictions [Cox71, Cox72, Gol73, Mor72]. Predictive [Cra70, Dav79b, FT70, Gei75, GE79, GS74b, Laa79, Lau70, Yea73]. Predictor [Kin72b]. Preferences [Bar73]. Preliminary [Arn70a, BS74a, BYJ73, pH78, Har77a, MA79, ST75]. Presence [Gla79, GR74, LM78]. Present [Vea75]. Presentation [GOST75]. Preserved [Hob76]. Pretesting [Leo77, LP70]. Prevalence [Gor73b, Lev77]. Price [Ano77e, BNW70, BD73, Bha71b, BS74b, Car75, Epp76, FS72, Hsu79, LP75, Tei71, Tur70, Yea73, YFS71]. Prices [CR72, CM74a, Epp79, HMW74, LMSW70, Sla73, Sla74]. Primary [Sud70, Sud78]. Principal [FHJ78, Gre75a, Haw74, Krz79, Krz81]. Principle [Bir70, Bly72a, Jos76, Sav70]. Prior [Bro76b, Bro76c, Mee74a, Sim72b, Hil70b]. Priors [Bar76a, BT76, Dic76b, Fra76d, GM78, Lea78b, Lin76a, Pra76, Sto76a, Sto76b, Vil76]. Probabilistic [Gol73, Mor72, NJM72, WL77, Win71]. Probabilities [Bre70, CR74, DM74, DF75, Est72, FH73, Fin77b, GL71, Haf78, Har73, HB72, Hen71, Hom78, KS71a, Kwe75, Lin78c, MMS71, Nau74, OPRS77, Sav71, SC77, TG75b, WN74a, Wan71, WM77, Woo76, ZM71]. Probability [AS76, BR70, Bly72b, Bly72c, CM73, Cha71c, Cha76c, Cra75, CR75c, DeG73, Dun75b, Edw75, EG71, Goo72b, Her76, Hil79, Hog75a, Hog75b, Jes70, Jes73, Kwe75, Lin72, Lin73c, LK76, Mee79, Pra72, RV77, Rea71, Sam70, Sim73, SC75b, SL74c, Sud70, Ton70, Vij75, Wei72, Wes77, Win72b, Win75d, JTB80, JTB81, Sha76a]. Probit [Mut79, Ame74]. Problem [BB72, BN74, CG70, CG71, Cor71, Fry70, Gho75b, GR75, Hwa75, Jen73, Lau72a, LB72, MS76a, Mee74a, MR70, Nau74, Ofo74, PH76, Sap72, Sch70a, Sch73d, Sch74b, Sco78, Sin71c, SD75, Som70, Som74, Ven78, Wac77, Wan71, Woo79, YFS71, GR77, WD72]. Problems [AM75, BBS74, Ber71b, BT75, Box79, Bro76a, Har74b, Har77c, Har77d, HS75, Leo77, MM71a, Nea75, Rao77a, Rub74b, Sed77, Sha73b, Sri71a, Sto78]. Procedur [TB72]. Procedure [AS76, BS73a, BS74a, Bin76, BG77, Gho75b, GR74, HB73a, Kat77, Mad74, MR79, McL75, Nor76, Ofo75, Owe75, Rea75, RC74, Rot78b, RA76, Sax76, SW78a, WC70, Dud71, Kat78]. Procedures [Atw71, BJ77, BT78, BM73a, CS73, CD76, DG75, DG76, Dye73, Dye74, FH79, Gas74, Gee72, GG72, Gli72, Gol75b, HB70, Har74a, Har74b, HR77, Hog74a, Hog74b, Hub74a, Hub74b, Lin71, Lin73a, Loy76, MM71a, Moo73b, OK76, PH76, SW76, Sha74, SB79, Tam79, WB71, Wel77, Wol77a, WE79, ZF73, Isa74, Ano78f, SW77]. Process [Bas73, BHF77, CR72, CZ77, Cor71, CR75c, DF74, GS75, Rom72, Scl79, SW78a, Web71]. Processes [And78a, Ant72, Edw75, FM74, FM77, Hog75a, Hog75b, KK73, KK75, PW74, Wes77, Win75d, Dun75b]. Processing [Min72, Pro71]. Producer [BM73b]. Product [RS73a]. Production [RZ71]. Productivity [Man74d, YFS71]. Products [Har71, Jam72]. Program [CT76, Cra76d, Cra76c]. Programming [AG72a, BNW70, Rao71a, SH73]. Programs [Fra73a, MM71b, Wam70, WNIO77]. Progress [Sha73b]. Projection [Sen72b]. Prone [Gre76]. Properties [Atk78, BJ70, DP72, Fro75, Gre75a, GS72, HS74, Hob76, Jes73, JP74, JR71, Kra76, Lau74, MJ79, Mer73, Mik72b, Mik75, Raj75, RG72a, San74b, Spi79b, TS77, Tih72, Wei71b, Wei73]. Property [FHJ78, Uth70]. Proportional [Ame73, Far76, RV77]. Proportional-Variance [Far76]. Proportions [Ano78e, BG73, BS77a, Gla79, Gol75a, GW79, Kle73c, Pre78a, Pre78b, SL74a, SL74b, TG77, Vit79]. Proposed [Fuc79]. Prospects [UB70]. Proverbs [Eck70]. Proxies [Pes76]. PS [Rao71b, Rao73]. Pseudorandom [Gav78, Kie78a, Mal78, SM78a, SM78b, Sim78b]. Publication [Ano74i]. Publications [Ano70i, Ano70j, Ano70k, Ano70l, Ano71k, Ano71l, Ano71m, Ano71n, Ano72n, Ano72o, Ano72p, Ano72q, Ano73j, Ano73k, Ano73l, Ano73m, Ano74j, Ano74k, Ano74l, Ano75g, Ano75h, Ano75i, Ano75j, Ano76j, Ano76k, Ano76l, Ano76m, Ano77i, Ano77j, Ano77k, Ano78j, Ano78k, Ano78l, Ano78m, Ano79o, Ano79p, Ano79q, Ano79r]. Purchase [BB76a, BH08, CG73, Pau78]. Purchasing [CG73]. Purposive [Wyn77]. Pycnophylactic [DWW79, Tob79a, Tob79b]. Pyx [Sti77b].

Quadrat [CM74c]. Quadratic [Den71, Ela72, GG76c, Gle72, JS72, JS73, LaM73, RBRH78b, Wal79a]. Qualitative [Ame76, Goo70, Pre78c]. Quality [KQ70, Roc75, Sub72]. Quantal [Owe75]. Quantifying [AKC76]. Quantile [Kre76, RR76, RS72]. Quantiles [Can74, Cha70a, Che75, Kab76, Rob77, Wei78d]. Quantitative [GKAH71, HP70, LC76, PB76]. Quantities [LMSW70]. Quantity [Adi73]. Quarterly [Cra70, Den71, SJL76]. Quasi [Fie70, Goo75a]. Quasi-Independence [Fie70, Goo75a]. Queensland [BL79]. Queries [Iji71]. Question [Fel72a, Fel72b, Lan75b, Lau72b, Moo71]. Questioning [Bor70b]. Questions [FGHA73]. Quick [KvE72, Nea75]. Quotients [Har71].

R [Ano79u, Hoa74, Kle77a, Geo74]. Race [Ceb74, Gor73b, MB74b]. Rail [Wes79]. Random [Bly72b, Bly72c, CMS76, CS72, CR75a, DH77, Fro73, Ger75, Gla76c, Goo72b, Gri72, How74a, Kel73, KR76, KR90, LM78, Lem77, Lev77, Lin72, LK76, Mee74b, Pra72, Raj75, SA73, Sah76, Sau74, Sho70, Sho72a, Sho73, SU74b, SL74a, SL74b, Ste78a, SM75, The74, Wak78, WM77, Win72b, CMS87, Che71b, Poo74, Rel72]. Random-Effects [Lem77]. Randomization [Zer79]. Randomized [DS75, Fen79, FGHA73, GG75, GKAH71, Lan75b, LW76c, LCM75, LC76, Loy76, Moo71, MS76d, PH72, PH74b, PB76, Rob75, Sam70, SB78b, SW78b, War71, WD78, WF79, Zer79, Poo74, Sen74]. Randomizing [LCM75]. Randomly [Bin76]. Randomness [Ala74, Lee77b]. Range [Haw71, JR71, Lev75b, Sin76, Sto78]. Ranges [Hah72]. Rank [AB78, AGA79, Bau72, BJ70, CS74, Cla70c, CH74a, CH74b, Con73, Dav71, Gil72, Gro72, Hil74c, HS75, Kra74, LSW78, Mil70b, Mil72, Ott73, Rah74, RH71, Sie78, WB71, Wes75, Zar72a]. Rank-Frequency [Hil74c]. Ranked [AST73, Dud72, Par79b]. Ranking [Kat77, Kat78, Pat76, Rob72]. Rankings [Qua79b]. Ranks [KvE72, Mie72, RBRH78a, War72]. Rare [Eri76, Lev77, Sud72]. Rate [AM73, Bas71, Ger77a, Har77b, PR75b, SW78a, Sri75, Web71]. Rates [Blu74, CS74, Coh76c, Hol70c, Hol71b, Key71, Nel70, Pie77, Sly72, Wil71]. Ratio [And71, Bha77, Cha71c, DS74, DeG73, Gal75b, Gla76c, GGM70, GGM71, GC78, Koo73, Lis72, MP71, Rao75, Rea71, Sam70, SD73a, SD73b, TB72, Tin70, Wei75, WT71, Wor79, Zac70]. Rational [Pes76, The74]. Ratios [Ano78e, HF70, MTC76, SA73, SL74a, SL74b, Sly72, TG77, Ano78e]. Reaching [DeG74]. Real [Bha71c]. Realizations [GGE74]. Reasons [Con74b, Con74c, Man74c, Mie74, SGS74]. Recapture [Wit74]. Received [Ano70i, Ano70j, Ano70k, Ano70l, Ano71k, Ano71l, Ano71m, Ano71n, Ano72n, Ano72o, Ano72p, Ano72q, Ano73j, Ano73k, Ano73l, Ano73m, Ano74j, Ano74i, Ano74k, Ano74l, Ano75g, Ano75h, Ano75i, Ano75j, Ano76j, Ano76k, Ano76l, Ano76m, Ano77i, Ano77j, Ano77k, Ano78j, Ano78k, Ano78l, Ano78m, Ano79o, Ano79p, Ano79q, Ano79r]. Reciprocal [Hei75, Tho76]. Reciprocation [Sau74]. Reclassification [McL75]. Recognition [AR71]. Recognizes [Gup76, Gup77]. Reconstructing [Dev73a, Dev73b]. Record [WC70]. Recording [Lue71, NMC71]. Recovery [BM77b, Por73, Por74]. Rectangular [Dri78, Dri79]. Recurrence [AE72]. Recursive [Bli74, DH72, OL71]. Reduced [BM76b]. Reducible [SW79, You76a]. Reduction [Ger77a, SGO71, Sim77]. Reevaluation [FT70]. Reexamination [BD73]. Reexamined [Ros72]. Reference [Pie77]. Refinement [Bjö78, Fle75]. Refining [AG72a]. Reflexive [Ban72]. Reforms [Reg73]. Reformulation [Ols77, Sch77a, Sch77b]. Regimes [Fer75]. Regional [TF73]. Regions [DWW79, SA73, Tob79a, Tob79b, UB70]. Regresion [DH72]. Regression [AK71, Ame73, Ano72e, Ano77c, AM75, Atw71, Bai79, Bal70, BB72, BK78, BRB76, BRB77, BT74, BW70, BA77, Ber77, Bha73a, BYJ73, Bow70, BM76a, BG70, Cha71a, Cha71b, CD72b, Chr73, Cle79, Con74a, Coo79, CMH73, Cox71, Cox72, Dag71, DC77a, DC77b, Efr75, Efr78, Ela72, Ell73, Eri74, EF76, Esp77, Fer75, FEJM73, Fro73, FKB73, FH78, FH81, GF73, Gal75b, Gal75c, GG76b, Gal77, Gar77, GS73a, GM78, Gre75a, GW78, Gro77, GM75, GM77b, Gut74, HG71a, Hal73, Ham76, Har77a, HB77, HH77, HT75, Hin71, Hoa70, HQ75, Hog75c, HB74, JW78, KW74b, KW74a, Kin72b, LM78, Lea78b, LMH73, Lit78, LSW78, LP70, MY76, Maz72, MC70, MM71b, MS72b, PD74, Poi73, Poi74, Pot72, PS76b, PW78, Pyn79, Qua79a, Raj75, Ram77, RC76, Rev74a]. Regression [Rev74b, Ros72, Rub79, SK76, Sch76, Shu70, Sie78, Sim74b, SS78a, Sti71, Str78, Str72b, Str72c, Str73, SM73, SM75, TS77, Vin76b, Vin76c, Web71, Wri78, YBJ72, Zel71, Zel75, Zel76a, Zel76b, dO74, PH74a, FEJM75, HH78]. Regressions [Bin78a, Bry78, CH78c, Fow78, GLT79, GS73c, Har78, Hos78, Joh78c, Kie78b, KG70, KNP79, MS76b, Qua72, QR78a, QR78b, QR79, Sam74, SU74b, SM75]. Regret [Bro76b, Bro76c]. Rejoinder [BF78b, Bly72c, Bra79b, Con74c, DSW77b, DC77b, Efr71b, Fra76c, GS77, Har77d, Hog75b, Hog74b, Hol71b, KZ72, Kie77b, McG73b, Met71c, Mul78b, NMC71, Par76, Par79d, QR78b, QR79, Ram78b, Sch77b, SM78b, Som74, Ste73, Sto76b, Str72c, SD73b, Tob79b, Zel79b]. Related [GG71, Har77c, Har77d, Hen79, Hen80, Hog72b, JW78, Rao77a, Sho70, Sho72a, Sho73, Wol74a]. Relation [AE72, Blo77b, Dol76, Dur70, Mis72]. Relations [Lee70a, Wal74]. Relationship [Goo75b, HP74, Lau72a]. Relationships [Low74, NNO76, Pie77, Smi75a, SJL76]. Relative [AGA79, Baw72, DS75, Gol75a, Her76, KMS75, LMH73, McG77, McK75c, MS76b, Vit79]. Relevance [Tuk79b]. Reliability [BM78, Ber74, Eas72, EP70, Har71, HS73, HS74, HB73b, Man74b, Win74b, BJ74]. Remark [Fre77, SC75a, SC76]. Remarks [BS73b, PS76a]. Removed [Wes75]. Renewal [Lee74]. Renyi [Yu75]. Renyi-Type [Yu75]. Repeated [And78a, HS73, HF70, KF70, MTC76, SS74, You72a]. Replacement [AS76, Fre77]. Replicated [Geo74]. Replication [JG72a, Lee72, Lee73a, McC76b]. Replications [Cor76, GJ75, KF70]. Report [Wam70]. Reported [DD77]. Reporting [Lau72b, MB74b].

Reports [Ano72r, Ano72s, Ano72t, Ano73n, Ano74m, Ano74n, Ano76n, Ano76o, BN73]. Represent [Che73b]. Representation [Gle76, Krz76, Vil77b]. Representations [Bul71, CR72, Har79, Kri72]. Representing [CR75a]. Research [Est72, GO72, SZ70]. Residential [Sla73, Sla74]. Residual [BT74, BP70, Dol76, Efr78, Hau76, McL79, Obe75, Pie72, Chi74, Pie73]. Residuals [AS77, Bro78c, Ell73, GS72, HM79a, HR70, Hem74, Qua79a, RS76, Rev76]. Resistance [TY74, Ylv77]. Resistant [Gre76]. Resource [NB71, NB72]. Respect [RB71]. Respondent [LW76c]. Respondents [PY74]. Response [Agn74, Ame76, Atw75, BHF77, Bor70b, BN73, DS75, Fel74b, Fel74c, FGHA73, Goo75a, GG75, GKAH71, Gri72, Kar70, Klo76, Koc73, KM74, Lan75b, LW76c, Lit74a, Lit74b, LCM75, LC76, LSB76, Loy76, MB74a, Moo71, Nat76, Owe75, PB76, San74b, SB78b, SB73, TA78, War71, WG72, WF79, Zer79, Poo74, Sen74]. Response-Duration [Klo76]. Response-Time [MB74a]. Responses [Can70]. Responsibilities [Fra76a]. Responsibility [Tuk79b]. Restoration [Bas73]. Restraints [RB71]. Restricted [BL77b, BL77c, DM76b, FHJ78, Pat75a, Rev76, Sam79]. Restriction [Rob78a]. Restrictions [GP75, Kad70, KN77, Sim72b]. Results [Bha76, CM74b, EHB78, KG71, Lea75, Low77, NBH73, OW72, RS76, Rev74a, Rev74b, Rev76, Som75, Win71, Wol73, Wol74a, YBJ72, Pri74, Ano72e]. Retail [BS74b, Liu70]. Retaining [KS71a]. Retention [Hen71, Hen72b]. Retirement [Rei76]. Retrospective [Mil76, MB74b]. Return [Blu74]. Returns [Mer71, Off72]. Reuse [Gei75]. Revealed [Lea78b]. Review [A.73, A.74a, A.74b, A.74c, AY76, Abo78, Abr79, Ada76, Ade78, Agr76a, Agy78, Aig71, Aig72, Aig75a, Aig75b, Aig75c, Ait72, Alb76, Alb79, Ame70, AZ79, Anb76, And70b, And77, And79, And74b, And78b, Ang78, Ant73a, Ant79, Ant73b, Arc73, Arn75, Arn76, Arn78, Arr70, Atk76, Atk79, Atw76, BJ72, Bal74, Bal75, Ban78, Bar79, Bar70a, Bar70b, Bar74, Bas78, Bau71, Bea76, Bec76, Bec73a, Bec73b, Bec74, Beh73, Ben75, Ben72, Ber71a, Ber73a, Ber78a, Ber73b, Ber70, BF78a, BF78b, Ber76, Ber78c, Ber78b, Bha71a, Bha79b, Bha79a, Bic73, Bin78b, Bir78a, Bir78b, Bla71, Blo77a, Blo71, Boa73, Boe71, Boe72, Boh75, Boh71, Bon79, Boo77, Boo79, Bor70a, Bou77, Boz73, Bra70, Bri71, Bri72, Bro78a, Bro78b, Bro71c, Bro71b, Bro71d, Bro72a, Bro72b]. Review [Bro72c, Bro79, Bro73, Bru73, Bru78, Bue76, Bue79, Bul72, But71, But79, But76, But77, C.71, C.72a, C.72b, C.72c, C.74a, C.74b, Cam79, Cas78, Cha76a, Cha79, Cha73, Cha75a, Cha74a, Cha78, Cha76b, Cha75b, Cha70b, Cha71d, Che73a, Chi71, Cho76, Cla70a, Cla70b, Cla71, Cle70a, Cle71, Coc70c, Coc70a, Coh71, Coh74b, Coh78, Col75, Col72, Con76a, Con76b, Coo74, Coo71b, Cox73, Cro79a, Cro71, Cru75, Cza70, D.72a, D.72b, D.73a, D.73b, D.74a, D.74b, D.74c, D.74d, D.74e, D.74f, D.74g, D.74h, D.75a, D.75b, D.75c, D.75d, Dal71, Dal77, Dal78, Dal79, Dav72a, Dav72b, Dav73b, Dav76, Dav77b, Dav79a, Dav79c, Daw76, Day79, DeG70, DeG71a, DeG71b, DeG75, DeG76a, DeG76b, DeL79, DeW71, Dea79, Dem73a, Den70, Den73, Des70, Dev78a, Dev78b, Dev79, Dia78, Dic70, Dic73]. Review [Dic77a, Dix78, Dok71, Dok72, Dou71, Dra73, Dra71, DG72, Dud78, Dun75a, Dun75b, Dun76a, Dun77, Dun79d, Dun79a, Dun79b, Dun79c, Dun74b, Dwy73, E.74, Eas77, Eat73, ET74, Eis73, EJ75, Ela75, Ela77, Els77, Eng72, Eni71, F.74, F.75, Fab70, Fai77, Far70, Far73, Far72, Fed70, Fei75, Fel79, Fie75a, Fie75b, Fie76a, Fie76b, Fie76c, Fie77a, Fie77b, Fie77c, Fis77, Fis70, FL76, Fox76, Fra70, FF78, Fra78, Fra79, Fra71, Fre78, Fri70, Fri78, Fry71, Fry72a, Fry72b, G.71, Gal73, Gan79, Gan71, Gar78, Gas74, Gav70, Gav73a, Gav73b, Ger72, Ger77b, Gib73, Gil74a, GE75, Gil76, Gil78, Gle73, Gle74, Gle75, Gli78, God74, God70, Goe74, Goe77, Goe78, Goe79, Gok79, Gol70, Gol76, Gol77, Gol72, Gol78, Goo72a, Goo78a, Goo78b]. Review [Goo78c, Goo78d, Goo79a, Goo79b, Goo79c, Gor73a, Gor76, Gov70, Gre73, Gre75b, Gre74, Gui78, Gul75, Gun79, H.71, H.72a, H72b, H.75a, H.75b, Hab77, Hag78, Hai70, Haj74, Hal74, Ham75, Han75, Har70, Har74a, Has76, Hay78, Hay76, Hea79, Hen72a, Hes71, Hey71, Hey78, Hil70a, Hil74a, Hil72, Hil73, Hil76, Hil78, Hin77a, Hin76a, Hin76b, Hin77b, Hix76, Hoa74, Hoe76a, Hoe76b, Hog79, Hog72a, Hog74a, Hog74b, Hog78, Hol75, Hol74, Hoo74, Hor76, How74b, Hsi70, Hub74a, Hud76, Hud75, Hum75, Hun72, Hun75, Hun70, Hur73, Hur74, Ign74, Iji70, Imr76, Imr78, Imr79, Ing77, Ism76, Ize78, Ize79, Jay79, Jed73, Jen76a, Jen76b, Joh72, Joh73b, Joh73c, Joh78a, Joi73, Jon73, K.71, K.72a, K.72b, K.72c, K.73, K.74, K.75, Kad73, Kad75a, Kad75b, Kai72, Kal78b, Kap70]. Review [Kas74, Kat70, Kau73, Kei72, Kem70, Kem72, Kem73, Kha77, Kim77, Kin72a, Kin76, Kin77, Kin78, Kin79, Kle77a, Kle77b, Kle70, Kle74, Kle73b, Kme74, Kme77a, Kme77b, Kme77c, Kme77d, Kme79a, Kme79b, Koo77, Koo72a, Koo72b, Koo76, Kot70, Kot72, Kot73, Kou72, Kro71, Kup76, Kup78, Kup79, Kur72, LaC79, LaV75, Lac74, Lac76, Lac77, Lah75, LS71, Lan75a, Lan79, Lap75, Lar74, Lar75, Lav79, Lay75, Lea78a, Leb74, Leh77, Lev75a, Li72, Lil75, Lin78a, Lin73b, Lin77b, Lin78b, Lin79b, Lin79c, Lin79d, Lin76b, Lov76, Lov78, Luc72, Lyt72, M.72, M.74, M.75c, M.75a, M.75b, Mad73, Mag72, Man72, Man74a, Man70, Man71, MYH76, Man75, Mar73, Mar78, Mar70a, Mar70b, Mas79, Mas72, McC74a, McC75a, McC76a, McC73b, McC74c, McC78a, McC78b, McC74d, McC75b, McC79a]. Review [McC79b, McK75a, McK77, McK75b, Mee78, Met71a, Mey72, Mey75a, Mey75b, Mic76, Mie71, Mil70a, Mil75, Mil79, MW79, Min71, Mir73, Mit71, Moo72b, Moo72a, Moo72c, Moo73a, Moo74b, Moo74a, Moo75b, Moo75c, Moo75a, Moo76b, Moo76a, Moo76c, Moo77b, Moo78a, Moo78b, Moo78c, Moo79, Mor77, Mor71a, Mor73, Mor70, Muh73, Mul78a, Mul78b, Mun77, Nag79, Nas79, Nel78, New76, New79, NFJN78, Noe70, Nol73, Nol74, Nor71, Van78, O.70b, O.70c, O.70d, O.70e, O.70a, O.71, Oak74, Ode71, Ode79, Ore79, Ork72, Ork75, Owe71, P.72, P.75a, P.75b, P.75c, Pac70, Pag70, Pai77, PR75a, Par79a, Par73, Pas72, Pas73, Pas76, Pat73, Pec71, Pet77b, Pet78b, Pet79, PA76, Pir73, Plo71, Poi79, Pol73a, Pol75, Pov79, Pra74, Pre76, Pre74, Pre78d, Pro77, Put78, Put79, Q.74a, Q.74b, Qua77a, Qua78]. Review [Qua77b, Rag71, Rag75, Rag78, Ram74, Rao77b, Rav78, Rel76a, Rel76b, Rho75, Rho77, Ric76, Ric72, Rob74, Rob78b, Rod76, Rog78, Roh77, Roh78, Rol72a, Rol72b, Rol76, Ros74a, Ros74b, Ros73, Rot78a, Rot71, Rub74a, Rya78, Rym71, S.73, S.74a, S.74b, S.74c, S.75, Saa71, San78, San74c, San77b, Sär72, Sär79, Sch79a, Sch71, Sch73c, Sch78a, Sch79b, Sch70b, Sch78b, Sch75a, Sch75b, Scl73, Scu70, Sed78, Sel75, Sen71, Sen72a, Sen75, Sen76b, Sen76c, Sen77, Sen78a, Sen78b, Sen78c, Sen79a, Sen79b, Sen73, Sen79c, Sha76a, Sha75, Sha78a, Sha77b, Sha78b, She79, She72, She75, Sho76, Sho72c, Sho75, Sie72, Sim79a, Sim79b, Sim76, Sim72a, Smi71a, Smi71b, Smi77b, Smi77c, Smi79a, Smi79b, Smi79c, Sok72, SZ70, Sri72a, Sta78, Sta79, Ste75, Ste71, Ste73, Ste70b, Sti78a]. Review [Sti73a, Sti75, Sti77a, Sti78b, Sti78c, Sto70, Str72a, Str76, Str71, Suc74, Sun79, Sut78, Swi73, Swi76, Swi78, Syk74, Tae76, Tae77, Tak71a, Tam72, Tan78, Tat71, Tay74a, Tay74b, Tha79a, Thi77a, Tho73a, Tia75, Tro75, Tru70, Tru77, Tru78a, Tru78b, Tuc78, Tuf70, Tuf79, Urq75, Van72, Van73, Vin76a, Wac79, Wah71, Wal79b, War74, War79, War77, War78b, Was79, Wat79, Wat75, Wat77, Wat78, Web76, Web78a, Weg74, Wei74, Wei76a, Wei76c, Wei76b, Wei77b, Wei78b, Wei71a, Wel73, Wet78, Whe73, Whi76, Whi79a, Wic79a, Wic79b, Wig70, Wig77, Wil76, Wil74, Wil73, Win74a, Win75a, Win75b, Win75c, Win76a, Win77, Win78a, Win78b, Win79, Win78c, Win76b, Wio75, Wip71, Wol79a, Wol74b, Wol74c, Yac74, YS72, You70, You73, Z.72, Zab76, Zab78]. Review [Zah77, Zar70, Zar72b, Zys73, Zys75, McK76]. Reviews [Ano75f, Ano76h, Ano76i, Ano77g, Ano77h, Ano78g, Ano78h, Ano78i, Ano79k, Ano79l, Ano79m, Ano79n]. Revisions [Bro76a, DO72]. Revisited [The70]. Revusky [BJ70]. rho [Cle70b]. Ridge [GM75, Str78, Vin76b, MG75]. Ridge-Type [MG75]. Rights [Pel71]. Risk [Bar73, DD75, EM71, EM72, HP70, KMS75, McK75c, Rom72]. Risks [Sha73a]. River [GCZH73]. Road [BL79]. Robert [Kle77a]. Robust [Atw71, BM70a, BLM71, Bre79, BF74b, Car79, Cle79, DM79, Fin77a, GC70, Gas74, Ham79, Ham74, Har74a, Har77a, HH77, HH78, HT75, Hog74a, Hog74b, Hub74a, Lay73, Lem78, Mik71, O’B79, PH76, Pre78a, Pre78b, Ram77, RBRH78b, RR77, RH73a, RH73b, SBH78, Sti73b, Sti76, Tha79b, BF74a, Hoa74]. Robustness [BG77, CB71, DL77, Eve79, Gro76, Ham76, SSM75, TL70, Ols74]. Rogers [Hoa74]. Role [Ham74, Sim72b, WWF74]. Ronald [Kle77a]. Root [DF79b, JRH73]. Roots [PAA70, TG75a, TG79]. Rotation [Bai75]. Rounding [DF79a, MM71b]. Row [Bur79, Ell77]. Rows [Bur79]. Roy [TW77]. Rule [Bly71, Bro71a, Dan71, Efr71a, Efr71b, EM73, FS77, FEJM73, FEJM75, Goo71a, Hoe72, HL75, KW74c, McL75, PS74, Pra71, Pra73a, SW71, Stu71, WD72, WD78]. Rules [BA77, Boh79, CF72, CA79, Dev73b, DG78, Dix76, Nat76]. Run [Blu74, DY78, Epp79, Hua73]. Run-Contingency [DY78].

S [Ano79v]. Sacrifice [BBC79]. Salary [Ang79, Hog75c]. Sales [Liu70, Sla73, Sla74]. Same [RY78]. Sampford [AS76]. Sample [And70a, Ano72e, Ano77d, Atk78, BMB78, BBDS71, BBDS72, BOLK75, Ber74, BDG78, Can74, Can75, Cha70a, Cha71c, CG78, Cha72, CD72b, CA77, CH75, CM76, CWR78, DG73a, Dem73b, Ebb73b, EHB78, Eng75, Eve79, FK76, Fli79, FW79, Fra73b, Fry70, GG76a, GC70, GR71, Gei75, Gla76c, Gli72, HW74, Het73, HBA72, HS79, Hoa75, HFR75, Hor78, Hur72, Ken76, Kin72b, Kle73c, Kno70, KL72, Koc73, Kra76, Lar78, Lay73, Lee72, Lee73a, Lis72, LSW78, LB72, MS76a, ML74, MS72a, MJ78, McC76b, McG73a, McG73b, Meh72, MS74, Mer73, Mie72, Mik72b, NR71, Ofo74, PD74, Par75, Par76, PH76, PS74, Pra73b, Raj75, Ram71, RG72a, Rev74a, Rev74b, Rev76, RB71, Rub77, San77a, Sax76, SS79, SS78a, Sir72, Som70, Som74, Som75, Spi79b, SS73a]. Sample [SH75, Sti76, TW77, TG71b, TDL75, Wal70, WB71, Wes75, Whi79b, WN73, WN74b, Wol77a, Yak76, YBJ72, Lin74]. Sample-Based [Gli72]. Samples [BOLK75, CT73, Che71a, DL76, DuM75, Fel74a, Fer78, Gal75a, Hah70a, Hah70b, Hah72, HM73, Joh70, JR71, Kle73a, Lit74a, Lit74b, Lit75, McK75c, McL77, MW70, Nat72, Sha76b, Sin76, Spi78, Sud70, Spi79a]. Sampling [AR71, AS76, BR70, BW70, Ber74, BS73b, Bre79, BB77, BS79, CD72a, Che79, Coh76b, CM74c, DE73, Eri76, FM74, FM77, Fis79, Fre77, Gav78, GK70, Hal75, Han73, Her76, Hoc77, Hoe72, Jen73, Jes70, Jes73, JP74, Jol73, Kie78a, Lee73b, Lev77, Mal78, Min72, NR71, Nat76, NL77, Nor76, PS74, RS73a, Rao71b, RG72b, Rao73, RV77, Roy76, RC78, Rub79, Sam70, SM78a, SM78b, SS71, SBH78, Sed77, Sim78b, SW71, SZ70, Sri71a, Sti77b, SSM75, Sub72, Sud70, Sud72, ST75, SY72, Ten70, TT72, Vij75, Wak78, WEB78b, Wyn77, ZG72, Ano79h]. Satterthwaite [HM70]. Saving [Cra70]. Scale [Arn70b, Arn73, Bha77, Bre79, Cha70a, CC74b, FK76, Fli79, GG78, Het73, JK72, KN75, KV73, KV75, Meh72, Mer71, MM72, Mie72, Ram77, Sol72b, Tha79b, TDL75, Vin76b]. Scale-Free [GG78]. Scaling [Goo75a]. Scatterplots [Cle79]. Scheffé [KGR79, Gil72, KR78b]. Scheme [Ten70]. Schemes [Ber74, DE73, Hoc77, Chi74]. Schlesselman [Fuc79]. Schoo [Gor73b]. School [Har77b]. Schooling [HP74]. Science [Box76, Hut72, Jay79]. Scientific [Ban71]. Score [LSW78]. Scores [LS79, Mos70b, PH72, PH74b, Pol73b]. Scoring [Dun74a, Mic72, War75]. Screening [CB70, Sud78, You76b]. Search [EZ73, MM72, RS79]. Searches [Fro75]. Seasonal [CT76, Pes72, Plo79, SF72b, TT71, Wal74]. Seasonality [GM78, Sim74b]. Seat [BL79]. Second [Meh72, PY74]. Secretary [SD75]. Section [NS74, You72b]. Sectionally [SM77b]. Security [Ano77e, Epp76, Fel76]. Seemingly [Ano72e, GS73c, KG70, MS76b, Rev74a, Rev74b, SU74b]. Segmented [GF73]. Selected [CCM71, Dah74]. Selecting [Baw72, BS73a, Can70, Eck76, Enn75, Gee72, NBH73, Nor76, Ofo75, Rot78b, SW71, Wac77]. Selection [AM75, BT78, Dav79b, Dud71, FEJM73, FEJM75, GE79, GR75, Goo73, Har74b, HURD72, Lea78b, NB71, NB72, Pat76, Pre78b, RS73b, RA76, SS78a, SBY79, Ste78a, You72a, GR77, Pre78a, RS74a]. Selectively [DD77]. Semi [WT75]. Semi-Metric [WT75]. Sense [Fis72, Tha79b]. Sensitive [Yea73]. Sensitivity [Bak74, Wor74]. Separate [Dye73, Dye74]. Sequence [Bly71, Bro71a, Dan71, Efr71a, Efr71b, Enn75, Fis72, Goo71a, Haw77, Pra71, Pra73a, Ste78a, Stu71]. Sequences [Hsu79]. Sequential [Alv78, BJ77, Bil77, Bin76, BM73a, Bre70, BH72, Car78, Cha74b, Cha71c, Cho71, Coo71a, Coo73, Eck76, EZ73, Gee72, Har75, HB72, HL75, Kat77, Kat78, Mil70b, Mil72, Muk76, Owe75, Pet78a, Pha71, PS75, Rea71, RS74b, Rom72, Sam70, SW76, SW77, SW78a, Tsa76, Wac77, WB71, Wei77a, Wei78a, Wol77a, Zac73]. Serial [BBC79, Sim75b, PH74a]. Serially [Ebb73a, Lee74]. Series [BJ77, BH79, Bha71c, Bor77, BP70, CL76, CT76, Den71, DF79b, Dor74, Ebb73c, Ebb73a, GGE74, Geo74, GG76c, GO72, Hau76, HB77, Hin79, Hol71a, JS75, Jon76, Kri72, Lee70a, LR71, Lit75, McL79, NNO76, Pie72, Pie73, Pie77, Pie79, SS74, Shu70, SA70, SS73b, Win74b, Wis72, You72b, SU74a, Hol72a]. Service [You72b]. Services [Gil74b]. Set [Bro76b, Bro76c, Goo73]. Sets [Agn74, Bau72]. Several [Bea77, Bro75, EG75, Fra73a, Gar72, Gar73, HB73b, Lev75b, LSW78, Ram77, Smi75b, Smi76, Lin74]. Severity [ZZ72]. Sex [Ceb74, Gor73b, MB74b]. Shape [Mos70a, WE79]. Shapiro [Che71a]. Sharp [TB79]. Shift [Haw77, Kno70, MR79, Sch76, Wor79]. Shifted [Arn70b, Arn73]. Shifting [FH70]. Shifts [Hsu79]. Shipments [BM73b]. Shod [Eck70]. Shoemaker [Eck70]. Shoemakers [Eck70]. Shooting [FK70]. Short [Epp79, Hua73, Plo79]. Short-Run [Hua73]. Short-term [Plo79]. Shortest [WT71]. Should [AS77]. Shrinkage [MS71]. Shuffling [Tho73b]. Sibson [Roh74]. Sided [Bil77, BS77b, DE73, Dud72, GG76a, HQ75, Kno70, Lem77]. Sign [Gas71]. Signals [Sob71]. Signed [Cla70c, CH74a, CH74b, Con73, Mil70b, Mil72, Rah74, WB71]. Signed-Rank [Con73, Mil70b, Mil72]. Significance [Arn70a, Bro76b, Bro76c, Cha76c, Joh75b, SM78c, SSM75, ST75, Vin76c, Zar72a]. Signs [Boh79, Qua79a]. Simmons [Lan75b]. Simon [Sti73b]. Simple [Bly71, BR77, Bro71a, Bur79, DC73, Dan71, Efr71a, Efr71b, Eng75, Gab78b, GS73a, Goo71a, Goo79d, Hay71a, HT75, JB73, LMH73, Nea75, Noe73, Pra71, Pra73a, Sam78, Sie78, Stu71, TB79, WG71, Woo71, PH74a, Rel72]. Simplification [GK72]. Simplified [Man74b]. Simplifying [Low77]. Simpson [Bly72a, Bly73]. Simulated [WC70, Yak76]. Simulating [CMS76, Mik72b, CMS87]. Simulation [All77, BL77a, Can75, CF72, DSW77a, DSW77b, EM77, Gav78, Gle76, Hoe77, Kel73, Kie78a, Mal78, Nel70, SM78a, SM78b, Sim78b, Smi75b, Smi77a, Thi77b]. Simultaneous [And74a, And76, Ang79, Ano78f, Atk78, Bow70, BG74, BB76b, Chr73, CZ75, DG75, DG76, DM76a, Far76, FH73, Gam77, Gre75c, Hah72, HQ75, HR77, Sri72c, Sri73, TB77b, DL74]. Simultaneous-Equations [Gre75c]. Single [Ban71, FM74, FM77, GH74, Gup76, Gup77, Hua73, Hub74b, Hwa74, JG75, Kle73c, Koh72, KS71b, Min72, Pyn79, Sam70, Sax76]. Single-Family [Hua73]. Single-Link [Hub74b]. Single-Sample [Sax76]. Single-Sampling [Sam70]. Single-Variable [Pyn79]. Singly [Kri72, Sim75a, Sim74a]. Singly-Ordered [Sim75a, Sim74a]. Singular [Ban72, HB73a, KN77].

Situations [MM71a, Tik71]. six [Ols74]. Sixth [Reg73]. Size [BM75, BDG78, Cha71c, CM76, Fab77, Fer78, GS77, Hor78, Kin72b, LG75, Mos70a, NR71, Ofo74, PD74, PF78, PS74, RV77, RB71, RH73b, San77a, San73, SS78a, Som70, Som74, Som75, WN74a, WM71]. Sizes [Ber74, CG78, Sha76b, SS73a, THP78, Wal70]. Skewness [SB75]. Slippage [Nea75]. Slope [DF72, FH70, Jol73, KW74b, KW74a, MS72b, Ram77]. Slopes [Ham76]. SLSEs [Low74]. Small [Atk78, Cha76c, CWR78, FH79, FM77, GC70, GH78b, Hoa75, HST79, Hom78, Hur72, Kad70, Ken76, Kin72b, Kno70, Kra76, Lar78, Lis72, Lit74a, Lit74b, Lit75, LSW78, ML74, McL77, Mer73, Mik72b, Mor78, MW70, Par75, Par76, Ram71, RG72a, Spi78, Spi79a, Spi79b, Sri72c, TH73, TG71b, TDL75, Yak76, Ano77d, Sri73]. Small-Area [GH78b]. Small-Sample [Atk78, CWR78, Hoa75, Hur72, Ken76, Lar78, LSW78, Mik72b, Par75, Par76, Spi79b, TDL75, Yak76]. Smaller [Fel74a]. Smallest [Ofo75, Som75, WN74a, Wal70]. Smear [Sco78]. Smear-and-Sweep [Sco78]. Smirnov [WN74b, Can75, GG76a, Wei78c, WN73]. Smoking [Pre70, Ste71, Ste73, Wil73]. Smooth [DWW79, KP79, TP79, Tob79a, Tob79b]. Smoothing [Cle79, Hob76, WT75]. Smoothness [GM78]. Social [Fel76, WNIO77]. Socioeconomic [BN73]. Software [FS77]. Solution [Fou77]. Solutions [BBS74, BRB76, Bjö78, Bli74, Esp77, Fle75, Sch70a]. Some [Agr76b, Ala74, Ano72e, Bha76, BJN70, BD78, Bly72b, Bly72c, Box79, BS75, CM74b, Coc70b, Con74b, Con74c, DG78, EHB78, EF76, Fow79, Fro73, Fro75, GC70, GR71, Gas74, GG72, Gho79, Gol75b, Goo72b, GY70, HS73, Har74a, Har74b, Har79, HS79, Hob76, HS75, Hog74a, Hog74b, Hub74a, HB73b, Jes73, JB73, JW78, JR71, KG71, Lau72a, Lau74, Leo77, Lin72, Lin78c, Low74, Low77, Man74c, MS76a, MG75, MJ79, Mie74, Noe73, OW72, Ofo74, OK76, Pra72, Pri74, RS73a, RG72a, RS76, RC76, Rev74b, Rev76, San74b, SM79, Sch76, ST76a, Som74, Sta75, SGS74, Ste74, TC72, TA78, TG71b, TS77, Tin70, Ton70, TDL75, Wam70, Wei71b, Wei73, Wei75, Win71, Win72b, Wol73, YBJ72, Yu75, SSM75, Rev74a]. Somerville [Ofo74, Som74]. Space [Che73b]. Spaces [And70a]. Spacing [Bof75]. Spatial [Mur70, Ord75, Sla73, Sla74]. Spatial-Temporal [Sla74]. Spearman [Hen80, CC73, CC74b, Hen79, Kra74, Ott73, Zar72a]. Special [BL77b, BL77c, CR75b, Pie77]. Species [Hil79, RS79]. Specific [Gup76, Key71, Lon70b, Gup77]. Specification [Cog73, Gal77, RZ71, RG72a, RS76, TS77, Thu79]. Specified [BT78, Bro75, Kwe75]. Specimens [Lit75]. Spectral [Ano77c, Bha73a, Lee75]. Spline [DGL77, EF76, Poi73]. Splines [BL77b, BL77c, Sam79, Poi74]. Split [LK78, McG77, Sch78c]. Split-Cable [McG77]. Split-Plot [LK78, Sch78c]. SPRT [CM76, KW79, Mad74, WT76]. SPSS [And78b, BF78a, BF78b, Dix78, McC78b, Mul78a, Mul78b, Nel78, NFJN78, Rya78, San78, Sch78b]. Spurious [Vea75]. Square [Ame73, Ame76, BE75, BR77, BM76b, Bur73, Bur74, Cha72, DG73b, DS74, DS73, Goo71b, GY70, HT72, HF70, KM74, Lis72, Lit74a, MG74, ML74, Odo70, RB71, SS77b, ZM71, Ame74]. Squared [Bai79, Coh76b, GM77b, Hoa77, KP78, Lar78, Moo77a, Sah76]. Squares [All77, Ame76, AS77, Ano79h, Bal70, Bea77, BL77a, Bjö78, Bra73, BL77c, Cha71a, CFY76, Dav77a, DSW77a, DSW77b, DP72, EM77, Ell73, Ell77, Fle71, Fle75, FHJ78, GK76, HM79a, Has72, Her76, HH77, HH78, Hoe77, Kel71, Lie76, McE76, Mik72a, Mis72, MM71b, Obe75, RW70, Roy76, Sam79, Sin71a, Smi77a, Spi79b, Sri72b, TT71, Thi77b, Wam70, BL77b]. Stabilities [BR70]. Stability [Bak74, Gar77]. Stable [CMS76, DuM73, DuM75, FR71, FS72, GLT79, Gup76, Gup77, HMW74, LP75, Pre72, SM79, TT71, CMS87]. Stably [DuM75]. Stacy [JK72]. Stage [AS77, Ano79h, BG77, BS79, CM76, Coo73, DP72, Gho75b, HT72, HS79, Kel71, Mik72a, Ofo75, Roy76, Sah76, SH76, Ven78, WK71]. Standard [BT78, DC73, Hah72, KF70, Lan73, Pet78a, SM79, TF73, WT71]. Standardized [Ell73, Ell77, MY76]. Standards [GOST75, WD76]. Starting [Obr79]. Starts [Hua73]. State [MMS71, NL79]. Statements [Bar77, Bro77, Bue77, Dem77, Kem77, Kie77a, Kie77b, Lin77a, Wol77b]. States [Coh76c, Lee70b, MB74a, Kob73, Lon73, SP76]. Stationary [Bor77, BG70, Fis72, Hau76, HT79, Jon76, Ros72, Shu70, Sob71, SU74a]. Statist [CMS87, FH81, JTB81, Krz81]. Statistic [Bai77, BE75, CG78, CH75, Cla70c, CH74a, Coh76b, GG76a, GGM70, GGM71, GC78, Hay71a, KTS75, LS79, MS74, Rah74, RR71, RB71, Wei78c, WN73, Cro74, CH74b, WN74b]. Statistical [And70c, Ang74, Ano70m, Ban71, BK79, BH08, Bir72, BYJ73, BG74, Chr79, Chr73, Cro79b, DuM73, DuM75, Edd79, Fel72b, Fie72, GCZH73, Has72, Kle77a, Lin79a, NL77, Par79c, Par79d, Pel71, Reg73, Rob79, Rot79, She78, SF72b, Tar79a, Vil77a, Wel79, ZF73, Zel79a, Zel79b, Fel72a, Tuk79a]. Statistically [Gav78, Kie78a, Mal78, SM78a, SM78b, Sim78b]. Statistician [Tuk79b, Cor75, Fra76a]. Statisticians [Eck70, Kis78, RR77]. Statistics [BOLK75, Bau72, Bof75, Box76, Box79, BS75, Can75, CCM71, Dav73a, Fra76a, Gal75a, Goo75b, Haw71, Hog72b, Hut72, KN75, Kis78, KV73, Lar78, MS76c, MW70, Pat75a, RH71, Roh76, SB75, Sie78, Ste74, Sti76, Sti77c, Uth70, Van76, Vit79, Wal70, Wol73, Kle77a, KV75]. Status [BN73, Mil76, SL79]. Steady [MMS71]. Steady-State [MMS71]. Stein [BG77, EM73, EM75, FH79]. Step [Bic75]. Stepwise [BA77, CA79, HR70, Wel77]. Stochastic [Ali74, Ano77e, BM77a, Bli74, BG70, CR72, CF72, DL72, Epp76, Koh72, Lee74, LW76b, Lis72, Lor71, Obr79, PW74, Qua74, Ros72, Sho70, Sho72a, Sho73, UB70, VG78]. Stock [BD73, Bha73b, Epp79, FS72, HMW74, Hsu79, LP75, Off72, Tei71, ZF73]. Stock-Market [HMW74]. Stopping [BA77, CA79, FS77, FEJM73, FEJM75, Hoe72, RS79]. Straight [HG71a, War72]. Strata [AKC76, KS71a]. Strategies [Ber78d, Lea78b]. Strategy [GO72]. Stratification [AKC76, BHF77, KA78, RH73b, Sin71b, Tho77, Ano79g]. Stratified [Lee73b, Lev77, McK75c, Nat72, SS71, Sir72, ST75, SY72]. Stratifying [Tho77]. Straw [Fie72]. Strength [Hal75, BJ74]. stress [BJ74]. stress- [BJ74]. Strictly [Wri78]. Stringent [Bea74, Kwe75]. Strong [Bar76a, BT76, Dic76b, Fra76d, Lin76a, Pra76, Sto76a, Sto76b, Vil76]. Stronger [CH78b, HW75]. Structural [BBDS71, BBDS72, Dol76, Hur72, Mal70, MM79, Max73, McD74, Mut79, RR71, Roh76, Sch76, ZP79]. Structure [Coh76c, FB73, Hin79, Mer73, Nel70, Pel71, SB78a, WSB73]. Structures [BM78, SM77a]. Student [LL72, Zel76b, LL71, Zel76a]. Student- [Zel76b, Zel76a]. Studentized [BT74, Ste70a, Sto78]. Studies [GS73a, GGS73, McC76b, Rub79, SBY79, TL70, WSB73, Cro74, JTB80, JTB81]. Study [All77, BMB78, BR70, BL79, BL77a, Can75, DSW77a, DSW77b, EM77, EG75, Eri76, Fie75b, Gar72, Gar73, GS73b, Gil72, Hoe77, KMS75, LMSW70, LS75, Lin78c, MM73, McK75b, McK76, Mik75, MR70, Nat76, Ram77, RG72a, Smi77a, Thi77b, Thu79, Vin76b]. Studying [Fow79]. Subaggregate [DWD76]. Subclass [Koc73]. Subclasses [SS71]. Subdomain [Laa79]. Subgroups [HST79]. Subject [Ano79s, Ano79t, Ano79u, Ano79v, CA77, Dav78, HG71a, McC72, Pra70, War72]. Subjective [Edw75, HB72, Hog75a, Hog75b, Rub77, Win75d]. Subjects [HL73]. Subpopulation [Car78]. Subsample [Sha76b, THP78]. Subsequent [ST75]. Subset [AM75, Gal75c, Rot78b]. Substitution [Wei77c]. Subsurvival [Pet77a]. Subsystem [Man74b, Win74b]. Sufficiency [And70a, Dur70]. Sufficient [Mal72, MTC76]. Suggestion [Man74c]. Suggestions [Dag71, Gas74, Har74a, Hog74a, Hog74b, Hub74a]. Sum [BN75, Ham71, Jam77, Mar76, McC73a, Nau73, SS77b]. Sum-Constrained [BN75]. Summaries [Ano70m]. Sums [Ell77, Gre70, Gre71]. Sun [Fie72]. Sunday [Tho75]. Sunspots [PW74]. Superpopulation [BS79]. Supply [Hol72b, Sed77]. Sure [Bly72a]. Sure-Thing [Bly72a]. Surface [Atw75]. Surfaces [HQ75, Kar70, KM74]. Surveillance [Che78]. Survey [BMB78, Bha76, GOST75, Hoa74, KL72, Koc73, Lau72b, Min72, Pro71, SB78b, Wyn77, OS70]. Surveys [Bai75, Bor70b, Bre79, CH78a, FGHA73, MR70, Raj77, Rub77, SS74, Sen76a, Sir70, Sir72, SB73, WM70]. Survival [BBC79, BH72, BT72, CH77, GS73a, HW74, KJT72, PF78, Sly72, SV76, Tar79b, TG75b]. Survivorship [Tur74, TBH74]. Sweep [Sco78]. Switching [Bin78a, Bry78, CH78c, Fer75, Fow78, Har78, Hos78, Joh78c, Kie78b, KNP79, Qua72, QR78a, QR78b, QR79, SM75]. Symmetric [DuM73, FR71, Gov71, Gov76, Gro76, JR71, Meh72, Noe73, Sch73d, Sch74b, TL70, WD72]. Symmetry [EE70, Fin77a, Gas71, WB71, Wol74a]. Synthesis [You72b]. System [And74a, Ber74, Che78, Eas72, GS74a, GCZH73, Ham78, Man74b, Par71, Vin76b, WNIO77, Win74b, WM71, DL74, And76]. System- [Man74b]. Systematic [FH76, MY78, Mik71, WM70]. Systems [Bar76b, EP70, HS73, Kel71, Kwe75, LR71, Mar71, Qua74, Rev76, ZF73, Mar72].

t [Lin78c, Roh76]. Table [Ano78e, BM77a, Cla70c, FG70, Gam77, Gol75a, Gup73, IJK76, Joh75b, LL71, LL72, SL79, TG77, Tho70, Sim74a, Sim75a].

Tabled [How74a]. Tables [Bai77, BM76b, BB77, CA77, Chi72, CR75b, Coh76b, Con74b, Con74c, DS73, Fie70, GM77a, GJ78, Goo71b, Goo73, Goo75b, Gre70, Hom78, Joh75a, KJT72, KVK71, Lee77a, LK76, Man74c, Mie74, Nat72, Odo70, Pat74, Rah74, Sim78a, SGS74, Sto78, TB72, Tik72, Wag70, Whi70, WG72, WT71, WN73, ZR71, ZR74, Chi73a, Gre71, WN74b]. Tail [DL76, DeG73, Dic77b, GL71, Hom78, Lin78c, Som76, Woo76, ZM71]. Tail-Censored [DL76]. Tailed [DuM73, Gro76, Smi75b]. Tailored [Lor71]. Tale [Sam74]. Tau [Sim77]. Teams [Har77b]. Technical [Ano72r, Ano72s, Ano72t, Ano73n, Ano74m, Ano74n, Ano76n, Ano76o]. Technique [Bor70b, Fro75, GP75, GG75, GKAH71, LCM75, MM72, Nea75, Sch73b, SW73, SB78b]. Techniques [Bak74, GM75, Hub74b, LMSW70]. Telephone [Sud78]. Temporal [Sla73, Sla74]. Temporally [Ano77e, Epp76]. Ten [CS73]. Term [Mis72, Nel70, Plo79]. Terms [Lee77a, McC73a, Zel75, Zel76a, Zel76b]. Terry [Bea77, Dav70]. Test [Arn70a, BG73, BS74a, BJ73, BJ70, BYJ73, Bre70, BR77, BB77, Cha71c, CG78, Cha72, Che71a, CA77, Cho71, Con72b, Con73, DG73b, EE70, FH70, Fin77a, GG78, Gal75b, Gas71, Ger75, Gla76a, Gla76b, GG75, GR74, GP74, HW74, Ham76, pH78, HM79a, Has72, HD77, Het73, HT72, HG70, HG71b, HL73, HFR75, Hol70a, Hol70b, HM70, Jen70, Jon79, Ken76, KW71, LM78, LW76b, Lem78, MY78, MA79, MS74, Mil70b, MS76c, Ols73, Par75, Par76, PH73, Pat74, Pat75b, PL76, Pha71, PH72, PH74b, PS75, Pyn79, RA75, Rea71, Rob75, Sam76b, SF72a, Sho72b, SW79, Som77, ST75, TB70, TH73, TB77a, TP79, Tik71, Tik72, THP78, Uth70, WC70, WN73, WN74b, Wol77a, Wor79, Ylv77, PH74a, HBA72, Mil72]. Testing [Arn70a, BBH72, BBC79, Ber74, BS77a, BS74a, Coh74a, Con72a, CG74, Dix76, Dye74, FS77, Gal75c, Gal77, GH74, GP75, GR74, Har71, Haw77, HR70, HB74, Hwa72, Hwa74, Hwa75, Joh75b, Kad70, KS71b, Lin73a, LS75, Lor71, LP70, McG77, Mee74a, Mee74b, MS70, NNO76, Nel70, Owe75, PE78, Rob78a, RY78, Rom72, SZB79, Sin71c, Vit79, YS77, Zar72a, Ano79f]. Tests [AGA79, Ala74, And71, Ano77d, Bea74, BC79, BJN70, Bil77, BS77b, Bro76b, Bro76c, BF74a, BF74b, Cha74b, Cha76c, CS74, CWR78, Dav71, DE73, DS73, Dem73b, DY78, Dun74a, DC71, Eve79, FK76, Fli79, Fra73b, Fry70, GR71, Gil72, Goo75b, GH76b, Gro72, GY70, Hal75, HJ76, Her71, HS79, JB73, Kal78a, Kno70, KP79, KP78, Lau70, Lay73, LG75, Lee77b, Lee79, Lev75b, Lis72, LF71, LF72, LF73, LSW78, MSD73, MA73, MS76a, McG73a, McG73b, Mie72, Moo77a, Nat72, O’B79, Odo70, PF78, PAA70, Pra73b, RZ71, Ram71, RG72a, RS76, RS74b, Sam70, SM79, She78, Sim75b, SW78b, Smi75b, Smi76, SH75, SH76, Sri75, Ste70a, Ste72, SM78c, SSM75, Tho79, TS77, Thu79, TDL75, TY74, Upt76, Vin76c, Wan71, WB71, Wis72, Yak76]. Tests [Yar70, Yu75, Joh78b, Ols74]. Theil [GS72, MS74, YBJ72]. Their [Ano78e, Jag73, TF73, TG77]. Theorem [Con74a, Dur70, LS77a, WD76]. Theoretic [BH76, FM74, Ven78, Win72a]. Theories [Adi73]. Theory [BK78, Dia78, Est72, Gas74, HS74, Har74a, Hog74a, Hog74b, Hol70b, Hub74a, Iji71, Jos76, Lin73c, LK76, RS73a, SW76, SW77, Sub72, The74, Ano79e, GK70].

There [Dun71, SS78a]. Thereof [Pie77]. Thing [Bly72a]. Thompson [Rao73, Rao71b]. Threat [LSB76]. Three [Boh79, Eck76, Goo75b, Klo73, McD74, Muk76, MW70, Pat74, PB76, Sah76, Vit79]. Three-Decision [Boh79]. Three-Dimensional [Pat74]. Three-Factor [Goo75b]. Three-Parameter [MW70]. Three-Stage [Sah76]. Tiebout [Ceb74]. Ties [Bof76, Con73, Dav70, KT77, Rob72]. Tiku [TW77]. Time [Ano79f, BJ77, BH79, Bor77, BP70, CR72, CG73, CL76, CT76, Dan73, Dav73a, DF79b, Dor74, GGE74, GS73a, GO72, Hau76, HB77, Hin79, Jon76, Lee70a, MB74a, McL79, NNO76, Pie72, Pie73, Pie77, Pie79, SS74, Shu70, SB73, SM77b, YS77, You72b, SU74a]. Time-Series [BJ77, You72b]. Time-Truncated [Ano79f, YS77]. Times [BB76a, BH08, Dob78, FK70, Odo70, Fie72]. Tolerance [Ber77, Kap77, Lem77, TB79, Zac70]. Toll [BL79]. Total [Bin76, FNL77]. Totally [SW79, You76a]. Totals [Den71, ST76a]. Tournaments [CH78b]. Toxicological [Gla79]. Trade [Lue71, NMC71]. Trading [Car70]. Traffic [Hsu79]. Trailer [DS78]. Training [Gor73b]. Tram [Jag73]. Transaction [Ano77e, Epp76]. Transformation [Fuc79, GL71, Poi78, Sch73a, Spi78, Spi79a]. Transformations [FB73, Lan74, Lan76]. Transient [MB74a]. Transition [Hen71, Lee74, Wor74]. Transitory [Hol71a, Hol72a]. Translation [BOLK75]. Transplant [CH77, MB74a, TBH74, Cro74]. Transportation [DS78, Wes79, You72b]. Treatment [CD72a, CD74, EHB78, Pet78a]. Treatments [Add70, Can70, MI78]. Trend [AGA79]. Trends [Con72a]. Trial [Sti77b]. Trials [Ber78d, GS73b, HL75, KK79, Klo76, Lad75, Pet78a, Wei77a, Wei78a, WD78]. Trigonometric [Tar79b]. Trimmed [Pre78a, Pre78b]. Trimming [Pre78a, Pre78b]. Truncated [AE72, Ano79f, DG73a, KW71, KW74c, Klo76, San77a, Sel74, TG71b, YS77]. Truncating [Mad74]. Truncation [Alv78, DQ72, Lau74, SMRL70]. Tukey [KGR79, Fen79, Hoa74, JR71, KTS75, KR78b, McD72]. Turning [Lon70b]. TV [McG77]. Two [AS77, Ano78e, Ano79h, AF76, BE75, Bak74, BT78, BS73a, Ber78d, Bil77, BG77, BG74, BB77, Bur79, BS79, Can70, Can75, Car77, CA77, Cle70b, CR75b, CM76, Coo73, CF72, DE73, Dem73b, DP72, DG75, Dud72, Ell77, EM78, Eve79, FG70, FK76, Fli79, FGHA73, Fow79, Fra73b, Fry70, GG76a, Gar77, Geo74, Gho75a, Gho75b, Gol75a, GR75, GR77, Goo75b, GW79, HW74, Ham71, HG73, Hau76, HB77, HJ76, HT72, HS79, HH77, Hin71, HFR75, How74a, Jen70, Joh75a, JG72a, JG72b, Kau72, Kel71, KTS75, KW71, KW74c, LG75, LF72, Lit75, LB72, MS76a, MS72a, McD74, McG73a, McG73b, Mee79, MTC76, MI78, Met71b, Met71c, Mie72, Mik72a, Nor76, Ofo75, PL76, PAA70, PS76a, Pra73b, Qua74, Rao75, Rev74a, Rev74b, RR71]. Two [RS74b, Roh76, Roy76, Sam74, Sch73a, Scl79, SH73, SW71, SH76, SA70, Sri74, SM78c, TC72, TA78, TG77, Tho77, Uth70, Vit79, WK71, Wol77a, WEB78b, Zac73, ZG72, Fox74, Isa74, HH78]. Two- [SA70]. two-arm [Fox74]. Two-Armed [Ber78d]. Two-Associate [Cle70b]. Two-Compartment [Met71b, Met71c]. Two-Dimensional [Scl79]. Two-Factor [GW79]. Two-Group [GR75, GR77]. Two-Level [TA78]. Two-Moment [BE75]. Two-Parameter [WEB78b]. Two-Party [Qua74]. Two-Phase [Hin71, Rao75]. Two-Sample [Can75, CA77, Eve79, FK76, Fli79, Fry70, GG76a, HW74, HFR75, LB72, MS76a, Mie72, Wol77a]. Two-Sided [Bil77, DE73, Dud72]. Two-Stage [AS77, Ano79h, BG77, BS79, CM76, Coo73, DP72, Gho75b, HT72, HS79, Kel71, Mik72a, Ofo75, Roy76, SH76, WK71]. Two-Way [BG74, BB77, Bur79, CR75b, Ell77, HG73, JG72a, JG72b, Sch73a]. Type [DY78, MG75, PE78, SS77a, WL77, Woo76, Yu75, Poo74, BS73b, Wan71, WEB78b]. Typed [HL73].

u [Lee77a]. U.S [AG72a, Mer71, Wei77c]. U.S. [BW76, LS73]. U.S.A. [Mos70b]. U.S.S.R [Maz72]. UMP [Bre70]. Unbalanced [CG74, GH78a, GH79, Hem79, SHH78, Wal79a]. Unbiased [AE72, Blu74, CCM71, CL76, Far76, KN75, LaM73, LaM77, PS75, Sha74, SH73, Str74, Str75, War75, ZM71, Sen74]. Uncertain [Atw75, Hil74b]. Uncertainty [Ger77a]. Unclassified [McL77, O’N78]. Unconditional [Law73]. Uncontrolled [Kin72b]. Uncorrelated [RH71, Wol73]. Undecided [PY74]. Underlying [Ste78a, HURD72]. Underreporting [War78a]. Understanding [Kle77a]. Unemployment [CM74a, GH78b, MB74b]. Unequal [AS76, BR70, CG78, DS73, Ham76, HJ73, Hoc76, KTS75, LG75, MD74, MS76d, SS73a, Tam79, Vij75]. Ungrouped [OW72]. Unguided [Goo73]. Unified [Che75]. Uniform [Bar76a, BT76, Coo71a, Coo73, Dic76b, Fra76d, Lin76a, Pra76, Sto76a, Sto76b, Vil76]. Uniformly [Tak71b, Zac70]. Unimodal [Kri79]. Unimodality [KG71]. Union [SW79]. Union-Intersection [SW79]. Unique [Fou77]. Uniqueness [TG75a, TG79]. Unit [DF79b, Sud78]. United [Kob73, Lon73, SP76]. Units [Add70, KS71a]. Univariate [Bea77, Fin77a, Hau76, McL79, Pre72, Ste79]. Universes [TW77]. Unknown [Het73, McE76]. Unrelated [Ano72e, GS73c, KG70, Lan75b, MS76b, Moo71, Rev74a, Rev74b, SU74b]. untitled [A.73, A.74a, A.74b, A.74c, AY76, Abo78, Abr79, Ada76, Ade78, Agr76a, Agy78, Aig71, Aig72, Aig75a, Aig75b, Aig75c, Ait72, Alb76, Alb79, Ame70, AZ79, Anb76, And70b, And77, And79, And74b, Ang78, Ano70n, Ano70o, Ano70p, Ano70q, Ant73a, Ant79, Ant73b, Arc73, Arn75, Arn76, Arn78, Arr70, Atk76, Atk79, Atw76, BJ72, Bal74, Bal75, Ban78, Bar79, Bar70a, Bar70b, Bar74, Bas78, Bau71, Bea76, Bec76, Bec73a, Bec73b, Bec74, Beh73, Ben75, Ben72, Ber71a, Ber73a, Ber78a, Ber73b, Ber70, Ber76, Ber78c, Ber78b, Bha71a, Bha79b, Bha79a, Bic73, Bin78b, Bir78a, Bir78b, Bla71, Blo77a, Blo71, Boa73, Boe71, Boe72, Boh75, Boh71, Bon79, Boo77, Boo79, Bor70a, Bou77, Boz73, Bra70, Bri71, Bri72, Bro78a, Bro78b, Bro71c, Bro71b, Bro71d, Bro72a]. untitled [Bro72b, Bro72c, Bro79, Bro73, Bru73, Bru78, Bue76, Bue79, Bul72, But71, But79, But76, But77, C.71, C.72a, C.72b, C.72c, C.74a, C.74b, Cam79, Cas78, Cha76a, Cha79, Cha73, Cha75a, Cha74a, Cha78, Cha76b, Cha75b, Cha70b, Cha71d, Che73a, Chi71, Cho76, Cla70a, Cla70b, Cla71, Cle70a, Cle71, Coc70c, Coc70a, Coh71, Coh74b, Coh78, Col75, Col72, Con76a, Con76b, Coo74, Coo71b, Cox73, Cro79a, Cro71, Cru75, Cza70, D.72a, D.72b, D.73a, D.73b, D.74a, D.74b, D.74c, D.74d, D.74e, D.74f, D.74g, D.74h, D.75a, D.75b, D.75c, D.75d, Dal71, Dal77, Dal78, Dal79, Dav72a, Dav72b, Dav73b, Dav76, Dav77b, Dav79a, Dav79c, Daw76, Day79, DeG70, DeG71a, DeG71b, DeG75, DeG76a, DeG76b, DeL79, DeW71, Dea79, Dem73a, Den70, Den73, Des70, Dev78a, Dev78b, Dev79, Dic70, Dic73]. untitled [Dic77a, Dok71, Dok72, Dou71, Dra73, Dra71, DG72, Dud78, Dun75a, Dun76a, Dun77, Dun79d, Dun79a, Dun79b, Dun79c, Dun74b, Dwy73, E.74, Eas77, Eat73, ET74, Eis73, EJ75, Ela75, Ela77, Els77, Eng72, Eni71, F.74, F.75, Fab70, Fai77, Far70, Far73, Far72, Fed70, Fei75, Fel79, Fie75a, Fie76a, Fie76b, Fie76c, Fie77a, Fie77b, Fie77c, Fis77, Fis70, FL76, Fox76, Fra70, FF78, Fra78, Fra79, Fra71, Fre78, Fri70, Fri78, Fry71, Fry72a, Fry72b, G.71, Gal73, Gan79, Gan71, Gar78, Gav70, Gav73a, Gav73b, Ger72, Ger77b, Gib73, Gil74a, GE75, Gil76, Gil78, Gle73, Gle74, Gle75, Gli78, God74, God70, Goe74, Goe77, Goe78, Goe79, Gok79, Gol70, Gol76, Gol77, Gol72, Gol78, Goo72a, Goo78a, Goo78b, Goo78c, Goo79a, Goo79b, Goo79c]. untitled [Gor73a, Gor76, Gov70, Gre73, Gre75b, Gre74, Gui78, Gul75, Gun79, H.71, H.72a, H72b, H.75a, H.75b, Hab77, Hag78, Hai70, Haj74, Hal74, Ham75, Han75, Har70, Has76, Hay78, Hay76, Hea79, Hen72a, Hes71, Hey71, Hey78, Hil70a, Hil74a, Hil72, Hil73, Hil76, Hil78, Hin77a, Hin76a, Hin76b, Hin77b, Hix76, Hoe76a, Hoe76b, Hog79, Hog72a, Hog78, Hol75, Hol74, Hoo74, Hor76, How74b, Hsi70, Hud76, Hud75, Hum75, Hun72, Hun75, Hun70, Hur73, Hur74, Ign74, Iji70, Imr76, Imr78, Imr79, Ing77, Ism76, Ize78, Ize79, Jed73, Jen76a, Jen76b, Joh72, Joh73b, Joh73c, Joh78a, Joi73, Jon73, K.71, K.72a, K.72b, K.72c, K.73, K.74, K.75, Kad73, Kad75a, Kad75b, Kai72, Kal78b, Kap70, Kas74, Kat70, Kau73, Kei72, Kem70, Kem72, Kem73, Kha77, Kim77, Kin72a]. untitled [Kin76, Kin77, Kin78, Kin79, Kle77b, Kle70, Kle74, Kle73b, Kme74, Kme77a, Kme77b, Kme77c, Kme77d, Kme79a, Kme79b, Koo77, Koo72a, Koo72b, Koo76, Kot70, Kot72, Kot73, Kou72, Kro71, Kup76, Kup78, Kup79, Kur72, LaC79, LaV75, Lac74, Lac76, Lac77, Lah75, LS71, Lan75a, Lan79, Lap75, Lar74, Lar75, Lav79, Lay75, Lea78a, Leb74, Leh77, Lev75a, Li72, Lil75, Lin78a, Lin73b, Lin77b, Lin78b, Lin79b, Lin79c, Lin79d, Lin76b, Lov76, Lov78, Luc72, Lyt72, M.72, M.74, M.75c, M.75a, M.75b, Mad73, Mag72, Man72, Man74a, Man70, Man71, MYH76, Man75, Mar73, Mar78, Mar70a, Mar70b, Mas79, Mas72, McC74a, McC75a, McC76a, McC73b, McC74c, McC78a, McC74d, McC75b, McC79a, McC79b, McK75a, McK77, Mee78, Met71a, Mey72, Mey75a, Mey75b, Mic76, Mie71, Mil70a]. untitled [Mil75, Mil79, MW79, Min71, Mir73, Mit71, Moo72b, Moo72a, Moo72c, Moo73a, Moo74b, Moo74a, Moo75b, Moo75c, Moo75a, Moo76b, Moo76a, Moo76c, Moo77b, Moo78a, Moo78b, Moo78c, Moo79, Mor77, Mor71a, Mor73, Mor70, Muh73, Mun77, Nag79, Nas79, New76, New79, Noe70, Nol73, Nol74, Nor71, Van78, O.70b, O.70c, O.70d, O.70e, O.70a, O.71, Oak74, Ode71, Ode79, Ore79, Ork72, Ork75, Owe71, P.72, P.75a, P.75b, P.75c, Pac70, Pag70, Pai77, PR75a, Par79a, Par73, Pas72, Pas73, Pas76, Pat73, Pec71, Pet77b, Pet78b, Pet79, PA76, Pir73, Plo71, Poi79, Pol73a, Pol75, Pov79, Pra74, Pre76, Pre74, Pre78d, Pro77, Put78, Put79, Q.74a, Q.74b, Qua77a, Qua78, Qua77b, Rag71, Rag75, Rag78, Ram74, Rao77b, Rav78, Rel76a, Rel76b, Rho75, Rho77, Ric76, Ric72, Rob74, Rob78b]. untitled [Rod76, Rog78, Roh77, Roh78, Rol72a, Rol72b, Rol76, Ros74a, Ros74b, Ros73, Rot78a, Rot71, Rub74a, Rym71, S.73, S.74a, S.74b, S.74c, S.75, Saa71, San74c, San77b, Sär72, Sär79, Sch79a, Sch71, Sch73c, Sch78a, Sch79b, Sch70b, Sch75a, Sch75b, Scl73, Scu70, Sed78, Sel75, Sen71, Sen72a, Sen75, Sen76b, Sen76c, Sen77, Sen78a, Sen78b, Sen78c, Sen79a, Sen79b, Sen73, Sen79c, Sha75, Sha78a, Sha77b, Sha78b, She79, She72, She75, Sho76, Sho72c, Sho75, Sie72, Sim79a, Sim79b, Sim76, Sim72a, Smi71a, Smi71b, Smi77b, Smi77c, Smi79a, Smi79b, Smi79c, Sok72, Sri72a, Sta78, Sta79, Ste75, Ste70b, Sti78a, Sti73a, Sti75, Sti77a, Sti78b, Sti78c, Sto70, Str72a, Str76, Str71, Suc74, Sun79, Sut78, Swi73, Swi76, Swi78, Syk74, Tae76, Tae77, Tak71a, Tam72, Tan78, Tat71, Tay74a]. untitled [Tay74b, Tha79a, Thi77a, Tho73a, Tia75, Tro75, Tru70, Tru77, Tru78a, Tru78b, Tuc78, Tuf70, Tuf79, Urq75, Van72, Van73, Vin76a, Wac79, Wah71, Wal79b, War74, War79, War77, War78b, Was79, Wat79, Wat75, Wat77, Wat78, Web76, Web78a, Weg74, Wei74, Wei76a, Wei76c, Wei76b, Wei77b, Wei78b, Wei71a, Wel73, Wet78, Whe73, Whi76, Whi79a, Wic79a, Wic79b, Wig70, Wig77, Wil76, Wil74, Win74a, Win75a, Win75b, Win75c, Win76a, Win77, Win78a, Win78b, Win79, Win78c, Win76b, Wio75, Wip71, Wol79a, Wol74b, Wol74c, Yac74, YS72, You70, You73, Z.72, Zab76, Zab78, Zah77, Zar70, Zar72b, Zys73, Zys75]. Unweighted [HG73]. Up-and-Down [Lit74a, Lit74b, Lit75]. Updating [BT74, Cha71b, GLT79, Har75]. Upon [Bro78d, Wyz78]. Upper [Mor72, SC75b, Wei72, Wes77, Zac70]. Urban [Gre75c, Haf78, Wes79]. Urn [Wei78a]. USDA [Bha71c]. Use [Bro76b, Bro76c, Che73b, CH78a, Geb71, Gla79, GG75, Har77b, Hoc77, LCM75, McC76b, Poi78, RS76, RO76, Rev76, RB71, Ste76a, TL70, Wei78c, Cro74, TB72]. Used [Ano79f, MM73, Tho77, Wam70, YS77]. Useful [Con74a, Dic77b, Gle76, Har79, KMS75, RS73a]. Usefulness [Met71b, Met71c]. Uses [Sud70]. Using [AKC76, Bau72, Bro78c, CS74, Cle70b, Con74b, Con74c, Cra73, Cra76b, EM75, EM78, GH74, GM77b, Haw74, Hem74, Hog75c, Krz75, Lee73a, MB74a, Man74c, Mar76, Maz72, Mie74, OW72, OK76, Poi73, Poi74, Qua79b, RBRH78b, Rub79, Sah76, SS74, SS79, SS78b, Sri71b, SZ73, SGS74, Sud78, Dev73b]. Usual [Arn70b, Arn73]. Utility [Hil74b, NL79]. Utilization [Dag71]. Utilizing [MSW74, Meh72].

Vacancy [OS70]. Validation [McC76b, PD74]. Value [Blo77b, Cle70b, KQ70, KQ72, Lit74b, RW77]. Valued [Sta75, TG75a, TG79]. Values [CD72a, CD74, CGD76, FK70, GG76a, Gla76a, Gla76b, GK76, Rah74, SJL76, WN73, WN74b, HM73]. Variability [Add70, BM76a, GW79, Pes72]. Variable [Ame73, DF72, Dor74, FEJM73, FEJM75, Hen71, JG75, Kin72b, Koh72, LS75, May73, NM70, Poi78, Pyn79, RH73b, Sin71b, Woo79, Poo74]. Variables [CMS76, CA74, CS72, Cor76, Cor71, Fel74a, FM74, FM77, Gho75a, GG76c, Gre70, HG71a, Han73, How74a, Jam72, Jam77, KW74a, Kel71, KR76, KR90, KW74d, Krz75, MS72a, McD74, Mee74b, Moo73b, Mos70a, Mut79, NR71, Pra70, PS76b, RW70, RR71, Sau74, Sho70, Sho72a, SL74b, SS77b, SJL76, Tho77, WS78, Wal79a, Wal74, War72, WWF74, CMS87, DL74, Rel72, Gre71, KW74b, SL74a]. Variance [AE72, Ame73, BMB78, BR70, BG74, Bro78c, BS78, BS79, CM74b, Coh72, Cra76c, Cra76d, CG74, DE73, Eri70, Far76, Fel74c, Fra73a, Gil72, GO72, GW79, Gre75a, GJ75, pH78, HN77, HG73, Har77c, Har77d, Hem74, Hoa75, HM70, Jon76, KS75, Kle73c, LaM73, Lan73, Lee72, Lee73a, Lee73b, LK78, LM71, LS77b, LS79, MM73, ML74, MJ78, McD74, Meh73, Ofo75, Rao72, RV77, Rao77a, RC78, SA73, Sah76, Sam76b, Sax71, Sch73a, SH79, SF72a, Shu70, Som75, Sri72c, Ste78b, THP78, Vij75, WT71, Woo71, WC76, Zac73, Ols74, Pri74, BW70, Fel74b, HN79, Sri73]. Variances [AM75, BF74b, Car79, CG78, Dol76, Gar72, Gla76b, GK72, HJ73, Hei75, Hoc76, HH75, HHD75, Joh75a, Lay73, LG75, Lee77a, Lev75b, Mal72, Mor71c, O’B79, PL76, Pot72, Rao70, Sam78, Sen72b, Sho73, Str72b, Str72c, Str73, Tam79, TW77, ZG72, Lin74, BF74a, Gar73, Gla76a, Str73]. Variate [Sta75]. Variates [GG71, Gle72, GR75, GR77, GP74, Ham71, Mor71b, Tho76]. Variation [Efr78, GH76a, Mur70]. Varietal [You76b]. Varying [Sly72, SM77b]. Vector [GS73c, SW79, SM78c, TG75a, TG79, You76a, Wor74]. Vector-Valued [TG75a, TG79]. Ventures [Hil74b]. Verdicts [GS73b, GS74a]. Verification [Min72]. Version [KW74c, Lan75b]. Versus [Bur79, DWD76, Law73, SW73]. Very [Epp79, Sud72]. Viewpoint [DH72]. Visibility [CM74c]. Vital [Coh76c, Raj77]. Vocabulary [McN73]. Volume [Ano70r, Ano71o, Ano72u, Ano73o, Ano74o, Ano75k, Ano76p, Ano77l, Ano78n, Ano79e, Ano79w]. Volumes [Ano77f, Ano77e, Epp76]. vs [Obe75].

W [Hoa74]. Wages [CM74a]. Waiting [Dav73a]. Wald [CM76, HD77, Mar76, Moo77a]. Walk [Kel73, LM78, WM77]. Wallis [KT77]. Walpole [Kle77a]. Warner [WF79]. Way [BG74, BB77, Bur79, CR75b, Ell77, HG73, Her71, Hoc76, JG72a, JG72b, KT77, Lem77, LS79, Sch73a]. Weakened [Sav70]. Wealth [Fel76]. Weather [Bra79a, Bra79b, CH79, DS79, Ell79, Flu79, Gab79, Kru79, Mie79, Mos79, Ney79, Sim79c]. Weibull [Bas73, BP73, Law73, PR75b, Pat75a, RAK74]. Weighing [Ban72]. Weight [BNW70, Can75]. Weighted [Ame76, Bea77, BW70, Bra73, Cle79, Obe75, Qua79b, Sie78, SS77b, Wol74a].

Weighted-Symmetry [Wol74a]. Weighting [HB73a]. Weights [Gil74b]. Welch [Wan71]. Well [Uth70]. Well-Known [Uth70]. Where [How74a]. Which [AS77, HF70, Jam75]. White [Chi76]. Whither [Kle71]. Who [Eck70]. Whose [GF73]. Wide [Fis72]. Wide-Sense [Fis72]. Widely [Wam70]. Width [Car77, CD76, SW76, SW77, Zac73]. Wife [Eck70]. Wiggins [Dun75b]. Wilcoxon [CH74b, CH75, Cla70c, CH74a, Con73, HW74, HL73, SH76]. Wilderness [RO76]. Wilk [Che71a]. Wilks [RR76]. William [Kle77a]. Win [CH78b]. Window [Nea72]. Winner [Hoe72, HL75, KW74c, Nor76, PS74, SW71, WD78]. Wishart [Gle76]. Withheld [San73]. Within [Cla70c, Sud78, Hei75]. Without [AS76, Fre77]. WLSEs [Low74]. Women [Gup76, Gup77]. Word [Sic75, Che71b]. Words [Par79b]. Work [Mil76]. Workweek [TG71a]. Worse [Eck70].

x [BE75, CT76]. X-11 [CT76].

Yates [Con74b, Con74c, Man74c, Mie74, SGS74]. Year [Bro76a]. York [Ano70m].

Z [Ano79v]. Zellner [MS76b]. Zero [DG73a, Efr78, Wag70]. Zero-One [Efr78]. Zipf [Hil70b, Hil74c, HW75].

References

A:1973:BRB

[A.73] D. J. A. Book review: [untitled]. Journal of the American Statistical Association, 68(342):500–501, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

         org/stable/2284124.

A:1974:BRBa

[A.74a] D. J. A. Book review: [untitled]. Journal of the American Statistical Association, 69(345):281–282, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

         org/stable/2285560.

A:1974:BRBb

[A.74b] D. J. A. Book review: [untitled]. Journal of the American Statistical Association, 69(345):282, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

         stable/2285561.

A:1974:BRBc

[A.74c] D. J. A. Book review: [untitled]. Journal of the American Statistical Association, 69(347):835, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286039.

Ager:1978:IAB

[AB78] Joel W. Ager, Jr. and Sandor B. Brent. An index of agreement between a hypothesized partial order and an empirical rank order. Journal of the American Statistical Association, 73(364):827–830, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286288.

Abowd:1978:BRB

[Abo78] John M. Abowd. Book review: [untitled]. Journal of the American Statistical Association, 73(363):682–683, September 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286633.

Abramson:1979:BRB

[Abr79] Lee R. Abramson. Book review: [untitled]. Journal of the American Statistical Association, 74(365):241–243, March 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286759.

Adams:1976:BRB

[Ada76] William James Adams. Book review: [untitled]. Journal of the American Statistical Association, 71(356):1004–1005, December 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286884.

Addelman:1970:VTE

[Add70] Sidney Addelman. Variability of treatments and experimental units in the design and analysis of experiments. Journal of the American Statistical Association, 65(331):1095–1108, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284277.

Addelman:1972:RDD

[Add72] Sidney Addelman. Recent developments in the design of factorial experiments. Journal of the American Statistical Association, 67(337):103–111, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284707.

Adelstein:1978:BRB

[Ade78] Richard P. Adelstein. Book review: [untitled]. Journal of the American Statistical Association, 73(361):226, March 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286567.

Adie:1973:ICQ

[Adi73] Douglas K. Adie. An international comparison of the quantity and income-expenditure theories. Journal of the American Statistical Association, 68(341):63–65, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284139.

Ahuja:1972:RRM

[AE72] J. C. Ahuja and E. A. Enneking. Recurrence relation for the minimum variance unbiased estimator of a parameter of a left-truncated Poisson distribution. Journal of the American Statistical Association, 67(337):232–233, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284734.

Armstrong:1976:CTA

[AF76] R. D. Armstrong and E. L. Frome. A comparison of two algorithms for absolute deviation curve fitting. Journal of the American Statistical Association, 71(354):328–330, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285306.

Aigner:1970:EPL

[AG70] Dennis J. Aigner and Arthur S. Goldberger. Estimation of Pareto’s law from grouped observations. Journal of the American Statistical Association, 65(330):712–723, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284582.

Adams:1972:ELP

[AG72a] F. Gerard Adams and James M. Griffin. An economic-linear programming model of the U.S. petroleum refining industry. Journal of the American Statistical Association, 67(339):542–551, September 1972.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284436.

Aigner:1972:CEP

[AG72b] Dennis J. Aigner and Arthur S. Goldberger. Corrigenda: Estimation of Pareto’s Law from Grouped Observations. Journal of the American Statistical Association, 67(337):252, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284755.

Aiyar:1979:ARE

[AGA79] Radhakrishnan J. Aiyar, Claude L. Guillier, and Willem Albers. Asymptotic relative efficiencies of rank tests for trend alternatives. Journal of the American Statistical Association, 74(365):226–231, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286755.

Agnew:1974:CSB

[Agn74] Robert A. Agnew. Confidence sets for binary response models. Journal of the American Statistical Association, 69(346):522–524, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285690.

Agresti:1976:BRB

[Agr76a] Alan Agresti. Book review: [untitled]. Journal of the American Statistical Association, 71(353):243–244, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285780.

Agresti:1976:ECC

[Agr76b] Alan Agresti. The effect of category choice on some ordinal measures of association. Journal of the American Statistical Association, 71(353):49–55, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285729.

Agresti:1977:CMP

[Agr77] Alan Agresti. Considerations in measuring partial association for ordinal categorical data. Journal of the American Statistical Association, 72(357):37–45, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286903.

Agyei:1978:BRB

[Agy78] William K. A. Agyei. Book review: [untitled]. Journal of the American Statistical Association, 73(363):681, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286630.

Aigner:1971:BRB

[Aig71] Dennis Aigner. Book review: [untitled]. Journal of the American Statistical Association, 66(335):653, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283551.

Aigner:1972:BRB

[Aig72] Dennis Aigner. Book review: [untitled]. Journal of the American Statistical Association, 67(340):956, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284674.

Aigner:1975:BRBa

[Aig75a] Dennis J. Aigner. Book review: [untitled]. Journal of the American Statistical Association, 70(352):967–968, December 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285488.

Aigner:1975:BRBb

[Aig75b] Dennis J. Aigner. Book review: [untitled]. Journal of the American Statistical Association, 70(352):968, December 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285489.

Aigner:1975:BRBc

[Aig75c] Dennis J. Aigner. Book review: [untitled]. Journal of the American Statistical Association, 70(352):968, December 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285490.

Aitkin:1972:BRB

[Ait72] Murray Aitkin. Book review: [untitled]. Journal of the American Statistical Association, 67(338):496–497, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284422.

Abrahamse:1971:NED

[AK71] A. P. J. Abrahamse and J. Koerts. New estimators of disturbances in regression analysis. Journal of the American Statistical Association, 66(333):71–74, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284850.

Anderson:1976:QGS

[AKC76] Dallas W. Anderson, Leslie Kish, and Richard G. Cornell. Quantifying gains from stratification for optimum and approximately optimum strata using a bivariate normal model. Journal of the American Statistical Association, 71(356):887–892, December 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286856.

Alam:1974:SNT

[Ala74] Khursheed Alam. Some nonparametric tests of randomness. Journal of the American Statistical Association, 69(347):738–739, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286010.

Albright:1976:BRB

[Alb76] S. Christian Albright. Book review: [untitled]. Journal of the American Statistical Association, 71(356):1007, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286888.

Albright:1979:BRB

[Alb79] S. Christian Albright. Book review: [untitled]. Journal of the American Statistical Association, 74(366):510–511, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286377.

Ali:1974:SOK

[Ali74] Mukhtar M. Ali. Stochastic ordering and kurtosis measure. Journal of the American Statistical Association, 69(346):543–545, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285694.

Allen:1977:SSA

[All77] David M. Allen. A simulation study of alternatives to ordinary least squares: Comment. Journal of the American Statistical Association, 72(357):95–96, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286912.

Alvo:1978:SET

[Alv78] Mayer Alvo. Sequential estimation of a truncation parameter. Journal of the American Statistical Association, 73(362):404–407, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286673.

Aldrich:1973:NBM

[AM73] Carole A. Aldrich and Thomas E. Morton. A note on Bayesian modification of the failure rate function. Journal of the American Statistical Association, 68(342):483–484, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284103.

Arvesen:1975:SSP

[AM75] J. N. Arvesen and G. P. McCabe, Jr. Subset selection problems for variances with applications to regression analysis. Journal of the American Statistical Association, 70(349):166–170, March 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285395.

Amemiya:1970:BRB

[Ame70] Takeshi Amemiya. Book review: [untitled]. Journal of the American Statistical Association, 65(332):1687–1689, December 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284357.

Amemiya:1973:RAW

[Ame73] Takeshi Amemiya. Regression analysis when the variance of the dependent variable is proportional to the square of its expectation. Journal of the American Statistical Association, 68(344):928–934, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284524.

Amemiya:1974:BPA

[Ame74] Takeshi Amemiya. Bivariate probit analysis: minimum chi-square methods. Journal of the American Statistical Association, 69 (348):940–944, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286167.

Amemiya:1976:MLM

[Ame76] Takeshi Amemiya. The maximum likelihood, the minimum chi-square and the nonlinear weighted least-squares estimator in the general qualitative response model. Journal of the American Statistical Association, 71(354):347–351, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285311.

Anbar:1976:BRB

[Anb76] Dan Anbar. Book review: [untitled]. Journal of the American Statistical Association, 71(355):775–776, September 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285635.

Andersen:1970:SEF

[And70a] Erling Bernhard Andersen. Sufficiency and exponential families for discrete sample spaces. Journal of the American Statistical Association, 65(331):1248–1255, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284291.

Anderson:1970:BRB

[And70b] Theodore R. Anderson. Book review: [untitled]. Journal of the American Statistical Association, 65(330):1003–1004, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284614.

Andrews:1970:CSI

[And70c] David F. Andrews. Calibration and statistical inference. Journal of the American Statistical Association, 65(331):1233–1242, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284289.

Andersen:1971:ADC

[And71] Erling B. Andersen. The asymptotic distribution of conditional likelihood ratio tests. Journal of the American Statistical Association, 66(335):630–633, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283544.

Anderson:1974:AED

[And74a] T. W. Anderson. An asymptotic expansion of the distribution of the limited information maximum likelihood estimate of a coefficient in a simultaneous equation system. Journal of the American Statistical Association, 69(346):565–573, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285698.

Andrews:1974:BRB

[And74b] David F. Andrews. Book review: [untitled]. Journal of the American Statistical Association, 69(345):279–280, March 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285554.

Anderson:1976:CAD

[And76] T. W. Anderson. Corrigenda: An Asymptotic of the Distribution of the Limited Information Maximum Likelihood Estimate of a Coefficient in a Simultaneous Equation System. Journal of the American Statistical Association, 71(356):1010, December 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286894.

Anderson:1977:BRB

[And77] Dallas W. Anderson. Book review: [untitled]. Journal of the American Statistical Association, 72(358):483, June 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286827.

Anderson:1978:RMA

[And78a] T. W. Anderson. Repeated measurements on autoregressive processes. Journal of the American Statistical Association, 73(362):371– 378, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286668.

Andrews:1978:RMB

[And78b] D. F. Andrews. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73(361):85, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286521.

Anderson:1979:BRB

[And79] W. H. Locke Anderson. Book review: [untitled]. Journal of the American Statistical Association, 74(366):505–506, June 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286370.

Angoff:1974:DSI

[Ang74] William H. Angoff. The development of statistical indices for detecting cheaters. Journal of the American Statistical Association, 69(345):44–49, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285498.

Angrist:1978:BRB

[Ang78] Shirley Angrist. Book review: [untitled]. Journal of the American Statistical Association, 73(362):443–444, June 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286696.

Angers:1979:SEP

[Ang79] Claude Angers. Simultaneous estimation of percentile curves with application to salary data. Journal of the American Statistical Association, 74(367):621–625, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286979.

Anonymous:1970:BMa

[Ano70a] Anonymous. Back matter. Journal of the American Statistical Association, 65(329):481–482, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283628.

Anonymous:1970:BMb

[Ano70b] Anonymous. Back matter. Journal of the American Statistical Association, 65(330):1045–1048, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284626.

Anonymous:1970:BMc

[Ano70c] Anonymous. Back matter. Journal of the American Statistical Association, 65(331):1424–1427, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284325.

Anonymous:1970:BMd

[Ano70d] Anonymous. Back matter. Journal of the American Statistical Association, 65(332):1696–1700, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284365.

Anonymous:1970:FMa

[Ano70e] Anonymous. Front matter. Journal of the American Statistical Association, 65(329):1–8, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283570.

Anonymous:1970:FMb

[Ano70f] Anonymous. Front matter. Journal of the American Statistical Association, 65(330):??, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284561.

Anonymous:1970:FMc

[Ano70g] Anonymous. Front matter. Journal of the American Statistical Association, 65(331):??, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284274.

Anonymous:1970:FMd

[Ano70h] Anonymous. Front matter. Journal of the American Statistical Association, 65(332):??, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284327.

Anonymous:1970:PRa

[Ano70i] Anonymous. Publications received. Journal of the American Statistical Association, 65(329):478–480, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283627.

Anonymous:1970:PRb

[Ano70j] Anonymous. Publications received. Journal of the American Statistical Association, 65(330):1010, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284619.

Anonymous:1970:PRc

[Ano70k] Anonymous. Publications received. Journal of the American Statistical Association, 65(331):1422–1423, September 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284324.

Anonymous:1970:PRd

[Ano70l] Anonymous. Publications received. Journal of the American Statistical Association, 65(332):1694–1695, December 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284364.

Anonymous:1970:SPD

[Ano70m] Anonymous. Summaries of papers delivered at the 129th Annual Meeting of the American Statistical Association, New York, August 19–22, 1969. Journal of the American Statistical Association, 65(330):1013–1044, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284625.

Anonymous:1970:Ua

[Ano70n] Anonymous. [untitled]. Journal of the American Statistical Association, 65(332):1693, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284360.

Anonymous:1970:Ub

[Ano70o] Anonymous. [untitled]. Journal of the American Statistical Association, 65(332):1694, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284361.

Anonymous:1970:Uc

[Ano70p] Anonymous. [untitled]. Journal of the American Statistical Association, 65(332):1694, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284362.

Anonymous:1970:Ud

[Ano70q] Anonymous. [untitled]. Journal of the American Statistical Association, 65(332):1694, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284363.

Anonymous:1970:VI

[Ano70r] Anonymous. Volume information. Journal of the American Statistical Association, 65(332):1701–1709, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284326.

Anonymous:1971:BMa

[Ano71a] Anonymous. Back matter. Journal of the American Statistical Association, 66(333):230, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284895.

Anonymous:1971:BMb

[Ano71b] Anonymous. Back matter. Journal of the American Statistical Association, 66(334):432, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283965.

Anonymous:1971:BMc

[Ano71c] Anonymous. Back matter. Journal of the American Statistical Association, 66(335):664, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283569.

Anonymous:1971:BMd

[Ano71d] Anonymous. Back matter. Journal of the American Statistical Association, 66(336):??, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284273.

Anonymous:1971:FAa

[Ano71e] Anonymous. Forthcoming articles. Journal of the American Statistical Association, 66(333):4, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284838.

Anonymous:1971:FAb

[Ano71f] Anonymous. Forthcoming articles. Journal of the American Statistical Association, 66(336):672, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284211.

Anonymous:1971:FMa

[Ano71g] Anonymous. Front matter. Journal of the American Statistical Association, 66(333):1–6, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284837.

Anonymous:1971:FMb

[Ano71h] Anonymous. Front matter. Journal of the American Statistical Association, 66(334):240–242, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283915.

Anonymous:1971:FMc

[Ano71i] Anonymous. Front matter. Journal of the American Statistical Association, 66(335):436–663, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283505.

Anonymous:1971:FMd

[Ano71j] Anonymous. Front matter. Journal of the American Statistical Association, 66(336):673–674, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284210.

Anonymous:1971:PRa

[Ano71k] Anonymous. Publications received. Journal of the American Statistical Association, 66(333):227–228, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284889.

Anonymous:1971:PRb

[Ano71l] Anonymous. Publications received. Journal of the American Statistical Association, 66(334):429–431, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283964.

Anonymous:1971:PRc

[Ano71m] Anonymous. Publications received. Journal of the American Statistical Association, 66(335):661–662, September 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283568.

Anonymous:1971:PRd

[Ano71n] Anonymous. Publications received. Journal of the American Statistical Association, 66(336):925–926, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284272.

Anonymous:1971:VI

[Ano71o] Anonymous. Volume information. Journal of the American Statistical Association, 66(336):927–934, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284209.

Anonymous:1972:BMa

[Ano72a] Anonymous. Back matter. Journal of the American Statistical Association, 67(337):256, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284768.

Anonymous:1972:BMb

[Ano72b] Anonymous. Back matter. Journal of the American Statistical Association, 67(338):503–506, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284430.

Anonymous:1972:BMc

[Ano72c] Anonymous. Back matter. Journal of the American Statistical Association, 67(339):726, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284496.

Anonymous:1972:BMd

[Ano72d] Anonymous. Back matter. Journal of the American Statistical Association, 67(340):976, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284692.

Anonymous:1972:CES

[Ano72e] Anonymous. Corrigenda: Estimators for Seemingly Unrelated Regression Equations: Some Exact Finite Sample Results. Journal of the American Statistical Association, 67(337):255, March 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284767.

Anonymous:1972:FAa

[Ano72f] Anonymous. Forthcoming articles. Journal of the American Statistical Association, 67(337):4, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284694.

Anonymous:1972:FAb

[Ano72g] Anonymous. Forthcoming articles. Journal of the American Statistical Association, 67(338):272, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284367.

Anonymous:1972:FAc

[Ano72h] Anonymous. Forthcoming articles. Journal of the American Statistical Association, 67(339):512, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284432.

Anonymous:1972:FAd

[Ano72i] Anonymous. Forthcoming articles. Journal of the American Statistical Association, 67(340):732–975, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284629.

Anonymous:1972:FMa

[Ano72j] Anonymous. Front matter. Journal of the American Statistical Association, 67(337):1–6, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284693.

Anonymous:1972:FMb

[Ano72k] Anonymous. Front matter. Journal of the American Statistical Association, 67(338):272–502, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284366.

Anonymous:1972:FMc

[Ano72l] Anonymous. Front matter. Journal of the American Statistical Association, 67(339):513–514, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284431.

Anonymous:1972:FMd

[Ano72m] Anonymous. Front matter. Journal of the American Statistical Association, 67(340):733–975, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284628.

Anonymous:1972:PRa

[Ano72n] Anonymous. Publications received. Journal of the American Statistical Association, 67(337):249–251, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284754.

Anonymous:1972:PRb

[Ano72o] Anonymous. Publications received. Journal of the American Statistical Association, 67(338):499–501, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284429.

Anonymous:1972:PRc

[Ano72p] Anonymous. Publications received. Journal of the American Statistical Association, 67(339):722–724, September 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284494.

Anonymous:1972:PRd

[Ano72q] Anonymous. Publications received. Journal of the American Statistical Association, 67(340):964–965, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284690.

Anonymous:1972:TRAa

[Ano72r] Anonymous. Technical reports available. Journal of the American Statistical Association, 67(338):273–508, June 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284368.

Anonymous:1972:TRAb

[Ano72s] Anonymous. Technical reports available. Journal of the American Statistical Association, 67(339):725, September 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284495.

Anonymous:1972:TRAc

[Ano72t] Anonymous. Technical reports available. Journal of the American Statistical Association, 67(340):972–974, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284691.

Anonymous:1972:VI

[Ano72u] Anonymous. Volume information. Journal of the American Statistical Association, 67(340):966–971, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284627.

Anonymous:1973:BMa

[Ano73a] Anonymous. Back matter. Journal of the American Statistical Association, 68(341):252–256, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284208.

Anonymous:1973:BMb

[Ano73b] Anonymous. Back matter. Journal of the American Statistical Association, 68(342):503–504, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284128.

Anonymous:1973:BMc

[Ano73c] Anonymous. Back matter. Journal of the American Statistical Association, 68(343):758–762, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284836.

Anonymous:1973:BMd

[Ano73d] Anonymous. Back matter. Journal of the American Statistical Association, 68(344):1039–1040, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284560.

Anonymous:1973:FA

[Ano73e] Anonymous. Forthcoming articles. Journal of the American Statistical Association, 68(341):4–255, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284130.

Anonymous:1973:FMa

[Ano73f] Anonymous. Front matter. Journal of the American Statistical Association, 68(341):1–6, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284129.

Anonymous:1973:FMb

[Ano73g] Anonymous. Front matter. Journal of the American Statistical Association, 68(342):267–270, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284061.

Anonymous:1973:FMc

[Ano73h] Anonymous. Front matter. Journal of the American Statistical Association, 68(343):513–514, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284769.

Anonymous:1973:FMd

[Ano73i] Anonymous. Front matter. Journal of the American Statistical Association, 68(344):769–770, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284498.

Anonymous:1973:PRa

[Ano73j] Anonymous. Publications received. Journal of the American Statistical Association, 68(341):247–248, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284198.

Anonymous:1973:PRb

[Ano73k] Anonymous. Publications received. Journal of the American Statistical Association, 68(342):501–502, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284127.

Anonymous:1973:PRc

[Ano73l] Anonymous. Publications received. Journal of the American Statistical Association, 68(343):756–757, September 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284831.

Anonymous:1973:PRd

[Ano73m] Anonymous. Publications received. Journal of the American Statistical Association, 68(344):1030–1032, December 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284559.

Anonymous:1973:TRA

[Ano73n] Anonymous. Technical reports available. Journal of the American Statistical Association, 68(341):249–250, March 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284199.

Anonymous:1973:VI

[Ano73o] Anonymous. Volume information. Journal of the American Statistical Association, 68(344):1033–1038, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284497.

Anonymous:1974:BMa

[Ano74a] Anonymous. Back matter. Journal of the American Statistical Association, 69(345):287–292, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285579.

Anonymous:1974:BMb

[Ano74b] Anonymous. Back matter. Journal of the American Statistical Association, 69(348):??, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286205.

Anonymous:1974:FAa

[Ano74c] Anonymous. Forthcoming articles. Journal of the American Statistical Association, 69(345):286–289, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285570.

Anonymous:1974:FAb

[Ano74d] Anonymous. Forthcoming articles. Journal of the American Statistical Association, 69(346):587–588, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285723.

Anonymous:1974:FMa

[Ano74e] Anonymous. Front matter. Journal of the American Statistical Association, 69(345):1–8, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285493.

Anonymous:1974:FMb

[Ano74f] Anonymous. Front matter. Journal of the American Statistical Association, 69(346):586–302, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285647.

Anonymous:1974:FMc

[Ano74g] Anonymous. Front matter. Journal of the American Statistical Association, 69(347):589–594, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285988.

Anonymous:1974:FMd

[Ano74h] Anonymous. Front matter. Journal of the American Statistical Association, 69(348):841–848, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286152.

Anonymous:1974:PRb

[Ano74i] Anonymous. Publication received. Journal of the American Statistical Association, 69(346):584–585, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285719.

Anonymous:1974:PRa

[Ano74j] Anonymous. Publications received. Journal of the American Statistical Association, 69(345):283–284, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285568.

Anonymous:1974:PRc

[Ano74k] Anonymous. Publications received. Journal of the American Statistical Association, 69(347):836–837, September 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286044.

Anonymous:1974:PRd

[Ano74l] Anonymous. Publications received. Journal of the American Statistical Association, 69(348):1053–1055, December 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286204.

Anonymous:1974:TRAa

[Ano74m] Anonymous. Technical reports available. Journal of the American Statistical Association, 69(345):285–286, March 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285569.

Anonymous:1974:TRAb

[Ano74n] Anonymous. Technical reports available. Journal of the American Statistical Association, 69(347):838–840, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286045.

Anonymous:1974:VI

[Ano74o] Anonymous. Volume information. Journal of the American Statistical Association, 69(348):1056–1064, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286151.

Anonymous:1975:BMa

[Ano75a] Anonymous. Back matter. Journal of the American Statistical Association, 70(350):495–496, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285870.

Anonymous:1975:BMb

[Ano75b] Anonymous. Back matter. Journal of the American Statistical Association, 70(351):24, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286150.

Anonymous:1975:BMc

[Ano75c] Anonymous. Back matter. Journal of the American Statistical Association, 70(351):738, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285987.

Anonymous:1975:BMd

[Ano75d] Anonymous. Back matter. Journal of the American Statistical Association, 70(352):971–974, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285492.

Anonymous:1975:FM

[Ano75e] Anonymous. Front matter. Journal of the American Statistical Association, 70(351):1–2, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286148.

Anonymous:1975:LBR

[Ano75f] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 70(352):955, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285464.

Anonymous:1975:PRa

[Ano75g] Anonymous. Publications received. Journal of the American Statistical Association, 70(349):259–260, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285431.

Anonymous:1975:PRb

[Ano75h] Anonymous. Publications received. Journal of the American Statistical Association, 70(350):493–494, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285864.

Anonymous:1975:PRc

[Ano75i] Anonymous. Publications received. Journal of the American Statistical Association, 70(351):736–737, September 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285986.

Anonymous:1975:PRd

[Ano75j] Anonymous. Publications received. Journal of the American Statistical Association, 70(352):968–970, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285491.

Anonymous:1975:VI

[Ano75k] Anonymous. Volume information. Journal of the American Statistical Association, 70(352):975–980, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285432.

Anonymous:1976:BMa

[Ano76a] Anonymous. Back matter. Journal of the American Statistical Association, 71(353):256–262, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285807.

Anonymous:1976:BMb

[Ano76b] Anonymous. Back matter. Journal of the American Statistical Association, 71(354):534, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285367.

Anonymous:1976:BMc

[Ano76c] Anonymous. Back matter. Journal of the American Statistical Association, 71(356):1011–1024, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286899.

Anonymous:1976:FMa

[Ano76d] Anonymous. Front matter. Journal of the American Statistical Association, 71(353):1–2, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285724.

Anonymous:1976:FMb

[Ano76e] Anonymous. Front matter. Journal of the American Statistical Association, 71(354):265–266, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285296.

Anonymous:1976:FMc

[Ano76f] Anonymous. Front matter. Journal of the American Statistical Association, 71(355):537–538, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285580.

Anonymous:1976:FMd

[Ano76g] Anonymous. Front matter. Journal of the American Statistical Association, 71(356):785–786, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286840.

Anonymous:1976:LBRa

[Ano76h] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 71(354):518, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285345.

Anonymous:1976:LBRb

[Ano76i] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 71(355):767, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285619.

Anonymous:1976:PRa

[Ano76j] Anonymous. Publications received. Journal of the American Statistical Association, 71(353):252–253, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285801.

Anonymous:1976:PRb

[Ano76k] Anonymous. Publications received. Journal of the American Statistical Association, 71(354):533, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285366.

Anonymous:1976:PRc

[Ano76l] Anonymous. Publications received. Journal of the American Statistical Association, 71(355):778–779, September 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285645.

Anonymous:1976:PRd

[Ano76m] Anonymous. Publications received. Journal of the American Statistical Association, 71(356):1009, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286893.

Anonymous:1976:TRAa

[Ano76n] Anonymous. Technical reports available. Journal of the American Statistical Association, 71(353):254–255, March 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285802.

Anonymous:1976:TRAb

[Ano76o] Anonymous. Technical reports available. Journal of the American Statistical Association, 71(355):780–782, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285646.

Anonymous:1976:VI

[Ano76p] Anonymous. Volume information. Journal of the American Statistical Association, 71(356):1015–1021, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286839.

Anonymous:1977:BMa

[Ano77a] Anonymous. Back matter. Journal of the American Statistical Association, 72(359):696–704, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286259.

Anonymous:1977:BMb

[Ano77b] Anonymous. Back matter. Journal of the American Statistical Association, 72(360):934–944, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286507.

Anonymous:1977:CMC

[Ano77c] Anonymous. Corrigenda: a Monte Carlo Comparison of the Regression and Spectral Methods of Prediction. Journal of the American Statistical Association, 72(359):703, September 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286256.

Anonymous:1977:CSS

[Ano77d] Anonymous. Corrigenda: Small Sample Power Functions for Nonparametric Tests of Location in the Double Exponential Family. Journal of the American Statistical Association, 72(359):703, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286258.

Anonymous:1977:CSD

[Ano77e] Anonymous. Corrigenda: The Stochastic Dependence of Security Price Changes and Transaction Volumes in a Model with Temporally Dependent Price Changes. Journal of the American Statistical Association, 72(359):703, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286257.

Anonymous:1977:CIV

[Ano77f] Anonymous. [cumulative index: Volumes 61–72]. Journal of the American Statistical Association, 72(360):1–55, December 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286506.

Anonymous:1977:LBRa

[Ano77g] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 72(357):226, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286943.

Anonymous:1977:LBRb

[Ano77h] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 72(360):922, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286488.

Anonymous:1977:PRa

[Ano77i] Anonymous. Publications received. Journal of the American Statistical Association, 72(357):240–242, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286967.

Anonymous:1977:PRb

[Ano77j] Anonymous. Publications received. Journal of the American Statistical Association, 72(359):693–695, September 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286255.

Anonymous:1977:PRc

[Ano77k] Anonymous. Publications received. Journal of the American Statistical Association, 72(360):932–933, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286505.

Anonymous:1977:VI

[Ano77l] Anonymous. Volume information. Journal of the American Statistical Association, 72(360):938–943, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286450.

Anonymous:1978:BMa

[Ano78a] Anonymous. Back matter. Journal of the American Statistical Association, 73(361):234, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286580.

Anonymous:1978:BMb

[Ano78b] Anonymous. Back matter. Journal of the American Statistical Association, 73(362):452–456, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286712.

Anonymous:1978:BMc

[Ano78c] Anonymous. Back matter. Journal of the American Statistical Association, 73(363):692, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286649.

Anonymous:1978:BMd

[Ano78d] Anonymous. Back matter. Journal of the American Statistical Association, 73(364):901–904, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286321.

Anonymous:1978:CTE

[Ano78e] Anonymous. Corrigenda: a Table of Exact Confidence Limits for Differences and Ratios of Two Proportions and Their Odds Ratios. Journal of the American Statistical Association, 73(361):233, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286579.

Anonymous:1978:CIS

[Ano78f] Anonymous. Corrigenda: Intermediate Simultaneous Inference Procedures. Journal of the American Statistical Association, 73(364):900, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286320.

Anonymous:1978:LBRa

[Ano78g] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 73(361):214, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286550.

Anonymous:1978:LBRb

[Ano78h] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 73(362):434, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286680.

Anonymous:1978:LBRc

[Ano78i] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 73(363):676, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286623.

Anonymous:1978:PRa

[Ano78j] Anonymous. Publications received. Journal of the American Statistical Association, 73(361):230–232, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286578.

Anonymous:1978:PRb

[Ano78k] Anonymous. Publications received. Journal of the American Statistical Association, 73(362):449–451, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286709.

Anonymous:1978:PRc

[Ano78l] Anonymous. Publications received. Journal of the American Statistical Association, 73(363):689–690, September 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286647.

Anonymous:1978:PRd

[Ano78m] Anonymous. Publications received. Journal of the American Statistical Association, 73(364):898–899, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286318.

Anonymous:1978:VI

[Ano78n] Anonymous. Volume information. Journal of the American Statistical Association, 73(364):905–910, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286260.

Anonymous:1979:AI

[Ano79a] Anonymous. Author index. Journal of the American Statistical Association, 74(366):1–42, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286390.

Anonymous:1979:BMa

[Ano79b] Anonymous. Back matter. Journal of the American Statistical Association, 74(365):258–260, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286785.

Anonymous:1979:BMb

[Ano79c] Anonymous. Back matter. Journal of the American Statistical Association, 74(367):743–746, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287031.

Anonymous:1979:BMc

[Ano79d] Anonymous. Back matter. Journal of the American Statistical Association, 74(368):942–946, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286449.

Anonymous:1979:CAM

[Ano79e] Anonymous. Corrigenda: Applications of Mathematics, Volume 7: Game Theory. Journal of the American Statistical Association, 74(366):516, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286387.

Anonymous:1979:CET

[Ano79f] Anonymous. Corrigenda: Estimation of a Time-Truncated Exponential Parameter Used in Life Testing. Journal of the American Statistical Association, 74(366):516, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286384.

Anonymous:1979:CMM

[Ano79g] Anonymous. Corrigenda: Multivariate and Multipurpose Stratification. Journal of the American Statistical Association, 74(366):516, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286385.

Anonymous:1979:CLL

[Ano79h] Anonymous. Corrigenda: The Linear Least-Squares Prediction Approach to Two-Stage Sampling. Journal of the American Statistical Association, 74(366):516, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286383.

Anonymous:1979:FM

[Ano79i] Anonymous. Front matter. Journal of the American Statistical Association, 74(366):i–iii, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286388.

Anonymous:1979:I

[Ano79j] Anonymous. Introduction. Journal of the American Statistical Association, 74(366):iv–v, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286389.

Anonymous:1979:LBRa

[Ano79k] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 74(365):240, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286758.

Anonymous:1979:LBRb

[Ano79l] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 74(366):500, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286362.

Anonymous:1979:LBRc

[Ano79m] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 74(367):723, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286998.

Anonymous:1979:LBRd

[Ano79n] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 74(368):921, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286424.

Anonymous:1979:PRa

[Ano79o] Anonymous. Publications received. Journal of the American Statistical Association, 74(365):256–257, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286784.

Anonymous:1979:PRb

[Ano79p] Anonymous. Publications received. Journal of the American Statistical Association, 74(366):514–515, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286382.

Anonymous:1979:PRc

[Ano79q] Anonymous. Publications received. Journal of the American Statistical Association, 74(367):741–742, September 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2287027.

Anonymous:1979:PRd

[Ano79r] Anonymous. Publications received. Journal of the American Statistical Association, 74(368):939–940, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286447.

Anonymous:1979:SI

[Ano79s] Anonymous. Subject index: [A–E]. Journal of the American Statistical Association, 74(366):43–99, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286391.

Anonymous:1979:SIF

[Ano79t] Anonymous. Subject index: [F–N]. Journal of the American Statistical Association, 74(366):99–151, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286392.

Anonymous:1979:SIR

[Ano79u] Anonymous. Subject index: [O–R]. Journal of the American Statistical Association, 74(366):151–182, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286393.

Anonymous:1979:SIZ

[Ano79v] Anonymous. Subject index: [S–Z]. Journal of the American Statistical Association, 74(366):182–225, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286394.

Anonymous:1979:VI

[Ano79w] Anonymous. Volume information. Journal of the American Statistical Association, 74(368):947–951, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286395.

Antelman:1972:IBP

[Ant72] Gordon R. Antelman. Interrelated Bernoulli processes. Journal of the American Statistical Association, 67(340):831–841, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284645.

Antelman:1973:BRB

[Ant73a] Gordon Antelman. Book review: [untitled]. Journal of the American Statistical Association, 68(344):1029, December 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284555.

Antoniak:1973:BRB

[Ant73b] C. Antoniak. Book review: [untitled]. Journal of the American Statistical Association, 68(341):247, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284197.

Antelman:1979:BRB

[Ant79] Gordon Antelman. Book review: [untitled]. Journal of the American Statistical Association, 74(367):729–731, September 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287008.

Antelman:1971:FSD

[AR71] Gordon Antelman and Harry V. Roberts. Finding sampling distributions by a recognition method. Journal of the American Statistical Association, 66(333):136–141, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284862.

Archibald:1973:BRB

[Arc73] Donald G. Archibald. Book review: [untitled]. Journal of the American Statistical Association, 68(341):246, March 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284193.

Arnold:1970:HTI

[Arn70a] Barry C. Arnold. Hypothesis testing incorporating a preliminary test of significance. Journal of the American Statistical Association, 65 (332):1590–1596, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284341.

Arnold:1970:IUS

[Arn70b] Barry C. Arnold. Inadmissibility of the usual scale estimate for a shifted exponential distribution. Journal of the American Statistical Association, 65(331):1260–1264, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284293.

Arnold:1973:CIU

[Arn73] Barry C. Arnold. Corrigenda: Inadmissibility of the Usual Scale Estimate for a Shifted Exponential Distribution. Journal of the American Statistical Association, 68(343):757, September 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284832.

Arney:1975:BRB

[Arn75] William Ray Arney. Book review: [untitled]. Journal of the American Statistical Association, 70(351):733, September 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285973.

Arney:1976:BRB

[Arn76] William Ray Arney. Book review: [untitled]. Journal of the American Statistical Association, 71(356):1006–1007, December 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286887.

Arnold:1978:BRB

[Arn78] Barry C. Arnold. Book review: [untitled]. Journal of the American Statistical Association, 73(361):219–220, March 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286555.

Arnold:1979:LME

[Arn79] Steven F. Arnold. Linear models with exchangeably distributed errors. Journal of the American Statistical Association, 74(365):194– 199, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286750.

Arrow:1970:BRB

[Arr70] K. J. Arrow. Book review: [untitled]. Journal of the American Statistical Association, 65(329):468–469, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283618.

Asok:1976:SPU

[AS76] C. Asok and B. V. Sukhatme. On Sampford’s procedure of unequal probability sampling without replacement. Journal of the American Statistical Association, 71(356):912–918, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286860.

Anderson:1977:TSL

[AS77] T. W. Anderson and Takamitsu Sawa. Two-stage least squares: In which direction should the residuals be minimized? Journal of the American Statistical Association, 72(357):187–191, March 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286935.

Agrawal:1978:CFD

[AS78] H. L. Agrawal and K. L. Sharma. On construction of four-dimensional incomplete block designs. Journal of the American Statistical Association, 73(364):844–849, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286291.

Alam:1973:OCI

[AST73] K. Alam, K. M. Lal Saxena, and Yung Liang Tong. Optimal confidence interval for a ranked parameter. Journal of the American Statistical Association, 68(343):720–725, September 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284806.

Atkin:1976:BRB

[Atk76] Robert S. Atkin. Book review: [untitled]. Journal of the American Statistical Association, 71(356):1007–1008, December 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286889.

Atkinson:1978:SSP

[Atk78] Scott E. Atkinson. Small-sample properties of simultaneous equation estimators with multicollinearity. Journal of the American Statistical Association, 73(364):719–723, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286264.

Atkins:1979:BRB

[Atk79] D. R. Atkins. Book review: [untitled]. Journal of the American Statistical Association, 74(365):247–248, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286768.

Atwood:1971:RPE

[Atw71] Corwin L. Atwood. Robust procedures for estimating polynomial regression. Journal of the American Statistical Association, 66 (336):855–860, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284241.

Atwood:1975:ERS

[Atw75] Corwin L. Atwood. Estimating a response surface with an uncertain number of parameters, assuming normal errors. Journal of the American Statistical Association, 70(351):613–617, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285942.

Atwood:1976:BRB

[Atw76] Corwin L. Atwood. Book review: [untitled]. Journal of the American Statistical Association, 71(353):249, March 1976. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285788.

Amemiya:1972:EAP

[AW72] Takeshi Amemiya and Roland Y. Wu. The effect of aggregation on prediction in the autoregressive model. Journal of the American Statistical Association, 67(339):628–632, September 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284454.

Abernathy:1976:BRB

[AY76] James R. Abernathy and Yasar Yesilcay. Book review: [untitled]. Journal of the American Statistical Association, 71(354):528–530, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285356.

Amershi:1979:BRB

[AZ79] Amin H. Amershi and William T. Ziemba. Book review: [untitled]. Journal of the American Statistical Association, 74(367):738–739, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2287024.

Bendel:1977:CSR

[BA77] Robert B. Bendel and A. A. Afifi. Comparison of stopping rules in forward “stepwise” regression. Journal of the American Statistical Association, 72(357):46–53, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286904.

Bacharach:1979:NBD

[Bac79] Michael Bacharach. Normal Bayesian dialogues. Journal of the American Statistical Association, 74(368):837–846, December 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286408.

Bailar:1975:ERG

[Bai75] Barbara A. Bailar. The effects of rotation group bias on estimates from panel surveys. Journal of the American Statistical Association, 70(349):23–30, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285370.

Bailey:1977:TBS

[Bai77] B. J. R. Bailey. Tables of the Bonferroni t statistic. Journal of the American Statistical Association, 72(358):469–478, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286821.

Baillie:1979:AMS

[Bai79] Richard T. Baillie. The asymptotic mean squared error of multistep prediction from the regression model with autoregressive errors. Journal of the American Statistical Association, 74(365):175–184, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286748.

Baker:1974:STH

[Bak74] Frank B. Baker. Stability of two hierarchical grouping techniques case 1: Sensitivity to data errors. Journal of the American Statistical Association, 69(346):440–445, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285675.

Balestra:1970:EOL

[Bal70] Pietro Balestra. On the efficiency of ordinary least-squares in regression models. Journal of the American Statistical Association, 65 (331):1330–1337, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284299.

Balaam:1974:BRB

[Bal74] L. N. Balaam. Book review: [untitled]. Journal of the American Statistical Association, 69(346):576, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285702.

Ball:1975:BRB

[Bal75] R. J. Ball. Book review: [untitled]. Journal of the American Statistical Association, 70(350):487–488, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285850.

Bancroft:1971:ASA

[Ban71] T. A. Bancroft. The American statistical association: a single scientific and educational community. Journal of the American Statistical Association, 66(333):7–12, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284839.

Banerjee:1972:SWD

[Ban72] K. S. Banerjee. Singular weighing designs and a reflexive generalized inverse. Journal of the American Statistical Association, 67(337):211– 212, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284726.

Bancroft:1978:BRB

[Ban78] T. A. Bancroft. Book review: [untitled]. Journal of the American Statistical Association, 73(361):229–230, March 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286575.

Bart:1970:BRBa

[Bar70a] William M. Bart. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1413, September 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284315.

Bart:1970:BRBb

[Bar70b] William M. Bart. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1413–1414, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284316.

Baron:1973:PER

[Bar73] David P. Baron. Point estimation and risk preferences. Journal of the American Statistical Association, 68(344):944–950, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284527.

Bartko:1974:BRB

[Bar74] John J. Bartko. Book review: [untitled]. Journal of the American Statistical Association, 69(345):277–278, March 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285550.

Barnard:1976:SIU

[Bar76a] George A. Barnard. Strong inconsistency from uniform priors: Comment. Journal of the American Statistical Association, 71(353):121– 122, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285748.

Barnett:1976:MLI

[Bar76b] William A. Barnett. Maximum likelihood and iterated Aitken estimation of nonlinear systems of equations. Journal of the American Statistical Association, 71(354):354–360, June 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285313.

Barnard:1977:CCS

[Bar77] George A. Barnard. Conditional confidence statements and confidence estimators: Comment. Journal of the American Statistical Association, 72(360):808–810, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286461.

Barry:1979:BRB

[Bar79] Christopher B. Barry. Book review: [untitled]. Journal of the American Statistical Association, 74(365):244, March 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286761.

Basu:1971:BFR

[Bas71] A. P. Basu. Bivariate failure rate. Journal of the American Statistical Association, 66(333):103–104, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284857.

Bassin:1973:BOO

[Bas73] William M. Bassin. A Bayesian optimal overhaul interval model for the Weibull restoration process case. Journal of the American Statistical Association, 68(343):575–578, September 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284778.

Basu:1977:ENP

[Bas77] Debabrata Basu. On the elimination of nuisance parameters. Journal of the American Statistical Association, 72(358):355–366, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286800.

Bassler:1978:BRB

[Bas78] John F. Bassler. Book review: [untitled]. Journal of the American Statistical Association, 73(363):678–679, September 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286625.

Bauer:1971:BRB

[Bau71] Bertrand N. Bauer. Book review: [untitled]. Journal of the American Statistical Association, 66(334):427, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283958.

Bauer:1972:CCS

[Bau72] David F. Bauer. Constructing confidence sets using rank statistics. Journal of the American Statistical Association, 67(339):687–690, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284469.

Bawa:1972:AEO

[Baw72] V. S. Bawa. Asymptotic efficiency of one R-factor experiment relative of R one-factor experiments for selecting the best normal population. Journal of the American Statistical Association, 67(339):660–661, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284461.

Bailey:1971:OMC

[BB71] Duncan Bailey and Stanley E. Boyle. The optimal measure of concentration. Journal of the American Statistical Association, 66 (336):702–706, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284215.

Barlow:1972:IRP

[BB72] R. E. Barlow and H. D. Brunk. The isotonic regression problem and its dual. Journal of the American Statistical Association, 67(337):140–147, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284712.

Block:1974:CBE

[BB74] Henry W. Block and A. P. Basu. A continuous bivariate exponential extension. Journal of the American Statistical Association, 69 (348):1031–1037, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286184.

Banerjee:1976:PIM

[BB76a] A. K. Banerjee and G. K. Bhattacharyya. A purchase incidence model with inverse Gaussian interpurchase times. Journal of the American Statistical Association, 71(356):823–829, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286845. See remarks [BH08].

Broemeling:1976:SCI

[BB76b] L. D. Broemeling and D. E. Bee. Simultaneous confidence intervals for parameters of a balanced incomplete block. Journal of the American Statistical Association, 71(354):425–428, June 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285328.

Brown:1977:SBT

[BB77] Morton B. Brown and Jacqueline K. Benedetti. Sampling behavior of test for correlation in two-way contingency tables. Journal of the American Statistical Association, 72(358):309–315, June 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286793.

Berlin:1979:TDD

[BBC79] Bengt Berlin, Joel Brodsky, and Peter Clifford. Testing disease dependence in survival experiments with serial sacrifice. Journal of the American Statistical Association, 74(365):5–14, March 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286714.

Basmann:1971:EFS

[BBDS71] R. L. Basmann, Franklin Lee Brown, William S. Dawes, and Gregory K. Schoepfle. Exact finite sample density functions of GCL estimators of structural coefficients in a leading exactly identifiable case. Journal of the American Statistical Association, 66(333):122– 126, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284860.

Basmann:1972:CEF

[BBDS72] R. L. Basmann, F. L. Brown, W. S. Dawes, and G. K. Schoepfle. Corrigenda: Exact Finite Sample Density Functions of GCL Estimators of Structural Coefficients in a Leading Exactly Identifiable Case. Journal of the American Statistical Association, 67(337):252, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284756.

Bemis:1972:EHT

[BBH72] Bruce M. Bemis, Lee J. Bain, and James J. Higgins. Estimation and hypothesis testing for the parameters of a bivariate exponential distribution. Journal of the American Statistical Association, 67(340):927–929, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284664.

Balmer:1974:OSP

[BBS74] D. W. Balmer, M. Boulton, and R. A. Sack. Optimal solutions in parameter estimation problems for the Cauchy distribution. Journal of the American Statistical Association, 69(345):238–242, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285534.

Berk:1979:AOM

[BC79] Robert H. Berk and Arthur Cohen. Asymptotically optimal methods of combining tests. Journal of the American Statistical Association, 74(368):812–814, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286404.

Barnea:1973:RED

[BD73] Amir Barnea and David H. Downes. A reexamination of the empirical distribution of stock price changes. Journal of the American Statistical Association, 68(342):348–350, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284074.

Bjerve:1978:ESN

[BD78] Steinar Bjerve and Kjell A. Doksum. Estimation in some nonlinear models. Journal of the American Statistical Association, 73(362):366– 370, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286667.

Blumenthal:1978:ECS

[BDG78] Saul Blumenthal, Ram C. Dahiya, and Alan J. Gross. Estimating the complete sample size from an incomplete Poisson sample. Journal of the American Statistical Association, 73(361):182–187, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286542.

Bulgren:1974:BDA

[BDH74] William G. Bulgren, Richard L. Dykstra, and John E. Hewett. A bivariate t-distribution with applications. Journal of the American Statistical Association, 69(346):525–532, June 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285691.

Bain:1975:TMC

[BE75] Lee J. Bain and Max Engelhardt. A two-moment chi-square approximation for the statistic log (x–/x?). Journal of the American Statistical Association, 70(352):948–950, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285462.

Beaver:1974:LAM

[Bea74] Robert J. Beaver. Locally asymptotically most stringent tests for paired comparison experiments. Journal of the American Statistical Association, 69(346):423–427, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285672.

Beaver:1976:BRB

[Bea76] William H. Beaver. Book review: [untitled]. Journal of the American Statistical Association, 71(354):530–531, June 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285358.

Beaver:1977:WLS

[Bea77] Robert J. Beaver. Weighted least-squares analysis of several univariate Bradley–Terry models. Journal of the American Statistical Association, 72(359):629–634, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286230.

Beckman:1973:BRBa

[Bec73a] Richard J. Beckman. Book review: [untitled]. Journal of the American Statistical Association, 68(341):247, March 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284195.

Beckman:1973:BRBb

[Bec73b] Richard J. Beckman. Book review: [untitled]. Journal of the American Statistical Association, 68(341):247, March 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284196.

Beckman:1974:BRB

[Bec74] Richard J. Beckman. Book review: [untitled]. Journal of the American Statistical Association, 69(346):577, June 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285704.

Beck:1976:BRB

[Bec76] E. M. Beck. Book review: [untitled]. Journal of the American Statistical Association, 71(355):772–773, September 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285628.

Behrman:1973:BRB

[Beh73] Jere R. Behrman. Book review: [untitled]. Journal of the American Statistical Association, 68(342):492, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284109.

Bennett:1972:BRB

[Ben72] Carl A. Bennett. Book review: [untitled]. Journal of the American Statistical Association, 67(337):247–248, March 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284746.

Benjamin:1975:BRB

[Ben75] Bernard Benjamin. Book review: [untitled]. Journal of the American Statistical Association, 70(351):729, September 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285966.

Berk:1970:BRB

[Ber70] Robert H. Berk. Book review: [untitled]. Journal of the American Statistical Association, 65(332):1680–1682, December 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284352.

Beran:1971:BRB

[Ber71a] R. J. Beran. Book review: [untitled]. Journal of the American Statistical Association, 66(333):223–224, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284878.

Berry:1971:PDO

[Ber71b] Brian J. L. Berry. Problems of data organization and analytical methods in geography. Journal of the American Statistical Association, 66(335):510–523, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283518.

Beran:1973:BRB

[Ber73a] Rudolf Beran. Book review: [untitled]. Journal of the American Statistical Association, 68(343):752–753, September 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284822.

Berger:1973:BRB

[Ber73b] Paul D. Berger. Book review: [untitled]. Journal of the American Statistical Association, 68(341):244, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284190.

Berry:1974:OSS

[Ber74] Donald A. Berry. Optimal sampling schemes for estimating system reliability by testing components–1: Fixed sample sizes. Journal of the American Statistical Association, 69(346):485–491, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285684.

Berry:1976:BRB

[Ber76] Brian J. L. Berry. Book review: [untitled]. Journal of the American Statistical Association, 71(355):769, September 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285622.

Berk:1977:TCR

[Ber77] Kenneth N. Berk. Tolerance and condition in regression computations. Journal of the American Statistical Association, 72 (360):863–866, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286476.

Berengut:1978:BRB

[Ber78a] David Berengut. Book review: [untitled]. Journal of the American Statistical Association, 73(364):892–893, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286308.

Berry:1978:BRBb

[Ber78b] Albert Berry. Book review: [untitled]. Journal of the American Statistical Association, 73(363):683–684, September 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286635.

Berry:1978:BRBa

[Ber78c] Donald A. Berry. Book review: [untitled]. Journal of the American Statistical Association, 73(361):218–219, March 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286553.

Berry:1978:MTA

[Ber78d] Donald A. Berry. Modified two-armed bandit strategies for certain clinical trials. Journal of the American Statistical Association, 73(362):339–345, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286662.

Betancourt:1971:NIH

[Bet71] Roger R. Betancourt. The normal income hypothesis in Chile. Journal of the American Statistical Association, 66(334):258–263, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283918.

Brown:1974:CRT

[BF74a] Morton B. Brown and Alan B. Forsythe. Corrigenda: Robust Tests for the Equality of Variances. Journal of the American Statistical Association, 69(347):840, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286046.

Brown:1974:RTE

[BF74b] Morton B. Brown and Alan B. Forsythe. Robust tests for the equality of variances. Journal of the American Statistical Association, 69(346):364–367, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285659.

Berk:1978:RMBa

[BF78a] Kenneth N. Berk and Ivor S. Francis. A review of the manuals for BMDP and SPSS. Journal of the American Statistical Association, 73(361):65–71, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286517.

Berk:1978:RMBb

[BF78b] Kenneth N. Berk and Ivor S. Francis. A review of the manuals for BMDP and SPSS: Rejoinder. Journal of the American Statistical Association, 73(361):94–95, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286527.

Burnett:1970:ESS

[BG70] Thomas D. Burnett and Donald Guthrie. Estimation of stationary stochastic regression parameters. Journal of the American Statistical Association, 65(332):1547–1553, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284335.

Berger:1973:NCT

[BG73] Agnes Berger and Ruth Z. Gold. Note on Cochran’s Q-test for the comparison of correlated proportions. Journal of the American Statistical Association, 68(344):989–993, December 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284537.

Bradu:1974:SSI

[BG74] Dan Bradu and K. R. Gabriel. Simultaneous statistical inference on interactions in two-way analysis of variance. Journal of the American Statistical Association, 69(346):428–436, June 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285673.

Blumenthal:1977:RST

[BG77] Saul Blumenthal and Z. Govindarajulu. Robustness of Stein’s twostage procedure for mixtures of normal populations. Journal of the American Statistical Association, 72(357):192–196, March 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286936.

Breslow:1972:SCE

[BH72] Norman Breslow and Charlene Haug. Sequential comparison of exponential survival curves. Journal of the American Statistical Association, 67(339):691–697, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284470.

Baker:1975:MPH

[BH75] Frank B. Baker and Lawrence J. Hubert. Measuring the power of hierarchical cluster analysis. Journal of the American Statistical Association, 70(349):31–38, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285371.

Baker:1976:GTA

[BH76] Frank B. Baker and Lawrence J. Hubert. A graph-theoretic approach to goodness-of-fit in complete-link hierarchical clustering. Journal of the American Statistical Association, 71(356):870–878, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286853.

Basilevsky:1979:KLA

[BH79] Alexander Basilevsky and Derek P. J. Hum. Karhunen–Loeve analysis of historical time series with an application to plantation births in Jamaica. Journal of the American Statistical Association, 74(366):284–290, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286324.

Bemmaor:2008:BBP

[BH08] Albert C. Bemmaor and Daniel Hoppe. On Banerjee and Bhattacharyya (1976), “A Purchase Incidence Model With Inverse Gaussian Interpurchase Times,” Journal of the American Statistical Association, 71, 823–829. Journal of the American Statistical Association, 103(482):894–895, June 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [BB76a].

Bhapkar:1971:BRB

[Bha71a] V. P. Bhapkar. Book review: [untitled]. Journal of the American Statistical Association, 66(336):921, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284261.

Bhatia:1971:PIN

[Bha71b] Kul B. Bhatia. A price index for nonfarm one-family houses, 1947– 64. Journal of the American Statistical Association, 66(333):23–32, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284842.

Bhatia:1971:USN

[Bha71c] Kul B. Bhatia. The USDA series on net investment in farm real estate. A critique. Journal of the American Statistical Association, 66(335):492–495, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283513.

Bhansali:1973:MCC

[Bha73a] R. J. Bhansali. A Monte Carlo comparison of the regression method and the spectral methods of prediction. Journal of the American Statistical Association, 68(343):621–625, September 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284789.

Bhatia:1973:EAC

[Bha73b] Kul B. Bhatia. The estimation of accrued capital gains on individuals’ corporate stock holdings. Journal of the American Statistical Association, 68(341):55–62, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284138.

Bhatia:1976:CGI

[Bha76] Kul B. Bhatia. Capital gains and inequality of personal income: Some results from survey data. Journal of the American Statistical Association, 71(355):575–580, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285585.

Bhattacharyya:1977:NER

[Bha77] Helen T. Bhattacharyya. Nonparametric estimation of ratio of scale parameters. Journal of the American Statistical Association, 72(358):459–463, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286819.

Bhattacharyya:1979:BRBb

[Bha79a] G. K. Bhattacharyya. Book review: [untitled]. Journal of the American Statistical Association, 74(367):736, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287019.

Bhattacharyya:1979:BRBa

[Bha79b] Helen T. Bhattacharyya. Book review: [untitled]. Journal of the American Statistical Association, 74(367):732–733, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287012.

Bielby:1977:REN

[BHF77] William T. Bielby, Robert M. Hauser, and David L. Featherman. Response errors of nonblack males in models of the stratification process. Journal of the American Statistical Association, 72(360):723–735, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286452.

Bicksler:1973:BRB

[Bic73] James L. Bicksler. Book review: [untitled]. Journal of the American Statistical Association, 68(343):750–751, September 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284818.

Bickel:1975:OSH

[Bic75] P. J. Bickel. One-step Huber estimates in the linear model. Journal of the American Statistical Association, 70(350):428–434, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285834.

Billard:1977:OPS

[Bil77] L. Billard. Optimum partial sequential tests for two-sided tests of the binomial parameter. Journal of the American Statistical Association, 72(357):197–201, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286937.

Binns:1976:SCP

[Bin76] M. R. Binns. A sequential counting procedure for estimating the total number of randomly distributed individuals. Journal of the American Statistical Association, 71(353):74–79, March 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285734.

Binder:1978:EMN

[Bin78a] David A. Binder. Estimating mixtures of normal distributions and switching regressions: Comment. Journal of the American Statistical Association, 73(364):746–747, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286270.

Bingham:1978:BRB

[Bin78b] Christopher Bingham. Book review: [untitled]. Journal of the American Statistical Association, 73(363):688, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286645.

Birnbaum:1970:DMP

[Bir70] Allan Birnbaum. On Durbin’s modified principle of conditionality. Journal of the American Statistical Association, 65(329):402–403, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283603.

Birnbaum:1972:MCS

[Bir72] Allan Birnbaum. More on concepts of statistical evidence. Journal of the American Statistical Association, 67(340):858–861, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284650.

Birch:1978:BRBa

[Bir78a] Jeffrey B. Birch. Book review: [untitled]. Journal of the American Statistical Association, 73(364):885–887, December 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286300.

Birch:1978:BRBb

[Bir78b] Jeffrey B. Birch. Book review: [untitled]. Journal of the American Statistical Association, 73(364):887, December 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286301.

Bhattacharyya:1970:PRT

[BJ70] G. K. Bhattacharyya and Richard A. Johnson. Properties of a rank test of cronholm and revusky. Journal of the American Statistical Association, 65(331):1308–1319, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284297.

Badalamenti:1972:BRB

[BJ72] Anthony F. Badalamenti and Joseph M. Jacisin. Book review: [untitled]. Journal of the American Statistical Association, 67(337):248, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284748.

Bhattacharyy:1973:TIB

[BJ73] G. K. Bhattacharyy and Richard A. Johnson. On a test of independence in a bivariate exponential distribution. Journal of the American Statistical Association, 68(343):704–706, September 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284802.

Bhattacharyya:1974:ERM

[BJ74] G. K. Bhattacharyya and Richard A. Johnson. Estimation of reliability in a multicomponent stress- strength model. Journal of the American Statistical Association, 69(348):966–970, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286173.

Bagshaw:1977:SPD

[BJ77] Michael Bagshaw and Richard A. Johnson. Sequential procedures for detecting parameter changes in a time-series model. Journal of the American Statistical Association, 72(359):593–597, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286222.

Bhattacharyya:1970:PPS

[BJN70] G. K. Bhattacharyya, Richard A. Johnson, and H. R. Neave. Percentage points of some non-parametric tests for independence and empirical power comparisons. Journal of the American Statistical Association, 65(330):976–983, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284602.

Bjorck:1978:CIR

[Bjö78] Å. Björck. Comment on the iterative refinement of least squares solutions. Journal of the American Statistical Association, 73(361):161– 166, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286538.

Bonacich:1976:EMA

[BK76] Phillip Bonacich and Douglas Kirby. Establishing a metric from assumptions about linearity. Journal of the American Statistical Association, 71(353):56–61, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285730.

Bassett:1978:ATL

[BK78] Gilbert Bassett, Jr. and Roger Koenker. Asymptotic theory of least absolute error regression. Journal of the American Statistical Association, 73(363):618–622, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286611.

Belsley:1979:SAE

[BK79] David A. Belsley and Edwin Kuh. Statistical analysis of econometric models: Comment. Journal of the American Statistical Association, 74(367):643–645, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286982.

Bingham:1977:SSA

[BL77a] Christopher Bingham and Kinley Larntz. A simulation study of alternatives to ordinary least squares: Comment. Journal of the American Statistical Association, 72(357):97–102, March 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286914.

Buse:1977:CCS

[BL77b] A. Buse and L. Lim. Corrigenda: Cubic Splines as a Special Case of Restricted Least Squares. Journal of the American Statistical Association, 72(358):488, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286837.

Buse:1977:CSS

[BL77c] A. Buse and L. Lim. Cubic splines as a special case of restricted least squares. Journal of the American Statistical Association, 72(357):64– 68, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286906.

Bhattacharyya:1979:ESB

[BL79] M. N. Bhattacharyya and Allan P. Layton. Effectiveness of seat belt legislation on the Queensland road toll — an Australian case study in intervention analysis. Journal of the American Statistical Association, 74(367):596–603, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286976.

Bland:1971:BRB

[Bla71] Richard P. Bland. Book review: [untitled]. Journal of the American Statistical Association, 66(335):658, September 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283558.

Blight:1974:RSE

[Bli74] B. J. N. Blight. Recursive solutions for the estimation of a stochastic parameter. Journal of the American Statistical Association, 69(346):477–481, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285682.

Birnbaum:1971:ORL

[BLM71] Allan Birnbaum, Eugene Laska, and Morris Meisner. Optimally robust linear estimators of location. Journal of the American Statistical Association, 66(334):302–310, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283927.

Blommers:1971:BRB

[Blo71] Paul Blommers. Book review: [untitled]. Journal of the American Statistical Association, 66(336):918–919, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284256.

Block:1977:BRB

[Blo77a] Henry W. Block. Book review: [untitled]. Journal of the American Statistical Association, 72(357):227–228, March 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286944.

Blomqvist:1977:RBC

[Blo77b] Nils Blomqvist. On the relation between change and initial value. Journal of the American Statistical Association, 72(360):746–749, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286454.

Blume:1974:UEL

[Blu74] Marshall E. Blume. Unbiased estimators of long-run expected rates of return. Journal of the American Statistical Association, 69 (347):634–638, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285993.

Blyth:1971:DOS

[Bly71] Colin R. Blyth. Does an observed sequence of numbers follow a simple rule? (another look at Bode’s law): Comment. Journal of the American Statistical Association, 66(335):566–567, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283527.

Blyth:1972:SPS

[Bly72a] Colin R. Blyth. On Simpson’s paradox and the sure-thing principle. Journal of the American Statistical Association, 67(338):364– 366, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284382.

Blyth:1972:SPPa

[Bly72b] Colin R. Blyth. Some probability paradoxes in choice from among random alternatives. Journal of the American Statistical Association, 67(338):366–373, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284383.

Blyth:1972:SPPb

[Bly72c] Colin R. Blyth. Some probability paradoxes in choice from among random alternatives: Rejoinder. Journal of the American Statistical Association, 67(338):379–381, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284388.

Blyth:1973:SPM

[Bly73] Colin R. Blyth. Simpson’s paradox and mutually favorable events. Journal of the American Statistical Association, 68(343):746, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284813.

Birnbaum:1970:ARE

[BM70a] Allan Birnbaum and Valerie Mike. Asymptotically robust estimators of location. Journal of the American Statistical Association, 65 (331):1265–1282, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284294.

Bradu:1970:ELL

[BM70b] Dan Bradu and Yair Mundlak. Estimation in lognormal linear models. Journal of the American Statistical Association, 65(329):198–211, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283587.

Blot:1973:SED

[BM73a] William J. Blot and Duane A. Meeter. Sequential experimental design procedures. Journal of the American Statistical Association, 68(343):586–593, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284782.

Bodkin:1973:OSM

[BM73b] Ronald G. Bodkin and K. S. R. Murthy. The orders-shipments mechanism in Canadian producer goods industries. Journal of the American Statistical Association, 68(342):297–305, June 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284066.

Blumenthal:1975:EPS

[BM75] Saul Blumenthal and Richard Marcus. Estimating population size with exponential failure. Journal of the American Statistical Association, 70(352):913–922, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285457.

Breiman:1976:GEI

[BM76a] L. Breiman and W. S. Meisel. General estimates of the intrinsic variability of data in nonlinear regression models. Journal of the American Statistical Association, 71(354):301–307, June 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285301.

Brown:1976:RMS

[BM76b] Charles C. Brown and Larry R. Muenz. Reduced mean square error estimation in contingency tables. Journal of the American Statistical Association, 71(353):176–182, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285764.

Beckwith:1977:SIC

[BM77a] Neil E. Beckwith and Donald G. Morrison. Stochastic interpretation of 2 × 2 classification table. Journal of the American Statistical Association, 72(358):303–308, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286792.

Bement:1977:RII

[BM77b] Thomas R. Bement and George A. Milliken. Recovery of interblock information for designs with groups of correlated blocks. Journal of the American Statistical Association, 72(357):157–159, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286928.

Basu:1978:ERS

[BM78] A. P. Basu and A. H. El Mawaziny. Estimates of reliability of k-out-of-m structures in the independent exponential case. Journal of the American Statistical Association, 73(364):850–854, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286292.

Bailey:1978:IVS

[BMB78] Leroy Bailey, Thomas F. Moore, and Barbara A. Bailar. An interviewer variance study for the eight impact cities of the national crime survey cities sample. Journal of the American Statistical Association, 73(361):16–23, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286510.

Borus:1973:RBR

[BN73] Michael E. Borus and Gilbert Nestel. Response bias in reports of Father’s education and socioeconomic status. Journal of the American Statistical Association, 68(344):816–820, December 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284505.

Berkson:1974:NME

[BN74] Joseph Berkson and B. N. Nagnur. A note on the minimum χ_{1}^{2} estimate and a L.A.M.S.T. χ^{2} in the “no interaction” problem. Journal of the American Statistical Association, 69(348):1038–1040, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286185.

Bultez:1975:CSC

[BN75] Alain V. Bultez and Philippe A. Naert. Consistent sum-constrained models. Journal of the American Statistical Association, 70(351): 529–535, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285928.

Balintfy:1970:ECB

[BNW70] Joseph L. Balintfy, John Neter, and William Wasserman. An experimental comparison between fixed weight and linear programming food price indexes. Journal of the American Statistical Association, 65(329):49–60, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283574.

Boardman:1973:BRB

[Boa73] Anthony E. Boardman. Book review: [untitled]. Journal of the American Statistical Association, 68(342):489–491, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284106.

Boen:1971:BRB

[Boe71] James R. Boen. Book review: [untitled]. Journal of the American Statistical Association, 66(334):426–427, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283957.

Boen:1972:BRB

[Boe72] James R. Boen. Book review: [untitled]. Journal of the American Statistical Association, 67(338):497, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284424.

Bofinger:1970:MCG

[Bof70] Eve Bofinger. Maximizing the correlation of grouped observations. Journal of the American Statistical Association, 65(332):1632–1638, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284346.

Bofinger:1975:OCD

[Bof75] Eve Bofinger. Optimal condensation of distributions and optimal spacing of order statistics. Journal of the American Statistical Association, 70(349):151–154, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285391.

Bofinger:1976:LFC

[Bof76] Eve Bofinger. Least favorable configuration when ties are broken. Journal of the American Statistical Association, 71(354):423–424, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285327.

Bohrer:1971:BRB

[Boh71] Robert Bohrer. Book review: [untitled]. Journal of the American Statistical Association, 66(336):922, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284262.

Bohning:1975:BRB

[Boh75] Daryl E. Bohning. Book review: [untitled]. Journal of the American Statistical Association, 70(350):487, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285849.

Bohrer:1979:MTD

[Boh79] Robert Bohrer. Multiple three-decision rules for parametric signs. Journal of the American Statistical Association, 74(366):432–437, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286349.

Basu:1975:ISM

[BOLK75] J. P. Basu, P. L. Odell, T. O. Lewis, and A. Kinderman. On independence of sample mean and translation invariant statistics of samples from multivariate normal populations. Journal of the American Statistical Association, 70(350):480–481, June 1975. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285844.

Bongaarts:1979:BRB

[Bon79] John Bongaarts. Book review: [untitled]. Journal of the American Statistical Association, 74(366):513, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286380.

Bookstein:1977:BRB

[Boo77] Fred L. Bookstein. Book review: [untitled]. Journal of the American Statistical Association, 72(357):230–231, March 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286948.

Booth:1979:BRB

[Boo79] Gordon D. Booth. Book review: [untitled]. Journal of the American Statistical Association, 74(367):731–732, September 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287009.

Borch:1970:BRB

[Bor70a] Karl Borch. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1404–1405, September 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284310.

Borus:1970:REQ

[Bor70b] Michael E. Borus. Response error and questioning technique in surveys of earnings information. Journal of the American Statistical Association, 65(330):566–575, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284567.

Borpujari:1977:EBA

[Bor77] Aroona S. Borpujari. An empirical Bayes approach for estimating the mean of N stationary time series. Journal of the American Statistical Association, 72(358):397–402, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286806.

Boulier:1977:BRB

[Bou77] Bryan L. Boulier. Book review: [untitled]. Journal of the American Statistical Association, 72(360):928–929, December 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286498.

Bowden:1970:SCB

[Bow70] David C. Bowden. Simultaneous confidence bands for linear regression models. Journal of the American Statistical Association, 65(329):413–421, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283605.

Box:1976:SS

[Box76] George E. P. Box. Science and statistics. Journal of the American Statistical Association, 71(356):791–799, December 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286841.

Box:1979:SPS

[Box79] George E. P. Box. Some problems of statistics and everyday life. Journal of the American Statistical Association, 74(365):1–4, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286713.

Boza:1973:BRB

[Boz73] Luis B. Boza. Book review: [untitled]. Journal of the American Statistical Association, 68(343):756, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284830.

Box:1970:DRA

[BP70] G. E. P. Box and David A. Pierce. Distribution of residual autocorrelations in autoregressive-integrated moving average time series models. Journal of the American Statistical Association, 65 (332):1509–1526, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284333.

Bogdanoff:1973:BFI

[BP73] David A. Bogdanoff and Donald A. Pierce. Bayes-fiducial inference for the Weibull distribution. Journal of the American Statistical Association, 68(343):659–664, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284795.

Bayless:1970:ESS

[BR70] D. L. Bayless and J. N. K. Rao. An empirical study of stabilities of estimators and variance estimators in unequal probability sampling (n = 3 or 4). Journal of the American Statistical Association, 65 (332):1645–1667, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284348.

Broffitt:1977:PAC

[BR77] James D. Broffitt and Ronald H. Randles. A power approximation for the chi-square goodness-of-fit test: Simple hypothesis case. Journal of the American Statistical Association, 72(359):604–607, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286227.

Bramblett:1970:BRB

[Bra70] Jerry E. Bramblett. Book review: [untitled]. Journal of the American Statistical Association, 65(330):995–996, June 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284607.

Bradley:1973:EML

[Bra73] Edwin L. Bradley. The equivalence of maximum likelihood and weighted least squares estimate in the exponential family. Journal of the American Statistical Association, 68(341):199–200, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284169.

Braham:1979:FEWa

[Bra79a] Roscoe R. Braham, Jr. Field experimentation in weather modification. Journal of the American Statistical Association, 74(365):57–68, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286722.

Braham:1979:FEWb

[Bra79b] Roscoe R. Braham, Jr. Field experimentation in weather modification: Rejoinder. Journal of the American Statistical Association, 74(365):97–104, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286733.

Beaton:1976:ARS

[BRB76] Albert E. Beaton, Donald B. Rubin, and John L. Barone. The acceptability of regression solutions: Another look at computational accuracy. Journal of the American Statistical Association, 71(353):158– 168, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285761.

Beaton:1977:MCA

[BRB77] Albert E. Beaton, Donald B. Rubin, and John L. Barone. More on computational accuracy in regression: Comment. Journal of the American Statistical Association, 72(359):600–601, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286224.

Breslow:1970:SMU

[Bre70] Norman Breslow. Sequential modification of the UMP test for binomial probabilities. Journal of the American Statistical Association, 65(330):639–648, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284576.

Brewer:1979:CRS

[Bre79] K. R. W. Brewer. A class of robust sampling designs for largescale surveys. Journal of the American Statistical Association, 74 (368):911–915, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286422.

Broffitt:1976:DFP

[BRH76] James D. Broffitt, Ronald H. Randles, and Robert V. Hogg. Distribution-free partial discriminant analysis. Journal of the American Statistical Association, 71(356):934–939, December 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286864.

Brill:1971:BRB

[Bri71] Edward A. Brill. Book review: [untitled]. Journal of the American Statistical Association, 66(336):921, December 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284260.

Brillinger:1972:BRB

[Bri72] David R. Brillinger. Book review: [untitled]. Journal of the American Statistical Association, 67(340):959, December 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284679.

Bross:1971:DOS

[Bro71a] I. D. J. Bross. Does an observed sequence of numbers follow a simple rule? (another look at Bode’s law): Comment. Journal of the American Statistical Association, 66(335):562–564, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283524.

Brown:1971:BRBb

[Bro71b] B. W. Brown. Book review: [untitled]. Journal of the American Statistical Association, 66(333):225–226, March 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284883.

Brown:1971:BRBa

[Bro71c] George F. Brown, Jr. Book review: [untitled]. Journal of the American Statistical Association, 66(333):222–223, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284877.

Brown:1971:BRBc

[Bro71d] George F. Brown, Jr. Book review: [untitled]. Journal of the American Statistical Association, 66(334):423–424, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283951.

Brown:1972:BRBa

[Bro72a] Byron Wm Brown. Book review: [untitled]. Journal of the American Statistical Association, 67(339):715, September 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284478.

Brown:1972:BRBb

[Bro72b] George F. Brown, Jr. Book review: [untitled]. Journal of the American Statistical Association, 67(339):716–717, September 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284481.

Brown:1972:BRBc

[Bro72c] George F. Brown, Jr. Book review: [untitled]. Journal of the American Statistical Association, 67(340):958–959, December 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284677.

Browne:1973:BRB

[Bro73] M. W. Browne. Book review: [untitled]. Journal of the American Statistical Association, 68(341):244–245, March 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284191.

Brown:1975:EIS

[Bro75] Lawrence D. Brown. Estimation with incompletely specified loss functions (the case of several location parameters). Journal of the American Statistical Association, 70(350):417–427, June 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285833.

Broder:1976:BYR

[Bro76a] Ivy E. Broder. Base year revisions in the FRB index: Problems in interpreting historical movements. Journal of the American Statistical Association, 71(353):62–67, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285731.

Brook:1976:CUR

[Bro76b] R. J. Brook. Corrigenda: On the Use of a Regret Function to Set Significance Points in Prior Tests of Estimation. Journal of the American Statistical Association, 71(356):1010, December 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286898.

Brook:1976:URF

[Bro76c] Richard J. Brook. On the use of a regret function to set significance points in prior tests of estimation. Journal of the American Statistical Association, 71(353):126–131, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285753.

Brown:1977:CCS

[Bro77] Lawrence D. Brown. Conditional confidence statements and confidence estimators: Comment. Journal of the American Statistical Association, 72(360):810–812, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286462.

Brockett:1978:BRBa

[Bro78a] Patrick L. Brockett. Book review: [untitled]. Journal of the American Statistical Association, 73(364):895–896, December 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286312.

Brockett:1978:BRBb

[Bro78b] Patrick L. Brockett. Book review: [untitled]. Journal of the American Statistical Association, 73(364):896, December 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286313.

Brown:1978:EVC

[Bro78c] Kenneth G. Brown. Estimation of variance components using residuals. Journal of the American Statistical Association, 73(361):141–146, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286535.

Brown:1978:EAP

[Bro78d] Byron W. Brown, Jr. The effect of air pollution upon mortality: a consideration of distributed lag models: Comment. Journal of the American Statistical Association, 73(363):472, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286582.

Brown:1979:BRB

[Bro79] James N. Brown. Book review: [untitled]. Journal of the American Statistical Association, 74(367):728–729, September 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287007.

Brubaker:1973:BRB

[Bru73] Earl R. Brubaker. Book review: [untitled]. Journal of the American Statistical Association, 68(343):754, September 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284826.

Bruckner:1978:BRB

[Bru78] Lawrence A. Bruckner. Book review: [untitled]. Journal of the American Statistical Association, 73(362):445–446, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286701.

Bryant:1978:EMN

[Bry78] Peter Bryant. Estimating mixtures of normal distributions and switching regressions: Comment. Journal of the American Statistical Association, 73(364):748–749, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286272.

Berry:1973:IPS

[BS73a] Donald A. Berry and Milton Sobel. An improved procedure for selecting the better of two Bernoulli populations. Journal of the American Statistical Association, 68(344):979–984, December 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284535.

Bowman:1973:RDS

[BS73b] K. O. Bowman and L. R. Shenton. Remarks on the distribution of \sqrt{b_{1}} in sampling from a normal mixture, and normal Type A distribution. Journal of the American Statistical Association, 68(344):998–1003, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284539.

Bhattacharya:1974:PTP

[BS74a] Samir K. Bhattacharya and Vishnu K. Srivastava. A preliminary test procedure in life testing. Journal of the American Statistical Association, 69(347):726–729, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286008.

Brown:1974:RPH

[BS74b] George F. Brown, Jr. and Lester P. Silverman. The retail price of heroin: Estimation and applications. Journal of the American Statistical Association, 69(347):595–606, September 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285989.

Brown:1975:SIC

[BS75] Robert A. Brown and Nozer D. Singpurwalla. Some inequalities for certain functions of order statistics from IFR distributions. Journal of the American Statistical Association, 70(349):245–247, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285411.

Bhapkar:1977:DWT

[BS77a] Vasant P. Bhapkar and Grant W. Somes. Distribution of Q when testing equality of matched proportions. Journal of the American Statistical Association, 72(359):658–661, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286235.

Boyett:1977:NOS

[BS77b] James M. Boyett and J. J. Shuster. Nonparametric one-sided tests in multivariate analysis with medical applications. Journal of the American Statistical Association, 72(359):665–668, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286237.

Burdick:1978:ECI

[BS78] Richard K. Burdick and Robert L. Sielken, Jr. Exact confidence intervals for linear combinations of variance components in nested classifications. Journal of the American Statistical Association, 73 (363):632–635, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286614.

Burdick:1979:VEB

[BS79] Richard K. Burdick and Robert L. Sielken, Jr. Variance estimation based on a superpopulation model in two-stage sampling. Journal of the American Statistical Association, 74(366):438–440, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286350.

Brindley:1972:DAA

[BT72] E. C. Brindley, Jr. and W. A. Thompson, Jr. Dependence and aging aspects of multivariate survival. Journal of the American Statistical Association, 67(340):822–830, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284644.

Beckman:1974:DAS

[BT74] R. J. Beckman and H. J. Trussell. The distribution of an arbitrary Studentized residual and the effects of updating in multiple regression. Journal of the American Statistical Association, 69(345):199– 201, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285524.

Box:1975:IAA

[BT75] G. E. P. Box and G. C. Tiao. Intervention analysis with applications to economic and environmental problems. Journal of the American Statistical Association, 70(349):70–79, March 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285379.

Box:1976:SIU

[BT76] G. E. P. Box and G. C. Tiao. Strong inconsistency from uniform priors: Comment. Journal of the American Statistical Association, 71(353):122, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285749.

Bechhofer:1978:TDS

[BT78] Robert E. Bechhofer and Bruce W. Turnbull. Two (k + 1)- decision selection procedures for comparing k normal means with a specified standard. Journal of the American Statistical Association, 73(362):385–392, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286670.

Buehler:1976:BRB

[Bue76] Robert J. Buehler. Book review: [untitled]. Journal of the American Statistical Association, 71(353):245–246, March 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285783.

Buehler:1977:CCS

[Bue77] Robert J. Buehler. Conditional confidence statements and confidence estimators: Comment. Journal of the American Statistical Association, 72(360):813–814, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286463.

Buehler:1979:BRB

[Bue79] Robert J. Buehler. Book review: [untitled]. Journal of the American Statistical Association, 74(365):255–256, March 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286783.

Bulgren:1971:RDN

[Bul71] W. G. Bulgren. On representations of the doubly non-central F distribution. Journal of the American Statistical Association, 66(333):184– 186, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284870.

Bulgren:1972:BRB

[Bul72] William G. Bulgren. Book review: [untitled]. Journal of the American Statistical Association, 67(338):498, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284427.

Burstein:1973:CAP

[Bur73] Herman Burstein. Close approximations of percentage points of the chi-square distribution an Poisson confidence limits. Journal of the American Statistical Association, 68(343):581–584, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284780.

Burstein:1974:CCA

[Bur74] Herman Burstein. Corrigenda: Close Approximations of Percentage Points of the Chi-Square Distribution and Poisson Confidence Limits. Journal of the American Statistical Association, 69(345):288, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285571.

Burstein:1975:FPC

[Bur75] Herman Burstein. Finite population correction for binomial confidence limits. Journal of the American Statistical Association, 70 (349):67–69, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285378.

Burdick:1979:RAC

[Bur79] Donald S. Burdick. Rows adjusted for columns versus simple row means in two-way ANOVA: When are the two hypotheses equivalent? Journal of the American Statistical Association, 74(366):457– 458, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286355.

Butler:1971:BRB

[But71] Calvin C. Butler. Book review: [untitled]. Journal of the American Statistical Association, 66(335):661, September 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283565.

Butterworth:1976:BRB

[But76] Richard W. Butterworth. Book review: [untitled]. Journal of the American Statistical Association, 71(355):773, September 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285630.

Butterworth:1977:BRB

[But77] Richard W. Butterworth. Book review: [untitled]. Journal of the American Statistical Association, 72(360):931, December 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286502.

Butler:1979:BRB

[But79] Gail Butler. Book review: [untitled]. Journal of the American Statistical Association, 74(367):727, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287005.

Bement:1970:CVW

[BW70] T. R. Bement and J. S. Williams. Corrigenda: Variance of Weighted Regression Estimators When Sampling Errors are Independent and Heteroscedastic. Journal of the American Statistical Association, 65(330):1011, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284620.

Bottoms:1972:AJC

[BW72] A. E. Bottoms and Monica A. Walker. The American jury: a critique. Journal of the American Statistical Association, 67(340):773–779, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284635.

Bezdek:1976:CCD

[BW76] Roger H. Bezdek and Robert M. Wendling. Current- and constantdollar input-output forecasts for the U.S. economy. Journal of the American Statistical Association, 71(355):543–551, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285581.

Bock:1973:SCP

[BYJ73] M. E. Bock, T. A. Yancey, and G. G. Judge. The statistical consequences of preliminary test estimators in regression. Journal of the American Statistical Association, 68(341):109–116, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284154.

C:1971:BRB

[C.71] E. L. C. Book review: [untitled]. Journal of the American Statistical Association, 66(336):924, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284270.

C:1972:BRBa

[C.72a] E. L. C. Book review: [untitled]. Journal of the American Statistical Association, 67(337):249, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284750.

C:1972:BRBb

[C.72b] E. L. C. Book review: [untitled]. Journal of the American Statistical Association, 67(337):249, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284752.

C:1972:BRBc

[C.72c] E. L. C. Book review: [untitled]. Journal of the American Statistical Association, 67(339):722, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284492.

C:1974:BRBa

[C.74a] E. L. C. Book review: [untitled]. Journal of the American Statistical Association, 69(345):282, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285562.

C:1974:BRBb

[C.74b] E. L. C. Book review: [untitled]. Journal of the American Statistical Association, 69(345):282, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285563.

Chang:1974:CBD

[CA74] P. C. Chang and A. A. Afifi. Classification based on dichotomous and continuous variables. Journal of the American Statistical Association, 69(346):336–339, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285653.

Chiacchierini:1977:TST

[CA77] Richard P. Chiacchierini and Jesse C. Arnold. A two-sample test for independence in 2 × 2 contingency tables with both margins subject to misclassification. Journal of the American Statistical Association, 72(357):170–174, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286932.

Costanza:1979:CSR

[CA79] Michael C. Costanza and A. A. Afifi. Comparison of stopping rules in forward stepwise discriminant analysis. Journal of the American Statistical Association, 74(368):777–785, December 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286399.

Campbell:1979:BRB

[Cam79] Cathy Campbell. Book review: [untitled]. Journal of the American Statistical Association, 74(365):247, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286767.

Canner:1970:SOT

[Can70] Paul L. Canner. Selecting one of two treatments when the responses are dichotomous. Journal of the American Statistical Association, 65(329):293–306, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283593.

Cane:1974:LEP

[Can74] Gwenda J. Cane. Linear estimation of parameters of the Cauchy distribution based on sample quantiles. Journal of the American Statistical Association, 69(345):243–245, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285535.

Canner:1975:SSO

[Can75] Paul L. Canner. A simulation study of one- and two-sample Kolmogorov–Smirnov statistics with a particular weight function. Journal of the American Statistical Association, 70(349):209–211, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285405.

Carney:1970:DTP

[Car70] M. K. Carney. Developments in trading patterns in the common market and EFTA. Journal of the American Statistical Association, 65(332):1455–1459, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284329.

Carlson:1975:PEN

[Car75] John A. Carlson. Are price expectations normally distributed? Journal of the American Statistical Association, 70(352):749–754, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285433.

Carroll:1977:CTA

[Car77] Raymond J. Carroll. A comparison of two approaches to fixed-width confidence interval estimation. Journal of the American Statistical Association, 72(360):901–907, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286484.

Carroll:1978:SCI

[Car78] Raymond J. Carroll. Sequential confidence intervals for the mean of a subpopulation of a finite population. Journal of the American Statistical Association, 73(362):408–413, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286674.

Carroll:1979:EVR

[Car79] Raymond J. Carroll. On estimating variances of robust estimators when the errors are asymmetric. Journal of the American Statistical Association, 74(367):674–679, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286990.

Castellan:1978:BRB

[Cas78] N. John Castellan, Jr. Book review: [untitled]. Journal of the American Statistical Association, 73(364):889–891, December 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286305.

Chatterjee:1970:NAC

[CB70] Samprit Chatterjee and Seymour Barcun. A nonparametric approach to credit screening. Journal of the American Statistical Association, 65(329):150–154, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283581.

Chase:1971:MCI

[CB71] G. R. Chase and William G. Bulgren. A Monte Carlo investigation of the robustness of T^{2}. Journal of the American Statistical Association, 66(335):499–502, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283515.

Church:1973:ESK

[CC73] J. D. Church and E. Benton Cobb. On the equivalence of Spearman– Karber and maximum likelihood estimates of the mean. Journal of the American Statistical Association, 68(341):201–202, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284170.

Chan:1974:ADA

[CC74a] Lai K. Chan and Smiley W. Cheng. An algorithm for determining the asymptotically best linear estimate of the mean from multiply censored logistic data. Journal of the American Statistical Association, 69(348):1027–1030, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286183.

Church:1974:MSE

[CC74b] J. D. Church and E. Benton Cobb. A modified Spearman estimator of a location or scale parameter. Journal of the American Statistical Association, 69(345):207–210, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285526.

Chan:1971:BLU

[CCM71] Lai K. Chan, N. N. Chan, and E. R. Mead. Best linear unbiased estimates of the parameters of the logistic distribution based on selected order statistics. Journal of the American Statistical Association, 66(336):889–892, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284248.

Chan:1972:TMV

[CD72a] Linda S. Chan and Olive Jean Dunn. The treatment of missing values in discriminant analysis-1. the sampling experiment. Journal of the American Statistical Association, 67(338):473–477, June 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284409.

Chen:1972:EPC

[CD72b] Edwin H. Chen and W. J. Dixon. Estimates of parameters of a censored regression sample. Journal of the American Statistical Association, 67(339):664–671, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284463.

Chan:1974:NAA

[CD74] Linda S. Chan and Olive Jean Dunn. A note on the asymptotic aspect of the treatment of missing values in discriminant analysis. Journal of the American Statistical Association, 69(347):672–673, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285999.

Chen:1976:PFW

[CD76] Hubert J. Chen and Edward J. Dudewicz. Procedures for fixed-width interval estimation of the largest normal mean. Journal of the American Statistical Association, 71(355):752–756, September 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285615.

Cebula:1974:IMT

[Ceb74] Richard J. Cebula. Interstate migration and the Tiebout hypothesis: An analysis according to race, sex and age. Journal of the American Statistical Association, 69(348):876–879, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286156.

Cooper:1972:SSM

[CF72] J. Phillip Cooper and Stanley Fischer. Stochastic simulation of monetary rules in two macroeconomic models. Journal of the American Statistical Association, 67(340):750–760, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284632.

Chhikara:1974:EIG

[CF74] Raj S. Chhikara and J. Leroy Folks. Estimation of the inverse Gaussian distribution function. Journal of the American Statistical Association, 69(345):250–254, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285537.

Charnes:1976:EGL

[CFY76] A. Charnes, E. L. Frome, and P. L. Yu. The equivalence of generalized least squares and maximum likelihood estimates in the exponential family. Journal of the American Statistical Association, 71(353):169– 171, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285762.

Cornell:1970:MPC

[CG70] John A. Cornell and I. J. Good. The mixture problem for categorized components. Journal of the American Statistical Association, 65(329):339–355, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283597.

Cornell:1971:CMP

[CG71] John A. Cornell and I. J. Good. Corrigenda: “The Mixture Problem for Categorized Components. Journal of the American Statistical Association, 66(333):229, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284890.

Chatfield:1973:CPM

[CG73] C. Chatfield and G. J. Goodhardt. A consumer purchasing model with Erlang inter-purchase time. Journal of the American Statistical Association, 68(344):828–835, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284508.

Cummings:1974:VCT

[CG74] W. B. Cummings and D. W. Gaylor. Variance component testing in unbalanced nested designs. Journal of the American Statistical Association, 69(347):765–771, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286015.

Chao:1978:EDB

[CG78] Min-Te Chao and Ronald E. Glaser. The exact distribution of Bartlett’s test statistic for homogeneity of variances with unequal sample sizes. Journal of the American Statistical Association, 73 (362):422–426, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286677.

Chan:1976:AAM

[CGD76] Linda S. Chan, June Aono Gilman, and Olive Jean Dunn. Alternative approaches to missing values in discriminant analysis. Journal of the American Statistical Association, 71(356):842–844, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286848.

Claypool:1974:ANE

[CH74a] P. L. Claypool and Donald Holbert. Accuracy of normal and Edgeworth approximations to the distribution of the Wilcoxon signed rank statistic. Journal of the American Statistical Association, 69(345):255–258, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285538.

Claypool:1974:CAN

[CH74b] P. L. Claypool and Donald Holbert. Corrigenda: Accuracy of Normal and Edgewood Approximations to the Distribution of the Wilcoxon Signed Rank Statistic. Journal of the American Statistical Association, 69(346):587, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285720.

Chow:1975:ADW

[CH75] Winston K. Chow and J. L. Hodges, Jr. An approximation for the distribution of the Wilcoxon one-sample statistic. Journal of the American Statistical Association, 70(351):648–655, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285949.

Crowley:1977:CAH

[CH77] John Crowley and Marie Hu. Covariance analysis of heart transplant survival data. Journal of the American Statistical Association, 72(357):27–36, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286902.

Chromy:1978:UMI

[CH78a] James R. Chromy and Daniel G. Horvitz. The use of monetary incentives in national assessment household surveys. Journal of the American Statistical Association, 73(363):473–478, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286583.

Chung:1978:DSP

[CH78b] F. R. K. Chung and F. K. Hwang. Do stronger players win more knockout tournaments? Journal of the American Statistical Association, 73(363):593–596, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286606.

Clark:1978:EMN

[CH78c] B. R. Clark and C. R. Heathcote. Estimating mixtures of normal distributions and switching regressions: Comment. Journal of the American Statistical Association, 73(364):749–750, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286273.

Cook:1979:FEW

[CH79] R. Dennis Cook and Norton Holschuh. Field experimentation in weather modification: Comment. Journal of the American Statistical Association, 74(365):68–70, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286723.

Chan:1970:LEL

[Cha70a] Lai K. Chan. Linear estimation of the location and scale parameters of the Cauchy distribution based on sample quantiles. Journal of the American Statistical Association, 65(330):851–859, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284592.

Chase:1970:BRB

[Cha70b] G. R. Chase. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1419–1420, September 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284321.

Chai:1971:CME

[Cha71a] John J. Chai. Correlated measurement errors and the least squares estimator of the regression coefficient. Journal of the American Statistical Association, 66(335):478–483, September 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283511.

Chambers:1971:RU

[Cha71b] John M. Chambers. Regression updating. Journal of the American Statistical Association, 66(336):744–748, December 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284221.

Chanda:1971:ADS

[Cha71c] K. C. Chanda. Asymptotic distribution of the sample size for a sequential probability ratio test. Journal of the American Statistical Association, 66(333):178–183, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284869.

Chatterjee:1971:BRB

[Cha71d] Samprit Chatterjee. Book review: [untitled]. Journal of the American Statistical Association, 66(334):425–426, June 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283955.

Chase:1972:CST

[Cha72] Gerald R. Chase. On the chi-square test when the parameters are estimated independently of the sample. Journal of the American Statistical Association, 67(339):609–611, September 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284448.

Chakravarti:1973:BRB

[Cha73] I. M. Chakravarti. Book review: [untitled]. Journal of the American Statistical Association, 68(342):493–494, June 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284112.

Chakravorti:1974:BRB

[Cha74a] S. R. Chakravorti. Book review: [untitled]. Journal of the American Statistical Association, 69(346):580, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285707.

Chand:1974:STC

[Cha74b] Nanak Chand. Sequential tests of composite hypotheses. Journal of the American Statistical Association, 69(346):394–397, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285667.

Chakravarti:1975:BRB

[Cha75a] I. M. Chakravarti. Book review: [untitled]. Journal of the American Statistical Association, 70(350):490, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285855.

Chandra:1975:BRB

[Cha75b] Satish Chandra. Book review: [untitled]. Journal of the American Statistical Association, 70(352):961–962, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285472.

Chaiken:1976:BRB

[Cha76a] Jan M. Chaiken. Book review: [untitled]. Journal of the American Statistical Association, 71(353):250, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285791.

Chambers:1976:BRB

[Cha76b] John M. Chambers. Book review: [untitled]. Journal of the American Statistical Association, 71(355):768–769, September 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285621.

Chapman:1976:CMP

[Cha76c] Judith-Anne W. Chapman. A comparison of the χ^{2}, −2 log R, and multinomial probability criteria for significance tests when expected frequencies are small. Journal of the American Statistical Association, 71(356):854–863, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286851.

Chalmers:1978:BRB

[Cha78] Colin P. Chalmers. Book review: [untitled]. Journal of the American Statistical Association, 73(364):894–895, December 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286311.

Chaiken:1979:BRB

[Cha79] Jan M. Chaiken. Book review: [untitled]. Journal of the American Statistical Association, 74(365):243–244, March 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286760.

Chew:1970:CME

[Che70] Victor Chew. Covariance matrix estimation in linear models. Journal of the American Statistical Association, 65(329):173–181, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283584.

Chen:1971:PSW

[Che71a] Edwin H. Chen. The power of the Shapiro–Wilk W test for normality in samples from contaminated normal distributions. Journal of the American Statistical Association, 66(336):760–762, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284225.

Chen:1971:RNN

[Che71b] Edwin H. Chen. A random normal number generator for 32bit-word computers. Journal of the American Statistical Association, 66(334):400–403, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283944.

Chernoff:1973:BRB

[Che73a] Herman Chernoff. Book review: [untitled]. Journal of the American Statistical Association, 68(343):747–748, September 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284814.

Chernoff:1973:UFR

[Che73b] Herman Chernoff. The use of faces to represent points in Kdimensional space graphically. Journal of the American Statistical Association, 68(342):361–368, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284077.

Cheng:1975:UAC

[Che75] Smiley W. Cheng. A unified approach to choosing optimum quantiles for the ABLE’s. Journal of the American Statistical Association, 70(349):155–159, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285392.

Chen:1978:SSC

[Che78] Rina Chen. A surveillance system for congenital malformations. Journal of the American Statistical Association, 73(362):323–327, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286660.

Chen:1979:LLM

[Che79] T. Timothy Chen. Log-linear models for categorical data with misclassification and double sampling. Journal of the American Statistical Association, 74(366):481–488, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286359.

Childs:1971:BRB

[Chi71] Gerald L. Childs. Book review: [untitled]. Journal of the American Statistical Association, 66(334):423, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283950.

Chiang:1972:CCL

[Chi72] Chin Long Chiang. On constructing current life tables. Journal of the American Statistical Association, 67(339):538–541, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284435.

Chiang:1973:CCC

[Chi73a] Chin Long Chiang. Corrigenda: On Constructing Current Life Tables. Journal of the American Statistical Association, 68(343):757, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284833.

Chiu:1973:CED

[Chi73b] W. K. Chiu. Comments on the economic design of X-charts. Journal of the American Statistical Association, 68(344):919–921, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284522.

Chitturi:1974:DRA

[Chi74] Ratnam V. Chitturi. Distribution of residual autocorrelations in multiple autoregressive schemes. Journal of the American Statistical Association, 69(348):928–934, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286165.

Chitturi:1976:DMW

[Chi76] Ratnam V. Chitturi. Distribution of multivariate white noise autocorrelations. Journal of the American Statistical Association, 71(353):223–226, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285773.

Choi:1971:STC

[Cho71] Sung C. Choi. Sequential test for correlation coefficients. Journal of the American Statistical Association, 66(335):575–576, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283531.

Chow:1976:BRB

[Cho76] Gregory C. Chow. Book review: [untitled]. Journal of the American Statistical Association, 71(355):771, September 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285625.

Chong:1977:CEG

[Cho77] Kong-Ming Chong. On characterizations of the exponential and geometric distributions by expectations. Journal of the American Statistical Association, 72(357):160–161, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286929.

Christensen:1973:SSI

[Chr73] Laurits R. Christensen. Simultaneous statistical inference in the normal multiple linear regression model. Journal of the American Statistical Association, 68(342):457–461, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284098.

Christ:1979:SAE

[Chr79] Carl F. Christ. Statistical analysis of econometric models: Comment. Journal of the American Statistical Association, 74(367):645–646, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286983.

Chow:1976:BLU

[CL76] Gregory C. Chow and An-Loh Lin. Best linear unbiased estimation of missing observations in an economic time series. Journal of the American Statistical Association, 71(355):719–721, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285608.

Clark:1970:BRB

[Cla70a] C. Scott Clark. Book review: [untitled]. Journal of the American Statistical Association, 65(332):1689–1691, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284358.

Clatworthy:1970:BRB

[Cla70b] Willard H. Clatworthy. Book review: [untitled]. Journal of the American Statistical Association, 65(330):991–993, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284605.

Claypool:1970:LIW

[Cla70c] P. L. Claypool. Linear interpolation within McCornack’s table of the Wilcoxon matched pair signed rank statistic. Journal of the American Statistical Association, 65(330):974–975, June 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284601.

Clatworthy:1971:BRB

[Cla71] Willard H. Clatworthy. Book review: [untitled]. Journal of the American Statistical Association, 66(336):920–921, December 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284259.

Clelland:1970:BRB

[Cle70a] Richard C. Clelland. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1401–1403, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284308.

Cleroux:1970:UIV

[Cle70b] Robert Cleroux. On using an incorrect value of —rho in group divisible two-associate PBIB designs. Journal of the American Statistical Association, 65(330):844–850, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284591.

Clelland:1971:BRB

[Cle71] Richard C. Clelland. Book review: [untitled]. Journal of the American Statistical Association, 66(335):655, September 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283554.

Cleveland:1979:RLW

[Cle79] William S. Cleveland. Robust locally weighted regression and smoothing scatterplots. Journal of the American Statistical Association, 74(368):829–836, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286407.

Coale:1972:DAF

[CM72] A. J. Coale and D. R. McNeil. The distribution by age of the frequency of first marriage in a female cohort. Journal of the American Statistical Association, 67(340):743–749, December 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284631.

Carter:1973:MLE

[CM73] Walter H. Carter, Jr. and Raymond H. Myers. Maximum likelihood estimation from linear combinations of discrete probability functions. Journal of the American Statistical Association, 68(341):203–206, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284171.

Cargill:1974:WPU

[CM74a] Thomas F. Cargill and Robert A. Meyer. Wages, prices and unemployment: Distributed lag estimates. Journal of the American Statistical Association, 69(345):98–107, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285506.

Carter:1974:EGV

[CM74b] Walter H. Carter, Jr. and Raymond H. Myers. The effect of grouping on the variance and bias of the maximum likelihood estimator of the Poisson parameter–some Monte Carlo results. Journal of the American Statistical Association, 69(346):482–484, June 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285683.

Cook:1974:MQS

[CM74c] R. Dennis Cook and Frank B. Martin. A model for quadrat sampling with “visibility bias”. Journal of the American Statistical Association, 69(346):345–349, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285655.

Colton:1976:TSP

[CM76] Theodore Colton and Klim McPherson. Two-stage plans compared with fixed-sample-size and Wald SPRT plans. Journal of the American Statistical Association, 71(353):80–86, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285735.

Cote:1973:MBA

[CMH73] R. Cote, A. R. Manson, and R. J. Hader. Minimum bias approximation of a general regression model. Journal of the American Statistical Association, 68(343):633–638, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284791.

Chambers:1976:MSS

[CMS76] J. M. Chambers, C. L. Mallows, and B. W. Stuck. A method for simulating stable random variables. Journal of the American Statistical Association, 71(354):340–344, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2285309. See correction [CMS87].

Chambers:1987:CMS

[CMS87] J. M. Chambers, C. L. Mallows, and B. W. Stuck. Correction to: “A method for simulating stable random variables” [J. Amer. Statist. Assoc. 71 (1976), no. 354, 340–344, MR 54 #4059]. Journal of the American Statistical Association, 82(398):704, ???? 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CMS76].

Cochran:1970:BRBb

[Coc70a] W. G. Cochran. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1396–1398, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284305.

Cochran:1970:SEE

[Coc70b] W. G. Cochran. Some effects of errors of measurement on multiple correlation. Journal of the American Statistical Association, 65(329): 22–34, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2283572.

Cochran:1970:BRBa

[Coc70c] William G. Cochran. Book review: [untitled]. Journal of the American Statistical Association, 65(329):456–457, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283611.

Cochran:1973:ENF

[Coc73] W. G. Cochran. Experiments for nonlinear functions. Journal of the American Statistical Association, 68(344):771–781, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284499.

Cogger:1973:SA

[Cog73] Kenneth O. Cogger. Specification analysis. Journal of the American Statistical Association, 68(344):899–905, December 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284519.

Cohen:1971:BRB

[Coh71] A. C. Cohen, Jr. Book review: [untitled]. Journal of the American Statistical Association, 66(333):224, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284879.

Cohen:1972:ICI

[Coh72] Arthur Cohen. Improved confidence intervals for the variance of a normal distribution. Journal of the American Statistical Association, 67(338):382–387, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284389.

Cohen:1974:PPH

[Coh74a] Arthur Cohen. To pool or not to pool in hypothesis testing. Journal of the American Statistical Association, 69(347):721–725, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286007.

Cohen:1974:BRB

[Coh74b] Joel E. Cohen. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1046–1047, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286189.

Cohen:1976:CEL

[Coh76a] Arthur Cohen. Combining estimates of location. Journal of the American Statistical Association, 71(353):172–175, March 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285763.

Cohen:1976:DCS

[Coh76b] Joel E. Cohen. The distribution of the chi-squared statistic under clustered sampling from contingency tables. Journal of the American Statistical Association, 71(355):665–670, September 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285597.

Cohen:1976:EAS

[Coh76c] Joel E. Cohen. Ergodicity of age structure in populations with Markovian vital rates, I: Countable states. Journal of the American Statistical Association, 71(354):335–339, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285308.

Cohn:1978:BRB

[Coh78] Richard Cohn. Book review: [untitled]. Journal of the American Statistical Association, 73(363):686–687, September 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286642.

Colton:1972:BRB

[Col72] Theodore Colton. Book review: [untitled]. Journal of the American Statistical Association, 67(338):491–492, June 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284415.

Collins:1975:BRB

[Col75] John R. Collins. Book review: [untitled]. Journal of the American Statistical Association, 70(351):727–728, September 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285964.

Connor:1972:GTT

[Con72a] Robert J. Connor. Grouping for testing trends in categorical data. Journal of the American Statistical Association, 67(339):601–604, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284446.

Conover:1972:KGF

[Con72b] W. J. Conover. A Kolmogorov goodness-of-fit test for discontinuous distributions. Journal of the American Statistical Association, 67(339):591–596, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284444.

Conover:1973:MHT

[Con73] William Jay Conover. On methods of handling ties in the Wilcoxon signed-rank test. Journal of the American Statistical Association, 68(344):985–988, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284536.

Conlisk:1974:NCF

[Con74a] John Conlisk. Notes on a canonical form theorem useful in regression designs. Journal of the American Statistical Association, 69(345):196–198, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285523.

Conover:1974:SRUa

[Con74b] W. J. Conover. Some reasons for not using the Yates continuity correction on 2 × 2 contingency tables. Journal of the American Statistical Association, 69(346):374–376, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285661.

Conover:1974:SRUb

[Con74c] W. J. Conover. Some reasons for not using the Yates continuity correction on 2 × 2 contingency tables: Rejoinder. Journal of the American Statistical Association, 69(346):382, June 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285665.

Constantinides:1976:BRBa

[Con76a] George M. Constantinides. Book review: [untitled]. Journal of the American Statistical Association, 71(354):526–527, June 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285351.

Constantinides:1976:BRBb

[Con76b] George M. Constantinides. Book review: [untitled]. Journal of the American Statistical Association, 71(354):527, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285352.

Cooke:1971:SEU

[Coo71a] Peter J. Cooke. Sequential estimation in the uniform density. Journal of the American Statistical Association, 66(335):614–617, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283540.

Cootner:1971:BRB

[Coo71b] Paul H. Cootner. Book review: [untitled]. Journal of the American Statistical Association, 66(336):919, December 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284257.

Cooke:1973:TSS

[Coo73] Peter J. Cooke. Two-stage sequential estimation in the uniform density. Journal of the American Statistical Association, 68(341):107– 108, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284153.

Cook:1974:BRB

[Coo74] R. Dennis Cook. Book review: [untitled]. Journal of the American Statistical Association, 69(347):832–833, September 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286035.

Cook:1979:IOL

[Coo79] R. Dennis Cook. Influential observations in linear regression. Journal of the American Statistical Association, 74(365):169–174, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286747.

Cornell:1971:PVM

[Cor71] John A. Cornell. Process variables in the mixture problem for categorized components. Journal of the American Statistical Association, 66(333):42–48, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284844.

Cornfield:1975:SA

[Cor75] Jerome Cornfield. A Statistician’s apology. Journal of the American Statistical Association, 70(349):7–14, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285368.

Cornelius:1976:ALD

[Cor76] P. L. Cornelius. The analysis of lattice designs with concomitant variables associated with replications. Journal of the American Statistical Association, 71(354):445–450, June 1976. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285331.

Cowell:1975:ELI

[Cow75] Frank A. Cowell. On the estimation of a lifetime income — a correction. Journal of the American Statistical Association, 70 (351):588–589, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285937.

Cox:1971:IEP

[Cox71] C. Philip Cox. Interval estimation for X-predictions from linear Y - on-X regression lines through the origin. Journal of the American Statistical Association, 66(336):749–751, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284222.

Cox:1972:CIE

[Cox72] C. Philip Cox. Corrigenda: Interval Estimation for X-Predictions from Linear Y -On-X Regression Lines Through the Origin. Journal of the American Statistical Association, 67(337):252, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284757.

Cox:1973:BRB

[Cox73] D. F. Cox. Book review: [untitled]. Journal of the American Statistical Association, 68(344):1029, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284554.

Cozzolino:1974:CDI

[Coz74] John M. Cozzolino. Conjugate distributions for incomplete observations. Journal of the American Statistical Association, 69(345):264– 266, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285540.

Cargill:1972:TFD

[CR72] Thomas F. Cargill and Gordon C. Rausser. Time and frequency domain representations of futures prices as a stochastic process. Journal of the American Statistical Association, 67(337):23–30, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284697.

Carter:1974:EBM

[CR74] Grace M. Carter and John E. Rolph. Empirical Bayes methods applied to estimating fire alarm probabilities. Journal of the American Statistical Association, 69(348):880–885, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286157.

Chernoff:1975:ECE

[CR75a] Herman Chernoff and M. Haseeb Rizvi. Effect on classification error of random permutations of features in representing multivariate data by faces. Journal of the American Statistical Association, 70(351):548–554, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285931.

Cleveland:1975:CIS

[CR75b] William S. Cleveland and Daniel A. Relles. Clustering by identification with special application to two-way tables of counts. Journal of the American Statistical Association, 70(351):626–630, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285945.

Cryer:1975:IEP

[CR75c] J. D. Cryer and Tim Robertson. Isotonic estimation of the probability of extinction of a branching process. Journal of the American Statistical Association, 70(352):905–912, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285456.

Craig:1970:PAQ

[Cra70] George D. Craig. Predictive accuracy of quarterly and annual aggregate saving functions. Journal of the American Statistical Association, 65(331):1131–1145, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284280.

Crain:1973:NDE

[Cra73] Bradford R. Crain. A note on density estimation using orthogonal expansions. Journal of the American Statistical Association, 68(344):964–965, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284531.

Cran:1975:NCL

[Cra75] G. W. Cran. A note on Chernoff and Lieberman’s generalized probability paper. Journal of the American Statistical Association, 70(349):229–232, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285408.

Crain:1976:EMM

[Cra76a] Bradford R. Crain. Exponential models, maximum likelihood estimation, and the Haar condition. Journal of the American Statistical Association, 71(355):737–740, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285612.

Crain:1976:MED

[Cra76b] Bradford R. Crain. More on estimation of distributions using orthogonal expansions. Journal of the American Statistical Association, 71(355):741–745, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285613.

Cramer:1976:CNA

[Cra76c] Eliot M. Cramer. Corrigenda: a Nonorthogonal Analysis of Variance Program. Journal of the American Statistical Association, 71(356):1010, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286895.

Cramer:1976:NAV

[Cra76d] Elliot M. Cramer. A nonorthogonal analysis of variance program. Journal of the American Statistical Association, 71(353):93–95, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285737.

Crow:1971:BRB

[Cro71] Edwin L. Crow. Book review: [untitled]. Journal of the American Statistical Association, 66(333):227, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284887.

Crowley:1974:ANN

[Cro74] John Crowley. Asymptotic normality of a new nonparametric statistic for use in organ transplant studies. Journal of the American Statistical Association, 69(348):1006–1011, December 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286180.

Cross:1979:BRB

[Cro79a] Anne R. Cross. Book review: [untitled]. Journal of the American Statistical Association, 74(366):512–513, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286379.

Crow:1979:NSD

[Cro79b] Edwin L. Crow. Nonparametric statistical data modeling: Comment. Journal of the American Statistical Association, 74(365):129, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286740.

Crump:1975:BRB

[Cru75] Kenny Crump. Book review: [untitled]. Journal of the American Statistical Association, 70(352):964, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285479.

Chao:1972:NMP

[CS72] M. T. Chao and W. E. Strawderman. Negative moments of positive random variables. Journal of the American Statistical Association, 67(338):429–431, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284399.

Carmer:1973:ETP

[CS73] S. G. Carmer and M. R. Swanson. An evaluation of ten pairwise multiple comparison procedures by Monte Carlo methods. Journal of the American Statistical Association, 68(341):66–74, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284140.

Chikkagoudar:1974:CFR

[CS74] M. S. Chikkagoudar and J. J. Shuster. Comparison of failure rates using rank tests. Journal of the American Statistical Association, 69(346):411–413, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285670.

Campbell:1973:ADM

[CT73] Janet W. Campbell and Chris P. Tsokos. The asymptotic distribution of maxima in bivariate samples. Journal of the American Statistical Association, 68(343):734–739, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284810.

Cleveland:1976:DST

[CT76] W. P. Cleveland and G. C. Tiao. Decomposition of seasonal time series: a model for the census X-11 program. Journal of the American Statistical Association, 71(355):581–587, September 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285586.

Carter:1971:MLE

[CVM71] Walter H. Carter, Jr., Jacob Van Bowen, Jr., and Raymond H. Myers. Maximum likelihood estimation from grouped Poisson data. Journal of the American Statistical Association, 66(334):351–353, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283935.

Conover:1978:NSS

[CWR78] W. J. Conover, Oscar Wehmanen, and Fred L. Ramsey. A note on the small-sample power functions for nonparametric tests of location in the double exponential family. Journal of the American Statistical Association, 73(361):188–190, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286543.

Clevenson:1975:SEM

[CZ75] M. Lawrence Clevenson and James V. Zidek. Simultaneous estimation of the means of independent Poisson laws. Journal of the American Statistical Association, 70(351):698–705, September 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285958.

Clevenson:1977:BLE

[CZ77] M. Lawrence Clevenson and James V. Zidek. Bayes linear estimators of the intensity function of the nonstationary Poisson process. Journal of the American Statistical Association, 72(357):112–120, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286918.

Czamanski:1970:BRB

[Cza70] Stan Czamanski. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1408–1411, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284313.

D:1972:BRBa

[D.72a] M. H. D. Book review: [untitled]. Journal of the American Statistical Association, 67(337):249, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284753.

D:1972:BRBb

[D.72b] M. H. D. Book review: [untitled]. Journal of the American Statistical Association, 67(340):963, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284686.

D:1973:BRBa

[D.73a] M. H. D. Book review: [untitled]. Journal of the American Statistical Association, 68(342):500, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284123.

D:1973:BRBb

[D.73b] M. H. D. Book review: [untitled]. Journal of the American Statistical Association, 68(344):1030, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284558.

D:1974:BRBa

[D.74a] M. H. D. Book review: [untitled]. Journal of the American Statistical Association, 69(345):282–283, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285566.

D:1974:BRBb

[D.74b] M. H. D. Book review: [untitled]. Journal of the American Statistical Association, 69(345):283, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285567.

D:1974:BRBc

[D.74c] M. H. D. Book review: [untitled]. Journal of the American Statistical Association, 69(346):583, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285715.

D:1974:BRBd

[D.74d] M. H. D. Book review: [untitled]. Journal of the American Statistical Association, 69(346):584, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285718.

D:1974:BRBe

[D.74e] M. H. D. Book review: [untitled]. Journal of the American Statistical Association, 69(347):835–836, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286041.

D:1974:BRBf

[D.74f] M. H. D. Book review: [untitled]. Journal of the American Statistical Association, 69(347):836, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286042.

D:1974:BRBg

[D.74g] M. H. D. Book review: [untitled]. Journal of the American Statistical Association, 69(347):836, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286043.

D:1974:BRBh

[D.74h] M. H. D. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1052–1053, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286200.

D:1975:BRBa

[D.75a] M. H. D. Book review: [untitled]. Journal of the American Statistical Association, 70(349):259, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285429.

D:1975:BRBb

[D.75b] M. H. D. Book review: [untitled]. Journal of the American Statistical Association, 70(349):259, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285430.

D:1975:BRBc

[D.75c] M. H. D. Book review: [untitled]. Journal of the American Statistical Association, 70(351):736, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285984.

D:1975:BRBd

[D.75d] M. H. D. Book review: [untitled]. Journal of the American Statistical Association, 70(351):736, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285985.

Dagenais:1971:FSC

[Dag71] Marcel G. Dagenais. Further suggestions concerning the utilization of incomplete observations in regression analysis. Journal of the American Statistical Association, 66(333):93–98, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284855.

Dahiya:1974:EMS

[Dah74] Ram C. Dahiya. Estimation of the mean of the selected population. Journal of the American Statistical Association, 69(345):226– 230, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285531.

Dallal:1971:BRB

[Dal71] G. E. Dallal. Book review: [untitled]. Journal of the American Statistical Association, 66(334):426, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283956.

Dallal:1977:BRB

[Dal77] Gerard E. Dallal. Book review: [untitled]. Journal of the American Statistical Association, 72(359):682–683, September 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286242.

Daly:1978:BRB

[Dal78] James C. Daly. Book review: [untitled]. Journal of the American Statistical Association, 73(363):680–681, September 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286629.

Daly:1979:BRB

[Dal79] James C. Daly. Book review: [untitled]. Journal of the American Statistical Association, 74(366):503, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286365.

Danby:1971:DOS

[Dan71] J. M. A. Danby. Does an observed sequence of numbers follow a simple rule? (another look at Bode’s law): Comment. Journal of the American Statistical Association, 66(335):565–566, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283526.

Daniel:1973:OTP

[Dan73] Cuthbert Daniel. One-at-a-time plans. Journal of the American Statistical Association, 68(342):353–360, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284076.

Davidson:1970:EBT

[Dav70] Roger R. Davidson. On extending the Bradley–Terry model to accommodate ties in paired comparison experiments. Journal of the American Statistical Association, 65(329):317–328, March 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283595.

Davies:1971:RTL

[Dav71] Robert B. Davies. Rank tests for “Lehmann’s alternative”. Journal of the American Statistical Association, 66(336):879–883, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284246.

David:1972:BRB

[Dav72a] H. T. David. Book review: [untitled]. Journal of the American Statistical Association, 67(339):717, September 1972. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284482.

Davis:1972:BRB

[Dav72b] E. W. Davis. Book review: [untitled]. Journal of the American Statistical Association, 67(337):242–243, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284737.

David:1973:WTP

[Dav73a] H. A. David. Waiting time paradoxes and order statistics. Journal of the American Statistical Association, 68(343):743–745, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284812.

Davis:1973:BRB

[Dav73b] Herbert T. Davis. Book review: [untitled]. Journal of the American Statistical Association, 68(341):241–242, March 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284184.

Davis:1976:BRB

[Dav76] William Davis. Book review: [untitled]. Journal of the American Statistical Association, 71(355):775, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285634.

Davis:1977:DEA

[Dav77a] A. W. Davis. A differential equation approach to linear combinations of independent chi-squares. Journal of the American Statistical Association, 72(357):212–214, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286940.

Davis:1977:BRB

[Dav77b] William W. Davis. Book review: [untitled]. Journal of the American Statistical Association, 72(360):927–928, December 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286495.

Davis:1978:BAL

[Dav78] William W. Davis. Bayesian analysis of the linear model subject to linear inequality constraints. Journal of the American Statistical Association, 73(363):573–579, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286603.

Davis:1979:BRB

[Dav79a] Herbert T. Davis. Book review: [untitled]. Journal of the American Statistical Association, 74(367):735, September 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2287016.

Davis:1979:ABP

[Dav79b] William W. Davis. Approximate Bayesian predictive distributions and model selection. Journal of the American Statistical Association, 74(366):312–317, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286329.

Davison:1979:BRB

[Dav79c] Mark L. Davison. Book review: [untitled]. Journal of the American Statistical Association, 74(368):935–936, December 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286443.

Dawid:1976:BRB

[Daw76] A. P. Dawid. Book review: [untitled]. Journal of the American Statistical Association, 71(356):998–999, December 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286878.

Dayal:1979:BRB

[Day79] Hari H. Dayal. Book review: [untitled]. Journal of the American Statistical Association, 74(368):933, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286438.

Dunn:1971:CTE

[DC71] Olive Jean Dunn and Virginia Clark. Comparison of tests of the equality of dependent correlation coefficients. Journal of the American Statistical Association, 66(336):904–908, December 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284252.

DAgostino:1973:CSL

[DC73] Ralph B. D’Agostino and Edward E. Cureton. A class of simple linear estimators of the standard deviation of the normal distribution. Journal of the American Statistical Association, 68(341):207–210, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284172.

Dent:1977:MCAa

[DC77a] Warren T. Dent and David C. Cavander. More on computational accuracy in regression. Journal of the American Statistical Association, 72(359):598–600, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286223.

Dent:1977:MCAb

[DC77b] Warren T. Dent and David C. Cavander. More on computational accuracy in regression: Rejoinder. Journal of the American Statistical Association, 72(359):601–602, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286225.

Dixon:1975:MBR

[DD75] Dennis O. Dixon and David B. Duncan. Minimum Bayes risk tintervals for multiple comparisons. Journal of the American Statistical Association, 70(352):822–831, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285443.

Dawid:1977:LBI

[DD77] A. P. Dawid and James M. Dickey. Likelihood and Bayesian inference from selectively reported data. Journal of the American Statistical Association, 72(360):845–850, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286472.

Dayananda:1973:BAS

[DE73] R. A. Dayananda and I. G. Evans. Bayesian acceptance-sampling schemes for two-sided tests of the mean of a normal distribution of known variance. Journal of the American Statistical Association, 68(341):131–136, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284156.

Deaton:1979:BRB

[Dea79] Angus Deaton. Book review: [untitled]. Journal of the American Statistical Association, 74(365):251, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286776.

DeGroot:1970:BRB

[DeG70] Morris H. DeGroot. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1398–1400, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284306.

DeGroot:1971:BRBa

[DeG71a] Morris H. DeGroot. Book review: [untitled]. Journal of the American Statistical Association, 66(334):428–429, June 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283962.

DeGroot:1971:BRBb

[DeG71b] Morris H. DeGroot. Book review: [untitled]. Journal of the American Statistical Association, 66(335):660–661, September 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283564.

DeGroot:1973:DWC

[DeG73] Morris H. DeGroot. Doing what comes naturally: Interpreting a tail area as a posterior probability or as a likelihood ratio. Journal of the American Statistical Association, 68(344):966–969, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284532.

DeGroot:1974:RC

[DeG74] Morris H. DeGroot. Reaching a consensus. Journal of the American Statistical Association, 69(345):118–121, March 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285509.

DeGroot:1975:BRB

[DeG75] Morris H. DeGroot. Book review: [untitled]. Journal of the American Statistical Association, 70(352):964–965, December 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285480.

DeGroot:1976:BRBa

[DeG76a] Morris H. DeGroot. Book review: [untitled]. Journal of the American Statistical Association, 71(353):252, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285797.

DeGroot:1976:BRBb

[DeG76b] Morris H. DeGroot. Book review: [untitled]. Journal of the American Statistical Association, 71(353):252, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285798.

DeLand:1979:BRB

[DeL79] Edward C. DeLand. Book review: [untitled]. Journal of the American Statistical Association, 74(366):509, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286375.

Deming:1973:BRB

[Dem73a] W. Edwards Deming. Book review: [untitled]. Journal of the American Statistical Association, 68(343):748–749, September 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284815.

Dempster:1973:PDT

[Dem73b] Arthur P. Dempster. Post-data two sample tests of location: Comments. Journal of the American Statistical Association, 68(341):100– 101, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284149.

Dempster:1977:CCS

[Dem77] Arthur P. Dempster. Conditional confidence statements and confidence estimators: Comment. Journal of the American Statistical Association, 72(360):814–816, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286464.

Denison:1970:BRB

[Den70] Edward F. Denison. Book review: [untitled]. Journal of the American Statistical Association, 65(329):465–468, March 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283617.

Denton:1971:AMQ

[Den71] Frank T. Denton. Adjustment of monthly or quarterly series to annual totals: An approach based on quadratic minimization. Journal of the American Statistical Association, 66(333):99–102, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284856.

Denny:1973:BRB

[Den73] J. L. Denny. Book review: [untitled]. Journal of the American Statistical Association, 68(343):755, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284828.

Desu:1970:BRB

[Des70] M. M. Desu. Book review: [untitled]. Journal of the American Statistical Association, 65(330):998–999, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284610.

Devore:1973:RNMa

[Dev73a] Jay Devore. Reconstructing a noisy Markov chain. Journal of the American Statistical Association, 68(342):394–398, June 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284082.

Devore:1973:RNMb

[Dev73b] Jay L. Devore. Reconstructing a noisy Markov chain using nearneighbor rules. Journal of the American Statistical Association, 68 (343):599–602, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284784.

Devore:1978:BRBa

[Dev78a] Jay L. Devore. Book review: [untitled]. Journal of the American Statistical Association, 73(361):220, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286556.

Devore:1978:BRBb

[Dev78b] Jay L. Devore. Book review: [untitled]. Journal of the American Statistical Association, 73(362):442, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286692.

Devore:1979:BRB

[Dev79] Jay L. Devore. Book review: [untitled]. Journal of the American Statistical Association, 74(366):506–507, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286372.

DeWald:1971:BRB

[DeW71] Charles G. DeWald. Book review: [untitled]. Journal of the American Statistical Association, 66(334):428, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283961.

Degracie:1972:ESA

[DF72] James S. Degracie and Wayne A. Fuller. Estimation of the slope and analysis of covariance when the concomitant variable is measured with error. Journal of the American Statistical Association, 67(340):930–937, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284665.

Deistler:1974:LMF

[DF74] Manfred Deistler and Gustav Feichtinger. The linear model formulation of a multitype branching process applied to population dynamics. Journal of the American Statistical Association, 69(347):662–664, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285997.

Dickey:1975:PDP

[DF75] James Dickey and Peter Freeman. Population-distributed personal probabilities. Journal of the American Statistical Association, 70(350):362–364, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285823.

Diaconis:1979:RP

[DF79a] Persi Diaconis and David Freedman. On rounding percentages. Journal of the American Statistical Association, 74(366):359–364, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284288;   http:/
/www.jstor.org/stable/2286335.

Dickey:1979:DEA

[DF79b] David A. Dickey and Wayne A. Fuller. Distribution of the estimators for autoregressive time series with a unit root. Journal of the American Statistical Association, 74(366):427–431, June 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286348.

Dudewicz:1972:BRB

[DG72] Edward J. Dudewicz and Nancy L. Geller. Book review: [untitled]. Journal of the American Statistical Association, 67(340):962–963, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284684.

Dahiya:1973:EZC

[DG73a] Ram C. Dahiya and Alan J. Gross. Estimating the zero class from a truncated Poisson sample. Journal of the American Statistical Association, 68(343):731–733, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284809.

Dahiya:1973:HMC

[DG73b] Ram C. Dahiya and John Gurland. How many classes in the Pearson chi-square test? Journal of the American Statistical Association, 68(343):707–712, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284803.

Draper:1975:TSM

[DG75] Norman Draper and Irwin Guttman. Two simultaneous measurement procedures: a Bayesian approach. Journal of the American Statistical Association, 70(349):43–46, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285373.

Draper:1976:KSM

[DG76] Norman Draper and Irwin Guttman. k simultaneous measurement procedures: a Bayesian approach. Journal of the American Statistical Association, 71(355):605–607, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285589.

Dillon:1978:PSM

[DG78] William R. Dillon and Matthew Goldstein. On the performance of some multinomial classification rules. Journal of the American Statistical Association, 73(362):305–313, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286657.

Draper:1977:ABD

[DGL77] Norman R. Draper, Irwin Guttman, and Peter Lipow. All-bias designs for spline functions joined at the axes. Journal of the American Statistical Association, 72(358):424–429, June 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286811.

Duncan:1972:LDR

[DH72] D. B. Duncan and S. D. Horn. Linear dynamic recursive estimation from the viewpoint of regresion analysis. Journal of the American Statistical Association, 67(340):815–821, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284643.

Dent:1977:MLE

[DH77] Warren T. Dent and Clifford Hildreth. Maximum likelihood estimation in random coefficient models. Journal of the American Statistical Association, 72(357):69–72, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286907.

Diaconis:1978:RBM

[Dia78] Persi Diaconis. Review of A Mathematical Theory of Evidence. Journal of the American Statistical Association, 73(363):677–678, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286624.

Dickey:1970:BRB

[Dic70] James M. Dickey. Book review: [untitled]. Journal of the American Statistical Association, 65(329):473–474, March 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283621.

Dickey:1973:BRB

[Dic73] James M. Dickey. Book review: [untitled]. Journal of the American Statistical Association, 68(341):235–237, March 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284178.

Dickey:1976:APD

[Dic76a] James M. Dickey. Approximate posterior distributions. Journal of the American Statistical Association, 71(355):680–689, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285601.

Dickey:1976:SIU

[Dic76b] James M. Dickey. Strong inconsistency from uniform priors: Comment. Journal of the American Statistical Association, 71(353):117– 119, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285745.

Dickey:1977:BRB

[Dic77a] James M. Dickey. Book review: [untitled]. Journal of the American Statistical Association, 72(357):228, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286945.

Dickey:1977:TAU

[Dic77b] James M. Dickey. Is the tail area useful as an approximate Bayes factor? Journal of the American Statistical Association, 72(357):138– 142, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286921.

Dixon:1976:IED

[Dix76] Dennis O. Dixon. Interval estimates derived from Bayes testing rules. Journal of the American Statistical Association, 71(354):406–408, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285323.

Dixon:1978:RMB

[Dix78] W. J. Dixon. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73(361):80–82, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286519.

Duncan:1972:IMC

[DL72] George T. Duncan and Lizbie G. Lin. Inference for Markov chains having stochastic entry and exit. Journal of the American Statistical Association, 67(340):761–767, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284633.

Dutta:1974:IIV

[DL74] M. Dutta and E. Lyttkens. Iterative instrumental variables method and estimation of a large simultaneous system. Journal of the American Statistical Association, 69(348):977–986, December 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286175.

DAgostino:1976:LEL

[DL76] Ralph B. D’Agostino and Austin F. S. Lee. Linear estimation of the logistic parameters for complete or tail-censored samples. Journal of the American Statistical Association, 71(354):462–464, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285334.

DAgostino:1977:RLE

[DL77] Ralph B. D’Agostino and Austin F. S. Lee. Robustness of location estimators under changes of population kurtosis. Journal of the American Statistical Association, 72(358):393–396, June 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286805.

DeRouen:1974:MEL

[DM74] T. A. DeRouen and Toby J. Mitchell. A G_{1}-minimax estimator for a linear combination of binomial probabilities. Journal of the American Statistical Association, 69(345):231–233, March 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285532.

Dreze:1976:BFI

[DM76a] Jacques H. Dreze and Juan-Antonio Morales. Bayesian full information analysis of simultaneous equations. Journal of the American Statistical Association, 71(356):919–923, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286861.

Dykstra:1976:RML

[DM76b] Richard L. Dykstra and Richard W. Madsen. Restricted maximum likelihood estimators for Poisson parameters. Journal of the American Statistical Association, 71(355):711–718, September 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285607.

Denby:1979:REF

[DM79] Lorraine Denby and R. Douglas Martin. Robust estimation of the first-order autoregressive parameter. Journal of the American Statistical Association, 74(365):140–146, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286743.

Denton:1972:MCA

[DO72] Frank T. Denton and Ernest H. Oksanen. A multi-country analysis of the effects of data revisions on an econometric model. Journal of the American Statistical Association, 67(338):286–291, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284370.

deOliveira:1974:RNB

[dO74] J. Tiago de Oliveira. Regression in the nondifferentiable bivariate extreme models. Journal of the American Statistical Association, 69(347):816–818, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286024.

Dobson:1978:ETL

[Dob78] Annette J. Dobson. Evolution times of languages. Journal of the American Statistical Association, 73(361):58–64, March 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286516.

Doksum:1971:BRB

[Dok71] Kjell Doksum. Book review: [untitled]. Journal of the American Statistical Association, 66(335):661, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283566.

Doksum:1972:BRB

[Dok72] Kjell Doksum. Book review: [untitled]. Journal of the American Statistical Association, 67(339):721, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284488.

Dolby:1976:NLS

[Dol76] Geoffrey R. Dolby. A note on the linear structural relation when both residual variances are known. Journal of the American Statistical Association, 71(354):352–353, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285312.

Doran:1974:PMO

[Dor74] H. E. Doran. Prediction of missing observations in the time series of an economic variable. Journal of the American Statistical Association, 69(346):546–554, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285695.

Douglas:1971:BRB

[Dou71] Paul H. Douglas. Book review: [untitled]. Journal of the American Statistical Association, 66(334):422–423, June 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283949.

Dhrymes:1972:API

[DP72] Phoebus J. Dhrymes and Vishwanath Pandit. Asymptotic properties of an iterate of the two-stage least squares estimator. Journal of the American Statistical Association, 67(338):444–447, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284402.

Dionne:1972:CMM

[DQ72] A. Dionne and C. P. Quesenberry. Comparisons of MVU and ML distribution function estimators for truncation parameter distributions. Journal of the American Statistical Association, 67(338):465– 467, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284407.

Darroch:1971:CDD

[DR71] J. N. Darroch and D. Ratcliff. A characterization of the Dirichlet distribution. Journal of the American Statistical Association, 66(335):641–643, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283547.

Draper:1971:BRB

[Dra71] Norman R. Draper. Book review: [untitled]. Journal of the American Statistical Association, 66(336):924, December 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284266.

Drake:1973:BRB

[Dra73] Alvin W. Drake. Book review: [untitled]. Journal of the American Statistical Association, 68(343):751–752, September 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284820.

Driscoll:1978:PMI

[Dri78] Michael F. Driscoll. On pairwise and mutual independence: Characterizations of rectangular distributions. Journal of the American Statistical Association, 73(362):432–433, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286679.

Driscoll:1979:CPM

[Dri79] Michael F. Driscoll. Corrigenda: On Pairwise and Mutual Independence: Characterizations of Rectangular Distributions. Journal of the American Statistical Association, 74(366):516, June 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286386.

Dayton:1973:ETC

[DS73] C. Mitchell Dayton and William D. Schafer. Extended tables of t and chi square for Bonferroni tests with unequal error allocation. Journal of the American Statistical Association, 68(341):78–83, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284142.

David:1974:BMS

[DS74] Isidoro P. David and B. V. Sukhatme. On the bias and mean square error of the ratio estimator. Journal of the American Statistical Association, 69(346):464–466, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285679.

Dowling:1975:RER

[DS75] T. A. Dowling and Richard H. Shachtman. On the relative efficiency of randomized response models. Journal of the American Statistical Association, 70(349):84–87, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285381.

Dent:1978:FLI

[DS78] W. T. Dent and J. A. Swanson. Forecasting with limited information: ARIMA models of the trailer on flatcar transportation market. Journal of the American Statistical Association, 73(362):293–299, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286655.

Dawkins:1979:FEW

[DS79] Shyrl M. Dawkins and Elizabeth L. Scott. Field experimentation in weather modification: Comment. Journal of the American Statistical Association, 74(365):70–77, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286724.

Dempster:1977:SSAa

[DSW77a] A. P. Dempster, Martin Schatzoff, and Nanny Wermuth. A simulation study of alternatives to ordinary least squares. Journal of the American Statistical Association, 72(357):77–91, March 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286909.

Dempster:1977:SSAb

[DSW77b] A. P. Dempster, Martin Schatzoff, and Nanny Wermuth. A simulation study of alternatives to ordinary least squares: Rejoinder. Journal of the American Statistical Association, 72(357):104–106, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286916.

Dudewicz:1971:NSP

[Dud71] Edward J. Dudewicz. A nonparametric selection Procedure’s efficiency: Largest location parameter case. Journal of the American Statistical Association, 66(333):152–161, March 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284866.

Dudewicz:1972:TSC

[Dud72] Edward J. Dudewicz. Two-sided confidence intervals for ranked means. Journal of the American Statistical Association, 67(338):462– 464, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284406.

Dudley:1978:BRB

[Dud78] R. M. Dudley. Book review: [untitled]. Journal of the American Statistical Association, 73(363):687, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286643.

DuMouchel:1973:SDS

[DuM73] William H. DuMouchel. Stable distributions in statistical inference: 1. symmetric stable distributions compared to other symmetric longtailed distributions. Journal of the American Statistical Association, 68(342):469–477, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284101.

DuMouchel:1975:SDS

[DuM75] W. H. DuMouchel. Stable distributions in statistical inference: 2. information from stably distributed samples. Journal of the American Statistical Association, 70(350):386–393, June 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285828.

Duncan:1971:EDC

[Dun71] Acheson J. Duncan. The economic design of X̄-charts when there is a multiplicity of assignable causes. Journal of the American Statistical Association, 66(333):107–121, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284859.

Duncan:1974:EBA

[Dun74a] George T. Duncan. An empirical Bayes approach to scoring multiplechoice tests in the misinformation model. Journal of the American Statistical Association, 69(345):50–57, March 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285499.

Dunn:1974:BRB

[Dun74b] D. M. Dunn. Book review: [untitled]. Journal of the American Statistical Association, 69(347):829–830, September 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286029.

Duncan:1975:BRBa

[Dun75a] George T. Duncan. Book review: [untitled]. Journal of the American Statistical Association, 70(351):731, September 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285969.

Duncan:1975:BRBb

[Dun75b] Otis Dudley Duncan. Book review: Panel Analysis: Latent Probability Models for Attitude and Behavior Processes, by Lee M. Wiggins. Journal of the American Statistical Association, 70(352):959–960, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285469.

Duncan:1976:BRB

[Dun76a] George T. Duncan. Book review: [untitled]. Journal of the American Statistical Association, 71(353):247–248, March 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285786.

Dunsmore:1976:NFM

[Dun76b] Ian R. Dunsmore. A note on Faulkenberry’s method of obtaining prediction intervals. Journal of the American Statistical Association, 71(353):193–194, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285767.

Duncan:1977:BRB

[Dun77] George T. Duncan. Book review: [untitled]. Journal of the American Statistical Association, 72(357):229, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286946.

Duncan:1979:BRBb

[Dun79a] George T. Duncan. Book review: [untitled]. Journal of the American Statistical Association, 74(366):504, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286367.

Duncan:1979:BRBc

[Dun79b] George T. Duncan. Book review: [untitled]. Journal of the American Statistical Association, 74(366):505, June 1979. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286369.

Duncan:1979:BRBd

[Dun79c] George T. Duncan. Book review: [untitled]. Journal of the American Statistical Association, 74(368):930, December 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286434.

Duncan:1979:BRBa

[Dun79d] Otis Dudley Duncan. Book review: [untitled]. Journal of the American Statistical Association, 74(365):253, March 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286778.

Durbin:1970:BTR

[Dur70] J. Durbin. On Birnbaum’s theorem on the relation between sufficiency, conditionality and likelihood. Journal of the American Statistical Association, 65(329):395–398, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283601.

Dunn:1976:AVS

[DWD76] D. M. Dunn, W. H. Williams, and T. L. DeChaine. Aggregate versus subaggregate models in local area forecasting. Journal of the American Statistical Association, 71(353):68–71, March 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285732.

Dyn:1979:SPI

[DWW79] Nira Dyn, Grace Wahba, and Wing-Hung Wong. Smooth pycnophylactic interpolation for geographical regions: Comment. Journal of the American Statistical Association, 74(367):530–535, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286969.

Dwyer:1973:BRB

[Dwy73] Paul S. Dwyer. Book review: [untitled]. Journal of the American Statistical Association, 68(341):239–240, March 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284181.

Denny:1978:ARC

[DY78] J. L. Denny and S. J. Yakowitz. Admissible run-contingency type tests for independence and Markov dependence. Journal of the American Statistical Association, 73(361):177–181, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286541.

Dyer:1973:DPS

[Dye73] Alan R. Dyer. Discrimination procedures for separate families of hypotheses. Journal of the American Statistical Association, 68 (344):970–974, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284533.

Dyer:1974:HTP

[Dye74] Alan R. Dyer. Hypothesis testing procedures for separate families of hypotheses. Journal of the American Statistical Association, 69(345):140–145, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285512.

E:1974:BRB

[E.74] W. A. E. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1053, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286203.

Easterling:1972:ACL

[Eas72] Robert G. Easterling. Approximate confidence limits for system reliability. Journal of the American Statistical Association, 67(337):220– 222, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284730.

Easterling:1977:BRB

[Eas77] Robert G. Easterling. Book review: [untitled]. Journal of the American Statistical Association, 72(358):482–483, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286826.

Eaton:1973:BRB

[Eat73] Morris L. Eaton. Book review: [untitled]. Journal of the American Statistical Association, 68(343):749, September 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284816.

Ebbeler:1973:MPCb

[Ebb73a] Donald H. Ebbeler. Measuring the permanent component of a series for serially correlated observations. Journal of the American Statistical Association, 68(343):579–580, September 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284779.

Ebbeler:1973:NLS

[Ebb73b] Donald H. Ebbeler. A note on large-sample approximations in lognormal linear models. Journal of the American Statistical Association, 68(341):231, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284176.

Ebbeler:1973:MPCa

[Ebb73c] Donald H. Ebbeler. On the measurement of the permanent component of a series. Journal of the American Statistical Association, 68(342):343–347, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284073.

Eckler:1970:SSW

[Eck70] A. Ross Eckler. Statisticians and shoemakers (“who is worse shod than the Shoemaker’s wife,” from Heywood’s proverbs.). Journal of the American Statistical Association, 65(329):9–21, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283571.

Eckardt:1976:SMS

[Eck76] Walter L. Eckardt, Jr. A sequential method for selecting the best of three binomial populations. Journal of the American Statistical Association, 71(354):473–474, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285336.

Eddy:1979:NSD

[Edd79] William F. Eddy. Nonparametric statistical data modeling: Comment. Journal of the American Statistical Association, 74(365):124– 126, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286737.

Edwards:1975:CPA

[Edw75] Ward Edwards. Cognitive processes and the assessment of subjective probability distributions: Comment. Journal of the American Statistical Association, 70(350):291–293, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285810.

Eaton:1970:HTU

[EE70] M. L. Eaton and Bradley Efron. Hotelling’s T^{2} test under symmetry conditions. Journal of the American Statistical Association, 65(330):702–711, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284581.

Ertel:1976:SAL

[EF76] James E. Ertel and Edward B. Fowlkes. Some algorithms for linear spline and piecewise multiple linear regression. Journal of the American Statistical Association, 71(355):640–648, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285594.

Efron:1971:DOSa

[Efr71a] Bradley Efron. Does an observed sequence of numbers follow a simple rule? (another look at Bode’s law). Journal of the American Statistical Association, 66(335):552–559, September 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283522.

Efron:1971:DOSb

[Efr71b] Bradley Efron. Does an observed sequence of numbers follow a simple rule? (another look at Bode’s law): Rejoinder. Journal of the American Statistical Association, 66(335):568, September 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283529.

Efron:1975:ELR

[Efr75] Bradley Efron. The efficiency of logistic regression compared to normal discriminant analysis. Journal of the American Statistical Association, 70(352):892–898, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285453.

Efron:1976:NAA

[Efr76] B. Efron. Necessary analysis and adaptive inference: Comment. Journal of the American Statistical Association, 71(353):111–112, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285742.

Efron:1977:ECL

[Efr77] Bradley Efron. The efficiency of Cox’s likelihood function for censored data. Journal of the American Statistical Association, 72 (359):557–565, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286217.

Efron:1978:RAZ

[Efr78] Bradley Efron. Regression and ANOVA with zero-one data: Measures of residual variation. Journal of the American Statistical Association, 73(361):113–121, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286531.

Enis:1971:EPL

[EG71] Peter Enis and Seymour Geisser. Estimation of the probability that Y < X. Journal of the American Statistical Association, 66(333):162–168, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284867.

Einot:1975:SPS

[EG75] Israel Einot and K. R. Gabriel. A study of the powers of several methods of multiple comparisons. Journal of the American Statistical Association, 70(351):574–583, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285935.

El-Helbawy:1978:TCP

[EHB78] Abdalla T. El-Helbawy and Ralph A. Bradley. Treatment contrasts in paired comparisons: Large-sample results, applications, and some optimal designs. Journal of the American Statistical Association, 73(364):831–839, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286289.

Eisenpress:1973:BRB

[Eis73] Harry Eisenpress. Book review: [untitled]. Journal of the American Statistical Association, 68(342):494, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284114.

Elandt-Johnson:1975:BRB

[EJ75] Regina C. Elandt-Johnson. Book review: [untitled]. Journal of the American Statistical Association, 70(350):489, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285853.

Ekblom:1973:NNM

[Ekb73] H. Ekblom. A note on nonlinear median estimators. Journal of the American Statistical Association, 68(342):431–432, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284092.

Elashoff:1972:MQO

[Ela72] Janet Dixon Elashoff. A model for quadratic outliers in linear regression. Journal of the American Statistical Association, 67(338):478– 485, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284410.

Elashoff:1975:BRB

[Ela75] Janet Dixon Elashoff. Book review: [untitled]. Journal of the American Statistical Association, 70(352):959, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285468.

Elashoff:1977:BRB

[Ela77] Janet D. Elashoff. Book review: [untitled]. Journal of the American Statistical Association, 72(357):235–236, March 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286956.

Ellenberg:1973:JDS

[Ell73] Jonas H. Ellenberg. The joint distribution of the standardized least squares residuals from a general linear regression. Journal of the American Statistical Association, 68(344):941–943, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284526.

Ellenberg:1977:JDS

[Ell77] Jonas H. Ellenberg. The joint distribution of the standardized row sums of squares from a balanced two-way layout. Journal of the American Statistical Association, 72(358):407–411, June 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286808.

Elliott:1979:FEW

[Ell79] R. D. Elliott. Field experimentation in weather modification: Comment. Journal of the American Statistical Association, 74(365):77, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286725.

Elston:1977:BRB

[Els77] Robert C. Elston. Book review: [untitled]. Journal of the American Statistical Association, 72(358):486, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286832.

Efron:1971:LRB

[EM71] Bradley Efron and Carl Morris. Limiting the risk of Bayes and empirical Bayes estimators — Part I: The Bayes case. Journal of the American Statistical Association, 66(336):807–815, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284231.

Efron:1972:LRB

[EM72] Bradley Efron and Carl Morris. Limiting the risk of Bayes and empirical Bayes estimators — Part II: The empirical Bayes case. Journal of the American Statistical Association, 67(337):130–139, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284711.

Efron:1973:SER

[EM73] Bradley Efron and Carl Morris. Stein’s estimation rule and its competitors — an empirical Bayes approach. Journal of the American Statistical Association, 68(341):117–130, March 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284155.

Efron:1975:DAU

[EM75] Bradley Efron and Carl Morris. Data analysis using Stein’s estimator and its generalizations. Journal of the American Statistical Association, 70(350):311–319, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285814.

Efron:1977:SSA

[EM77] Bradley Efron and Carl Morris. A simulation study of alternatives to ordinary least squares: Comment. Journal of the American Statistical Association, 72(357):91–93, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286910.

Evans:1978:OED

[EM78] James W. Evans and A. R. Manson. Optimal experimental designs in two dimensions using minimum bias estimation. Journal of the American Statistical Association, 73(361):171–176, March 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286540.

Engle:1972:BRB

[Eng72] Robert F. Engle. Book review: [untitled]. Journal of the American Statistical Association, 67(340):959, December 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284678.

Engelhardt:1975:SLE

[Eng75] Max Engelhardt. Simple linear estimation of the parameters of the logistic distribution from a complete or censored sample. Journal of the American Statistical Association, 70(352):899–902, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285454.

Enis:1971:BRB

[Eni71] Peter Enis. Book review: [untitled]. Journal of the American Statistical Association, 66(335):654–655, September 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283553.

Enns:1975:SMS

[Enn75] E. G. Enns. Selecting the maximum of a sequence with imperfect information. Journal of the American Statistical Association, 70(351):640–643, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285947.

Esary:1970:RBS

[EP70] J. D. Esary and F. Proschan. A reliability bound for systems of maintained, interdependent components. Journal of the American Statistical Association, 65(329):329–338, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283596.

Epps:1976:SDS

[Epp76] Thomas W. Epps. The stochastic dependence of security price changes and transaction volumes in a model with temporally dependent price changes. Journal of the American Statistical Association, 71(356):830–834, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286846.

Epps:1979:CSP

[Epp79] Thomas W. Epps. Comovements in stock prices in the very short run. Journal of the American Statistical Association, 74(366):291– 298, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286325.

Ericson:1970:PMV

[Eri70] W. A. Ericson. On the posterior mean and variance of a population mean. Journal of the American Statistical Association, 65(330):649– 652, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284577.

Ericksen:1974:RME

[Eri74] Eugene P. Ericksen. A regression method for estimating population changes of local areas. Journal of the American Statistical Association, 69(348):867–875, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286155.

Ericksen:1976:SRP

[Eri76] Eugene P. Ericksen. Sampling a rare population: a case study. Journal of the American Statistical Association, 71(356):816–822, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286844.

Evans:1974:NEL

[ES74] I. G. Evans and S. A. Shaban. A note on estimation in lognormal models. Journal of the American Statistical Association, 69 (347):779–781, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286017.

Espasa:1977:NAR

[Esp77] Antoni Espasa. A note on the acceptability of regression solutions: Another look at computational accuracy. Journal of the American Statistical Association, 72(359):603, September 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286226.

Estes:1972:RTL

[Est72] William K. Estes. Research and theory on the learning of probabilities. Journal of the American Statistical Association, 67(337):81–102, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284706.

Eckland:1974:BRB

[ET74] Bruce K. Eckland and Howard F. Taylor. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1041–1044, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286186.

Everitt:1979:MCI

[Eve79] Brian S. Everitt. A Monte Carlo investigation of the robustness of Hotelling’s one- and two-sample T^{2} tests. Journal of the American Statistical Association, 74(365):48–51, March 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286719.

Eichhorn:1973:SSO

[EZ73] B. H. Eichhorn and S. Zacks. Sequential search of an optimal dosage, I. Journal of the American Statistical Association, 68(343):594–598, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284783.

F:1974:BRB

[F.74] S. E. F. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1053, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286201.

F:1975:BRB

[F.75] S. E. F. Book review: [untitled]. Journal of the American Statistical Association, 70(350):491–492, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285859.

Fabian:1970:BRB

[Fab70] Vaclav Fabian. Book review: [untitled]. Journal of the American Statistical Association, 65(329):459–461, March 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283614.

Fabian:1977:EJS

[Fab77] Vaclav Fabian. On the effect of jury size. Journal of the American Statistical Association, 72(359):535–536, September 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286212.

Fair:1977:BRB

[Fai77] Ray C. Fair. Book review: [untitled]. Journal of the American Statistical Association, 72(358):485, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286830.

Farley:1970:BRB

[Far70] John U. Farley. Book review: [untitled]. Journal of the American Statistical Association, 65(330):1002–1003, June 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284613.

Farr:1972:BRB

[Far72] S. David Farr. Book review: [untitled]. Journal of the American Statistical Association, 67(339):715–716, September 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284480.

Farley:1973:BRB

[Far73] Reynolds Farley. Book review: [untitled]. Journal of the American Statistical Association, 68(342):493, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284111.

Farebrother:1976:MPV

[Far76] R. W. Farebrother. The minimum proportional-variance unbiased linear estimator of β and simultaneous confidence intervals. Journal of the American Statistical Association, 71(355):761–762, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285617.

Fase:1971:ELI

[Fas71] Martin M. G. Fase. On the estimation of lifetime income. Journal of the American Statistical Association, 66(336):686–692, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284213.

Faulkenberry:1973:MOP

[Fau73] G. David Faulkenberry. A method of obtaining prediction intervals. Journal of the American Statistical Association, 68(342):433–435, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284093.

Fuller:1973:TEL

[FB73] Wayne A. Fuller and George E. Battese. Transformations for estimation of linear models with nested-error structure. Journal of the American Statistical Association, 68(343):626–632, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284790.

Feder:1970:BRB

[Fed70] Paul I. Feder. Book review: [untitled]. Journal of the American Statistical Association, 65(330):993–995, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284606.

Feinleib:1975:BRB

[Fei75] Manning Feinleib. Book review: [untitled]. Journal of the American Statistical Association, 70(349):257, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285423.

Forsythe:1973:SRV

[FEJM73] Alan B. Forsythe, Laszlo Engleman, Robert Jennrich, and Philip R. A. May. A stopping rule for variable selection in multiple regression. Journal of the American Statistical Association, 68(341):75–77, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284141.

Forsythe:1975:CSR

[FEJM75] Alan B. Forsythe, Laszlo Engelman, Robert Jennrich, and Philip R. A. May. Corrigenda: a Stopping Rule for Variable Selection in Multiple Regression. Journal of the American Statistical Association, 70(350):494, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285865.

Fellagi:1972:QSC

[Fel72a] I. Fellagi. On the question of statistical confidentiality. Journal of the American Statistical Association, 61(337):??, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fellegi:1972:QSC

[Fel72b] I. P. Fellegi. On the question of statistical confidentiality. Journal of the American Statistical Association, 67(337):7–18, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284695.

Feldstein:1974:EVC

[Fel74a] Martin Feldstein. Errors in variables: a consistent estimator with smaller MSE in finite samples. Journal of the American Statistical Association, 69(348):990–996, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286177.

Fellegi:1974:CIM

[Fel74b] Ivan P. Fellegi. Corrigenda: An Improved Method of Etimating the Correlated Response Variance. Journal of the American Statistical Association, 69(347):840, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286047.

Fellegi:1974:IME

[Fel74c] Ivan P. Fellegi. An improved method of estimating the correlated response variance. Journal of the American Statistical Association, 69(346):496–501, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285686.

Feldstein:1976:SSD

[Fel76] Martin Feldstein. Social security and the distribution of wealth. Journal of the American Statistical Association, 71(356):800–807, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286842.

Fellegi:1979:BRB

[Fel79] Ivan P. Fellegi. Book review: [untitled]. Journal of the American Statistical Association, 74(368):922, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286425.

Fenech:1979:TMM

[Fen79] Alan Paul Fenech. Tukey’s method of multiple comparison in the randomized blocks model. Journal of the American Statistical Association, 74(368):881–884, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286417.

Ferreira:1975:BAS

[Fer75] Pedro E. Ferreira. A Bayesian analysis of a switching regression model: Known number of regimes. Journal of the American Statistical Association, 70(350):370–374, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285825.

Ferguson:1978:MLE

[Fer78] Thomas S. Ferguson. Maximum likelihood estimates of the parameters of the Cauchy distribution for samples of size 3 and 4. Journal of the American Statistical Association, 73(361):211–213, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286549.

Frane:1978:BRBa

[FF78] Margaret V. Frane and James W. Frane. Book review: [untitled]. Journal of the American Statistical Association, 73(361):217–218, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286552.

Fienberg:1970:GTT

[FG70] Stephen E. Fienberg and John P. Gilbert. The geometry of a two by two contingency table. Journal of the American Statistical Association, 65(330):694–701, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284580.

Folsom:1973:TAQ

[FGHA73] Ralph E. Folsom, Bernard G. Greenberg, Daniel G. Horvitz, and James R. Abernathy. The two alternate questions randomized response model for human surveys. Journal of the American Statistical Association, 68(343):525–530, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284771.

Farley:1970:TSS

[FH70] John U. Farley and Melvin J. Hinich. A test for a shifting slope coefficient in a linear model. Journal of the American Statistical Association, 65(331):1320–1329, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284298.

Fienberg:1973:SEM

[FH73] Stephen E. Fienberg and Paul W. Holland. Simultaneous estimation of multinomial cell probabilities. Journal of the American Statistical Association, 68(343):683–691, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284799.

Fellegi:1976:SAA

[FH76] I. P. Fellegi and D. Holt. A systematic approach to automatic edit and imputation. Journal of the American Statistical Association, 71(353): 17–35, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285726.

Fuller:1978:REA

[FH78] Wayne A. Fuller and Michael A. Hidiroglou. Regression estimation after correcting for attenuation. Journal of the American Statistical Association, 73(361):99–104, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2286529. See correction [FH81].

Fay:1979:EIS

[FH79] Robert E. Fay, III and Roger A. Herriot. Estimates of income for small places: An application of James–Stein procedures to census data. Journal of the American Statistical Association, 74(366):269– 277, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286322.

Fuller:1981:CRE

[FH81] Wayne A. Fuller and Michael A. Hidiroglou. Correction: “Regression estimation after correcting for attenuation” [J. Amer. Statist. Assoc. 73 (1978), no. 361, 99–104, MR 80k:62087]. Journal of the American Statistical Association, 76(376):1022, December 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287626. See [FH78, JTB81].

Fomby:1978:OPP

[FHJ78] Thomas B. Fomby, R. Carter Hill, and Stan R. Johnson. An optimal property of principal components in the context of restricted least squares. Journal of the American Statistical Association, 73(361):191–193, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286544.

Fienberg:1970:QIM

[Fie70] Stephen E. Fienberg. Quasi-independence and maximum likelihood estimation in incomplete contingency tables. Journal of the American Statistical Association, 65(332):1610–1616, December 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284344.

Fienberg:1972:STS

[Fie72] Stephen E. Fienberg. The Sun–Times straw poll, 1968 and 1970: A statistical appraisal. Journal of the American Statistical Association, 67(338):292–297, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284371.

Fienberg:1975:BRB

[Fie75a] Stephen E. Fienberg. Book review: [untitled]. Journal of the American Statistical Association, 70(351):734, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285975.

Fienberg:1975:DAO

[Fie75b] Stephen E. Fienberg. The design and analysis of the observational study — a review: Comment. Journal of the American Statistical Association, 70(351):521–523, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285926.

Fienberg:1976:BRBa

[Fie76a] Stephen E. Fienberg. Book review: [untitled]. Journal of the American Statistical Association, 71(354):532, June 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285363.

Fienberg:1976:BRBb

[Fie76b] Stephen E. Fienberg. Book review: [untitled]. Journal of the American Statistical Association, 71(356):1005–1006, December 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286885.

Fienberg:1976:BRBc

[Fie76c] Stephen E. Fienberg. Book review: [untitled]. Journal of the American Statistical Association, 71(356):1008, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286892.

Fienberg:1977:BRBa

[Fie77a] Stephen E. Fienberg. Book review: [untitled]. Journal of the American Statistical Association, 72(357):239, March 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286963.

Fienberg:1977:BRBb

[Fie77b] Stephen E. Fienberg. Book review: [untitled]. Journal of the American Statistical Association, 72(358):485–486, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286831.

Fienberg:1977:BRBc

[Fie77c] Stephen E. Fienberg. Book review: [untitled]. Journal of the American Statistical Association, 72(359):691, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286250.

Finch:1977:RUT

[Fin77a] Stephen J. Finch. Robust univariate test of symmetry. Journal of the American Statistical Association, 72(358):387–392, June 1977.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286804.

Finney:1977:PBC

[Fin77b] D. J. Finney. Probabilities based on circumstantial evidence. Journal of the American Statistical Association, 72(358):316–318, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286794.

Fishlow:1970:BRB

[Fis70] Albert Fishlow. Book review: [untitled]. Journal of the American Statistical Association, 65(329):463–465, March 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283616.

Fishman:1972:EMO

[Fis72] George S. Fishman. Estimating the mean of observations from a wide-sense stationary autoregressive sequence. Journal of the American Statistical Association, 67(338):402–406, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284392.

Fishburn:1977:BRB

[Fis77] Peter C. Fishburn. Book review: [untitled]. Journal of the American Statistical Association, 72(359):683–684, September 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286244.

Fishman:1979:SBD

[Fis79] George S. Fishman. Sampling from the binomial distribution on a computer. Journal of the American Statistical Association, 74 (366):418–423, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286346.

Fox:1970:VST

[FK70] Martin Fox and George S. Kimeldorf. Values and shooting times in noisy duels. Journal of the American Statistical Association, 65(329):422–430, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283606.

Fligner:1976:DFT

[FK76] Michael A. Fligner and Timothy J. Killeen. Distribution-free twosample tests for scale. Journal of the American Statistical Association, 71(353):210–213, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285771.

Frome:1973:RAP

[FKB73] Edward L. Frome, Michael H. Kutner, and John J. Beauchamp. Regression analysis of Poisson-distributed data. Journal of the American Statistical Association, 68(344):935–940, December 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284525.

Folkerts-Landau:1976:BRB

[FL76] David Folkerts-Landau. Book review: [untitled]. Journal of the American Statistical Association, 71(355):776, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285637.

Fleiss:1971:DLC

[Fle71] Joseph L. Fleiss. On the distribution of a linear combination of independent chi squares. Journal of the American Statistical Association, 66(333):142–144, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284863.

Fletcher:1975:IRL

[Fle75] R. H. Fletcher. On the iterative refinement of least squares solutions. Journal of the American Statistical Association, 70(349):109– 112, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285385.

Fligner:1979:CTS

[Fli79] Michael A. Fligner. A class of two-sample distribution-free tests for scale. Journal of the American Statistical Association, 74 (368):889–893, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286419.

Flueck:1979:FEW

[Flu79] John A. Flueck. Field experimentation in weather modification: Comment. Journal of the American Statistical Association, 74(365):

77–80, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286726.

Fertig:1974:DTA

[FM74] Kenneth W. Fertig and Nancy R. Mann. A decision-theoretic approach to defining variables sampling plans for finite lots: Single sampling for exponential and Gaussian processes. Journal of the American Statistical Association, 69(347):665–671, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285998.

Fertig:1977:PIA

[FM77] Kenneth W. Fertig and Nancy R. Mann. A prediction interval approach to obtaining variables sampling plans for small lots: Single sampling from Gaussian processes. Journal of the American Statistical Association, 72(359):585–592, September 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286221.

Fienberg:1977:ETO

[FNL77] Stephen E. Fienberg, John Neter, and R. A. Leitch. Estimating the total overstatement error in accounting populations. Journal of the American Statistical Association, 72(358):295–302, June 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286791.

Foutz:1977:UCS

[Fou77] Robert V. Foutz. On the unique consistent solution to the likelihood equations. Journal of the American Statistical Association, 72(357):147–148, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286925.

Fowlkes:1978:EMN

[Fow78] Edward B. Fowlkes. Estimating mixtures of normal distributions and switching regressions: Comment. Journal of the American Statistical Association, 73(364):747–748, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286271.

Fowlkes:1979:SMS

[Fow79] Edward B. Fowlkes. Some methods for studying the mixture of two normal (lognormal) distributions. Journal of the American Statistical Association, 74(367):561–575, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286973.

Fox:1974:FHB

[Fox74] B. L. Fox. Finite horizon behavior of policies for two-arm bandits. Journal of the American Statistical Association, 69(348):963–965, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286172.

Fox:1976:BRB

[Fox76] Karl A. Fox. Book review: [untitled]. Journal of the American Statistical Association, 71(355):769–770, September 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285623.

Fama:1971:PES

[FR71] Eugene F. Fama and Richard Roll. Parameter estimates for symmetric stable distributions. Journal of the American Statistical Association, 66(334):331–338, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283932.

Francis:1970:BRB

[Fra70] Ivor Francis. Book review: [untitled]. Journal of the American Statistical Association, 65(332):1677–1680, December 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284350.

Frankel:1971:BRB

[Fra71] Lester R. Frankel. Book review: [untitled]. Journal of the American Statistical Association, 66(335):656–657, September 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283556.

Francis:1973:CSA

[Fra73a] Ivor Francis. A comparison of several analysis of variance programs. Journal of the American Statistical Association, 68(344):860–865, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284513.

Fraser:1973:PDT

[Fra73b] D. A. Fraser. Post-data two sample tests of location: Comments. Journal of the American Statistical Association, 68(341):101–104, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284150.

Frankel:1976:SPS

[Fra76a] Lester R. Frankel. Statistics and people — the Statistician’s responsibilities. Journal of the American Statistical Association, 71(353): 9–16, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285725.

Fraser:1976:NAAa

[Fra76b] D. A. S. Fraser. Necessary analysis and adaptive inference. Journal of the American Statistical Association, 71(353):99–110, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285740.

Fraser:1976:NAAb

[Fra76c] D. A. S. Fraser. Necessary analysis and adaptive inference: Rejoinder. Journal of the American Statistical Association, 71(353):112– 113, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285743.

Fraser:1976:SIU

[Fra76d] D. A. S. Fraser. Strong inconsistency from uniform priors: Comment. Journal of the American Statistical Association, 71(353):122– 123, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285750.

Frane:1978:BRBb

[Fra78] James W. Frane. Book review: [untitled]. Journal of the American Statistical Association, 73(364):887–888, December 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286302.

Frank:1979:BRB

[Fra79] Jeff Frank. Book review: [untitled]. Journal of the American Statistical Association, 74(367):737–738, September 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2287023.

Freedman:1977:RDB

[Fre77] David Freedman. A remark on the difference between sampling with and without replacement. Journal of the American Statistical Association, 72(359):681, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286241.

Freedman:1978:BRB

[Fre78] David Freedman. Book review: [untitled]. Journal of the American Statistical Association, 73(363):687–688, September 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286644.

Friedman:1970:BRB

[Fri70] Herman P. Friedman. Book review: [untitled]. Journal of the American Statistical Association, 65(330):1001–1002, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284612.

Friedman:1978:BRB

[Fri78] James W. Friedman. Book review: [untitled]. Journal of the American Statistical Association, 73(364):896, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286314.

Froehlich:1973:SER

[Fro73] B. R. Froehlich. Some estimators for a random coefficient regression mode. Journal of the American Statistical Association, 68(342):329– 335, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284071.

Frost:1975:SPA

[Fro75] Peter A. Frost. Some properties of the Almon lag technique when one searches for degree of polynomial and lag. Journal of the American Statistical Association, 70(351):606–612, September 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285941.

Fryer:1970:NTD

[Fry70] J. G. Fryer. On the nonparametric tests of David and Fix for the bivariate two-sample location problem. Journal of the American Statistical Association, 65(331):1297–1307, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284296.

Fryer:1971:BRB

[Fry71] H. C. Fryer. Book review: [untitled]. Journal of the American Statistical Association, 66(333):225, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284882.

Fryer:1972:BRBa

[Fry72a] H. C. Fryer. Book review: [untitled]. Journal of the American Statistical Association, 67(338):493–494, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284419.

Fryer:1972:BRBb

[Fry72b] H. C. Fryer. Book review: [untitled]. Journal of the American Statistical Association, 67(338):494–495, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284420.

Fielitz:1972:ASD

[FS72] B. D. Fielitz and E. W. Smith. Asymmetric stable distributions of stock price changes. Journal of the American Statistical Association, 67(340):813–814, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284642.

Forman:1977:ESR

[FS77] Ernest H. Forman and Nozer D. Singpurwalla. An empirical stopping rule for debugging and testing computer software. Journal of the American Statistical Association, 72(360):750–757, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286455.

Friend:1970:RPA

[FT70] Irwin Friend and William Thomas. A reevaluation of the predictive ability of plant and equipment anticipations. Journal of the American Statistical Association, 65(330):510–519, June 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284563.

Fuchs:1979:CCT

[Fuc79] Camil Fuchs. Comments on a criterion of transformation proposed by schlesselman. Journal of the American Statistical Association, 74(365):238–239, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286757.

Fisher:1973:ADA

[FV73] Lloyd Fisher and John W. Van Ness. Admissible discriminant analysis. Journal of the American Statistical Association, 68(343):603–607, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284785.

Fligner:1979:NPI

[FW79] Michael A. Fligner and Douglas A. Wolfe. Nonparametric prediction intervals for a future sample median. Journal of the American Statistical Association, 74(366):453–456, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286354.

G:1971:BRB

[G.71] M. S. G. Book review: [untitled]. Journal of the American Statistical Association, 66(336):924, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284269.

Gabriel:1978:PDB

[Gab78a] K. Ruben Gabriel. Power differences between pairwise multiple comparisons: Comment. Journal of the American Statistical Association, 73(363):485–487, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286585.

Gabriel:1978:SMM

[Gab78b] K. Ruben Gabriel. A simple method of multiple comparisons of means. Journal of the American Statistical Association, 73 (364):724–729, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286265.

Gabriel:1979:FEW

[Gab79] K. Ruben Gabriel. Field experimentation in weather modification: Comment. Journal of the American Statistical Association, 74(365):

81–84, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286727.

Galambos:1973:BRB

[Gal73] J. Galambos. Book review: [untitled]. Journal of the American Statistical Association, 68(344):1029, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284553.

Galambos:1975:OSS

[Gal75a] Janos Galambos. Order statistics of samples from multivariate distributions. Journal of the American Statistical Association, 70 (351):674–680, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285954.

Gallant:1975:PLR

[Gal75b] A. R. Gallant. The power of the likelihood ratio test of location in nonlinear regression models. Journal of the American Statistical Association, 70(349):198–203, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285403.

Gallant:1975:TSP

[Gal75c] A. Ronald Gallant. Testing a subset of the parameters of a nonlinear regression model. Journal of the American Statistical Association, 70(352):927–932, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285459.

Gallant:1977:TNR

[Gal77] A. Ronald Gallant. Testing a nonlinear regression specification: a nonregular case. Journal of the American Statistical Association, 72(359):523–530, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286210.

Games:1977:ITS

[Gam77] Paul A. Games. An improved t table for simultaneous control on g contrasts. Journal of the American Statistical Association, 72 (359):531–534, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286211.

Gani:1971:BRB

[Gan71] J. Gani. Book review: [untitled]. Journal of the American Statistical Association, 66(336):919–920, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284258.

Ganeshalingam:1979:BRB

[Gan79] S. Ganeshalingam. Book review: [untitled]. Journal of the American Statistical Association, 74(368):928–929, December 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286431.

Gartside:1972:SMC

[Gar72] Peter S. Gartside. A study of methods for comparing several variances. Journal of the American Statistical Association, 67(338):342– 346, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284380.

Gartside:1973:CSM

[Gar73] Peter S. Gartside. Corrigenda: a Study of Methods for Comparing Several Variances. Journal of the American Statistical Association, 68(341):251, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284200.

Garbade:1977:TME

[Gar77] Kenneth Garbade. Two methods for examining the stability of regression coefficients. Journal of the American Statistical Association, 72(357):54–63, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286905.

Garber:1978:BRB

[Gar78] M. J. Garber. Book review: [untitled]. Journal of the American Statistical Association, 73(363):686, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286640.

Gastwirth:1971:STS

[Gas71] Joseph L. Gastwirth. On the sign test for symmetry. Journal of the American Statistical Association, 66(336):821–823, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284233.

Gastwirth:1974:ARP

[Gas74] Joseph L. Gastwirth. Adaptive robust procedures: a partial review and some suggestions for future applications and theory: Comment. Journal of the American Statistical Association, 69(348):925–926, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286162.

Gaver:1970:BRB

[Gav70] Donald P. Gaver, Jr. Book review: [untitled]. Journal of the American Statistical Association, 65(329):461–463, March 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283615.

Gaver:1973:BRBa

[Gav73a] Donald P. Gaver. Book review: [untitled]. Journal of the American Statistical Association, 68(342):491–492, June 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284108.

Gaver:1973:BRBb

[Gav73b] Kenneth M. Gaver. Book review: [untitled]. Journal of the American Statistical Association, 68(342):497, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284117.

Gaver:1978:PNA

[Gav78] Donald P. Gaver. Pseudorandom number assignment in statistically designed simulation and distribution sampling experiments: Comment. Journal of the American Statistical Association, 73(363):522, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286593;   http://www.jstor.org/stable/
pdfplus/2286593.pdf. See [SM78a].

Gastwirth:1970:SSB

[GC70] J. L. Gastwirth and M. L. Cohen. Small sample behavior of some robust linear estimators of location. Journal of the American Statistical Association, 65(330):946–973, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284600.

Good:1974:BNB

[GC74] Irving John Good and James Flinn Crook. The Bayes/non-Bayes compromise and the multinomial distribution. Journal of the American Statistical Association, 69(347):711–720, September 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286006.

Good:1978:CED

[GC78] I. J. Good and J. F. Crook. Corrigenda: Exact Distributions for χ^{2} and for the Likelihood-Ratio Statistic for the Equiprobable Multinomial Distribution. Journal of the American Statistical Association, 73(364):900, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286319.

Greenberg:1973:SDO

[GCZH73] Michael R. Greenberg, George W. Carey, Leonard Zobler, and Robert M. Hordon. A statistical dissolved oxygen model for a freeflowing river system. Journal of the American Statistical Association, 68(342):279–283, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284063.

Gilbert:1975:BRB

[GE75] John P. Gilbert and Jim Eastman. Book review: [untitled]. Journal of the American Statistical Association, 70(349):249–250, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285413.

Geisser:1979:PAM

[GE79] Seymour Geisser and William F. Eddy. A predictive approach to model selection. Journal of the American Statistical Association, 74(365):153–160, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286745.

Gebhardt:1971:IBI

[Geb71] Friedrich Gebhardt. Incomplete beta-integral B(x; 2/3, 2/3) and [p(1 − p)]^{−1/6} for use with Borges’ approximation of the binomial distribution. Journal of the American Statistical Association, 66(333):189–191, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284872.

Geertsema:1972:NSP

[Gee72] J. C. Geertsema. Nonparametric sequential procedures for selecting the best of K populations. Journal of the American Statistical Association, 67(339):614–616, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284450.

Geisser:1975:PSR

[Gei75] Seymour Geisser. The predictive sample reuse method with applications. Journal of the American Statistical Association, 70(350):320– 328, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285815.

George:1974:RPC

[Geo74] Stephen L. George. Replicated paired comparisons between two objects: The (k, r)-series. Journal of the American Statistical Association, 69(347):750–754, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286013.

Geraci:1972:BRB

[Ger72] Vincent J. Geraci. Book review: [untitled]. Journal of the American Statistical Association, 67(340):961–962, December 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284683.

Gerig:1975:MEF

[Ger75] Thomas M. Gerig. A multivariate extension of Friedman’s χ_{r}^{2}-test with random covariates. Journal of the American Statistical Association, 70(350):443–447, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285837.

Gerber:1977:UFC

[Ger77a] Hans U. Gerber. Uncertainty functions with a constant rate of reduction and comparison of experiments. Journal of the American Statistical Association, 72(360):899–900, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286483.

Gersovitz:1977:BRB

[Ger77b] Mark Gersovitz. Book review: [untitled]. Journal of the American Statistical Association, 72(358):483–484, June 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286828.

Gallant:1973:FSP

[GF73] A. R. Gallant and Wayne A. Fuller. Fitting segmented polynomial regression models whose join points have to be estimated. Journal of the American Statistical Association, 68(341):144–147, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284158.

Gideon:1971:IMA

[GG71] Rudy A. Gideon and John Gurland. An interpolatory method of approximating distribution functions with applications to the gamma and related variates. Journal of the American Statistical Association, 66(335):577–582, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283532.

Gessaman:1972:CSM

[GG72] M. P. Gessaman and P. H. Gessaman. A comparison of some multivariate discrimination procedures. Journal of the American Statistical Association, 67(338):468–472, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284408.

Goodstadt:1975:RRT

[GG75] Michael S. Goodstadt and Valerie Gruson. The randomized response technique: a test on drug use. Journal of the American Statistical Association, 70(352):814–818, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285441.

Gail:1976:CVO

[GG76a] Mitchell H. Gail and Sylvan B. Green. Critical values for the onesided two-sample Kolmogorov–Smirnov statistic. Journal of the American Statistical Association, 71(355):757–760, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285616.

Gallant:1976:NRA

[GG76b] A. Ronald Gallant and J. Jeffery Goebel. Nonlinear regression with autocorrelated errors. Journal of the American Statistical Association, 71(356):961–967, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286869.

Gideon:1976:SEQ

[GG76c] Rudy A. Gideon and John Gurland. Series expansions for quadratic forms in normal variables. Journal of the American Statistical Association, 71(353):227–232, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285774.

Gail:1978:SFG

[GG78] Mitchell H. Gail and Joseph L. Gastwirth. A scale-free goddness-of-fit test for the exponential distribution based on the Lorenz curve. Journal of the American Statistical Association, 73(364):787–793, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286281.

Gallant:1974:TSR

[GGE74] A. R. Gallant, Thomas M. Gerig, and J. W. Evans. Time series realizations obtained according to an experimental design. Journal of the American Statistical Association, 69(347):639–645, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285994.

Good:1970:EDL

[GGM70] I. J. Good, T. N. Gover, and G. J. Mitchell. Exact distributions for χ^{2} and for the likelihood-ratio statistic for the equiprobable multinomial distribution. Journal of the American Statistical Association, 65(329):267–283, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2283591. See corrigenda [GGM71].

Good:1971:CED

[GGM71] I. J. Good, T. N. Gover, and G. J. Mitchell. Corrigenda: Exact Distributions for χ^{2} and for the Likelihood-Ratio Statistic for the Equiprobable Multinomial Distribution. Journal of the American Statistical Association, 66(333):229, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284891. See [GGM70].

Ghosh:1973:NML

[GGS73] Malay Ghosh, James E. Grizzle, and Pranab Kumar Sen. Nonparametric methods in longitudinal studies. Journal of the American Statistical Association, 68(341):29–36, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284134.

Garey:1974:ISD

[GH74] M. R. Garey and F. K. Hwang. Isolating a single defective using group testing. Journal of the American Statistical Association, 69(345):151– 153, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285514.

Gleser:1976:EMN

[GH76a] Leon Jay Gleser and John D. Healy. Estimating the mean of a normal distribution with known coefficient of variation. Journal of the American Statistical Association, 71(356):977–981, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286872.

Green:1976:PME

[GH76b] J. R. Green and Y. A. S. Hegazy. Powerful modified-EDF goodnessof-fit tests. Journal of the American Statistical Association, 71 (353):204–209, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285770.

Genizi:1978:IEM

[GH78a] Abraham Genizi and Yosef Hochberg. On improved extensions of the T -method of multiple comparisons for unbalanced designs. Journal of the American Statistical Association, 73(364):879–884, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286299.

Gonzalez:1978:SAE

[GH78b] Maria Elena Gonzalez and Christine Hoza. Small-area estimation with application to unemployment and housing estimates. Journal of the American Statistical Association, 73(361):7–15, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286509.

Genizi:1979:CIE

[GH79] Abraham Genizi and Yosef Hochberg. Corrigenda: On Improved Extensions of the T = Method of Multiple Comparisons for Unbalanced Designs. Journal of the American Statistical Association, 74(367):744, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287029.

Ghosh:1975:DDT

[Gho75a] B. K. Ghosh. On the distribution of the difference of two t-variables. Journal of the American Statistical Association, 70(350):463–467, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285841.

Ghosh:1975:TSP

[Gho75b] B. K. Ghosh. A two-stage procedure for the Behrens–Fisher problem. Journal of the American Statistical Association, 70(350):457– 462, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285840.

Ghosh:1979:CSA

[Gho79] B. K. Ghosh. A comparison of some approximate confidence intervals for the binomial parameter. Journal of the American Statistical Association, 74(368):894–900, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286420.

Gibbs:1973:BRB

[Gib73] Jack P. Gibbs. Book review: [untitled]. Journal of the American Statistical Association, 68(341):242, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284185.

Gilbert:1972:MCS

[Gil72] Richard O. Gilbert. A Monte Carlo study of analysis of variance and competing rank tests for Scheffé’s mixed model. Journal of the American Statistical Association, 67(337):71–75, March 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284704.

Gilbert:1974:BRB

[Gil74a] Roy F. Gilbert. Book review: [untitled]. Journal of the American Statistical Association, 69(347):830–831, September 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286031.

Gillespie:1974:MDC

[Gil74b] Robert W. Gillespie. Measuring the demand for court services: a critique of the Federal District Courts case weights. Journal of the American Statistical Association, 69(345):38–43, March 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285497.

Gilbert:1976:BRB

[Gil76] John P. Gilbert. Book review: [untitled]. Journal of the American Statistical Association, 71(354):523–524, June 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285349.

Gilbert:1978:BRB

[Gil78] Neil Gilbert. Book review: [untitled]. Journal of the American Statistical Association, 73(362):444, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286697.

Gurney:1975:COR

[GJ75] Margaret Gurney and Robert S. Jewett. Constructing orthogonal replications for variance estimation. Journal of the American Statistical Association, 70(352):819–821, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285442.

Gokhale:1978:CAI

[GJ78] D. V. Gokhale and N. S. Johnson. A class of alternatives to independence in contingency tables. Journal of the American Statistical Association, 73(364):800–804, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286283.

Goodman:1970:CMA

[GK70] Leo A. Goodman and W. H. Kruskal. Corrigenda: Measures of Association for Cross Classifications. III: Approximate Sampling Theory. Journal of the American Statistical Association, 65(330):1011, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284622.

Goodman:1972:MAC

[GK72] Leo A. Goodman and William H. Kruskal. Measures of association for cross classifications, IV: Simplification of asymptotic variances. Journal of the American Statistical Association, 67(338):415–421, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284396.

Gastwirth:1975:BMG

[GK75] Joseph L. Gastwirth and Abba M. Krieger. On bounding moments from grouped data. Journal of the American Statistical Association, 70(350):468–471, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285842.

Goldstein:1976:LSF

[GK76] A. J. Goldstein and J. B. Kruskal. Least-squares fitting by monotonic functions having integer values. Journal of the American Statistical Association, 71(354):370–373, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285316.

Greenberg:1971:ARR

[GKAH71] Bernard G. Greenberg, Roy R. Kuebler, Jr., James R. Abernathy, and Daniel G. Horvitz. Application of the randomized response technique in obtaining quantitative data. Journal of the American Statistical Association, 66(334):243–250, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283916.

Gray:1971:ATP

[GL71] H. L. Gray and T. O. Lewis. Approximation of tail probabilities by means of the B_{n}-transformation. Journal of the American Statistical Association, 66(336):897–899, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284250.

Glaser:1976:CEC

[Gla76a] Ronald E. Glaser. Corrigenda: Exact Critical Values for Bartlett’s Test for Homogeneity of Variances. Journal of the American Statistical Association, 71(356):1010, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286896.

Glaser:1976:ECV

[Gla76b] Ronald E. Glaser. Exact critical values for Bartlett’s test for homogeneity of variances. Journal of the American Statistical Association, 71(354):488–490, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285339.

Glaser:1976:RGM

[Gla76c] Ronald E. Glaser. The ratio of the geometric mean to the arithmetic mean for a random sample from a gamma distribution. Journal of the American Statistical Association, 71(354):480–487, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285338.

Gladen:1979:UJE

[Gla79] Beth Gladen. The use of the jackknife to estimate proportions from toxicological data in the presence of litter effects. Journal of the American Statistical Association, 74(366):278–283, June 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286323.

Gleser:1972:NCB

[Gle72] Leon Jay Gleser. On a new class of bounds for the distribution of quadratic forms in normal variates. Journal of the American Statistical Association, 67(339):655–659, September 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284460.

Gleser:1973:BRB

[Gle73] Leon Jay Gleser. Book review: [untitled]. Journal of the American Statistical Association, 68(342):491, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284107.

Gleser:1974:BRB

[Gle74] Leon Jay Gleser. Book review: [untitled]. Journal of the American Statistical Association, 69(346):582, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285711.

Gleser:1975:BRB

[Gle75] Leon Jay Gleser. Book review: [untitled]. Journal of the American Statistical Association, 70(352):956, December 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285465.

Gleser:1976:CRN

[Gle76] Leon Jay Gleser. A canonical representation for the noncentral Wishart distribution useful for simulation. Journal of the American Statistical Association, 71(355):690–695, September 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285602.

Glick:1972:SBC

[Gli72] Ned Glick. Sample-based classification procedures derived from density estimators. Journal of the American Statistical Association, 67(337):116–122, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284709.

Glick:1978:BRB

[Gli78] Frederick P. Glick. Book review: [untitled]. Journal of the American Statistical Association, 73(361):223, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286562.

Gragg:1979:NSM

[GLT79] W. B. Gragg, R. J. Leveque, and J. A. Trangenstein. Numerically stable methods for updating regressions. Journal of the American Statistical Association, 74(365):161–168, March 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286746.

Guilkey:1975:DRR

[GM75] David K. Guilkey and James L. Murphy. Directed ridge regression techniques in cases of multicollinearity. Journal of the American Statistical Association, 70(352):769–775, December 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285435.

Gail:1977:CNC

[GM77a] Mitchell Gail and Nathan Mantel. Counting the number of r × c contingency tables with fixed margins. Journal of the American Statistical Association, 72(360):859–862, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286475.

Gunst:1977:BER

[GM77b] Richard F. Gunst and Robert L. Mason. Biased estimation in regression: An evaluation using mean squared error. Journal of the American Statistical Association, 72(359):616–628, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286229.

Gersovitz:1978:SRA

[GM78] Mark Gersovitz and James G. MacKinnon. Seasonality in regression: An application of smoothness priors. Journal of the American Statistical Association, 73(362):264–273, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286651.

Granger:1972:IVR

[GO72] Clive W. J. Granger and Daniel Orr. “infinite variance” and research strategy in time series analysis. Journal of the American Statistical Association, 67(338):275–285, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284369.

Godfrey:1970:BRB

[God70] Michael D. Godfrey. Book review: [untitled]. Journal of the American Statistical Association, 65(330):1006–1007, June 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284616.

Godambe:1974:BRB

[God74] V. P. Godambe. Book review: [untitled]. Journal of the American Statistical Association, 69(346):578–580, June 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285706.

Goel:1974:BRB

[Goe74] Prem K. Goel. Book review: [untitled]. Journal of the American Statistical Association, 69(345):280–281, March 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285556.

Goel:1977:BRB

[Goe77] Prem K. Goel. Book review: [untitled]. Journal of the American Statistical Association, 72(357):233, March 1977. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286954.

Goel:1978:BRB

[Goe78] Prem K. Goel. Book review: [untitled]. Journal of the American Statistical Association, 73(362):441, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286688.

Goel:1979:BRB

[Goe79] Prem K. Goel. Book review: [untitled]. Journal of the American Statistical Association, 74(366):508–509, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286374.

Gokhale:1973:ADD

[Gok73] D. V. Gokhale. Approximating disrete distributions, with applications. Journal of the American Statistical Association, 68 (344):1009–1012, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284541.

Gokhale:1979:BRB

[Gok79] D. V. Gokhale. Book review: [untitled]. Journal of the American Statistical Association, 74(367):724, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286999.

Goldberg:1970:BRB

[Gol70] Samuel Goldberg. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1420–1421, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284322.

Goldberger:1972:BRB

[Gol72] Arthur S. Goldberger. Book review: [untitled]. Journal of the American Statistical Association, 67(340):951–952, December 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284668.

Goldberger:1973:CBB

[Gol73] Arthur S. Goldberger. Correlations between binary outcomes and probabilistic predictions. Journal of the American Statistical Association, 68(341):84, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284143.

Goldberg:1975:EMB

[Gol75a] Judith D. Goldberg. The effects of misclassification on the bias in the difference between two proportions and the relative odds in the fourfold table. Journal of the American Statistical Association, 70 (351):561–567, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285933.

Goldstein:1975:CSD

[Gol75b] Matthew Goldstein. Comparison of some density estimate classification procedures. Journal of the American Statistical Association, 70(351):666–669, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285952.

Goldberg:1976:BRB

[Gol76] Michael A. Goldberg. Book review: [untitled]. Journal of the American Statistical Association, 71(354):527, June 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285353.

Goldberg:1977:BRB

[Gol77] Judith D. Goldberg. Book review: [untitled]. Journal of the American Statistical Association, 72(360):923, December 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286489.

Goldfeld:1978:BRB

[Gol78] Stephen M. Goldfeld. Book review: [untitled]. Journal of the American Statistical Association, 73(363):683, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286634.

Goodman:1970:MAQ

[Goo70] Leo A. Goodman. The multivariate analysis of qualitative data: Interactions among multiple classifications. Journal of the American Statistical Association, 65(329):226–256, March 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283589.

Good:1971:DOS

[Goo71a] I. J. Good. Does an observed sequence of numbers follow a simple rule? (another look at Bode’s law): Comment. Journal of the American Statistical Association, 66(335):559–562, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283523.

Goodman:1971:PCS

[Goo71b] Leo A. Goodman. Partitioning of chi-square, analysis of marginal contingency tables, and estimation of expected frequencies in multidimensional contingency tables. Journal of the American Statistical Association, 66(334):339–344, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283933.

Good:1972:BRB

[Goo72a] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 67(337):245–246, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284742.

Good:1972:SPP

[Goo72b] I. J. Good. Some probability paradoxes in choice from among random alternatives: Comment. Journal of the American Statistical Association, 67(338):374–375, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284385.

Goodman:1973:GUM

[Goo73] Leo A. Goodman. Guided and unguided methods for the selection of models for a set of T multidimensional contingency tables. Journal of the American Statistical Association, 68(341):165–175, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284162.

Goodman:1975:NMS

[Goo75a] Leo A. Goodman. A new model for scaling response patterns: An application of the quasi-independence concept. Journal of the American Statistical Association, 70(352):755–768, December 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285434.

Goodman:1975:RBT

[Goo75b] Leo A. Goodman. On the relationship between two statistics pertaining to tests of three-factor interaction in contingency tables. Journal of the American Statistical Association, 70(351):624–625, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285944.

Good:1978:BRBb

[Goo78a] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 73(362):441, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286689.

Good:1978:BRBc

[Goo78b] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 73(362):441–442, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286690.

Good:1978:BRBd

[Goo78c] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 73(362):442, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286691.

Good:1978:RBF

[Goo78d] I. J. Good. Review: Fractals: Form, Chance, and Dimension, by Benoît B. Mandelbrot. Journal of the American Statistical Association, 73(362):438, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286682.

Good:1979:BRBa

[Goo79a] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 74(366):501–502, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286363.

Good:1979:BRBb

[Goo79b] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 74(366):502–503, June 1979. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286364.

Good:1979:BRBc

[Goo79c] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 74(367):732, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287011.

Goodman:1979:SMA

[Goo79d] Leo A. Goodman. Simple models for the analysis of association in cross-classifications having ordered categories. Journal of the American Statistical Association, 74(367):537–552, September 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286971.

Gordon:1973:BRB

[Gor73a] Louis Gordon. Book review: [untitled]. Journal of the American Statistical Association, 68(341):242–243, March 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284186.

Gordon:1973:EEP

[Gor73b] Robert A. Gordon. An explicit estimation of the prevalence of commitment to a training schoo to age 18, by race and by sex. Journal of the American Statistical Association, 68(343):547–553, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284774.

Gordon:1976:BRB

[Gor76] A. D. Gordon. Book review: [untitled]. Journal of the American Statistical Association, 71(356):1000–1001, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286880.

Gonzalez:1975:SDP

[GOST75] Maria E. Gonzalez, Jack L. Ogus, Gary Shapiro, and Benjamin J. Tepping. Standards for discussion and presentation of errors in survey and census data. Journal of the American Statistical Association, 70(351):5–23, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286149.

Govindarajulu:1970:BRB

[Gov70] Z. Govindarajulu. Book review: [untitled]. Journal of the American Statistical Association, 65(330):996–997, June 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284608.

Govindarajulu:1971:CBL

[Gov71] Z. Govindarajulu. Corrigenda: Best Linear Estimates under Symmetric Censoring of the Parameters of a Double Exponential Population. Journal of the American Statistical Association, 66(333):229, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284892.

Govindarajulu:1976:CBL

[Gov76] Zakkula Govindarajulu. Corrigenda: Best Linear Estimates Under Symmetric Censoring of the Parameters of a Double Exponential Population. Journal of the American Statistical Association, 71(353): 255, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285803.

Gupta:1974:PNT

[GP74] Somesh Das Gupta and Michael D. Perlman. Power of the noncentral F -test: Effect of additional variates on Hotelling’s T^{2}-test. Journal of the American Statistical Association, 69(345):174–180, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285519.

Godfrey:1975:TRA

[GP75] L. G. Godfrey and D. S. Poskitt. Testing the restrictions of the Almon lag technique. Journal of the American Statistical Association, 70(349):105–108, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285384.

Gastwirth:1971:EDL

[GR71] Joseph L. Gastwirth and Herman Rubin. Effect of dependence on the level of some one-sample tests. Journal of the American Statistical Association, 66(336):816–820, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284232.

Graff:1974:GTP

[GR74] Lois E. Graff and Robert Roeloffs. A group-testing procedure in the presence of test error. Journal of the American Statistical Association, 69(345):159–163, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285516.

Goldstein:1975:SVT

[GR75] Matthew Goldstein and Manus Rabinowitz. Selection of variates for the two-group multinomial classification problem. Journal of the American Statistical Association, 70(352):776–781, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285436.

Goldstein:1977:CSV

[GR77] Matthew Goldstein and Manus Rabinowitz. Corrigenda: Selection of Variates for the Two-Group Multinomial Classification Problem. Journal of the American Statistical Association, 72(358):488, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286836.

Graybill:1977:WLA

[Gra77] Franklin A. Graybill. What lies ahead? Journal of the American Statistical Association, 72(357):7–10, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286900.

Greenberg:1970:TPS

[Gre70] Irwin Greenberg. Tables of Poisson sums of normal variables. Journal of the American Statistical Association, 65(330):984–989, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284603.

Greenberg:1971:CTP

[Gre71] Irwin Greenberg. Corrigenda: Tables of Poisson Sums of Normal Variables. Journal of the American Statistical Association, 66(333): 229, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284893.

Green:1973:BRB

[Gre73] Bert F. Green. Book review: [untitled]. Journal of the American Statistical Association, 68(341):238–239, March 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284180.

Grenander:1974:BRB

[Gre74] Ulf Grenander. Book review: [untitled]. Journal of the American Statistical Association, 69(347):829, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286028.

Greenberg:1975:MVP

[Gre75a] Edward Greenberg. Minimum variance properties of principal component regression. Journal of the American Statistical Association, 70(349):194–197, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285402.

Greenhouse:1975:BRB

[Gre75b] Samuel W. Greenhouse. Book review: [untitled]. Journal of the American Statistical Association, 70(350):482–484, June 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285845.

Greenwood:1975:SEM

[Gre75c] Michael J. Greenwood. A simultaneous-equations model of urban growth and migration. Journal of the American Statistical Association, 70(352):797–810, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285439.

Green:1976:OPO

[Gre76] Richard F. Green. Outlier-prone and outlier-resistant distributions. Journal of the American Statistical Association, 71(354):502–505, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285341.

Griffiths:1972:EAR

[Gri72] W. E. Griffiths. Estimation of actual response coefficients in the Hildreth–Houck random coefficient model. Journal of the American Statistical Association, 67(339):633–635, September 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284455.

Groeneveld:1972:AOG

[Gro72] Richard A. Groeneveld. Asymptotically optimal group rank tests for location. Journal of the American Statistical Association, 67 (340):847–849, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284647.

Gross:1976:CIR

[Gro76] Alan M. Gross. Confidence interval robustness with long-tailed symmetric distributions. Journal of the American Statistical Association, 71(354):409–416, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285324.

Gross:1977:CIB

[Gro77] Alan M. Gross. Confidence intervals for bisquare regression estimates. Journal of the American Statistical Association, 72(358):341– 354, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286799.

Grossman:1972:OPT

[GS72] Stanley I. Grossman and George P. H. Styan. Optimality properties of Theil’s BLUS residuals. Journal of the American Statistical Association, 67(339):672–673, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284464.

Gehan:1973:SRM

[GS73a] Edmund A. Gehan and M. M. Siddiqui. Simple regression methods for survival time studies. Journal of the American Statistical Association, 68(344):848–856, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284511.

Gelfand:1973:SPM

[GS73b] Alan E. Gelfand and Herbert Solomon. A study of Poisson’s models for jury verdicts in criminal and civil trials. Journal of the American Statistical Association, 68(342):271–278, June 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284062.

Guilkey:1973:ESU

[GS73c] David K. Guilkey and Peter Schmidt. Estimation of seemingly unrelated regressions with vector autoregressive errors. Journal of the American Statistical Association, 68(343):642–647, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284793.

Gelfand:1974:MJV

[GS74a] Alan E. Gelfand and Herbert Solomon. Modeling jury verdicts in the American legal system. Journal of the American Statistical Association, 69(345):32–37, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285496.

Gillo:1974:PMM

[GS74b] Martin W. Gillo and Maynard W. Shelly. Predictive modeling of multivariable and multivariate data. Journal of the American Statistical Association, 69(347):646–653, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285995.

Gelfand:1975:ADM

[GS75] Alan E. Gelfand and Herbert Solomon. Analyzing the decisionmaking process of the American jury. Journal of the American Statistical Association, 70(350):305–310, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285813.

Gelfand:1977:EJS

[GS77] Alan E. Gelfand and Herbert Solomon. On the effect of jury size: Rejoinder. Journal of the American Statistical Association, 72(359):536–537, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286213.

Guinn:1978:BRB

[Gui78] Ted Guinn. Book review: [untitled]. Journal of the American Statistical Association, 73(361):229, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286573.

Gulati:1975:BRB

[Gul75] Chandra M. Gulati. Book review: [untitled]. Journal of the American Statistical Association, 70(352):963, December 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285474.

Gunst:1979:BRB

[Gun79] Richard F. Gunst. Book review: [untitled]. Journal of the American Statistical Association, 74(367):733, September 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2287013.

Gupta:1973:BTC

[Gup73] Biswatosh Sen Gupta. The birth table and its construction for Belgium. Journal of the American Statistical Association, 68(341):44–45, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284136.

Gupta:1976:APN

[Gup76] Prithwis Das Gupta. Age-parity-nuptiality-specific stable population model that recognizes births to single women. Journal of the American Statistical Association, 71(354):308–314, June 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285302.

Gupta:1977:CAP

[Gup77] Prithwis Das Gupta. Corrigenda: Age-Parity-Nuptiality-Specific Stable Population Model that Recognizes Births to Single Women. Journal of the American Statistical Association, 72(358):488, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286835.

Guthery:1974:PR

[Gut74] Scott B. Guthery. Partition regression. Journal of the American Statistical Association, 69(348):945–947, December 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286168.

Griffiths:1978:PRP

[GW78] David Griffiths and Mary Willcox. Percentile regression: a parametric approach. Journal of the American Statistical Association, 73(363):496–498, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286588.

Graybill:1979:CIP

[GW79] Franklin A. Graybill and Chih-Ming Wang. Confidence intervals for proportions of variability in two-factor nested variance component models. Journal of the American Statistical Association, 74(366):368– 374, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286337.

Guthrie:1970:EES

[GY70] Donald Guthrie and Mary N. Youssef. Empirical evaluation of some chi-square tests for the order of a Markov chain. Journal of the American Statistical Association, 65(330):631–634, June 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284574.

H:1971:BRB

[H.71] J. A. H. Book review: [untitled]. Journal of the American Statistical Association, 66(336):924, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284268.

H:1972:BRBa

[H.72a] M. H. Book review: [untitled]. Journal of the American Statistical Association, 67(337):248–249, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284749.

H:1972:BRBb

[H72b] M. H. Book review: [untitled]. Journal of the American Statistical Association, 67(339):722, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284491.

H:1975:BRBa

[H.75a] D. V. H. Book review: [untitled]. Journal of the American Statistical Association, 70(350):492, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285860.

H:1975:BRBb

[H.75b] D. V. H. Book review: [untitled]. Journal of the American Statistical Association, 70(350):492, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285863.

Hamdan:1971:CEC

[HAB71] M. A. Hamdan and H. A. Al-Bayyati. Canonical expansion of the compound correlated bivariate Poisson distribution. Journal of the American Statistical Association, 66(334):390–393, June 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283942.

Haberman:1977:BRB

[Hab77] Shelby J. Haberman. Book review: [untitled]. Journal of the American Statistical Association, 72(360):923–924, December 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286490.

Haff:1978:MNM

[Haf78] L. R. Haff. The multivariate normal mean with intraclass correlated components: Estimation of urban fire alarm probabilities. Journal of the American Statistical Association, 73(364):767–774, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286278.

Haggstrom:1978:BRB

[Hag78] Gus W. Haggstrom. Book review: [untitled]. Journal of the American Statistical Association, 73(363):688, September 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286646.

Hahn:1970:AFC

[Hah70a] Gerald J. Hahn. Additional factors for calculating prediction intervals for samples from a normal distribution. Journal of the American Statistical Association, 65(332):1668–1676, December 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284349.

Hahn:1970:CFC

[Hah70b] Gerald J. Hahn. Corrigenda: Factors for Calculating Samples from a Normal Distribution. Journal of the American Statistical Association, 65(330):1011, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284621.

Hahn:1972:SPI

[Hah72] Gerald J. Hahn. Simultaneous prediction intervals to contain the standard deviations or ranges of future samples from a normal distribution. Journal of the American Statistical Association, 67 (340):938–942, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284666.

Haines:1970:BRB

[Hai70] George H. Haines, Jr. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1406–1407, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284311.

Hajek:1974:BRB

[Haj74] Jaroslav Hajek. Book review: [untitled]. Journal of the American Statistical Association, 69(345):279, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285552.

Halpern:1973:PRB

[Hal73] Elkan F. Halpern. Polynomial regression from a Bayesian approach. Journal of the American Statistical Association, 68(341):137–143, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284157.

Halperin:1974:BRB

[Hal74] Max Halperin. Book review: [untitled]. Journal of the American Statistical Association, 69(345):281, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285558.

Hald:1975:ODS

[Hal75] A. Hald. Optimum double sampling tests of given strength I. the normal distribution. Journal of the American Statistical Association, 70(350):451–456, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285839.

Hamdan:1971:LST

[Ham71] M. A. Hamdan. The logarithm of the sum of two correlated lognormal variates. Journal of the American Statistical Association, 66(333):105–106, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284858.

Hampel:1974:ICR

[Ham74] Frank R. Hampel. The influence curve and its role in robust estimation. Journal of the American Statistical Association, 69(346):383– 393, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285666.

Hamermesh:1975:BRB

[Ham75] Daniel S. Hamermesh. Book review: [untitled]. Journal of the American Statistical Association, 70(349):251, March 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285415.

Hamilton:1976:MCT

[Ham76] Basil L. Hamilton. A Monte Carlo test of the robustness of parametric and nonparametric analysis of covariance against unequal regression slopes. Journal of the American Statistical Association, 71(356):864–869, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286852.

Ham:1978:NEE

[Ham78] John C. Ham. A note on the efficient estimation of the linear expenditure system. Journal of the American Statistical Association, 73(361):208–210, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286548.

Hamilton:1979:REE

[Ham79] Martin A. Hamilton. Robust estimates of the ED50. Journal of the American Statistical Association, 74(366):344–354, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286333.

Han:1973:DSP

[Han73] Chien-Pai Han. Double sampling with partial information on auxiliary variables. Journal of the American Statistical Association, 68(344):914–918, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284521.

Hannula:1975:BRB

[Han75] Russell E. Hannula. Book review: [untitled]. Journal of the American Statistical Association, 70(349):259, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285427.

Han:1978:EMW

[Han78] Chien-Pai Han. Estimating means when a group of observations is classified by a linear discriminant function. Journal of the American Statistical Association, 73(363):661–665, September 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286620.

Harkness:1970:BRB

[Har70] W. L. Harkness. Book review: [untitled]. Journal of the American Statistical Association, 65(332):1686–1687, December 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284356.

Harris:1971:HTC

[Har71] Bernard Harris. Hypothesis testing and confidence intervals for products and quotients of Poisson parameters with applications to reliability. Journal of the American Statistical Association, 66(335):609–613, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2283539.

Hartigan:1972:DCD

[Har72] J. A. Hartigan. Direct clustering of a data matrix. Journal of the American Statistical Association, 67(337):123–129, March 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284710.

Harville:1973:APO

[Har73] David A. Harville. Assigning probabilities to the outcomes of multientry competitions. Journal of the American Statistical Association, 68(342):312–316, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284068.

Harter:1974:ARP

[Har74a] H. Leon Harter. Adaptive robust procedures: a partial review and some suggestions for future applications and theory: Comment. Journal of the American Statistical Association, 69(348):923–925, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286161.

Harville:1974:OPS

[Har74b] David A. Harville. Optimal procedures for some constrained selection problems. Journal of the American Statistical Association, 69(346):446–452, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285676.

Harkema:1975:ACC

[Har75] R. Harkema. An analytical comparison of certainty equivalence and sequential updating. Journal of the American Statistical Association, 70(350):348–350, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285820.

Harvey:1977:CPE

[Har77a] A. C. Harvey. A comparison of preliminary estimators for robust regression. Journal of the American Statistical Association, 72(360):910–913, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286486.

Harville:1977:ULM

[Har77b] David Harville. The use of linear-model methodology to rate high school or college football teams. Journal of the American Statistical Association, 72(358):278–289, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286789.

Harville:1977:MLAa

[Har77c] David A. Harville. Maximum likelihood approaches to variance component estimation and to related problems. Journal of the American Statistical Association, 72(358):320–338, June 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286796.

Harville:1977:MLAb

[Har77d] David A. Harville. Maximum likelihood approaches to variance component estimation and to related problems: Rejoinder. Journal of the American Statistical Association, 72(358):339–340, June 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286798.

Hartley:1978:EMN

[Har78] Michael J. Hartley. Estimating mixtures of normal distributions and switching regressions: Comment. Journal of the American Statistical Association, 73(364):738–741, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286267.

Harville:1979:SUR

[Har79] David A. Harville. Some useful representations for constrained mixed-model estimation. Journal of the American Statistical Association, 74(365):200–206, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286751.

Hastings:1972:TDS

[Has72] W. Keith Hastings. Test data for statistical algorithms: Least squares and ANOVA. Journal of the American Statistical Association, 67 (340):874–879, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284653.

Hasenkamp:1976:BRB

[Has76] Georg Hasenkamp. Book review: [untitled]. Journal of the American Statistical Association, 71(353):249–250, March 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285790.

Haugh:1976:CIT

[Hau76] Larry D. Haugh. Checking the independence of two covariancestationary time series: a univariate residual cross-correlation approach. Journal of the American Statistical Association, 71(354):378– 385, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285318.

Hawkins:1971:BRO

[Haw71] Douglas M. Hawkins. On the bounds of the range of order statistics. Journal of the American Statistical Association, 66(335):644–645, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2283548.

Hawkins:1974:DEM

[Haw74] Douglas M. Hawkins. The detection of errors in multivariate data using principal components. Journal of the American Statistical Association, 69(346):340–344, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285654.

Hawkins:1977:TSO

[Haw77] Douglas M. Hawkins. Testing a sequence of observations for a shift in location. Journal of the American Statistical Association, 72(357):180–186, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286934.

Hayes:1971:ESO

[Hay71a] Robert H. Hayes. Efficiency of simple order statistic estimates when losses are piecewise linear. Journal of the American Statistical Association, 66(333):127–135, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284861.

Haynes:1971:CAP

[Hay71b] J. D. Haynes. Comment on article on pharmacokinetics. Journal of the American Statistical Association, 66(333):53–54, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284846.

Hays:1976:BRB

[Hay76] William L. Hays. Book review: [untitled]. Journal of the American Statistical Association, 71(356):1002–1003, December 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286882.

Hayes:1978:BRB

[Hay78] Richard E. Hayes. Book review: [untitled]. Journal of the American Statistical Association, 73(362):448, June 1978. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286707.

Hager:1970:IPG

[HB70] Harold W. Hager and Lee J. Bain. Inferential procedures for the generalized gamma distribution. Journal of the American Statistical Association, 65(332):1601–1609, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284343.

Hendrickson:1972:ESP

[HB72] Arlo D. Hendrickson and Robert J. Buehler. Elicitation of subjective probabilities by sequential choices. Journal of the American Statistical Association, 67(340):880–883, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284654.

Hazra:1973:APS

[HB73a] P. K. Hazra and K. S. Banerjee. On the augmentation procedure in singular weighting designs. Journal of the American Statistical Association, 68(342):392–393, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284081.

Hwang:1973:CIS

[HB73b] Dar-Shong Hwang and Robert J. Buehler. Confidence intervals for some functions of several Bernoulli parameters with reliability applications. Journal of the American Statistical Association, 68(341):211–217, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284173.

Huang:1974:TRD

[HB74] Cliff J. Huang and Ben W. Bolch. On the testing of regression disturbances for normality. Journal of the American Statistical Association, 69(346):330–335, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285652.

Haugh:1977:IDR

[HB77] Larry D. Haugh and G. E. P. Box. Identification of dynamic regression (distributed lag) models connecting two time series. Journal of the American Statistical Association, 72(357):121–130, March 1977.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286919.

Hewett:1972:CND

[HBA72] John E. Hewett, William G. Bulgren, and D. E. Amos. Corrigenda: a Note on a Double Sample Test. Journal of the American Statistical Association, 67(337):252–253, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284759.

Hauck:1977:WTA

[HD77] Walter W. Hauck, Jr. and Allan Donner. Wald’s test as applied to hypotheses in logit analysis. Journal of the American Statistical Association, 72(360):851–853, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286473.

Hansen:1978:CLC

[HD78] Pierre Hansen and Michel Delattre. Complete-link cluster analysis by graph coloring. Journal of the American Statistical Association, 73(362):397–403, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286672.

Heath:1979:BRB

[Hea79] David Heath. Book review: [untitled]. Journal of the American Statistical Association, 74(365):247, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286766.

Heiberger:1975:LMC

[Hei75] Richard M. Heiberger. Linear models for cell means and reciprocal within cell variances. Journal of the American Statistical Association, 70(351):618–623, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285943.

Hemmerle:1974:NAV

[Hem74] William J. Hemmerle. Nonorthogonal analysis of variance using iterative improvement and balanced residuals. Journal of the American Statistical Association, 69(347):772–778, September 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286016.

Hemmerle:1976:INA

[Hem76] William J. Hemmerle. Iterative nonorthogonal analysis of covariance. Journal of the American Statistical Association, 71(353):195– 199, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285768.

Hemmerle:1979:BHU

[Hem79] William J. Hemmerle. Balanced hypotheses and unbalanced data. Journal of the American Statistical Association, 74(368):794–798, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286401.

Henry:1971:RMM

[Hen71] Neil W. Henry. The retention model: a Markov chain with variable transition probabilities. Journal of the American Statistical Association, 66(334):264–267, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283919.

Hendricks:1972:BRB

[Hen72a] Walter J. Hendricks. Book review: [untitled]. Journal of the American Statistical Association, 67(338):498–499, June 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284428.

Henry:1972:CRM

[Hen72b] Neil W. Henry. Corrigenda: The Retention Model. Journal of the American Statistical Association, 67(337):252, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284758.

Henze:1979:END

[Hen79] Fritz H.-H. Henze. The exact noncentral distributions of Spearman’s r and other related correlation coefficients. Journal of the American Statistical Association, 74(366):459–464, June 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286356. See corrigenda [Hen80].

Henze:1980:CEN

[Hen80] Fritz H.-H. Henze. Corrigenda: The Exact Noncentral Distribution of Spearman’s r and Other Related Correlation Coefficients. Journal of the American Statistical Association, 75(371):765, September 1980.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287701. See [Hen79].

Herrey:1971:PPD

[Her71] Erna M. J. Herrey. Percentage points of H-distribution for computing confidence limits or performing t-tests by way of the mean absolute deviation. Journal of the American Statistical Association, 66(333):187–188, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284871.

Herson:1976:IRE

[Her76] Jay Herson. An investigation of relative efficiency of least-squares prediction to conventional probability sampling plans. Journal of the American Statistical Association, 71(355):700–703, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285604.

Hester:1971:BRB

[Hes71] Donald D. Hester. Book review: [untitled]. Journal of the American Statistical Association, 66(335):657, September 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283557.

Hettmansperger:1973:LSC

[Het73] Thomas P. Hettmansperger. A large sample conservative test for location with unknown scale parameters. Journal of the American Statistical Association, 68(342):466–468, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284100.

Heyman:1971:BRB

[Hey71] Daniel P. Heyman. Book review: [untitled]. Journal of the American Statistical Association, 66(336):923, December 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284264.

Heyns:1978:BRB

[Hey78] Barbara Heyns. Book review: [untitled]. Journal of the American Statistical Association, 73(361):224–225, March 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286564.

Huynh:1970:CUW

[HF70] Huynh Huynh and Leonard S. Feldt. Conditions under which mean square ratios in repeated measurements designs have exact F - distributions. Journal of the American Statistical Association, 65 (332):1582–1589, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284340.

Hogg:1975:TSA

[HFR75] Robert V. Hogg, Doris M. Fisher, and Ronald H. Randles. A twosample adaptive distribution-free test. Journal of the American Statistical Association, 70(351):656–661, September 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285950.

Hinz:1970:TFN

[HG70] Paul Hinz and John Gurland. A test of fit for the negative binomial and other contagious distributions. Journal of the American Statistical Association, 65(330):887–903, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284595.

Halperin:1971:NES

[HG71a] Max Halperin and Joan Gurian. A note on estimation in straight line regression when both variables are subject to error. Journal of the American Statistical Association, 66(335):587–589, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283534.

Hinz:1971:CTF

[HG71b] Paul Hinz and John Gurland. Corrigenda: a Test of Fit for the Negative Binomial and Other Contagious Distributions. Journal of the American Statistical Association, 66(333):229, March 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284894.

Hartwell:1973:EVC

[HG73] T. D. Hartwell and D. W. Gaylor. Estimating variance components for two-way disproportionate data with missing cells by the method of unweighted mean. Journal of the American Statistical Association, 68(342):379–383, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284079.

Horn:1975:CEH

[HH75] Susan D. Horn and Roger A. Horn. Comparison of estimators of heteroscedastic variances in linear models. Journal of the American Statistical Association, 70(352):872–879, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285450.

Hill:1977:TRA

[HH77] Richard W. Hill and Paul W. Holland. Two robust alternatives to least-squares regression. Journal of the American Statistical Association, 72(360):828–833, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286469.

Hill:1978:CTR

[HH78] Richard W. Hill and Paul W. Holland. Corrigenda: Two Robust Alternatives to Least-Squares Regression. Journal of the American Statistical Association, 73(362):455, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286710.

Horn:1975:EHV

[HHD75] Susan D. Horn, Roger A. Horn, and David B. Duncan. Estimating heteroscedastic variances in linear models. Journal of the American Statistical Association, 70(350):380–385, June 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285827.

Higgins:1970:DAE

[Hig70] Gerald F. Higgins. A discriminant analysis of employment in defense and nondefense industries. Journal of the American Statistical Association, 65(330):613–622, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284572.

Hildebrand:1970:BRB

[Hil70a] David K. Hildebrand. Book review: [untitled]. Journal of the American Statistical Association, 65(329):457–458, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283612.

Hill:1970:ZLP

[Hil70b] Bruce M. Hill. Zipf’s law and prior distributions for the composition of a population. Journal of the American Statistical Association, 65 (331):1220–1232, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284288.

Hilderbrand:1972:BRB

[Hil72] David K. Hilderbrand. Book review: [untitled]. Journal of the American Statistical Association, 67(340):953, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284669.

Hill:1973:BRB

[Hil73] Bruce M. Hill. Book review: [untitled]. Journal of the American Statistical Association, 68(342):487–489, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284105.

Hildebrand:1974:BRB

[Hil74a] David K. Hildebrand. Book review: [untitled]. Journal of the American Statistical Association, 69(345):276, March 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285548.

Hildreth:1974:EUU

[Hil74b] Clifford Hildreth. Expected utility of uncertain ventures. Journal of the American Statistical Association, 69(345):9–17, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285494.

Hill:1974:RFF

[Hil74c] Bruce M. Hill. The rank-frequency form of Zipf’s law. Journal of the American Statistical Association, 69(348):1017–1026, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286182.

Hill:1975:ABL

[Hil75] Bruce M. Hill. Aberrant behavior of the likelihood function in discrete cases. Journal of the American Statistical Association, 70 (351):717–719, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285961.

Hill:1976:BRB

[Hil76] Bruce M. Hill. Book review: [untitled]. Journal of the American Statistical Association, 71(356):999–1000, December 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286879.

Hill:1978:BRB

[Hil78] Ken Hill. Book review: [untitled]. Journal of the American Statistical Association, 73(362):445, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286700.

Hill:1979:PMN

[Hil79] Bruce M. Hill. Posterior moments of the number of species in a finite population and the posterior probability of finding a new species. Journal of the American Statistical Association, 74(367):668–673, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286989.

Hinkley:1971:ITP

[Hin71] David V. Hinkley. Inference in two-phase regression. Journal of the American Statistical Association, 66(336):736–743, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284220.

Hinkley:1976:BRBa

[Hin76a] David V. Hinkley. Book review: [untitled]. Journal of the American Statistical Association, 71(355):777–778, September 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285642.

Hinkley:1976:BRBb

[Hin76b] David V. Hinkley. Book review: [untitled]. Journal of the American Statistical Association, 71(355):778, September 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285643.

Hinich:1977:BRB

[Hin77a] Melvin J. Hinich. Book review: [untitled]. Journal of the American Statistical Association, 72(358):482, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286825.

Hinz:1977:BRB

[Hin77b] Paul N. Hinz. Book review: [untitled]. Journal of the American Statistical Association, 72(360):929, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286499.

Hinich:1979:ELS

[Hin79] Melvin J. Hinich. Estimating the lag structure of a nonlinear time series model. Journal of the American Statistical Association, 74(366):449–452, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286353.

Hixson:1976:BRB

[Hix76] Marilyn Hixson. Book review: [untitled]. Journal of the American Statistical Association, 71(355):771–772, September 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285626.

Hartley:1973:ELM

[HJ73] H. O. Hartley and K. S. E. Jayatillake. Estimation for linear models with unequal variances. Journal of the American Statistical Association, 68(341):189–192, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284166.

Hegemann:1976:PTT

[HJ76] Victor Hegemann and Dallas E. Johnson. The power of two tests for nonadditivity. Journal of the American Statistical Association, 71(356):945–948, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286866.

Hodges:1973:WTM

[HL73] J. L. Hodges, Jr. and E. L. Lehmann. Wilcoxon and t test for matched pairs of typed subjects. Journal of the American Statistical Association, 68(341):151–158, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284160.

Hsi:1975:MPW

[HL75] Bartholomew P. Hsi and Thomas A. Louis. A modified play-the-winner rule for sequential trials. Journal of the American Statistical Association, 70(351):644–647, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285948.

Howe:1970:AST

[HM70] Richard B. Howe and Raymond H. Myers. An alternative to Satterthwaite’s test involving positive linear combinations of variance components. Journal of the American Statistical Association, 65(329):404– 412, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2283604.

Harter:1973:CML

[HM73] Leon H. Harter and Albert H. Moore. Corrigenda: Maximum-Likelihood Estimation, from Doubly Censored Samples, of the Parameters of the First Asymptotic Distribution of Extreme Values. Journal of the American Statistical Association, 68(343):757, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284834.

Harrison:1979:THB

[HM79a] M. J. Harrison and B. P. M. McCabe. A test for heteroscedasticity based on ordinary least squares residuals. Journal of the American Statistical Association, 74(366):494–499, June 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286361.

Hartz:1979:BAN

[HM79b] Mary A. Hartz and Donald L. Meyer. The Bayesian analysis of a nonadditive model. Journal of the American Statistical Association, 74(366):383–388, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286340.

Hsu:1974:SPB

[HMW74] Der-Ann Hsu, Robert B. Miller, and Dean W. Wichern. On the stable Paretian behavior of stock-market prices. Journal of the American Statistical Association, 69(345):108–113, March 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285507.

Hannan:1977:EPE

[HN77] E. J. Hannan and D. F. Nicholls. The estimation of the prediction error variance. Journal of the American Statistical Association, 72(360):834–840, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286470.

Hannan:1979:CEP

[HN79] E. J. Hannan and D. F. Nicholls. Corrigenda: The Estimation of the Prediction Error Variance. Journal of the American Statistical Association, 74(368):941, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286448.

Hocking:1971:MLE

[HO71] R. R. Hocking and H. H. Oxspring. Maximum likelihood estimation with incomplete multinomial data. Journal of the American Statistical Association, 66(333):65–70, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284849.

Hoadley:1970:BLI

[Hoa70] Bruce Hoadley. A Bayesian look at inverse linear regression. Journal of the American Statistical Association, 65(329):356–369, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283598.

Hoaglin:1974:RBE

[Hoa74] David C. Hoaglin. Review: Robust Estimates of Location: Survey and Advances, by D. F. Andrews, P. J. Bickel, F. R. Hampel, P. J. Huber, W. H. Rogers, and J. W. Tukey. Journal of the American Statistical Association, 69(345):272–274, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285544.

Hoaglin:1975:SSV

[Hoa75] David C. Hoaglin. The small-sample variance of the Pitman location estimators. Journal of the American Statistical Association, 70(352):880–888, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285451.

Hoaglin:1977:DAC

[Hoa77] David C. Hoaglin. Direct approximations for chi-squared percentage points. Journal of the American Statistical Association, 72 (359):508–515, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286208.

Hobson:1976:PPS

[Hob76] Richard Hobson. Properties preserved by some smoothing functions. Journal of the American Statistical Association, 71(355):763–766, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285618.

Hochberg:1976:MMM

[Hoc76] Yosef Hochberg. A modification of the T -method of multiple comparisons for a one-way layout with unequal variances. Journal of the American Statistical Association, 71(353):200–203, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285769.

Hochberg:1977:UDS

[Hoc77] Yosef Hochberg. On the use of double sampling schemes in analyzing categorical data with misclassification errors. Journal of the American Statistical Association, 72(360):914–921, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286487.

Hoel:1972:ISR

[Hoe72] David G. Hoel. An inverse stopping rule for play-the-winner sampling. Journal of the American Statistical Association, 67(337):148– 151, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284713.

Hoel:1976:BRBa

[Hoe76a] David G. Hoel. Book review: [untitled]. Journal of the American Statistical Association, 71(353):252, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285799.

Hoel:1976:BRBb

[Hoe76b] David G. Hoel. Book review: [untitled]. Journal of the American Statistical Association, 71(353):252, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285800.

Hoerl:1977:SSA

[Hoe77] Arthur E. Hoerl. A simulation study of alternatives to ordinary least squares: Comment. Journal of the American Statistical Association, 72(357):94–95, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286911.

Hogg:1972:BRB

[Hog72a] Robert V. Hogg. Book review: [untitled]. Journal of the American Statistical Association, 67(338):497, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284423.

Hogg:1972:MLK

[Hog72b] Robert V. Hogg. More light on the kurtosis and related statistics. Journal of the American Statistical Association, 67(338):422–424, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284397.

Hogg:1974:ARPa

[Hog74a] Robert V. Hogg. Adaptive robust procedures: a partial review and some suggestions for future applications and theory. Journal of the American Statistical Association, 69(348):909–923, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286160.

Hogg:1974:ARPb

[Hog74b] Robert V. Hogg. Adaptive robust procedures: a partial review and some suggestions for future applications and theory: Rejoinder. Journal of the American Statistical Association, 69(348):927, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286164.

Hogarth:1975:CPAa

[Hog75a] Robin M. Hogarth. Cognitive processes and the assessment of subjective probability distributions. Journal of the American Statistical Association, 70(350):271–289, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285808.

Hogarth:1975:CPAb

[Hog75b] Robin M. Hogarth. Cognitive processes and the assessment of subjective probability distributions: Rejoinder. Journal of the American Statistical Association, 70(350):294, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285811.

Hogg:1975:EPR

[Hog75c] Robert V. Hogg. Estimates of percentile regression lines using salary data. Journal of the American Statistical Association, 70(349):56–59, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285375.

Hogg:1976:NAA

[Hog76] Robert V. Hogg. Necessary analysis and adaptive inference: Comment. Journal of the American Statistical Association, 71(353):110– 111, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285741.

Hogg:1978:BRB

[Hog78] Robert V. Hogg. Book review: [untitled]. Journal of the American Statistical Association, 73(364):891–892, December 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286307.

Hogarth:1979:BRB

[Hog79] Robin M. Hogarth. Book review: [untitled]. Journal of the American Statistical Association, 74(366):504–505, June 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286368.

Hollander:1970:DFT

[Hol70a] Myles Hollander. A distribution-free test for parallelism. Journal of the American Statistical Association, 65(329):387–394, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283600.

Holmes:1970:DTF

[Hol70b] James M. Holmes. A direct test of Friedman’s permanent income theory. Journal of the American Statistical Association, 65 (331):1159–1162, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284282.

Holmes:1970:DBR

[Hol70c] R. A. Holmes. Discriminatory bias in rates charged by the Canadian Automobile Insurance Industry. Journal of the American Statistical Association, 65(329):108–122, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283578.

Holmes:1971:CIP

[Hol71a] James M. Holmes. A condition for independence of permanent and transitory components of a series. Journal of the American Statistical Association, 66(333):13–15, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284840.

Holmes:1971:DBR

[Hol71b] R. A. Holmes. Discriminatory bias in rates charged by the Canadian Automobile Insurance Industry: Rejoinder. Journal of the American Statistical Association, 66(334):290–291, June 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283925.

Holmes:1972:CCI

[Hol72a] James M. Holmes. Corrigenda: a Condition for Independence of Permanent and Transitory Components of a Series. Journal of the American Statistical Association, 67(337):253, March 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284760.

Holmes:1972:KAS

[Hol72b] James M. Holmes. The Keynesian aggregate supply function for labor. Journal of the American Statistical Association, 67(340): 797–802, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284639.

Holgate:1974:BRB

[Hol74] Philip Holgate. Book review: [untitled]. Journal of the American Statistical Association, 69(347):831, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286032.

Holbrook:1975:BRB

[Hol75] Robert S. Holbrook. Book review: [untitled]. Journal of the American Statistical Association, 70(349):257, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285424.

Hommel:1978:TPC

[Hom78] Gerhard Hommel. Tail probabilities for contingency tables with small expectations. Journal of the American Statistical Association, 73(364):764–766, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286277.

Hooke:1974:BRB

[Hoo74] Robert Hooke. Book review: [untitled]. Journal of the American Statistical Association, 69(345):276–277, March 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285549.

Horowitz:1976:BRB

[Hor76] Ira Horowitz. Book review: [untitled]. Journal of the American Statistical Association, 71(354):532, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285362.

Hora:1978:SSD

[Hor78] Stephen C. Hora. Sample size determination in Bayesian discriminant analysis. Journal of the American Statistical Association, 73 (363):569–572, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286602.

Hosmer:1978:EMN

[Hos78] David W. Hosmer. Estimating mixtures of normal distributions and switching regressions: Comment. Journal of the American Statistical Association, 73(364):741–744, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286268.

Howe:1974:ACL

[How74a] W. G. Howe. Approximate confidence limits on the mean of X + Y where X and Y are two tabled independent random variables. Journal of the American Statistical Association, 69(347):789–794, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286019.

Howrey:1974:BRB

[How74b] E. Philip Howrey. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1051–1052, December 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286197.

Hammonds:1970:QMR

[HP70] T. M. Hammonds and D. I. Padberg. A quantitative measure of the risk factor in investment decisions. Journal of the American Statistical Association, 65(330):602–612, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284571.

Heckman:1974:EEF

[HP74] James Heckman and Solomon Polachek. Empirical evidence on the functional form of the earnings-schooling relationship. Journal of the American Statistical Association, 69(346):350–354, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285656.

Hochberg:1975:OSS

[HQ75] Yosef Hochberg and Dana Quade. One-sided simultaneous confidence bounds on regression surfaces with intercepts. Journal of the American Statistical Association, 70(352):889–891, December 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285452.

Hedayat:1970:ISR

[HR70] A. Hedayat and D. S. Robson. Independent stepwise residuals for testing homoscedasticity. Journal of the American Statistical Association, 65(332):1573–1581, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284339.

Hochberg:1977:ISI

[HR77] Yosef Hochberg and German Rodriguez. Intermediate simultaneous inference procedures. Journal of the American Statistical Association, 72(357):220–225, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286942.

Harris:1973:RSI

[HS73] Bernard Harris and Andrew P. Soms. The reliability of systems of independent parallel components when some components are repeated. Journal of the American Statistical Association, 68(344):894–898, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284518.

Harris:1974:PGI

[HS74] Bernard Harris and Andrew P. Soms. Properties of the generalized incomplete modified Bessel distributions with applications to reliability theory. Journal of the American Statistical Association, 69(345):259– 263, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285539.

Hocking:1975:FRA

[HS75] R. R. Hocking and F. M. Speed. A full rank analysis of some linear model problems. Journal of the American Statistical Association, 70(351):706–712, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285959.

Hewett:1979:STS

[HS79] John E. Hewett and John D. Spurrier. Some two-stage K-sample tests. Journal of the American Statistical Association, 74(366):398– 404, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286343.

Hsi:1970:BRB

[Hsi70] Bartholomew P. Hsi. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1418–1419, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284320.

Hsiao:1979:AMC

[Hsi79] Cheng Hsiao. Autoregressive modeling of Canadian money and income data. Journal of the American Statistical Association, 74 (367):553–560, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286972.

Holt:1979:MBA

[HST79] D. Holt, T. M. F. Smith, and T. J. Tomberlin. A model-based approach to estimation for small subgroups of a population. Journal of the American Statistical Association, 74(366):405–410, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286344.

Hsu:1979:DSP

[Hsu79] D. A. Hsu. Detecting shifts of parameter in gamma sequences with applications to stock price and air traffic flow analysis. Journal of the American Statistical Association, 74(365):31–40, March 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286717.

Hewett:1972:TSC

[HT72] John E. Hewett and R. K. Tsutakawa. Two-stage chi square goodness-of-fit test. Journal of the American Statistical Association, 67(338):395–401, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284391.

Hinich:1975:SMR

[HT75] Melvin J. Hinich and Prem P. Talwar. A simple method for robust regression. Journal of the American Statistical Association, 70(349):113–119, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285386.

Hillmer:1979:LFS

[HT79] Steven C. Hillmer and George C. Tiao. Likelihood function of stationary multiple autoregressive moving average models. Journal of the American Statistical Association, 74(367):652–660, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286987.

Huang:1973:SRI

[Hua73] David S. Huang. Short-run instability in single-family housing starts. Journal of the American Statistical Association, 68(344):788–792, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284501.

Huber:1974:ARP

[Hub74a] Peter J. Huber. Adaptive robust procedures: a partial review and some suggestions for future applications and theory: Comment. Journal of the American Statistical Association, 69(348):926–927, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286163.

Hubert:1974:AET

[Hub74b] Lawrence Hubert. Approximate evaluation techniques for the singlelink and complete-link hierarchical clustering procedures. Journal of the American Statistical Association, 69(347):698–704, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286004.

Hudleston:1975:BRB

[Hud75] Peter J. Hudleston. Book review: [untitled]. Journal of the American Statistical Association, 70(352):963, December 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285475.

Huddleston:1976:BRB

[Hud76] Harold F. Huddleston. Book review: [untitled]. Journal of the American Statistical Association, 71(355):774, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285632.

Hume:1975:BRB

[Hum75] M. W. Hume. Book review: [untitled]. Journal of the American Statistical Association, 70(352):963–964, December 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285476.

Huntsberger:1970:BRB

[Hun70] David V. Huntsberger. Book review: [untitled]. Journal of the American Statistical Association, 65(330):997–998, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284609.

Hunt:1972:BRB

[Hun72] B. R. Hunt. Book review: [untitled]. Journal of the American Statistical Association, 67(337):248, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284747.

Hunt:1975:BRB

[Hun75] B. R. Hunt. Book review: [untitled]. Journal of the American Statistical Association, 70(349):252, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285417.

Hurd:1972:SSE

[Hur72] Michael D. Hurd. Small-sample estimation of a structural equation with autocorrelated errors. Journal of the American Statistical Association, 67(339):567–573, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284439.

Hurd:1973:BRB

[Hur73] Michael D. Hurd. Book review: [untitled]. Journal of the American Statistical Association, 68(344):1025–1027, December 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284549.

Hurd:1974:BRB

[Hur74] Michael D. Hurd. Book review: [untitled]. Journal of the American Statistical Association, 69(347):833–834, September 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286036.

Hogg:1972:SUD

[HURD72] Robert V. Hogg, Vincent A. Uthoff, Ronald H. Randles, and Alan S. Davenport. On the selection of the underlying distribution and adaptive estimation. Journal of the American Statistical Association, 67(339):597–600, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284445.

Hutter:1972:SPS

[Hut72] James L. Hutter. Statistics and political science. Journal of the American Statistical Association, 67(340):735–742, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284630.

Halperin:1974:EDC

[HW74] Max Halperin and James Ware. Early decision in a censored Wilcoxon two-sample test for accumulating survival data. Journal of the American Statistical Association, 69(346):414–422, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285671.

Hill:1975:SFZ

[HW75] Bruce M. Hill and Michael Woodroofe. Stronger forms of Zipf’s law. Journal of the American Statistical Association, 70(349):212– 219, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285406.

Hwang:1972:MDA

[Hwa72] F. K. Hwang. A method for detecting all defective members in a population by group testing. Journal of the American Statistical Association, 67(339):605–608, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284447.

Hwang:1974:FSD

[Hwa74] F. K. Hwang. On finding a single defective in binomial group testing. Journal of the American Statistical Association, 69(345):146– 150, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285513.

Hwang:1975:GBG

[Hwa75] F. K. Hwang. A generalized binomial group testing problem. Journal of the American Statistical Association, 70(352):923–926, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285458.

Ifram:1970:MDA

[Ifr70] Adnan F. Ifram. On mixtures of distributions with applications to estimation. Journal of the American Statistical Association, 65(330):749–754, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284584.

Ignall:1974:BRB

[Ign74] Edward Ignall. Book review: [untitled]. Journal of the American Statistical Association, 69(346):580–581, June 1974. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285708.

Ijiri:1970:BRB

[Iji70] Yuji Ijiri. Book review: [untitled]. Journal of the American Statistical Association, 65(332):1685–1686, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284355.

Ijiri:1971:FQA

[Iji71] Yuji Ijiri. Fundamental queries in aggregation theory. Journal of the American Statistical Association, 66(336):766–782, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284228.

Imrey:1976:ICT

[IJK76] Peter B. Imrey, William D. Johnson, and Gary G. Koch. An incomplete contingency table approach to paired-comparison experiments. Journal of the American Statistical Association, 71(355):614–623, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285591.

Imrey:1976:BRB

[Imr76] Peter B. Imrey. Book review: [untitled]. Journal of the American Statistical Association, 71(353):245, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285782.

Imrey:1978:BRB

[Imr78] Peter B. Imrey. Book review: [untitled]. Journal of the American Statistical Association, 73(363):679–680, September 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286628.

Imrey:1979:BRB

[Imr79] Peter B. Imrey. Book review: [untitled]. Journal of the American Statistical Association, 74(367):725–726, September 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2287003.

Inglis:1977:BRB

[Ing77] James Inglis. Book review: [untitled]. Journal of the American Statistical Association, 72(360):932, December 1977. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286504.

Isaki:1974:BTC

[Isa74] Cary T. Isaki. Bias of two classification procedures in a dichotomous population. Journal of the American Statistical Association, 69(348):957–960, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286170.

Ismail:1976:BRB

[Ism76] A. H. Ismail. Book review: [untitled]. Journal of the American Statistical Association, 71(354):531–532, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285360.

Izenman:1978:BRB

[Ize78] Alan J. Izenman. Book review: [untitled]. Journal of the American Statistical Association, 73(362):435–438, June 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286681.

Izenman:1979:BRB

[Ize79] Alan Julian Izenman. Book review: [untitled]. Journal of the American Statistical Association, 74(368):924–926, December 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286428.

Jagers:1973:HMP

[Jag73] Peter Jagers. How many people pay their tram fares? Journal of the American Statistical Association, 68(344):801–804, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284503.

James:1972:PIB

[Jam72] Ian R. James. Products of independent beta variables with applications to connor and Mosimann’s generalized Dirichlet distribution. Journal of the American Statistical Association, 67(340):910–912, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284660.

James:1975:MDW

[Jam75] Ian R. James. Multivariate distributions which have beta conditional distributions. Journal of the American Statistical Association, 70(351):681–684, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285955.

James:1977:INB

[Jam77] Ian R. James. On F -independence for nonnegative, bounded-sum variables when the bound is large. Journal of the American Statistical Association, 72(357):215–219, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286941.

Jaynes:1979:BRB

[Jay79] Edwin T. Jaynes. Book review: Inference, method and decision: Towards a Bayesian philosophy of science. Journal of the American Statistical Association, 74(367):740–741, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287026.

Jogdeo:1973:SSE

[JB73] Kumar Jogdeo and Robert Bohrer. Some simple examples and counterexamples about the existence of optimum tests. Journal of the American Statistical Association, 68(343):679–682, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284798.

Jedamus:1973:BRB

[Jed73] Paul Jedamus. Book review: [untitled]. Journal of the American Statistical Association, 68(343):752, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284821.

Jennrich:1970:ATE

[Jen70] Robert I. Jennrich. An asymptotic χ^{2} test for the equality of two correlation matrices. Journal of the American Statistical Association, 65(330):904–912, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284596.

Jennergren:1973:OSP

[Jen73] L. Peter Jennergren. An optimal sampling policy for a matched sampling problem. Journal of the American Statistical Association, 68(341):148–150, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284159.

Jensen:1976:BRBa

[Jen76a] Louis Jensen. Book review: [untitled]. Journal of the American Statistical Association, 71(354):532, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285361.

Jensen:1976:BRBb

[Jen76b] Louis Jensen. Book review: [untitled]. Journal of the American Statistical Association, 71(355):778, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285644.

Jessen:1970:PSM

[Jes70] R. J. Jessen. Probability sampling with marginal constraints. Journal of the American Statistical Association, 65(330):776–796, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284587.

Jessen:1973:SPP

[Jes73] R. J. Jessen. Some properties of probability lattice sampling. Journal of the American Statistical Association, 68(341):20–28, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284133.

Johnson:1972:ATW

[JG72a] Dallas E. Johnson and Franklin A. Graybill. An analysis of a twoway model with interaction and no replication. Journal of the American Statistical Association, 67(340):862–868, December 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284651.

Johnson:1972:ETW

[JG72b] Dallas E. Johnson and Franklin A. Graybill. Estimation of σ^{2} in a two-way classification model with interaction. Journal of the American Statistical Association, 67(338):388–394, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284390.

Joreskog:1975:EMM

[JG75] Karl G. Joreskog and Arthur S. Goldberger. Estimation of a model with multiple indicators and multiple causes of a single latent variable. Journal of the American Statistical Association, 70 (351):631–639, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285946.

Johnson:1972:CAE

[JK72] N. L. Johnson and S. Kotz. Comparisons among estimators of a scale parameter of the beta-stacy distribution. Journal of the American Statistical Association, 67(337):203–205, March 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284724.

John:1970:AMS

[Joh70] Sleeba John. On analyzing mixed samples. Journal of the American Statistical Association, 65(330):755–762, June 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284585.

John:1972:BRB

[Joh72] Peter W. M. John. Book review: [untitled]. Journal of the American Statistical Association, 67(340):954, December 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284671.

Johansson:1973:GLF

[Joh73a] Johny K. Johansson. A generalized logistic function with an application to the effect of advertising. Journal of the American Statistical Association, 68(344):824–827, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284507.

Johnson:1973:BRBa

[Joh73b] Norman L. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 68(341):246–247, March 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284194.

Johnson:1973:BRBb

[Joh73c] Richard A. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 68(344):1027–1028, December 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284550.

Johns:1974:NEL

[Joh74] M. V. Johns, Jr. Nonparametric estimation of location. Journal of the American Statistical Association, 69(346):453–460, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285677.

John:1975:TCT

[Joh75a] S. John. Tables for comparing two normal variances or two gamma means. Journal of the American Statistical Association, 70(350):344– 347, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285819.

Johnson:1975:MTS

[Joh75b] N. S. Johnson. C_{α} method for testing for significance in the r × c contingency table. Journal of the American Statistical Association, 70(352):942–947, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285461.

Johnson:1978:BRB

[Joh78a] N. L. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 73(361):229, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286574.

Johnson:1978:MTC

[Joh78b] Norman J. Johnson. Modified t tests and confidence intervals for asymmetrical populations. Journal of the American Statistical Association, 73(363):536–544, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286597.

Johnson:1978:EMN

[Joh78c] Norman L. Johnson. Estimating mixtures of normal distributions and switching regressions: Comment. Journal of the American Statistical Association, 73(364):750, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286274.

Joiner:1973:BRB

[Joi73] Brian L. Joiner. Book review: [untitled]. Journal of the American Statistical Association, 68(343):754–755, September 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284827.

Jolicoeur:1973:ICL

[Jol73] Pierre Jolicoeur. Imaginary confidence limits of the slope of the major axis of a bivariate normal distribution: a sampling experiment. Journal of the American Statistical Association, 68(344):866–871, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284514.

Jones:1973:BRB

[Jon73] Robert C. Jones. Book review: [untitled]. Journal of the American Statistical Association, 68(343):753–754, September 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284825.

Jones:1975:FA

[Jon75] Richard H. Jones. Fitting autoregressions. Journal of the American Statistical Association, 70(351):590–592, September 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285938.

Jones:1976:EIG

[Jon76] Richard H. Jones. Estimation of the innovation generalized variance of a multivariate stationary time series. Journal of the American Statistical Association, 71(354):386–388, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285319.

Jones:1979:EAD

[Jon79] Douglas H. Jones. An efficient adaptive distribution-free test for location. Journal of the American Statistical Association, 74 (368):822–828, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286406.

Joshi:1976:NBT

[Jos76] V. M. Joshi. A note on Birnbaum’s theory of the likelihood principle. Journal of the American Statistical Association, 71(354):345– 346, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285310.

Joanes:1974:SPB

[JP74] D. N. Joanes and H. W. Peers. On the sampling properties of Bayesian point estimators. Journal of the American Statistical Association, 69(346):560–564, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285697.

Joiner:1971:SPR

[JR71] Brian L. Joiner and Joan R. Rosenblatt. Some properties of the range in samples from Tukey’s symmetric lambda distributions. Journal of the American Statistical Association, 66(334):394–399, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283943.

Jenkins:1973:RE

[JRH73] O. C. Jenkins, L. J. Ringer, and H. O. Hartley. Root estimators. Journal of the American Statistical Association, 68(342):414–419, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284088.

Jensen:1972:GAD

[JS72] D. R. Jensen and Herbert Solomon. A Gaussian approximation to the distribution of a definite quadratic form. Journal of the American Statistical Association, 67(340):898–902, December 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284657.

Jensen:1973:CGA

[JS73] D. R. Jensen and H. Solomon. Corrigenda: a Gaussian Approximation to the Distribution of a Definite Quadratic Form. Journal of the American Statistical Association, 68(343):757, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284835.

Jain:1975:BPS

[JS75] Gopi C. Jain and Naunihal Singh. On bivariate power series distributions associated with Lagrange expansion. Journal of the American Statistical Association, 70(352):951–954, December 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285463.

Johnson:1980:NFP

[JTB80] Mark E. Johnson, Gary L. Tietjen, and Richard J. Beckman. A new family of probability distributions with applications to Monte Carlo studies. Journal of the American Statistical Association, 75(370):276– 279, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2287446.

See correction [JTB81].

Johnson:1981:CNF

[JTB81] Mark E. Johnson, Gary L. Tietjen, and Richard J. Beckman. Correction: “A new family of probability distributions with applications to Monte Carlo studies” [J. Amer. Statist. Assoc. 75 (1980), no. 370, 276–279, MR 81f:62032]. Journal of the American Statistical Association, 76(376):1022, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287627. See [FH81, JTB80].

Johnson:1978:SAL

[JW78] Richard A. Johnson and Thomas E. Wehrly. Some angular-linear distributions and related regression models. Journal of the American Statistical Association, 73(363):602–606, September 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286608.

K:1971:BRB

[K.71] J. B. K. Book review: [untitled]. Journal of the American Statistical Association, 66(336):924, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284271.

K:1972:BRBa

[K.72a] J. B. K. Book review: [untitled]. Journal of the American Statistical Association, 67(339):721, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284490.

K:1972:BRBb

[K.72b] J. B. K. Book review: [untitled]. Journal of the American Statistical Association, 67(339):722, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284493.

K:1972:BRBc

[K.72c] J. B. K. Book review: [untitled]. Journal of the American Statistical Association, 67(340):964, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284688.

K:1973:BRB

[K.73] G. M. K. Book review: [untitled]. Journal of the American Statistical Association, 68(342):501, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284126.

K:1974:BRB

[K.74] J. K. Book review: [untitled]. Journal of the American Statistical Association, 69(346):583, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285716.

K:1975:BRB

[K.75] J. K. Book review: [untitled]. Journal of the American Statistical Association, 70(349):259, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285428.

Kish:1978:MMS

[KA78] Leslie Kish and Dallas W. Anderson. Multivariate and multipurpose stratification. Journal of the American Statistical Association, 73(361):24–34, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286511.

Kabe:1976:CBQ

[Kab76] D. G. Kabe. On confidence bands for quantiles of a normal population. Journal of the American Statistical Association, 71(354):417– 419, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285325.

Kadane:1970:TOR

[Kad70] Joseph B. Kadane. Testing overidentifying restrictions when the disturbances are small. Journal of the American Statistical Association, 65(329):182–185, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283585.

Kadane:1973:BRB

[Kad73] Joseph B. Kadane. Book review: [untitled]. Journal of the American Statistical Association, 68(344):1025, December 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284548.

Kadane:1975:BRBa

[Kad75a] Joseph B. Kadane. Book review: [untitled]. Journal of the American Statistical Association, 70(349):248–249, March 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285412.

Kadane:1975:BRBb

[Kad75b] Joseph B. Kadane. Book review: [untitled]. Journal of the American Statistical Association, 70(352):964, December 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285478.

Kaiser:1972:BRB

[Kai72] Henry F. Kaiser. Book review: [untitled]. Journal of the American Statistical Association, 67(338):493, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284418.

Kalbfleisch:1978:LMN

[Kal78a] John D. Kalbfleisch. Likelihood methods and nonparametric tests. Journal of the American Statistical Association, 73(361):167–170, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286539.

Kallianpur:1978:BRB

[Kal78b] G. Kallianpur. Book review: [untitled]. Journal of the American Statistical Association, 73(362):440, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286686.

Kapadia:1970:BRB

[Kap70] C. H. Kapadia. Book review: [untitled]. Journal of the American Statistical Association, 65(329):475–476, March 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283623.

Kappenman:1977:TID

[Kap77] Russell F. Kappenman. Tolerance intervals for the double exponential distribution. Journal of the American Statistical Association, 72(360):908–909, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286485.

Karson:1970:DCM

[Kar70] M. J. Karson. Design criterion for minimum bias estimation of response surfaces. Journal of the American Statistical Association, 65 (332):1565–1572, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284338.

Kass:1974:BRB

[Kas74] Gordon V. Kass. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1050–1051, December 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286195.

Katz:1970:BRB

[Kat70] Leo Katz. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1400–1401, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284307.

Katz:1977:ESR

[Kat77] Leo Katz. An efficient sequential ranking procedure. Journal of the American Statistical Association, 72(360):841–844, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286471.

Katz:1978:CES

[Kat78] Leo Katz. Corrigenda: An Efficient Sequential Ranking Procedure. Journal of the American Statistical Association, 73(362):455, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286711.

Kaufman:1972:TDI

[Kau72] G. M. Kaufman. Two distributions involving a Bessel function. Journal of the American Statistical Association, 67(340):906–909, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284659.

Kaufman:1973:BRB

[Kau73] Gordon M. Kaufman. Book review: [untitled]. Journal of the American Statistical Association, 68(342):492–493, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284110.

Keith:1972:BRB

[Kei72] E. Gordon Keith. Book review: [untitled]. Journal of the American Statistical Association, 67(339):720, September 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284486.

Kelejian:1971:TSL

[Kel71] Harry H. Kelejian. Two-stage least squares and econometric systems linear in parameters but nonlinear in the endogenous variables. Journal of the American Statistical Association, 66(334):373–374, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283938.

Kelker:1973:RWE

[Kel73] Douglas Kelker. A random walk epidemic simulation. Journal of the American Statistical Association, 68(344):821–823, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284506.

Kempthorne:1970:BRB

[Kem70] Oscar Kempthorne. Book review: [untitled]. Journal of the American Statistical Association, 65(329):455–456, March 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283610.

Kempthorne:1972:BRB

[Kem72] Oscar Kempthorne. Book review: [untitled]. Journal of the American Statistical Association, 67(339):712, September 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284475.

Kempthorne:1973:BRB

[Kem73] Oscar Kempthorne. Book review: [untitled]. Journal of the American Statistical Association, 68(341):245–246, March 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284192.

Kempthorne:1977:CCS

[Kem77] Oscar Kempthorne. Conditional confidence statements and confidence estimators: Comment. Journal of the American Statistical Association, 72(360):816–819, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286465.

Kenkel:1976:CSS

[Ken76] James L. Kenkel. Comment on the small-sample power of Durbin’s htest. Journal of the American Statistical Association, 71(353):96–97, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285738.

Keyfitz:1971:LIA

[Key71] Nathan Keyfitz. Linkages of intrinsic to age-specific rates. Journal of the American Statistical Association, 66(334):275–281, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283921.

Keyfitz:1972:FP

[Key72] Nathan Keyfitz. On future population. Journal of the American Statistical Association, 67(338):347–363, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284381.

Kish:1970:BRR

[KF70] Leslie Kish and Martin R. Frankel. Balanced repeated replications for standard errors. Journal of the American Statistical Association, 65(331):1071–1094, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284276.

Kmenta:1970:ESU

[KG70] Jan Kmenta and Roy F. Gilbert. Estimation of seemingly unrelated regressions with autoregressive disturbances. Journal of the American Statistical Association, 65(329):186–197, March 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283586.

Keilson:1971:SRD

[KG71] J. Keilson and H. Gerber. Some results for discrete unimodality. Journal of the American Statistical Association, 66(334):386–389, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283941.

Keselman:1979:ACM

[KGR79] H. J. Keselman, Paul A. Games, and Joanne C. Rogan. An addendum to “A Comparison of the Modified-Tukey and Scheffé Methods of Multiple Comparisons for Pairwise Contrasts”. Journal of the American Statistical Association, 74(367):626–627, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286980. See [KR78b].

Klatzky:1971:CCA

[KH71] Sheila R. Klatzky and Robert W. Hodge. A canonical correlation analysis of occupational mobility. Journal of the American Statistical Association, 66(333):16–22, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284841.

Khan:1977:BRB

[Kha77] Rasul A. Khan. Book review: [untitled]. Journal of the American Statistical Association, 72(357):231, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286949.

Kiefer:1977:CCSa

[Kie77a] J. Kiefer. Conditional confidence statements and confidence estimators. Journal of the American Statistical Association, 72 (360):789–808, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286460.

Kiefer:1977:CCSb

[Kie77b] J. Kiefer. Conditional confidence statements and confidence estimators: Rejoinder. Journal of the American Statistical Association, 72(360):822–827, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286468.

Kiefer:1978:PNA

[Kie78a] J. Kiefer. Pseudorandom number assignment in statistically designed simulation and distribution sampling experiments: Comment. Journal of the American Statistical Association, 73 (363):523–524, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286594;   http://www.jstor.org/stable/
pdfplus/2286594.pdf. See [SM78a].

Kiefer:1978:EMN

[Kie78b] Nicholas M. Kiefer. Estimating mixtures of normal distributions and switching regressions: Comment. Journal of the American Statistical Association, 73(364):744–745, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286269.

Kim:1970:CAI

[Kim70] Sungwoo Kim. Capital appropriations and investment behavior. Journal of the American Statistical Association, 65(331):1180–1194, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284284.

Kim:1977:BRB

[Kim77] Geung Ho Kim. Book review: [untitled]. Journal of the American Statistical Association, 72(358):486–487, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286833.

King:1972:BRB

[Kin72a] Benjamin King. Book review: [untitled]. Journal of the American Statistical Association, 67(339):713–714, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284477.

King:1972:ALR

[Kin72b] Nathan King. An alternative for the linear regression equation when the predictor variable is uncontrolled and the sample size is small. Journal of the American Statistical Association, 67(337):217–219, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284729.

King:1976:BRB

[Kin76] Benjamin F. King. Book review: [untitled]. Journal of the American Statistical Association, 71(353):246–247, March 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285784.

Kinory:1977:BRB

[Kin77] Shimshon Kinory. Book review: [untitled]. Journal of the American Statistical Association, 72(358):487, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286834.

Kinory:1978:BRB

[Kin78] Shimshon Kinory. Book review: [untitled]. Journal of the American Statistical Association, 73(362):443, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286695.

Kinory:1979:BRB

[Kin79] Shimshon Kinory. Book review: [untitled]. Journal of the American Statistical Association, 74(366):513, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286381.

Kish:1978:CSS

[Kis78] Leslie Kish. Chance, statistics, and statisticians. Journal of the American Statistical Association, 73(361):1–6, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286508.

Koch:1972:LMA

[KJT72] Gary G. Koch, William D. Johnson, and H. Dennis Tolley. A linear models approach to the analysis of survival and extent of disease in multidimensional contingency tables. Journal of the American Statistical Association, 67(340):783–796, December 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284638.

Kaufman:1973:BAN

[KK73] G. M. Kaufman and Benjamin King. A Bayesian analysis of nonresponse in dichotomous processes. Journal of the American Statistical Association, 68(343):670–678, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284797.

Kedem-Kimelfeld:1975:ELL

[KK75] Benjamin Kedem-Kimelfeld. Estimating the lags of lag processes. Journal of the American Statistical Association, 70(351):603–605, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285940.

Kadane:1979:FNP

[KK79] Joseph B. Kadane and David Kairys. Fair numbers of peremptory challenges in jury trials. Journal of the American Statistical Association, 74(368):747–753, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286396.

Koch:1972:AMA

[KL72] Gary G. Koch and Stanley Lemeshow. An application of multivariate analysis to complex sample survey data. Journal of the American Statistical Association, 67(340):780–782, December 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284637.

Kau:1976:FFE

[KL76] James B. Kau and Cheng F. Lee. The functional form in estimating the density gradient: An empirical investigation. Journal of the American Statistical Association, 71(354):326–327, June 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285305.

Klein:1970:BRB

[Kle70] L. R. Klein. Book review: [untitled]. Journal of the American Statistical Association, 65(329):469–471, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283619.

Klein:1971:WE

[Kle71] Lawrence R. Klein. Whither econometrics? Journal of the American Statistical Association, 66(334):415–421, June 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283948.

Klein:1973:CEO

[Kle73a] Roger W. Klein. K-class estimators: The optimum normalization for finite samples. Journal of the American Statistical Association, 68(342):445–451, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284096.

Kleiner:1973:BRB

[Kle73b] Beat Kleiner. Book review: [untitled]. Journal of the American Statistical Association, 68(343):751, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284819.

Kleinman:1973:PEV

[Kle73c] Joel C. Kleinman. Proportions with extraneous variance: Single and independent sample. Journal of the American Statistical Association, 68(341):46–54, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284137.

Klein:1974:BRB

[Kle74] Roger W. Klein. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1045–1046, December 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286188.

Kleber:1977:BRBa

[Kle77a] Richard S. Kleber. Book review: Elementary Statistical Concepts, by Ronald E. Walpole; Understanding Statistics, 2nd edition, by William Mendenhall; Lyman Ott; Elementary Statistics, 2nd edition, by Robert R. Johnson; Statistics, by Norma Gilbert. Journal of the American Statistical Association, 72(359):684–687, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286245.

Kleber:1977:BRBb

[Kle77b] Richard S. Kleber. Book review: [untitled]. Journal of the American Statistical Association, 72(359):687, September 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286246.

Klotz:1973:OGT

[Klo73] Benjamin P. Klotz. Oscillatory growth in three nations. Journal of the American Statistical Association, 68(343):562–567, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284776.

Klotz:1976:TGR

[Klo76] Jerome H. Klotz. Truncated geometric response-duration of response models for clinical trials. Journal of the American Statistical Association, 71(354):331–334, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285307.

Kupper:1974:EDC

[KM74] Lawrence L. Kupper and Edward F. Meydrech. Experimental design considerations based on a new approach to mean square error estimation of response surfaces. Journal of the American Statistical Association, 69(346):461–463, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285678.

Khorasani:1979:ECB

[KM79] Fariborz Khorasani and George A. Milliken. On the exactness of confidence bands about a linear model. Journal of the American Statistical Association, 74(366):446–448, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286352.

Kmenta:1974:BRB

[Kme74] J. Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 69(347):835, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286038.

Kmenta:1977:BRBa

[Kme77a] Jan Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 72(357):237, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286959.

Kmenta:1977:BRBb

[Kme77b] Jan Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 72(357):238, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286960.

Kmenta:1977:BRBc

[Kme77c] Jan Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 72(357):238, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286961.

Kmenta:1977:BRBd

[Kme77d] Jan Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 72(357):238–239, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286962.

Kmenta:1979:BRBa

[Kme79a] Jan Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 74(365):245, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286763.

Kmenta:1979:BRBb

[Kme79b] Jan Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 74(365):248–249, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286770.

Kupper:1975:HES

[KMS75] L. L. Kupper, A. J. McMichael, and R. Spirtas. A hybrid epidemiologic study design useful in estimating relative risk. Journal of the American Statistical Association, 70(351):524–528, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285927.

Kaminsky:1975:BLU

[KN75] Kenneth S. Kaminsky and Paul I. Nelson. Best linear unbiased prediction of order statistics in location and scale families. Journal of the American Statistical Association, 70(349):145–150, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285390.

Kreijger:1977:ELR

[KN77] Rudolf G. Kreijger and Heinz Neudecker. Exact linear restrictions on parameters in the general linear model with a singular covariance matrix. Journal of the American Statistical Association, 72(358):430– 432, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286812.

Knott:1970:SSP

[Kno70] M. Knott. The small sample power of one-sided Kolmogorov tests for a shift in location of the normal distribution. Journal of the American Statistical Association, 65(331):1384–1391, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284303.

Knoke:1976:MCD

[Kno76] James D. Knoke. Multiple comparisons with dichotomous data. Journal of the American Statistical Association, 71(356):849–853, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286850.

Kumar:1979:CEM

[KNP79] K. D. Kumar, E. H. Nicklin, and A. S. Paulson. Comment on “estimating mixtures of normal distributions and switching regressions”. Journal of the American Statistical Association, 74(365):52– 55, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286720.

Kobrin:1973:HHC

[Kob73] Frances E. Kobrin. Household headship and its changes in the United States, 1940–1960, 1970. Journal of the American Statistical Association, 68(344):793–800, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284502.

Koch:1973:AAM

[Koc73] Gary G. Koch. An alternative approach to multivariate response error models for sample survey data with applications to estimators involving subclass means. Journal of the American Statistical Association, 68(344):906–913, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284520.

Kohn:1972:CEM

[Koh72] Robert E. Kohn. A cost-effectiveness model for air pollution control with a single stochastic variable. Journal of the American Statistical Association, 67(337):19–22, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284696.

Koopmans:1972:BRBa

[Koo72a] L. H. Koopmans. Book review: [untitled]. Journal of the American Statistical Association, 67(339):718–719, September 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284484.

Koopmans:1972:BRBb

[Koo72b] L. H. Koopmans. Book review: [untitled]. Journal of the American Statistical Association, 67(340):960, December 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284680.

Koop:1973:NKA

[Koo73] J. C. Koop. A note on Koop’s approach for finding the bias of the ratio estimate: Comment. Journal of the American Statistical Association, 68(342):407, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284085.

Koopmans:1976:BRB

[Koo76] L. H. Koopmans. Book review: [untitled]. Journal of the American Statistical Association, 71(354):521–522, June 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285347.

Koop:1977:BRB

[Koo77] J. C. Koop. Book review: [untitled]. Journal of the American Statistical Association, 72(357):231–232, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286950.

Kotz:1970:BRB

[Kot70] Samuel Kotz. Book review: [untitled]. Journal of the American Statistical Association, 65(330):990–991, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284604.

Kotz:1972:BRB

[Kot72] Samuel Kotz. Book review: [untitled]. Journal of the American Statistical Association, 67(337):244–245, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284740.

Kotz:1973:BRB

[Kot73] Samuel Kotz. Book review: [untitled]. Journal of the American Statistical Association, 68(343):749–750, September 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284817.

Kousser:1972:BRB

[Kou72] J. Morgan Kousser. Book review: [untitled]. Journal of the American Statistical Association, 67(338):493, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284417.

Koziol:1978:CIC

[KP78] James A. Koziol and Michael D. Perlman. Combining independent chi-squared tests. Journal of the American Statistical Association, 73(364):753–763, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286276.

Kopecky:1979:ESG

[KP79] Kenneth J. Kopecky and Donald A. Pierce. Efficiency of smooth goodness-of-fit tests. Journal of the American Statistical Association, 74(366):393–397, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286342.

Kain:1970:MVH

[KQ70] John F. Kain and John M. Quigley. Measuring the value of housing quality. Journal of the American Statistical Association, 65(330):532– 548, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284565.

Kain:1972:NOE

[KQ72] John F. Kain and John M. Quigley. Note on Owner’s estimate of housing value. Journal of the American Statistical Association, 67(340):803–806, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284640.

Kinderman:1976:CGN

[KR76] A. J. Kinderman and J. G. Ramage. Computer generation of normal random variables. Journal of the American Statistical Association, 71(356):893–896, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2286857. See correction [KR90].

Kaminsky:1978:PIL

[KR78a] Kenneth S. Kaminsky and Lennart S. Rhodin. The prediction information in the latest failure. Journal of the American Statistical Association, 73(364):863–866, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286295.

Keselman:1978:CMT

[KR78b] H. J. Keselman and Joanne C. Rogan. A comparison of the modified-Tukey and Scheffé methods of multiple comparisons for pairwise contrasts. Journal of the American Statistical Association, 73(361):47– 52, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286514.

See addendum [KGR79].

Kinderman:1990:CCG

[KR90] A. J. Kinderman and J. G. Ramage. Correction: Computer generation of normal random variables. Journal of the American Statistical Association, 85(409):272, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289582. See [KR76].

Kraemer:1973:IAN

[Kra73] Helena Chmura Kraemer. Improved approximation to the non-null distribution of the correlation coefficient. Journal of the American Statistical Association, 68(344):1004–1008, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284540.

Kraemer:1974:NND

[Kra74] Helena Chmura Kraemer. The non-null distribution of the Spearman rank correlation coefficient. Journal of the American Statistical Association, 69(345):114–117, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285508.

Kraemer:1975:CIA

[Kra75] Helena Chmura Kraemer. Corrigenda: Improved Approximation to the Non-Null Distribution of the Correlation Coefficient. Journal of the American Statistical Association, 70(350):494, June 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285866.

Kraemer:1976:SSN

[Kra76] Helena Chmura Kraemer. The small sample nonnull properties of Kendall’s coefficient of concordance for normal populations. Journal of the American Statistical Association, 71(355):608–613, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285590.

Krewski:1976:DFC

[Kre76] Daniel Krewski. Distribution-free confidence intervals for quantile intervals. Journal of the American Statistical Association, 71(354):420– 422, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285326.

Krishnan:1972:SRB

[Kri72] Marakatha Krishnan. Series representations of a bivariate singly noncentral t-distribution. Journal of the American Statistical Association, 67(337):228–231, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284733.

Krieger:1979:BMG

[Kri79] Abba M. Krieger. Bounding moments, the Gini index and Lorenz curve from grouped data for unimodal density functions. Journal of the American Statistical Association, 74(366):375–378, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286338.

Krotki:1971:BRB

[Kro71] Karol J. Krotki. Book review: [untitled]. Journal of the American Statistical Association, 66(333):220–221, March 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284874.

Kruskal:1979:FEW

[Kru79] William Kruskal. Field experimentation in weather modification: Comment. Journal of the American Statistical Association, 74(365): 84–86, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286728.

Krzanowski:1975:DCU

[Krz75] W. J. Krzanowski. Discrimination and classification using both binary and continuous variables. Journal of the American Statistical Association, 70(352):782–790, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285437.

Krzanowski:1976:CRL

[Krz76] W. J. Krzanowski. Canonical representation of the location model for discrimination or classification. Journal of the American Statistical Association, 71(356):845–848, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286849.

Krzanowski:1979:BGC

[Krz79] W. J. Krzanowski. Between-groups comparison of principal components. Journal of the American Statistical Association, 74(367): 703–707, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2286995. See correction [Krz81].

Krzanowski:1981:CBG

[Krz81] W. J. Krzanowski. Correction: “Between-groups comparison of principal components” [J. Amer. Statist. Assoc. 74 (1979), no. 367, 703– 707, MR 81a:62049]. Journal of the American Statistical Association, 76(376):1022, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287629. See [Krz79].

Kish:1971:RUA

[KS71a] Leslie Kish and Alastair Scott. Retaining units after changing strata and probabilities. Journal of the American Statistical Association, 66(335):461–470, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283509.

Kumar:1971:FSD

[KS71b] Satindar Kumar and Milton Sobel. Finding a single defective in binomial group-testing. Journal of the American Statistical Association, 66(336):824–828, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284234.

Kimeldorf:1973:CCI

[KS73] George Kimeldorf and Allan Sampson. A class of covariance inequalities. Journal of the American Statistical Association, 68(341):228– 230, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284175.

Khatri:1975:EVC

[KS75] C. G. Khatri and K. R. Shah. Exact variance of combined interand intra-block estimates in incomplete block designs. Journal of the American Statistical Association, 70(350):402–406, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285831.

Klotz:1977:OWL

[KT77] Jerome Klotz and James Teng. One-way layout for counts and the exact enumeration of the Kruskal–Wallis H distribution with ties. Journal of the American Statistical Association, 72(357):165–169, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286931.

Keselman:1975:ETU

[KTS75] H. J. Keselman, Larry E. Toothaker, and M. Shooter. An evaluation of two unequal n_{k} forms of the Tukey multiple comparison statistic. Journal of the American Statistical Association, 70(351):584–587, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285936.

Kupper:1976:BRB

[Kup76] Lawrence L. Kupper. Book review: [untitled]. Journal of the American Statistical Association, 71(353):250–251, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285792.

Kupper:1978:BRB

[Kup78] Lawrence L. Kupper. Book review: [untitled]. Journal of the American Statistical Association, 73(361):221, March 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286559.

Kupper:1979:BRB

[Kup79] Lawrence L. Kupper. Book review: [untitled]. Journal of the American Statistical Association, 74(368):932–933, December 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286437.

Kurtz:1972:BRB

[Kur72] Thomas E. Kurtz. Book review: [untitled]. Journal of the American Statistical Association, 67(338):488–490, June 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284412.

Kulldorff:1973:ELS

[KV73] Gunnar Kulldorff and Kerstin Vannman. Estimation of the location and scale parameters of a Pareto distribution by linear functions of order statistics. Journal of the American Statistical Association, 68(341):218–227, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284174.

Kulldorff:1975:CEL

[KV75] Gunnar Kulldorff and Kerstin Vannman. Corrigenda: Estimation of the Location and Scale Parameters of a Pareto Distribution by Linear Functions of Order Statistics. Journal of the American Statistical Association, 70(350):494, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285867.

Kraft:1972:AEQ

[KvE72] Charles H. Kraft and Constance van Eeden. Asymptotic efficiencies of quick methods of computing efficient estimates based on ranks. Journal of the American Statistical Association, 67(337):199–202, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284723.

Ku:1971:AMC

[KVK71] Harry H. Ku, Ruth N. Varner, and S. Kullback. On the analysis of multidimensional contingency tables. Journal of the American Statistical Association, 66(333):55–64, March 1971. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284848.

Kiefer:1971:TTC

[KW71] James E. Kiefer and George H. Weiss. A truncated test for choosing the better of two binomial populations. Journal of the American Statistical Association, 66(336):867–871, December 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284243.

Karni:1974:CES

[KW74a] Edi Karni and Ishay Weissman. A consistent estimator of the slope in a regression model with errors in the variables. Journal of the American Statistical Association, 69(345):211–213, March 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285527.

Karni:1974:CCE

[KW74b] Edi Karni and Ishay Weissman. Corrigenda: a Consistent Estimator of the Slope in a Regression Model with Errors in the Variables. Journal of the American Statistical Association, 69(347):840, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286048.

Kiefer:1974:TVP

[KW74c] James E. Kiefer and George H. Weiss. Truncated version of a playthe-winner rule for choosing the better of two binomial populations. Journal of the American Statistical Association, 69(347):807–809, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286022.

Kruskal:1974:IPC

[KW74d] J. B. Kruskal and H. S. Witsenhausen. An inequality for positively correlated variables. Journal of the American Statistical Association, 69(346):540–542, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285693.

Kollerstrom:1979:SNC

[KW79] Julian Kollerstrom and G. Barrie Wetherill. SPRT’s for the normal correlation coefficient. Journal of the American Statistical Association, 74(368):815–821, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286405.

Kwerel:1975:MSB

[Kwe75] Seymour M. Kwerel. Most stringent bounds on aggregated probabilities of partially specified dependent probability systems. Journal of the American Statistical Association, 70(350):472–479, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285843.

Kalven:1972:AJC

[KZ72] Harry Kalven, Jr. and Hans Zeisel. The American jury: a critique: Rejoinder. Journal of the American Statistical Association, 67(340):779, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284636.

Laake:1979:PAS

[Laa79] Petter Laake. A predictive approach to subdomain estimation in finite populations. Journal of the American Statistical Association, 74(366):355–358, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286334.

Lachenbruch:1974:BRB

[Lac74] Peter A. Lachenbruch. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1048–1049, December 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286192.

Lachenbruch:1976:BRB

[Lac76] Peter A. Lachenbruch. Book review: [untitled]. Journal of the American Statistical Association, 71(353):249, March 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285789.

Lachenbruch:1977:BRB

[Lac77] P. A. Lachenbruch. Book review: [untitled]. Journal of the American Statistical Association, 72(359):690–691, September 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286249.

LaCava:1979:BRB

[LaC79] Jerry LaCava. Book review: [untitled]. Journal of the American Statistical Association, 74(365):249, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286771.

Ladd:1975:ABD

[Lad75] Daniel W. Ladd. An algorithm for the binomial distribution with dependent trials. Journal of the American Statistical Association, 70(350):333–340, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285817.

Laha:1975:BRB

[Lah75] R. G. Laha. Book review: [untitled]. Journal of the American Statistical Association, 70(351):728–729, September 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285965.

Laird:1978:NML

[Lai78] Nan Laird. Nonparametric maximum likelihood estimation of a mixing distribution. Journal of the American Statistical Association, 73(364):805–811, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286284.

LaMotte:1973:NNQ

[LaM73] Lynn Roy LaMotte. On non-negative quadratic unbiased estimation of variance components. Journal of the American Statistical Association, 68(343):728–730, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284808.

LaMotte:1977:ACL

[LaM77] Lynn Roy LaMotte. On admissibility and completeness of linear unbiased estimators in a general linear model. Journal of the American Statistical Association, 72(358):438–441, June 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286814.

Land:1973:SCL

[Lan73] Charles E. Land. Standard confidence limits for linear functions of the normal mean and variance. Journal of the American Statistical Association, 68(344):960–963, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284530.

Land:1974:CIE

[Lan74] Charles E. Land. Confidence interval estimation for means after data transformations to normality. Journal of the American Statistical Association, 69(347):795–802, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286020.

Land:1975:BRB

[Lan75a] Kenneth C. Land. Book review: [untitled]. Journal of the American Statistical Association, 70(349):256–257, March 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285422.

Lanke:1975:CUQ

[Lan75b] Jan Lanke. On the choice of the unrelated question in Simmons’ version of randomized response. Journal of the American Statistical Association, 70(349):80–83, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285380.

Land:1976:CCI

[Lan76] Charles E. Land. Corrigenda: Confidence Interval Estimation for Means After Data Transformations to Normality. Journal of the American Statistical Association, 71(353):255, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285804.

Landaw:1979:BRB

[Lan79] Elliot M. Landaw. Book review: [untitled]. Journal of the American Statistical Association, 74(368):931–932, December 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286436.

Lapan:1975:BRB

[Lap75] Harvey E. Lapan. Book review: [untitled]. Journal of the American Statistical Association, 70(350):489–490, June 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285854.

Larntz:1974:BRB

[Lar74] Kinley Larntz. Book review: [untitled]. Journal of the American Statistical Association, 69(346):578, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285705.

Larntz:1975:BRB

[Lar75] Kinley Larntz. Book review: [untitled]. Journal of the American Statistical Association, 70(351):730–731, September 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285968.

Larntz:1978:SSC

[Lar78] Kinley Larntz. Small-sample comparisons of exact levels for chisquared goodness-of-fit statistics. Journal of the American Statistical Association, 73(362):253–263, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286650.

Laughhunn:1970:DED

[Lau70] D. J. Laughhunn. Distributional effects in demand analysis: Observations and predictive tests. Journal of the American Statistical Association, 65(330):576–585, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284568.

Laub:1972:SAA

[Lau72a] P. Michael Laub. Some aspects of the aggregation problem in the dividend-earnings relationship. Journal of the American Statistical Association, 67(339):552–559, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284437.

Laurent:1972:EQL

[Lau72b] Andre Laurent. Effects of question length on reporting behavior in the survey interview. Journal of the American Statistical Association, 67(338):298–305, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284372.

Laurent:1974:CSD

[Lau74] Andre G. Laurent. On characterization of some distributions by truncation properties. Journal of the American Statistical Association, 69(347):823–827, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286026.

LaValle:1975:BRB

[LaV75] Irving H. LaValle. Book review: [untitled]. Journal of the American Statistical Association, 70(349):254–255, March 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285419.

Lavenhar:1979:BRB

[Lav79] Marvin A. Lavenhar. Book review: [untitled]. Journal of the American Statistical Association, 74(367):726–727, September 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287004.

Lawless:1973:CVU

[Law73] J. F. Lawless. Conditional versus unconditional confidence intervals for the parameters o the Weibull distribution. Journal of the American Statistical Association, 68(343):665–669, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284796.

Layard:1973:RLS

[Lay73] M. W. J. Layard. Robust large-sample tests for homogeneity of variances. Journal of the American Statistical Association, 68(341):195– 198, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284168.

Layard:1975:BRB

[Lay75] M. W. J. Layard. Book review: [untitled]. Journal of the American Statistical Association, 70(350):486–487, June 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285848.

Lochner:1972:BAT

[LB72] Robert H. Lochner and A. P. Basu. Bayesian analysis of the twosample problem with incomplete data. Journal of the American Statistical Association, 67(338):432–438, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284400.

Liu:1976:NDQ

[LC76] P. T. Liu and L. P. Chow. A new discrete quantitative randomized response model. Journal of the American Statistical Association, 71(353):72–73, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285733.

Liu:1975:URR

[LCM75] P. T. Liu, L. P. Chow, and W. H. Mosley. Use of the randomized response technique with a new randomizing device. Journal of the American Statistical Association, 70(350):329–332, June 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285816.

Leamer:1974:FMP

[Lea74] Edward E. Leamer. False models and post-data model construction. Journal of the American Statistical Association, 69(345):122– 131, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285510.

Leamer:1975:EYR

[Lea75] Edward E. Leamer. “explaining your results” as access-biased memory. Journal of the American Statistical Association, 70(349):88–93, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285382.

Leadbetter:1978:BRB

[Lea78a] M. R. Leadbetter. Book review: [untitled]. Journal of the American Statistical Association, 73(362):439–440, June 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286685.

Leamer:1978:RSS

[Lea78b] Edward E. Leamer. Regression selection strategies and revealed priors. Journal of the American Statistical Association, 73(363): 580–587, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286604.

Lebanon:1974:BRB

[Leb74] Alexander Lebanon. Book review: [untitled]. Journal of the American Statistical Association, 69(347):834–835, September 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286037.

Lee:1970:EDD

[Lee70a] Feng-Yao Lee. Estimation of dynamic demand relations from a time series of family budget data. Journal of the American Statistical Association, 65(330):586–597, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284569.

Lee:1970:MDF

[Lee70b] Maw Lin Lee. A model of the distribution of Federal expenditures among states. Journal of the American Statistical Association, 65(329):136–149, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283580.

Lee:1972:PBD

[Lee72] Kok-Huat Lee. Partially balanced designs for half sample replication method of variance estimation. Journal of the American Statistical Association, 67(338):324–334, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284377.

Lee:1973:UPB

[Lee73a] Kok-Huat Lee. Using partially balanced designs for the half sample replication method of variance estimation. Journal of the American Statistical Association, 68(343):612–614, September 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284787.

Lee:1973:VES

[Lee73b] Kok-Huat Lee. Variance estimation in stratified sampling. Journal of the American Statistical Association, 68(342):336–342, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284072.

Lee:1974:FBP

[Lee74] Ronald Demos Lee. Forecasting births in post-transition populations: Stochastic renewal with serially correlated fertility. Journal of the American Statistical Association, 69(347):607–617, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285990.

Lee:1975:NFP

[Lee75] Ronald Demos Lee. Natural fertility, population cycles and the spectral analysis of births and marriages. Journal of the American Statistical Association, 70(350):295–304, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285812.

Lee:1977:AVU

[Lee77a] S. Keith Lee. On the asymptotic variances of u terms in loglinear models of multidimensional contingency tables. Journal of the American Statistical Association, 72(358):412–419, June 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286809.

Lee:1977:MTR

[Lee77b] Young Jack Lee. Maximin tests of randomness against ordered alternatives: The multinomial distribution case. Journal of the American Statistical Association, 72(359):673–675, September 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286239.

Lee:1979:MTC

[Lee79] Kerry L. Lee. Multivariate tests for clusters. Journal of the American Statistical Association, 74(367):708–714, September 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286996.

Lehman:1977:BRB

[Leh77] Eugene-H. Lehman. Book review: [untitled]. Journal of the American Statistical Association, 72(357):232, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286952.

Lemon:1977:FOS

[Lem77] Glen H. Lemon. Factors for one-sided tolerance limits for balanced one-way-ANOVA random-effects model. Journal of the American Statistical Association, 72(359):676–680, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286240.

Lemmer:1978:RTD

[Lem78] Hermanus H. Lemmer. A robust test for dispersion. Journal of the American Statistical Association, 73(362):419–421, June 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286676.

Leonard:1977:BAS

[Leo77] Tom Leonard. A Bayesian approach to some multinomial estimation and pretesting problems. Journal of the American Statistical Association, 72(360):869–874, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286478.

Levenbach:1973:EHM

[Lev73] Hans Levenbach. The estimation of heteroscedasticity from a marginal likelihood function. Journal of the American Statistical Association, 68(342):436–439, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284094.

Lever:1975:BRB

[Lev75a] W. E. Lever. Book review: [untitled]. Journal of the American Statistical Association, 70(349):256, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285421.

Levy:1975:ECS

[Lev75b] Kenneth J. Levy. An empirical comparison of several multiple range tests for variances. Journal of the American Statistical Association, 70(349):180–183, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285398.

Levy:1977:OAS

[Lev77] Paul S. Levy. Optimum allocation in stratified random network sampling for estimating the prevalence of attributes in rare populations. Journal of the American Statistical Association, 72(360):758–763, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286456.

Littell:1971:AOF

[LF71] Ramon C. Littell and J. Leroy Folks. Asymptotic optimality of Fisher’s method of combining independent tests. Journal of the American Statistical Association, 66(336):802–806, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284230.

Littell:1972:CTM

[LF72] Ramon C. Littell and J. Leroy Folks. On the comparison of two methods of combining independent tests. Journal of the American Statistical Association, 67(337):223, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284731.

Littell:1973:AOF

[LF73] Ramon C. Littell and J. Leroy Folks. Asymptotic optimality of Fisher’s method of combining independent tests II. Journal of the American Statistical Association, 68(341):193–194, March 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284167.

Lee:1975:SPT

[LG75] Austin F. S. Lee and John Gurland. Size and power of tests for equality of means of two normal populations with unequal variances. Journal of the American Statistical Association, 70(352):933–941, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285460.

Li:1972:BRB

[Li72] C. C. Li. Book review: [untitled]. Journal of the American Statistical Association, 67(337):241–242, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284736.

Liew:1976:ICL

[Lie76] Chong Kiew Liew. Inequality constrained least-squares estimation. Journal of the American Statistical Association, 71(355):746–751, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285614.

Lilien:1975:BRB

[Lil75] Gary L. Lilien. Book review: [untitled]. Journal of the American Statistical Association, 70(349):258, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285426.

Lin:1971:EPD

[Lin71] Pi-Erh Lin. Estimation procedures for difference of means with missing data. Journal of the American Statistical Association, 66 (335):634–636, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283545.

Lindley:1972:SPP

[Lin72] D. V. Lindley. Some probability paradoxes in choice from among random alternatives: Comment. Journal of the American Statistical Association, 67(338):373–374, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284384.

Lin:1973:PTD

[Lin73a] Pi-Erh Lin. Procedures for testing the difference of means with incomplete data. Journal of the American Statistical Association, 68(343):699–703, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284801.

Ling:1973:BRB

[Lin73b] Robert F. Ling. Book review: [untitled]. Journal of the American Statistical Association, 68(344):1023, December 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284546.

Ling:1973:PTC

[Lin73c] Robert F. Ling. A probability theory of cluster analysis. Journal of the American Statistical Association, 68(341):159–164, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284161.

Ling:1974:CSA

[Lin74] Robert F. Ling. Comparison of several algorithms for computing sample means and variances. Journal of the American Statistical Association, 69(348):859–866, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286154;   https://www.tandfonline.com/
doi/abs/10.1080/01621459.1974.10480219;   https://www.tandfonline.
com/doi/pdf/10.1080/01621459.1974.10480219.

Lindley:1976:SIU

[Lin76a] D. V. Lindley. Strong inconsistency from uniform priors: Comment. Journal of the American Statistical Association, 71(353):120–121, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285747.

Lingner:1976:BRB

[Lin76b] Joan W. Lingner. Book review: [untitled]. Journal of the American Statistical Association, 71(355):777, September 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285638.

Lindley:1977:CCS

[Lin77a] Dennis V. Lindley. Conditional confidence statements and confidence estimators: Comment. Journal of the American Statistical Association, 72(360):819–820, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286466.

Ling:1977:BRB

[Lin77b] Robert F. Ling. Book review: [untitled]. Journal of the American Statistical Association, 72(360):930–931, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286501.

Lindsey:1978:BRB

[Lin78a] James Keith Lindsey. Book review: [untitled]. Journal of the American Statistical Association, 73(363):685, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286638.

Ling:1978:BRB

[Lin78b] Robert F. Ling. Book review: [untitled]. Journal of the American Statistical Association, 73(361):225, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286565.

Ling:1978:SAS

[Lin78c] Robert F. Ling. A study of the accuracy of some approximations for t, χ^{2}, and F tail probabilities. Journal of the American Statistical Association, 73(362):274–283, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286652.

Lindley:1979:NSD

[Lin79a] D. V. Lindley. Nonparametric statistical data modeling: Comment. Journal of the American Statistical Association, 74(365):127, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286738.

Ling:1979:BRBa

[Lin79b] Robert F. Ling. Book review: [untitled]. Journal of the American Statistical Association, 74(367):724–725, September 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287001.

Ling:1979:BRBb

[Lin79c] Robert F. Ling. Book review: [untitled]. Journal of the American Statistical Association, 74(367):737, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287021.

Ling:1979:BRBc

[Lin79d] Robert F. Ling. Book review: [untitled]. Journal of the American Statistical Association, 74(368):929–930, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286433.

Lissitz:1972:CSS

[Lis72] Robert W. Lissitz. Comparison of the small sample power of the chi-square and likelihood ratio tests of the assumptions for stochastic models. Journal of the American Statistical Association, 67 (339):574–577, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284440.

Little:1974:MSE

[Lit74a] R. E. Little. A mean square error comparison of certain median response estimates for the up-and-down method with small samples. Journal of the American Statistical Association, 69(345):202–206, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285525.

Little:1974:MSS

[Lit74b] R. E. Little. The up-and-down method for small samples with extreme value response distributions. Journal of the American Statistical Association, 69(347):803–806, September 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286021.

Little:1975:MSS

[Lit75] R. E. Little. The up-and-down method for small samples with two specimens “in series”. Journal of the American Statistical Association, 70(352):846–851, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285446.

Little:1978:CRM

[Lit78] Roderick J. A. Little. Consistent regression methods for discriminant analysis with incomplete data. Journal of the American Statistical Association, 73(362):319–322, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286659.

Liu:1970:DRS

[Liu70] Ben-Chieh Liu. Determinants of retail sales in large metropolitan areas, 1954 and 1963. Journal of the American Statistical Association, 65(332):1460–1473, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284330.

Ling:1976:PTC

[LK76] Robert F. Ling and George G. Killough. Probability tables for cluster analysis based on a theory of random graphs. Journal of the American Statistical Association, 71(354):293–300, June 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285300.

Li:1978:CVE

[LK78] Shou-Hua Li and Jerome H. Klotz. Components of variance estimation for the split-plot design. Journal of the American Statistical Association, 73(361):147–152, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286536.

Lempers:1971:ETS

[LL71] Fred B. Lempers and Adri S. Louter. An extension of the table of the Student distribution. Journal of the American Statistical Association, 66(335):503, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283516.

Lempers:1972:CNE

[LL72] Fred B. Lempers and Adri S. Louter. Corrigenda: a Note on an Extension of the Table of the Student Distribution. Journal of the American Statistical Association, 67(337):253, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284761.

Light:1971:AVC

[LM71] Richard J. Light and Barry H. Margolin. An analysis of variance for categorical data. Journal of the American Statistical Association, 66(335):534–544, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283520.

LaMotte:1978:ETP

[LM78] Lynn Roy LaMotte and Archer McWhorter, Jr. An exact test for the presence of random walk coefficients in a linear regression model. Journal of the American Statistical Association, 73(364):816–820, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286286.

Leone:1973:REB

[LMH73] Fred C. Leone and Effat Moussa-Hamouda. Relative efficiencies of ‘O–BLUE’ estimators in simple linear regression. Journal of the American Statistical Association, 68(344):953–959, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284529.

Leuthold:1970:FDH

[LMSW70] R. M. Leuthold, A. J. A. MacCormick, A. Schmitz, and D. G. Watts. Forecasting daily hog prices and quantities: a study of alternative forecasting techniques. Journal of the American Statistical Association, 65(329):90–107, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283577.

Long:1970:MGM

[Lon70a] Larry H. Long. On measuring geographic mobility. Journal of the American Statistical Association, 65(331):1195–1203, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284285.

Long:1970:FST

[Lon70b] Richard Long. Forecasting specific turning points. Journal of the American Statistical Association, 65(330):520–531, June 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284564.

Long:1973:NEM

[Lon73] Larry H. Long. New estimates of migration expectancy in the United States. Journal of the American Statistical Association, 68(341):37– 43, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284135.

Lord:1971:TTA

[Lor71] Frederic M. Lord. Tailored testing, an application of stochastic approximation. Journal of the American Statistical Association, 66 (336):707–711, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284216.

Lovell:1976:BRB

[Lov76] Michael Lovell. Book review: [untitled]. Journal of the American Statistical Association, 71(353):251–252, March 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285795.

Lovell:1978:BRB

[Lov78] Michael C. Lovell. Book review: [untitled]. Journal of the American Statistical Association, 73(361):228, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286571.

Lowerre:1974:SRB

[Low74] James Lowerre. Some relationships between BLUEs, WLSEs and SLSEs. Journal of the American Statistical Association, 69(345):223– 225, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285530.

Lowerre:1977:SSR

[Low77] James M. Lowerre. Some simplifying results on BLUEs. Journal of the American Statistical Association, 72(358):433–437, June 1977.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286813.

Loynes:1976:AOR

[Loy76] R. M. Loynes. Asymptotically optimal randomized response procedures. Journal of the American Statistical Association, 71(356): 924–928, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286862.

Lovell:1970:MRI

[LP70] Michael C. Lovell and Edward Prescott. Multiple regression with inequality constraints: Pretesting bias, hypothesis testing and efficiency. Journal of the American Statistical Association, 65(330):913– 925, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284597.

Leitch:1975:ESL

[LP75] R. A. Leitch and A. S. Paulson. Estimation of stable law parameters: Stock price behavior application. Journal of the American Statistical Association, 70(351):690–697, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285957.

Lieberman:1971:CII

[LR71] Gerald J. Lieberman and Sheldon M. Ross. Confidence intervals for independent exponential series systems. Journal of the American Statistical Association, 66(336):837–840, December 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284237.

Lianos:1972:ADP

[LR72] Theodore P. Lianos and Gordon C. Rausser. Approximate distribution of parameters in a distributed lag model. Journal of the American Statistical Association, 67(337):64–67, March 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284702.

Lancaster:1971:BRB

[LS71] H. O. Lancaster and Harry Smith, Jr. Book review: [untitled]. Journal of the American Statistical Association, 66(336):917–918, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284255.

Lave:1973:AAB

[LS73] Lester B. Lave and Eugene P. Seskin. An analysis of the association between U.S. mortality and air pollution. Journal of the American Statistical Association, 68(342):284–290, June 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284064.

Lin:1975:TEM

[LS75] Pi-Erh Lin and Lawrence E. Stivers. Testing for equality of means with incomplete data on one variable: a Monte Carlo study. Journal of the American Statistical Association, 70(349):190–193, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285401.

Lehmann:1977:FTM

[LS77a] E. L. Lehmann and Juliet Popper Shaffer. On a fundamental theorem in multiple comparisons. Journal of the American Statistical Association, 72(359):576–578, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286219.

Liu:1977:CMV

[LS77b] Lon-Mu Liu and Jerome Senturia. Computation of MINQUE variance component estimates. Journal of the American Statistical Association, 72(360):867–868, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286477.

Lu:1979:DNS

[LS79] Hollington T. Lu and Paul J. Smith. Distribution of the normal scores statistic for nonparametric one-way analysis of variance. Journal of the American Statistical Association, 74(367):715–722, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286997.

Locander:1976:IIM

[LSB76] William Locander, Seymour Sudman, and Norman Bradburn. An investigation of interview method, threat and response distortion. Journal of the American Statistical Association, 71(354):269–275, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285297.

Lo:1978:SSC

[LSW78] Louise C. Lo, Mark G. Simkin, and Reginald G. Worthley. A smallsample comparison of rank score tests for parallelism of several regression lines. Journal of the American Statistical Association, 73 (363):666–669, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286621.

Lucas:1972:BRB

[Luc72] Robert E. Lucas, Jr. Book review: [untitled]. Journal of the American Statistical Association, 67(337):245, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284741.

Luey:1971:NMC

[Lue71] Paul Luey. On naya and Morgan’s criterion of accuracy in trade recording. Journal of the American Statistical Association, 66 (336):763–765, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284226.

Larntz:1976:MMD

[LW76a] Kinley Larntz and Sanford Weisberg. Multiplicative models for dyad formation. Journal of the American Statistical Association, 71(354):455–461, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285333.

Lee:1976:DFT

[LW76b] Young Jack Lee and Douglas A. Wolfe. A distribution-free test for stochastic ordering. Journal of the American Statistical Association, 71(355):722–727, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285609.

Leysieffer:1976:RJO

[LW76c] Frederick W. Leysieffer and Stanley L. Warner. Respondent jeopardy and optimal designs in randomized response models. Journal of the American Statistical Association, 71(355):649–656, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285595.

Lyttkens:1972:BRB

[Lyt72] E. Lyttkens. Book review: [untitled]. Journal of the American Statistical Association, 67(340):960–961, December 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284682.

M:1972:BRB

[M.72] R. H. M. Book review: [untitled]. Journal of the American Statistical Association, 67(337):249, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284751.

M:1974:BRB

[M.74] D. S. M. Book review: [untitled]. Journal of the American Statistical Association, 69(345):282, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285565.

M:1975:BRBb

[M.75a] D. S. M. Book review: [untitled]. Journal of the American Statistical Association, 70(351):735, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285980.

M:1975:BRBc

[M.75b] D. S. M. Book review: [untitled]. Journal of the American Statistical Association, 70(351):735, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285981.

M:1975:BRBa

[M.75c] R. H. M. Book review: [untitled]. Journal of the American Statistical Association, 70(350):492, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285861.

Malkovich:1973:TMN

[MA73] J. F. Malkovich and A. A. Afifi. On tests for multivariate normality. Journal of the American Statistical Association, 68(341):176–179, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284163.

Meeden:1979:APT

[MA79] Glen Meeden and Barry C. Arnold. The admissibility of a preliminary test estimator when the loss incorporates a complexity cost. Journal of the American Statistical Association, 74(368):872–874, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286414.

Maddala:1973:BRB

[Mad73] G. S. Maddala. Book review: [untitled]. Journal of the American Statistical Association, 68(344):1028, December 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284551.

Madsen:1974:PTS

[Mad74] Richard Madsen. A procedure for truncating SPRT’s. Journal of the American Statistical Association, 69(346):403–410, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285669.

Magee:1972:BRB

[Mag72] Stephen P. Magee. Book review: [untitled]. Journal of the American Statistical Association, 67(339):712–713, September 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284476.

Mallios:1970:ASE

[Mal70] William S. Mallios. The analysis of structural effects in experimental design. Journal of the American Statistical Association, 65(330):808– 827, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284589.

Mallela:1972:NSC

[Mal72] Parthasaradhi Mallela. Necessary and sufficient conditions for MINQU-estimation of heteroskedastic variances in linear models. Journal of the American Statistical Association, 67(338):486–487, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284411.

Mallows:1978:PNA

[Mal78] Colin L. Mallows. Pseudorandom number assignment in statistically designed simulation and distribution sampling experiments: Comment. Journal of the American Statistical Association, 73(363):520, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286591;   http://www.jstor.org/stable/
pdfplus/2286591.pdf. See [SM78a].

Mandeville:1970:BRB

[Man70] Garrett K. Mandeville. Book review: [untitled]. Journal of the American Statistical Association, 65(332):1682–1683, December 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284353.

Mandeville:1971:BRB

[Man71] Garrett K. Mandeville. Book review: [untitled]. Journal of the American Statistical Association, 66(333):224, March 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284880.

Mandel:1972:BRB

[Man72] John Mandel. Book review: [untitled]. Journal of the American Statistical Association, 67(339):711–712, September 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284474.

Mandel:1974:BRB

[Man74a] John Mandel. Book review: [untitled]. Journal of the American Statistical Association, 69(347):828, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286027.

Mann:1974:SEO

[Man74b] Nancy R. Mann. Simplified expressions for obtaining approximately optimum system- reliability confidence bounds from exponential subsystem data. Journal of the American Statistical Association, 69(346):492–495, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285685.

Mantel:1974:SRU

[Man74c] Nathan Mantel. Some reasons for not using the Yates continuity correction on 2×2 contingency tables: Comment and a suggestion. Journal of the American Statistical Association, 69(346):378–380, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285663.

Mantell:1974:FAL

[Man74d] Edmund H. Mantell. Factors affecting labor productivity in post offices. Journal of the American Statistical Association, 69(346):303– 309, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285648.

Manski:1975:BRB

[Man75] C. F. Manski. Book review: [untitled]. Journal of the American Statistical Association, 70(352):964, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285477.

Marshall:1970:BRB

[Mar70a] K. T. Marshall. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1416–1418, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284319.

Martin:1970:BRB

[Mar70b] Margaret E. Martin. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1403–1404, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284309.

Marschak:1971:EIS

[Mar71] Jacob Marschak. Economics of information systems. Journal of the American Statistical Association, 66(333):192–219, March 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284873.

Marschak:1972:CEI

[Mar72] Jacob Marschak. Corrigenda: Economics of Information Systems. Journal of the American Statistical Association, 67(337):253, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284762.

Marascuilo:1973:BRB

[Mar73] Leonard A. Marascuilo. Book review: [untitled]. Journal of the American Statistical Association, 68(342):495–496, June 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284115.

Marcus:1976:PSD

[Mar76] Allan H. Marcus. Power sum distributions: An easier approach using the Wald distribution. Journal of the American Statistical Association, 71(353):237–238, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285776.

Markham:1978:BRB

[Mar78] Jesse W. Markham. Book review: [untitled]. Journal of the American Statistical Association, 73(362):448–449, June 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286708.

Masson:1972:BRB

[Mas72] Robert Tempest Masson. Book review: [untitled]. Journal of the American Statistical Association, 67(337):243–244, March 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284739.

Masnick:1979:BRB

[Mas79] George Masnick. Book review: [untitled]. Journal of the American Statistical Association, 74(368):923–924, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286427.

Maxwell:1973:ASM

[Max73] E. A. Maxwell. Application of the structural model to censored data. Journal of the American Statistical Association, 68(342):440–444, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284095.

Mayer:1973:ECC

[May73] Lawrence S. Mayer. Estimating a correlation coefficient when one variable is not directly observed. Journal of the American Statistical Association, 68(342):420–421, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284089.

Mazur:1972:URM

[Maz72] D. Peter Mazur. Using regression models to estimate the expectation of life for the U.S.S.R. Journal of the American Statistical Association, 67(337):31–36, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284698.

Mantel:1974:ERT

[MB74a] Nathan Mantel and David P. Byar. Evaluation of response-time data involving transient states: An illustration using heart-transplant data. Journal of the American Statistical Association, 69(345):81–86, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285503.

Morgenstern:1974:RBU

[MB74b] Richard D. Morgenstern and Nancy S. Barrett. The retrospective bias in unemployment reporting by sex, race and age. Journal of the American Statistical Association, 69(346):355–357, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285657.

McGee:1970:PR

[MC70] Victor E. McGee and Willard T. Carleton. Piecewise regression. Journal of the American Statistical Association, 65(331):1109–1124, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284278.

McCall:1971:MMI

[McC71] John J. McCall. A Markovian model of income dynamics. Journal of the American Statistical Association, 66(335):439–447, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283506.

McCarthy:1972:EDI

[McC72] Philip J. McCarthy. The effects of discarding inliars when binomial data are subject to classification errors. Journal of the American Statistical Association, 67(339):515–529, September 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284433.

McCabe:1973:ENT

[McC73a] George P. McCabe, Jr. Estimation of the number of terms in a sum. Journal of the American Statistical Association, 68(342):452– 456, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284097.

McCarthy:1973:BRB

[McC73b] Michael D. McCarthy. Book review: [untitled]. Journal of the American Statistical Association, 68(344):1020–1021, December 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284544.

McCabe:1974:BRB

[McC74a] G. P. McCabe. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1045, December 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286187.

McCall:1974:CMM

[McC74b] John J. McCall. Corrigenda: a Markovian Model of Income Dynamics. Journal of the American Statistical Association, 69(347):840, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286049.

McCarthy:1974:BRB

[McC74c] Michael D. McCarthy. Book review: [untitled]. Journal of the American Statistical Association, 69(345):275, March 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285546.

McCloskey:1974:BRB

[McC74d] Donald N. McCloskey. Book review: [untitled]. Journal of the American Statistical Association, 69(345):274–275, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285545.

McCabe:1975:BRB

[McC75a] G. P. McCabe. Book review: [untitled]. Journal of the American Statistical Association, 70(352):962, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285473.

McClung:1975:BRB

[McC75b] Nelson McClung. Book review: [untitled]. Journal of the American Statistical Association, 70(351):734–735, September 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285977.

McCabe:1976:BRB

[McC76a] George P. McCabe. Book review: [untitled]. Journal of the American Statistical Association, 71(354):530, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285357.

McCarthy:1976:UBH

[McC76b] Philip J. McCarthy. The use of balanced half-sample replication in cross-validation studies. Journal of the American Statistical Association, 71(355):596–604, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285588.

McCarthy:1978:BRB

[McC78a] James F. McCarthy. Book review: [untitled]. Journal of the American Statistical Association, 73(363):682, September 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286632.

McCarthy:1978:RMB

[McC78b] John L. McCarthy. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73(361): 86–89, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286522.

McClave:1978:EOA

[McC78c] James T. McClave. Estimating the order of autoregressive models: The max χ^{2} method. Journal of the American Statistical Association, 73(361):122–128, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286532.

McCullagh:1979:BRBa

[McC79a] Peter McCullagh. Book review: [untitled]. Journal of the American Statistical Association, 74(367):724, September 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2287000.

McCullagh:1979:BRBb

[McC79b] Peter McCullagh. Book review: [untitled]. Journal of the American Statistical Association, 74(368):934–935, December 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286440.

McDonald:1972:MET

[McD72] Lyman L. McDonald. A multivariate extension of Tukey’s one degree of freedom for non-additivity. Journal of the American Statistical Association, 67(339):674–675, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284465.

McDonald:1974:EEM

[McD74] James B. McDonald. On the existence of the exact moments of three LIML structural variance estimators for the case of two included endogenous variables. Journal of the American Statistical Association, 69(346):514–521, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285689.

McElroy:1976:OLS

[McE76] F. W. McElroy. Optimality of least squares in linear models with unknown error covariance matrix. Journal of the American Statistical Association, 71(354):374–377, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285317.

McGilchrist:1973:PDTa

[McG73a] Clyde A. McGilchrist. Post-data two sample tests of location. Journal of the American Statistical Association, 68(341):97–99, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284148.

McGilchrist:1973:PDTb

[McG73b] Clyde A. McGilchrist. Post-data two sample tests of location: Rejoinder. Journal of the American Statistical Association, 68(341):105– 106, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284152.

McGuire:1977:MTR

[McG77] Timothy W. McGuire. Measuring and testing relative advertising effectiveness with split-cable TV panel data. Journal of the American Statistical Association, 72(360):736–745, December 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286453.

McKay:1975:BRB

[McK75a] Michael D. McKay. Book review: [untitled]. Journal of the American Statistical Association, 70(351):733–734, September 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285974.

McKinlay:1975:DAO

[McK75b] Sonja M. McKinlay. The design and analysis of the observational study — a review. Journal of the American Statistical Association, 70(351):503–520, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285925.

McKinlay:1975:EBE

[McK75c] Sonja M. McKinlay. The effect of bias on estimators of relative risk for pair-matched and stratified samples. Journal of the American Statistical Association, 70(352):859–864, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285448.

McKinlay:1976:CDA

[McK76] Sonja McKinlay. Corrigenda: The Design and Analysis of the Observational Study — a Review. Journal of the American Statistical Association, 71(353):255, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285805.

McKean:1977:BRB

[McK77] Joseph McKean. Book review: [untitled]. Journal of the American Statistical Association, 72(359):691–692, September 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286251.

McLachlan:1975:IRP

[McL75] G. J. McLachlan. Iterative reclassification procedure for constructing an asymptotically optimal rule of allocation in discriminant analysis. Journal of the American Statistical Association, 70(350):365–369, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285824.

McLachlan:1977:ELD

[McL77] Geoffrey John McLachlan. Estimating the linear discriminant function from initial samples containing a small number of unclassified observations. Journal of the American Statistical Association, 72(358):403–406, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286807.

McLeod:1979:DRC

[McL79] A. Ian McLeod. Distribution of the residual cross-correlation in univariate ARMA time series models. Journal of the American Statistical Association, 74(368):849–855, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286410.

McNeil:1973:EAV

[McN73] Donald R. McNeil. Estimating an Author’s vocabulary. Journal of the American Statistical Association, 68(341):92–96, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284147.

Marks:1974:DFW

[MD74] Sidney Marks and Olive Jean Dunn. Discriminant functions when covariance matrices are unequal. Journal of the American Statistical Association, 69(346):555–559, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285696.

Meeden:1974:CPB

[Mee74a] Glen Meeden. Choosing a prior for a binomial testing problem with incomplete knowledge. Journal of the American Statistical Association, 69(347):740–743, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286011.

Meelis:1974:THK

[Mee74b] E. Meelis. Testing for homogeneity of k independent negative binomial distributed random variables. Journal of the American Statistical Association, 69(345):181–186, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285520.

Meeden:1978:BRB

[Mee78] Glen Meeden. Book review: [untitled]. Journal of the American Statistical Association, 73(362):440–441, June 1978. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286687.

Meeden:1979:CTP

[Mee79] Glen Meeden. Comparing two probability appraisers. Journal of the American Statistical Association, 74(366):299–302, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286326.

Mehta:1972:UIS

[Meh72] J. S. Mehta. On utilizing information from a second sample in estimating the scale parameter for a family of symmetric distributions. Journal of the American Statistical Association, 67(338):448–452, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284403.

Mehran:1973:VML

[Meh73] Farhad Mehran. Variance of the MVUE for the lognormal mean. Journal of the American Statistical Association, 68(343):726–727, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284807.

Mehran:1975:BGI

[Meh75] Farhad Mehran. Bounds on the Gini index based on observed points of the Lorenz curve. Journal of the American Statistical Association, 70(349):64–66, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285377.

Merewitz:1971:CRS

[Mer71] Leonard Merewitz. Costs and returns to scale in U.S. post offices. Journal of the American Statistical Association, 66(335):504–509, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2283517.

Merriwether:1973:SSP

[Mer73] Jacob D. Merriwether. Small sample properties of distributed lag estimators with misspecified lag structure. Journal of the American Statistical Association, 68(343):568–574, September 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284777.

Metcalf:1971:BRB

[Met71a] Charles E. Metcalf. Book review: [untitled]. Journal of the American Statistical Association, 66(333):221–222, March 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284875.

Metzler:1971:UTCa

[Met71b] Carl M. Metzler. Usefulness of the two-compartment open model in pharmacokinetics. Journal of the American Statistical Association, 66(333):49–53, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284845.

Metzler:1971:UTCb

[Met71c] Carl M. Metzler. Usefulness of the two-compartment open model in pharmacokinetics: Rejoinder. Journal of the American Statistical Association, 66(333):54, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284847.

Meyer:1972:BRB

[Mey72] Donald L. Meyer. Book review: [untitled]. Journal of the American Statistical Association, 67(340):955–956, December 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284673.

Meyer:1975:BRBa

[Mey75a] Donald L. Meyer. Book review: [untitled]. Journal of the American Statistical Association, 70(351):731–732, September 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285970.

Meyer:1975:BRBb

[Mey75b] Donald L. Meyer. Book review: [untitled]. Journal of the American Statistical Association, 70(351):735, September 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285978.

Milliken:1970:EGL

[MG70] George A. Milliken and Franklin A. Graybill. Extensions of the general linear hypothesis model. Journal of the American Statistical Association, 65(330):797–807, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284588.

Mann:1974:CSA

[MG74] Nancy R. Mann and Frank E. Grubbs. Chi-square approximations for exponential parameters, prediction intervals and beta percentiles. Journal of the American Statistical Association, 69(347):654–661, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285996.

McDonald:1975:MCE

[MG75] Gary C. McDonald and Diane I. Galarneau. A Monte Carlo evaluation of some Ridge-type estimators. Journal of the American Statistical Association, 70(350):407–416, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285832.

Mendoza:1978:ECM

[MI78] Gaston Mendoza and Boris Iglewicz. An extension of Colton’s model for comparing two medical treatments. Journal of the American Statistical Association, 73(363):646–649, September 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286617.

Michelini:1972:CPS

[Mic72] C. Michelini. Convergence pattern of the scoring method in estimating parameters of a log-normal function. Journal of the American Statistical Association, 67(338):319–323, June 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284376.

Mickelsen:1976:BRB

[Mic76] Richard Mickelsen. Book review: [untitled]. Journal of the American Statistical Association, 71(356):1008, December 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286890.

Mielke:1971:BRB

[Mie71] Paul W. Mielke, Jr. Book review: [untitled]. Journal of the American Statistical Association, 66(336):923, December 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284265.

Mielke:1972:ABT

[Mie72] Paul W. Mielke, Jr. Asymptotic behavior of two-sample tests based on powers of ranks for detecting scale and location alternatives. Journal of the American Statistical Association, 67(340):850–854, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284648.

Miettinen:1974:SRU

[Mie74] Olli S. Miettinen. Some reasons for not using the Yates continuity correction on 2 × 2 contingency tables: Comment. Journal of the American Statistical Association, 69(346):380–382, June 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285664.

Mielke:1979:FEW

[Mie79] Paul W. Mielke, Jr. Field experimentation in weather modification: Comment. Journal of the American Statistical Association, 74(365): 87–88, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286729.

Mike:1971:ERS

[Mik71] Valerie Mike. Efficiency-robust systematic linear estimators of location. Journal of the American Statistical Association, 66(335): 594–601, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283536.

Mikhail:1972:BTS

[Mik72a] W. M. Mikhail. The bias of the two-stage least squares estimator. Journal of the American Statistical Association, 67(339):625–627, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284453.

Mikhail:1972:SSS

[Mik72b] W. M. Mikhail. Simulating the small-sample properties of econometric estimators. Journal of the American Statistical Association, 67(339):620–624, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284452.

Mikhail:1975:CMC

[Mik75] W. M. Mikhail. A comparative Monte Carlo study of the properties of econometric estimators. Journal of the American Statistical Association, 70(349):94–104, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285383.

Miller:1970:BRB

[Mil70a] M. Clinton Miller. Book review: [untitled]. Journal of the American Statistical Association, 65(329):475, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283622.

Miller:1970:SSR

[Mil70b] Rupert G. Miller, Jr. Sequential signed-rank test. Journal of the American Statistical Association, 65(332):1554–1561, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284336.

Miller:1972:CSS

[Mil72] Rupert G. Miller, Jr. Corrigenda: a Sequential Signed-Rank Test. Journal of the American Statistical Association, 67(337):253, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284763.

Miller:1975:BRB

[Mil75] Robert B. Miller. Book review: [untitled]. Journal of the American Statistical Association, 70(352):956–957, December 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285466.

Miller:1976:RDW

[Mil76] Ann R. Miller. Retrospective data on work status in the 1970 census of population: An attempt at evaluation. Journal of the American Statistical Association, 71(354):286–292, June 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285299.

Miller:1977:DMC

[Mil77] Rupert G. Miller, Jr. Developments in multiple comparisons 1966–1976. Journal of the American Statistical Association, 72 (360):779–788, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286459.

Miller:1979:BRBa

[Mil79] Robert B. Miller. Book review: [untitled]. Journal of the American Statistical Association, 74(368):927, December 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286429.

Minton:1971:BRB

[Min71] Paul D. Minton. Book review: [untitled]. Journal of the American Statistical Association, 66(334):428, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283959.

Minton:1972:VES

[Min72] George Minton. Verification error in single sampling inspection plans for processing survey data. Journal of the American Statistical Association, 67(337):46–54, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284700.

Mirer:1973:BRB

[Mir73] Thad W. Mirer. Book review: [untitled]. Journal of the American Statistical Association, 68(343):753, September 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284823.

Misra:1972:RBP

[Mis72] P. N. Misra. Relation between Pearsonian coefficients of distributions of least squares estimators and the disturbance term. Journal of the American Statistical Association, 67(339):662–663, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284462.

Mitchel:1971:BRB

[Mit71] T. J. Mitchel. Book review: [untitled]. Journal of the American Statistical Association, 66(333):226, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284884.

Maritz:1978:NEV

[MJ78] J. S. Maritz and R. G. Jarrett. A note on estimating the variance of the sample median. Journal of the American Statistical Association, 73(361):194–196, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286545.

McDonald:1979:ASP

[MJ79] James B. McDonald and Bartell C. Jensen. An analysis of some properties of alternative measures of income inequality based on the gamma distribution function. Journal of the American Statistical Association, 74(368):856–860, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286411.

Margolin:1974:AVC

[ML74] Barry H. Margolin and Richard J. Light. An analysis of variance for categorical data, II: Small sample comparisons with chi square and other competitors. Journal of the American Statistical Association, 69(347):755–764, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286014.

Mantel:1971:PCM

[MM71a] Nathan Mantel and Max Myers. Problems of convergence of maximum likelihood iterative procedures in multiparameter situations. Journal of the American Statistical Association, 66(335):484–491, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2283512.

Mullet:1971:NME

[MM71b] Gary M. Mullet and Tracy W. Murray. A new method for examining rounding error in least-squares regression computer programs. Journal of the American Statistical Association, 66(335):496–498, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2283514.

Messenger:1972:MST

[MM72] Robert Messenger and Lewis Mandell. A modal search technique for predictibe nominal scale multivariate analys. Journal of the American Statistical Association, 67(340):768–772, December 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284634.

Maddala:1973:CSA

[MM73] G. S. Maddala and T. D. Mount. A comparative study of alternative estimators for variance components models used in econometric applications. Journal of the American Statistical Association, 68(342):324–328, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284070.

Mariano:1979:NDF

[MM79] Roberto S. Mariano and James B. McDonald. A note on the distribution functions of LIML and 2SLS structural coefficient in the exactly identified case. Journal of the American Statistical Association, 74(368):847–848, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286409.

Morrison:1971:ENP

[MMS71] Donald G. Morrison, William F. Massy, and Fred N. Silverman. The effect of nonhomogeneous populations on Markov steady-state probabilities. Journal of the American Statistical Association, 66(334):268– 274, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2283920.

Moors:1971:OUQ

[Moo71] J. J. A. Moors. Optimization of the unrelated question randomized response model. Journal of the American Statistical Association, 66(335):627–629, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283543.

Moore:1972:BRBb

[Moo72a] David S. Moore. Book review: [untitled]. Journal of the American Statistical Association, 67(338):492–493, June 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284416.

Moore:1972:BRBa

[Moo72b] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 67(338):491, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284414.

Moore:1972:BRBc

[Moo72c] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 67(338):498, June 1972. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284426.

Moore:1973:BRB

[Moo73a] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 68(342):494, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284113.

Moore:1973:EFD

[Moo73b] Dan H. Moore, II. Evaluation of five discrimination procedures for binary variables. Journal of the American Statistical Association, 68(342):399–404, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284083.

Moore:1974:BRBb

[Moo74a] David S. Moore. Book review: [untitled]. Journal of the American Statistical Association, 69(346):575–576, June 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285701.

Moore:1974:BRBa

[Moo74b] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 69(345):281, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285559.

Moore:1975:BRBc

[Moo75a] David S. Moore. Book review: [untitled]. Journal of the American Statistical Association, 70(352):967, December 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285486.

Moore:1975:BRBa

[Moo75b] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 70(352):966, December 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285484.

Moore:1975:BRBb

[Moo75c] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 70(352):966–967, December 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285485.

Moore:1976:BRBb

[Moo76a] David S. Moore. Book review: [untitled]. Journal of the American Statistical Association, 71(355):773, September 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285629.

Moore:1976:BRBa

[Moo76b] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 71(355):772, September 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285627.

Moore:1976:BRBc

[Moo76c] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 71(355):773–774, September 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285631.

Moore:1977:GIW

[Moo77a] David S. Moore. Generalized inverses, Wald’s method, and the construction of chi-squared tests of fit. Journal of the American Statistical Association, 72(357):131–137, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286920.

Moore:1977:BRB

[Moo77b] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 72(357):232–233, March 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286953.

Moore:1978:BRBa

[Moo78a] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 73(362):446, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286702.

Moore:1978:BRBb

[Moo78b] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 73(362):446, June 1978. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286703.

Moore:1978:BRBc

[Moo78c] S. Craig Moore. Book review: [untitled]. Journal of the American Statistical Association, 73(362):447, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286706.

Moore:1979:BRB

[Moo79] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 74(366):506, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286371.

Morton:1970:BRB

[Mor70] J. E. Morton. Book review: [untitled]. Journal of the American Statistical Association, 65(330):999–1001, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284611.

Morrison:1971:BRB

[Mor71a] Donald F. Morrison. Book review: [untitled]. Journal of the American Statistical Association, 66(335):653–654, September 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283552.

Morrison:1971:DLF

[Mor71b] Donald F. Morrison. The distribution of linear functions of independent F variates. Journal of the American Statistical Association, 66(334):383–385, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283940.

Morrison:1971:EVM

[Mor71c] Donald F. Morrison. Expectations and variances of maximum likelihood estimates of the multivariate normal distribution parameters with missing data. Journal of the American Statistical Association, 66(335):602–604, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283537.

Morrison:1972:UBC

[Mor72] Donald G. Morrison. Upper bounds for correlations between binary outcomes and probabilistic predictions. Journal of the American Statistical Association, 67(337):68–70, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284703.

Morrison:1973:BRB

[Mor73] Donald F. Morrison. Book review: [untitled]. Journal of the American Statistical Association, 68(344):1019–1020, December 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284543.

Moreno:1977:BRB

[Mor77] Jerry Moreno. Book review: [untitled]. Journal of the American Statistical Association, 72(357):236–237, March 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286958.

Morimune:1978:ILI

[Mor78] Kimio Morimune. Improving the limited information maximum likelihood estimator when the disturbances are small. Journal of the American Statistical Association, 73(364):867–871, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286296.

Mosimann:1970:SAS

[Mos70a] James E. Mosimann. Size allometry: Size and shape variables with characterizations of the lognormal and generalized gamma distributions. Journal of the American Statistical Association, 65(330):930– 945, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284599.

Mosteller:1970:CFS

[Mos70b] Frederick Mosteller. Collegiate football scores, U.S.A. Journal of the American Statistical Association, 65(329):35–48, March 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283573.

Mosteller:1979:FEW

[Mos79] Frederick Mosteller. Field experimentation in weather modification: Comment. Journal of the American Statistical Association, 74(365):

88–90, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286730.

Mitchell:1971:CCI

[MP71] Ann F. S. Mitchell and Clive D. Payne. A conservative confidence interval for a likelihood ratio. Journal of the American Statistical Association, 66(336):861–866, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284242.

Murthy:1970:PIS

[MR70] M. N. Murthy and A. S. Roy. A problem in integration of surveys. A case study. Journal of the American Statistical Association, 65(329):123–135, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283579.

Marks:1979:EPD

[MR79] Ronald G. Marks and P. V. Rao. An estimation procedure for data containing outliers with a one-directional shift in the mean. Journal of the American Statistical Association, 74(367):614–620, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286978.

Mehra:1970:NET

[MS70] K. L. Mehra and G. E. J. Smith. On nonparametric estimation and testing for interactions in factorial experiments. Journal of the American Statistical Association, 65(331):1283–1296, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284295.

Mehta:1971:EMS

[MS71] J. S. Mehta and R. Srinivasan. Estimation of the mean by shrinkage to a point. Journal of the American Statistical Association, 66(333): 86–90, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284853.

Mariano:1972:EFS

[MS72a] Roberto S. Mariano and Takamitsu Sawa. The exact finite-sample distribution of the limited-information maximum likelihood estimator in the case of two included endogenous variables. Journal of the American Statistical Association, 67(337):159–163, March 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284715.

Murty:1972:ODE

[MS72b] V. N. Murty and W. J. Studden. Optimal designs for estimating the slope of a polynomial regression. Journal of the American Statistical Association, 67(340):869–873, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284652.

Mehta:1973:BAB

[MS73] J. S. Mehta and P. A. V. B. Swamy. Bayesian analysis of a bivariate normal distribution with incomplete observations. Journal of the American Statistical Association, 68(344):922–927, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284523.

Mehta:1974:EFS

[MS74] J. S. Mehta and P. A. V. B. Swamy. The exact finite sample distribution of Theil’s compatibility test statistic and its application. Journal of the American Statistical Association, 69(345):154–158, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285515.

Mardia:1976:STB

[MS76a] Kanti V. Mardia and Barrie D. Spurr. On some tests for the bivariate two-sample location problem. Journal of the American Statistical Association, 71(356):994–995, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286876.

Mehta:1976:FER

[MS76b] J. S. Mehta and P. A. V. B. Swamy. Further evidence on the relative efficiencies of Zellner’s seemingly unrelated regressions estimator. Journal of the American Statistical Association, 71(355):634–639, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285593.

Monti:1976:LOC

[MS76c] Katherine L. Monti and Pranab K. Sen. The locally optimal combination of independent test statistics. Journal of the American Statistical Association, 71(356):903–911, December 1976. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286859.

Mudholkar:1976:UPM

[MS76d] Govind S. Mudholkar and Perla Subbaiah. Unequal precision multiple comparisons for randomized block designs under nonstandard conditions. Journal of the American Statistical Association, 71(354):429– 434, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285329.

Maag:1973:GFT

[MSD73] Urs R. Maag, Franz Streit, and Pierre A. Drouilly. Goodness-of-fit tests for grouped data. Journal of the American Statistical Association, 68(342):462–465, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284099.

Mayer:1974:UIE

[MSW74] Lawrence S. Mayer, Jagbir Singh, and Thomas A. Willke. Utilizing intial estimates in estimating the coefficients in a linear model. Journal of the American Statistical Association, 69(345):219–222, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285529.

Mendoza:1976:NSC

[MTC76] Jorge L. Mendoza, Larry E. Toothaker, and Bradford R. Crain. Necessary and sufficient conditions for F ratios in the L×J ×K factorial design with two repeated factors. Journal of the American Statistical Association, 71(356):992–993, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286875.

Muhsam:1973:BRB

[Muh73] H. V. Muhsam. Book review: [untitled]. Journal of the American Statistical Association, 68(343):753, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284824.

Mukhopadhyay:1976:SEL

[Muk76] Nitis Mukhopadhyay. Sequential estimation of a linear function of means of three normal populations. Journal of the American Statistical Association, 71(353):149–153, March 1976. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285759.

Muller:1978:RMBa

[Mul78a] Mervin E. Muller. A review of the manuals for BMDP and SPSS. Journal of the American Statistical Association, 73(361):71– 80, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286518.

Muller:1978:RMBb

[Mul78b] Mervin E. Muller. A review of the manuals for BMDP and SPSS: Rejoinder. Journal of the American Statistical Association, 73(361): 95–98, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286528.

Munch:1977:BRB

[Mun77] Patricia Munch. Book review: [untitled]. Journal of the American Statistical Association, 72(360):929–930, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286500.

Murray:1970:CIV

[Mur70] Barbara B. Murray. Coefficient of interarea variation as a measure of spatial income inequality. Journal of the American Statistical Association, 65(330):598–601, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284570.

Murty:1971:MD

[Mur71] V. N. Murty. Minimax designs. Journal of the American Statistical Association, 66(334):319–320, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283929.

Murray:1972:EEC

[Mur72] Tracy W. Murray. An empirical examination of the classical assumptions concerning errors in data. Journal of the American Statistical Association, 67(339):530–537, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284434.

Muthen:1979:SPM

[Mut79] Bengt Muthen. A structural probit model with latent variables. Journal of the American Statistical Association, 74(368):807–811, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286403.

Munro:1970:EBO

[MW70] A. H. Munro and R. A. J. Wixley. Estimators based on order statistics of small samples from a three-parameter lognormal distribution. Journal of the American Statistical Association, 65(329):212–225, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283588.

Miller:1979:BRBb

[MW79] Thomas W. Miller and Stanley S. Wasserman. Book review: [untitled]. Journal of the American Statistical Association, 74(368): 936–937, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286444.

Mayer:1976:ESR

[MY76] Lawrence S. Mayer and Mary Sue Younger. Estimation of standardized regression coefficients. Journal of the American Statistical Association, 71(353):154–157, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285760.

Maronna:1978:BTD

[MY78] Ricardo Maronna and Victor J. Yohai. A bivariate test for the detection of a systematic change in mean. Journal of the American Statistical Association, 73(363):640–645, September 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286616.

Mann:1976:BRB

[MYH76] Nancy R. Mann, A. J. Yorgiadis, and H. Himelblau. Book review: [untitled]. Journal of the American Statistical Association, 71(355):774–775, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285633.

Nagatani:1979:BRB

[Nag79] Keizo Nagatani. Book review: [untitled]. Journal of the American Statistical Association, 74(368):933–934, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286439.

Nash:1979:BRB

[Nas79] Stanley W. Nash. Book review: [untitled]. Journal of the American Statistical Association, 74(365):254–255, March 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286781.

Nathan:1972:APT

[Nat72] Gad Nathan. On the asymptotic power of tests for independence in contingency tables from stratified samples. Journal of the American Statistical Association, 67(340):917–920, December 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284662.

Nathan:1976:ESR

[Nat76] Gad Nathan. An empirical study of response and sampling errors for multiplicity estimates with different counting rules. Journal of the American Statistical Association, 71(356):808–815, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286843.

Naus:1973:PSD

[Nau73] Joseph I. Naus. Power sum distributions. Journal of the American Statistical Association, 68(343):740–742, September 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284811.

Naus:1974:PGB

[Nau74] Joseph Naus. Probabilities for a generalized birthday problem. Journal of the American Statistical Association, 69(347):810–815, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286023.

Neuhardt:1971:SMF

[NB71] John B. Neuhardt and Hugh E. Bradley. On the selection of multifactor experimental arrangements with resource constraints. Journal of the American Statistical Association, 66(335):618–621, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283541.

Neuhardt:1972:CSM

[NB72] John B. Neuhardt and Hugh E. Bradley. Corrigenda: On the Selection of Multi-Factor Experimental Arrangements with Resource Constraints. Journal of the American Statistical Association, 67(337): 253, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284764.

Neuhardt:1973:CRS

[NBH73] John B. Neuhardt, Hugh E. Bradley, and Richard W. Henning. Computational results in selecting multi-factor experimental arrangements. Journal of the American Statistical Association, 68 (343):608–611, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284786.

Neave:1972:CLW

[Nea72] Henry R. Neave. A comparison of lag window generators. Journal of the American Statistical Association, 67(337):152–158, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284714.

Neave:1975:QST

[Nea75] Henry R. Neave. A quick and simple technique for general slippage problems. Journal of the American Statistical Association, 70(351):721–726, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285963.

Nelson:1970:TMT

[Nel70] Charles R. Nelson. Testing a model of the term structure of interest rates by simulation of market forecasts. Journal of the American Statistical Association, 65(331):1163–1179, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284283.

Nelder:1978:RMB

[Nel78] J. A. Nelder. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73(361):89–90, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286523.

Newbold:1976:BRB

[New76] Paul Newbold. Book review: [untitled]. Journal of the American Statistical Association, 71(353):247, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285785.

Newman:1977:MM

[New77] Peter Newman. Malaria and mortality. Journal of the American Statistical Association, 72(358):257–263, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286786.

Newbold:1979:BRB

[New79] Paul Newbold. Book review: [untitled]. Journal of the American Statistical Association, 74(367):734–735, September 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2287015.

Neyman:1979:FEW

[Ney79] Jerzy Neyman. Field experimentation in weather modification: Comment. Journal of the American Statistical Association, 74(365):90–94, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286731.

Nie:1978:RMB

[NFJN78] Norman H. Nie, Jonathan B. Fry, Jean G. Jenkins, and Marija J. Norusis. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73(361):82–84, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286520.

Naus:1972:PMI

[NJM72] Joseph I. Naus, Thomas G. Johnson, and Ramiro Montalvo. A probabilistic model for identifying errors in data editing. Journal of the American Statistical Association, 67(340):943–950, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284667.

Neter:1977:BSE

[NL77] John Neter and James K. Loebbecke. On the behavior of statistical estimators when sampling accounting populations. Journal of the American Statistical Association, 72(359):501–507, September 1977.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286207.

Novick:1979:FSA

[NL79] Melvin R. Novick and Dennis V. Lindley. Fixed-state assessment of utility functions. Journal of the American Statistical Association, 74(366):306–311, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286328.

Neter:1970:ACC

[NM70] John Neter and E. Scott Maynes. On the appropriateness of the correlation coefficient with a 0, 1 dependent variable. Journal of the American Statistical Association, 65(330):501–509, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284562.

Naya:1971:NMC

[NMC71] Seiji Naya, Theodore Morgan, and Joseph A. Colosi. On naya and Morgan’s criterion of accuracy in trade recording: Rejoinder. Journal of the American Statistical Association, 66(336):765, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284227.

Nakamura:1976:TRB

[NNO76] Alice Orcutt Nakamura, Masao Nakamura, and Guy H. Orcutt. Testing for relationships between time series. Journal of the American Statistical Association, 71(353):214–222, March 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285772.

Noether:1970:BRB

[Noe70] Gottfried E. Noether. Book review: [untitled]. Journal of the American Statistical Association, 65(332):1680, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284351.

Noether:1973:SSD

[Noe73] Gottfried E. Noether. Some simple distribution-free confidence intervals for the center of a symmetric distribution. Journal of the American Statistical Association, 68(343):716–719, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284805.

Nold:1973:BRB

[Nol73] F. C. Nold. Book review: [untitled]. Journal of the American Statistical Association, 68(343):755, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284829.

Nold:1974:BRB

[Nol74] F. C. Nold. Book review: [untitled]. Journal of the American Statistical Association, 69(346):583, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285713.

Northam:1971:BRB

[Nor71] Jack I. Northam. Book review: [untitled]. Journal of the American Statistical Association, 66(335):655–656, September 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283555.

Nordbrock:1976:IPW

[Nor76] Earl Nordbrock. An improved play-the-winner sampling procedure for selecting the better of two binomial populations. Journal of the American Statistical Association, 71(353):137–139, March 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285755.

Nagendra:1971:OSS

[NR71] Y. Nagendra and G. Rai. Optimum sample size and sampling interval for controlling the mean of non-normal variables. Journal of the American Statistical Association, 66(335):637–640, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283546.

Nelson:1974:EPD

[NS74] Charles R. Nelson and G. William Schwert. Estimating the parameters of a distributed lag model from cross-section data: The case of hospital admissions and discharges. Journal of the American Statistical Association, 69(347):627–633, September 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285992.

O:1970:BRBe

[O.70a] D. B. O. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1421, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284323.

O:1970:BRBa

[O.70b] I. O. Book review: [untitled]. Journal of the American Statistical Association, 65(329):476–477, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283624.

O:1970:BRBb

[O.70c] I. O. Book review: [untitled]. Journal of the American Statistical Association, 65(329):477, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283625.

O:1970:BRBc

[O.70d] I. O. Book review: [untitled]. Journal of the American Statistical Association, 65(329):477, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283626.

O:1970:BRBd

[O.70e] I. O. Book review: [untitled]. Journal of the American Statistical Association, 65(330):1009, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284618.

O:1971:BRB

[O.71] D. B. O. Book review: [untitled]. Journal of the American Statistical Association, 66(333):227, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284888.

Oakes:1974:BRB

[Oak74] David Oakes. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1051, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286196.

OBrien:1979:GAM

[O’B79] Ralph G. O’Brien. A general ANOVA method for robust tests of additive models for variances. Journal of the American Statistical Association, 74(368):877–880, December 1979. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286416.

Obenchain:1975:ROO

[Obe75] R. L. Obenchain. Residual optimality: Ordinary vs. weighted vs. biased least squares. Journal of the American Statistical Association, 70(350):375–379, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285826.

Obremski:1979:CSP

[Obr79] Thomas E. Obremski. On the choice of starting point in stochastic approximation. Journal of the American Statistical Association, 74(367):700–702, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286994.

Odeh:1971:BRB

[Ode71] R. E. Odeh. Book review: [untitled]. Journal of the American Statistical Association, 66(334):428, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283960.

Odell:1979:BRB

[Ode79] Patrick L. Odell. Book review: [untitled]. Journal of the American Statistical Association, 74(365):253, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286779.

Odoroff:1970:CML

[Odo70] Charles L. Odoroff. A comparison of minimum logit chi-square estimation and maximum likelihood estimation in 2—times 2—times 2 and 3—times 2—times 2 contingency tables: Tests for interaction. Journal of the American Statistical Association, 65(332):1617–1631, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284345.

Officer:1972:DSR

[Off72] R. R. Officer. The distribution of stock returns. Journal of the American Statistical Association, 67(340):807–812, December 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284641.

Ofosu:1974:OSS

[Ofo74] J. B. Ofosu. Optimum sample size for a problem in choosing the population with the largest mean: Some comments on Somerville’s paper. Journal of the American Statistical Association, 69(345):270, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285542.

Ofosu:1975:TSM

[Ofo75] John B. Ofosu. A two-stage minimax procedure for selecting the normal population with the smallest variance. Journal of the American Statistical Association, 70(349):171–174, March 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285396.

Ott:1976:SCP

[OK76] Jurg Ott and Richard A. Kronmal. Some classification procedures for multivariate binary data using orthogonal functions. Journal of the American Statistical Association, 71(354):391–399, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285321.

Odell:1971:BLR

[OL71] P. L. Odell and T. O. Lewis. Best linear recursive estimation. Journal of the American Statistical Association, 66(336):893–896, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284249.

Olshen:1973:CLT

[Ols73] Richard A. Olshen. The conditional level of the F -test. Journal of the American Statistical Association, 68(343):692–698, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284800.

Olson:1974:CRS

[Ols74] Chester L. Olson. Comparative robustness of six tests in multivariate analysis of variance. Journal of the American Statistical Association, 69(348):894–908, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286159.

Olshen:1977:NRM

[Ols77] Richard A. Olshen. A note on a reformulation of the S-method of multiple comparison: Comment. Journal of the American Statistical Association, 72(357):144–146, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286923.

ONeill:1978:NDU

[O’N78] Terence J. O’Neill. Normal discrimination with unclassified observations. Journal of the American Statistical Association, 73 (364):821–826, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286287.

Osterburg:1977:DMF

[OPRS77] James W. Osterburg, T. Parthasarathy, T. E. S. Raghavan, and Stanley L. Sclove. Development of a mathematical formula for the calculation of fingerprint probabilities based on individual characteristics. Journal of the American Statistical Association, 72(360):772–778, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286458.

Ord:1975:EMM

[Ord75] Keith Ord. Estimation methods for models of spatial interaction. Journal of the American Statistical Association, 70(349):120–126, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285387.

Orey:1979:BRB

[Ore79] Steven Orey. Book review: [untitled]. Journal of the American Statistical Association, 74(367):725, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287002.

Orkin:1972:BRB

[Ork72] Michael Orkin. Book review: [untitled]. Journal of the American Statistical Association, 67(340):963, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284685.

Orkin:1975:BRB

[Ork75] Michael Orkin. Book review: [untitled]. Journal of the American Statistical Association, 70(350):489, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285852.

Ostry:1970:DDA

[OS70] Sylvia Ostry and Alan Sunter. Definitional and design aspects of the Canadian Job Vacancy Survey. Journal of the American Statistical Association, 65(331):1059–1070, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284275.

Otten:1973:NSR

[Ott73] A. Otten. Note on the Spearman rank correlation coefficient. Journal of the American Statistical Association, 68(343):585, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284781.

ONeill:1972:SRR

[OW72] Brian O’Neill and W. T. Wells. Some recent results in lognormal parameter estimation using grouped and ungrouped data. Journal of the American Statistical Association, 67(337):76–80, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284705.

Owen:1971:BRB

[Owe71] Donald B. Owen. Book review: [untitled]. Journal of the American Statistical Association, 66(333):227, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284886.

Owen:1975:BSP

[Owe75] Roger J. Owen. A Bayesian sequential procedure for quantal response in the context of adaptive mental testing. Journal of the American Statistical Association, 70(350):351–356, June 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285821.

P:1972:BRB

[P.72] B. S. P. Book review: [untitled]. Journal of the American Statistical Association, 67(340):964, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284689.

P:1975:BRBa

[P.75a] B. S. P. Book review: [untitled]. Journal of the American Statistical Association, 70(350):492, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285862.

P:1975:BRBb

[P.75b] B. S. P. Book review: [untitled]. Journal of the American Statistical Association, 70(351):735, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285979.

P:1975:BRBc

[P.75c] B. S. P. Book review: [untitled]. Journal of the American Statistical Association, 70(351):735–736, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285982.

Pillai:1976:BRB

[PA76] K. C. S. Pillai and Corwin L. Atwood. Book review: [untitled]. Journal of the American Statistical Association, 71(353):248–249, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285787.

Pillai:1970:PCT

[PAA70] K. C. Sreedharan Pillai and Sabri Al-Ani. Power comparisons of tests of equality of two covariance matrices based on individual characteristic roots. Journal of the American Statistical Association, 65(329):438–446, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283608.

Pack:1970:BRB

[Pac70] Howard Pack. Book review: [untitled]. Journal of the American Statistical Association, 65(330):1004–1006, June 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284615.

Page:1970:BRB

[Pag70] E. S. Page. Book review: [untitled]. Journal of the American Statistical Association, 65(329):458–459, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283613.

Paige:1977:BRB

[Pai77] Chris Paige. Book review: [untitled]. Journal of the American Statistical Association, 72(359):692, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286253.

Parks:1971:MLE

[Par71] Richard W. Parks. Maximum likelihood estimation of the linear expenditure system. Journal of the American Statistical Association, 66(336):900–903, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284251.

Parzen:1973:BRB

[Par73] Emanuel Parzen. Book review: [untitled]. Journal of the American Statistical Association, 68(341):237–238, March 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284179.

Park:1975:SSP

[Par75] Soo-Bin Park. On the small-sample power of Durbin’s h test. Journal of the American Statistical Association, 70(349):60–63, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285376.

Park:1976:CSS

[Par76] Soo-Bin Park. Comment on the small-sample power of Durbin’s h-test: Rejoinder. Journal of the American Statistical Association, 71(353):97–98, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285739.

Parke:1979:BRB

[Par79a] Robert Parke. Book review: [untitled]. Journal of the American Statistical Association, 74(366):511–512, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286378.

Parunak:1979:GAR

[Par79b] Anita Parunak. Graphical analysis of ranked counts (of words). Journal of the American Statistical Association, 74(365):25–30, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286716.

Parzen:1979:NSDa

[Par79c] Emanuel Parzen. Nonparametric statistical data modeling. Journal of the American Statistical Association, 74(365):105–121, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286734. See com-

ments [Tuk79a].

Parzen:1979:NSDb

[Par79d] Emanuel Parzen. Nonparametric statistical data modeling: Rejoinder. Journal of the American Statistical Association, 74(365):129– 131, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286741.

Pasternack:1972:BRB

[Pas72] Bernard S. Pasternack. Book review: [untitled]. Journal of the American Statistical Association, 67(339):721, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284489.

Pasternack:1973:BRB

[Pas73] B. S. Pasternack. Book review: [untitled]. Journal of the American Statistical Association, 68(344):1030, December 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284557.

Pasternack:1976:BRB

[Pas76] Bernard S. Pasternack. Book review: [untitled]. Journal of the American Statistical Association, 71(354):532, June 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285364.

Patton:1973:BRB

[Pat73] R. A. Patton. Book review: [untitled]. Journal of the American Statistical Association, 68(341):241, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284183.

Patil:1974:ITT

[Pat74] Kashinath D. Patil. Interaction test for three-dimensional contingency tables. Journal of the American Statistical Association, 69(345):164–168, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285517.

Patel:1975:BML

[Pat75a] Jagdish K. Patel. Bounds on moments of linear functions of order statistics from Weibull and other restricted families. Journal of the American Statistical Association, 70(351):670–673, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285953.

Patil:1975:CTE

[Pat75b] Kashinath D. Patil. Cochran’s Q test: Exact distribution. Journal of the American Statistical Association, 70(349):186–189, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285400.

Patel:1976:RSI

[Pat76] Jagdish K. Patel. Ranking and selection of IFR populations based on means. Journal of the American Statistical Association, 71(353):143– 146, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285757.

Paull:1978:GCP

[Pau78] Allan E. Paull. A generalized compound Poisson model for consumer purchase panel data analysis. Journal of the American Statistical Association, 73(364):706–713, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286262.

Pollock:1976:CTR

[PB76] K. H. Pollock and Yuksel Bek. A comparison of three randomized response models for quantitative data. Journal of the American Statistical Association, 71(356):884–886, December 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286855.

Park:1974:CVA

[PD74] Colin N. Park and Arthur L. Dudycha. A cross-validation approach to sample size determination for regression models. Journal of the American Statistical Association, 69(345):214–218, March 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285528.

Pfeifer:1978:DTG

[PE78] Charles G. Pfeifer and Peter Enis. Dorfman-type group testing for a modified binomial model. Journal of the American Statistical Association, 73(363):588–592, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286605.

Peck:1971:BRB

[Pec71] Jon K. Peck. Book review: [untitled]. Journal of the American Statistical Association, 66(334):424–425, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283953.

Pelto:1971:SSB

[Pel71] Chester R. Pelto. The statistical structure of bidding for oil and mineral rights. Journal of the American Statistical Association, 66 (335):456–460, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283508.

Pesando:1972:SVD

[Pes72] James E. Pesando. Seasonal variability in distributed lag models. Journal of the American Statistical Association, 67(338):311–312, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284374.

Pesando:1976:RED

[Pes76] James E. Pesando. Rational expectations and distributed lag expectations proxies. Journal of the American Statistical Association, 71(353):36–42, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285727.

Peterson:1977:EKM

[Pet77a] Arthur V. Peterson, Jr. Expressing the Kaplan–Meier estimator as a function of empirical subsurvival functions. Journal of the American Statistical Association, 72(360):854–858, December 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286474.

Petkau:1977:BRB

[Pet77b] John Petkau. Book review: [untitled]. Journal of the American Statistical Association, 72(360):931–932, December 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286503.

Petkau:1978:SMT

[Pet78a] A. John Petkau. Sequential medical trials for comparing an experimental with a standard treatment. Journal of the American Statistical Association, 73(362):328–338, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286661.

Petkau:1978:BRB

[Pet78b] John Petkau. Book review: [untitled]. Journal of the American Statistical Association, 73(363):684–685, September 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286637.

Petkau:1979:BRB

[Pet79] A. John Petkau. Book review: [untitled]. Journal of the American Statistical Association, 74(367):735–736, September 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287018.

Peace:1978:SPA

[PF78] Karl E. Peace and Roger E. Flora. Size and power assessments of tests of hypotheses on survival parameters. Journal of the American Statistical Association, 73(361):129–132, March 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286533.

Pirie:1972:DFN

[PH72] Walter R. Pirie and Myles Hollander. A distribution-free normal scores test for ordered alternatives in the randomized block design. Journal of the American Statistical Association, 67(340):855–857, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284649.

Patel:1973:NTI

[PH73] Kantilal M. Patel and David G. Hoel. A nonparametric test for interaction in factorial experiments. Journal of the American Statistical Association, 68(343):615–620, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284788.

Phillips:1974:STS

[PH74a] G. D. A. Phillips and A. C. Harvey. A simple test for serial correlation in regression analysis. Journal of the American Statistical Association, 69(348):935–939, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286166.

Pirie:1974:CDF

[PH74b] Walter R. Pirie and M. Hollander. Corrigenda: a Distribution-Free Normal Scores Test for Ordered Alternatives in the Randomized Block Design. Journal of the American Statistical Association, 69(345):288, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285572.

Policello:1976:ARP

[PH76] George E. Policello, II and Thomas P. Hettmansperger. Adaptive robust procedures for the one-sample location problem. Journal of the American Statistical Association, 71(355):624–633, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285592.

Han:1978:NPT

[pH78] Chien pai Han. Nonnegative and preliminary test estimators of variance components. Journal of the American Statistical Association, 73(364):855–858, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286293.

Phatarfod:1971:STG

[Pha71] R. M. Phatarfod. A sequential test for gamma distributions. Journal of the American Statistical Association, 66(336):876–878, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284245.

Pierce:1972:RCD

[Pie72] David A. Pierce. Residual correlations and diagnostic checking in dynamic-disturbance time series models. Journal of the American Statistical Association, 67(339):636–640, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284456.

Pierce:1973:CRC

[Pie73] David A. Pierce. Corrigenda: Residual Correlations and Diagnostic Checking in Dynamic- Disturbance Time Series Models. Journal of the American Statistical Association, 68(341):251, March 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284201.

Pierce:1977:RLT

[Pie77] David A. Pierce. Relationships — and the lack thereof — between economic time series, with special reference to money and interest rates. Journal of the American Statistical Association, 72(357):11–26, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286901.

Pierce:1979:MTS

[Pie79] David A. Pierce. R^{2} measures for time series. Journal of the American Statistical Association, 74(368):901–910, December 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286421.

Pirie:1973:BRB

[Pir73] Walter R. Pirie. Book review: [untitled]. Journal of the American Statistical Association, 68(344):1028–1029, December 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284552.

Pillai:1970:EDP

[PJ70] K. C. Sreedharan Pillai and Kanta Jayachandran. On the exact distribution of Pillai’s V^{(s)} criterion. Journal of the American Statistical Association, 65(329):447–454, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283609.

Perng:1976:TET

[PL76] S. K. Perng and Ramon C. Littell. A test of equality of two normal population means and variances. Journal of the American Statistical Association, 71(356):968–971, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286870.

Ploch:1971:BRB

[Plo71] Donald R. Ploch. Book review: [untitled]. Journal of the American Statistical Association, 66(334):424, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283952.

Plosser:1979:STF

[Plo79] Charles I. Plosser. Short-term forecasting and seasonal adjustment. Journal of the American Statistical Association, 74(365):15–24, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286715.

Poirier:1973:PRU

[Poi73] Dale J. Poirier. Piecewise regression using cubic spline. Journal of the American Statistical Association, 68(343):515–524, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284770.

Poirier:1974:CPR

[Poi74] Dale J. Poirier. Corrigenda: Piecewise Regression Using Cubic Splines. Journal of the American Statistical Association, 69(345): 288, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285573.

Poirier:1978:UBC

[Poi78] Dale J. Poirier. The use of the Box–Cox transformation in limited dependent variable models. Journal of the American Statistical Association, 73(362):284–287, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286653.

Poirier:1979:BRB

[Poi79] Dale J. Poirier. Book review: [untitled]. Journal of the American Statistical Association, 74(368):938–939, December 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286446.

Pollak:1973:BRB

[Pol73a] Edward Pollak. Book review: [untitled]. Journal of the American Statistical Association, 68(344):1023–1025, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284547.

Pollard:1973:CFS

[Pol73b] R. Pollard. Collegiate football scores and the negative binomial distribution. Journal of the American Statistical Association, 68(342):351– 352, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284075.

Pollak:1975:BRB

[Pol75] Edward Pollak. Book review: [untitled]. Journal of the American Statistical Association, 70(350):484, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285846.

Poole:1974:EDF

[Poo74] W. Kenneth Poole. Estimation of the distribution function of a continuous type random variable through randomized response. Journal of the American Statistical Association, 69(348):1002–1005, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286179.

Portnoy:1973:RIB

[Por73] Stephen Portnoy. On recovery of intra-block information. Journal of the American Statistical Association, 68(342):384–391, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284080.

Portnoy:1974:CRI

[Por74] Stephen Portnoy. Corrigenda: On Recovery of Intra-Block Information. Journal of the American Statistical Association, 69(345):288, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285574.

Potthoff:1972:CCM

[Pot72] Richard F. Potthoff. Comparing conditional means and variances in a regression model with measurement errors of known variances: Comment. Journal of the American Statistical Association, 67(338):412– 413, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284394.

Povey:1979:BRB

[Pov79] George Povey. Book review: [untitled]. Journal of the American Statistical Association, 74(365):253–254, March 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286780.

Paris:1975:BRB

[PR75a] Quirino Paris and Gordon Rausser. Book review: [untitled]. Journal of the American Statistical Association, 70(350):485–486, June 1975.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285847.

Patel:1975:BCM

[PR75b] Jagdish K. Patel and Campbell B. Read. Bounds on conditional moments of Weibull and other monotone failure rate families. Journal of the American Statistical Association, 70(349):238–244, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285410.

Prato:1970:MCD

[Pra70] Anthony A. Prato. Measurement of citrus demands when all variables are subject to error. Journal of the American Statistical Association, 65(331):1146–1158, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284281.

Pratt:1971:DOS

[Pra71] John W. Pratt. Does an observed sequence of numbers follow a simple rule? (another look at Bode’s law): Comment. Journal of the American Statistical Association, 66(335):567–568, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283528.

Pratt:1972:SPP

[Pra72] John W. Pratt. Some probability paradoxes in choice from among random alternatives: Comment. Journal of the American Statistical Association, 67(338):378–379, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284387.

Pratt:1973:CDO

[Pra73a] John W. Pratt. Corrigenda: Does an Observed Sequence of Numbers Follow a Simple Rule? (Another Look at Bode’s Law): Comment. Journal of the American Statistical Association, 68(341):251, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284202.

Pratt:1973:PDT

[Pra73b] John W. Pratt. Post-data two sample tests of location: Comments. Journal of the American Statistical Association, 68(341):104–105, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284151.

Prabhu:1974:BRB

[Pra74] N. U. Prabhu. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1048, December 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286191.

Pratt:1976:SIU

[Pra76] John W. Pratt. Strong inconsistency from uniform priors: Comment. Journal of the American Statistical Association, 71(353):119– 120, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285746.

Preston:1970:AID

[Pre70] Samuel H. Preston. The age-incidence of death from smoking. Journal of the American Statistical Association, 65(331):1125–1130, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284279.

Press:1972:EUM

[Pre72] S. James Press. Estimation in univariate and multivariate stable distributions. Journal of the American Statistical Association, 67 (340):842–846, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284646.

Preston:1974:BRB

[Pre74] Samuel H. Preston. Book review: [untitled]. Journal of the American Statistical Association, 69(346):574–575, June 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285699.

Prescott:1976:BRB

[Pre76] James R. Prescott. Book review: [untitled]. Journal of the American Statistical Association, 71(353):243, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285779.

Prescott:1978:CST

[Pre78a] P. Prescott. Corrigenda: Selection of Trimming Proportions for Robust Adaptive Trimmed Means. Journal of the American Statistical Association, 73(363):691, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286648.

Prescott:1978:STP

[Pre78b] P. Prescott. Selection of trimming proportions for robust adaptive trimmed means. Journal of the American Statistical Association, 73(361):133–140, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286534.

Press:1978:QCF

[Pre78c] S. James Press. Qualitative controlled feedback for forming group judgments and making decisions. Journal of the American Statistical Association, 73(363):526–535, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286596.

Preston:1978:BRB

[Pre78d] D. B. Preston. Book review: [untitled]. Journal of the American Statistical Association, 73(362):438–439, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286683.

Pringle:1974:SRE

[Pri74] R. M. Pringle. Some results on the estimation of variance components by MINQUE. Journal of the American Statistical Association, 69(348):987–989, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286176.

Proctor:1971:PSD

[Pro71] Charles H. Proctor. Processing survey data with ordered alternative hypotheses. Journal of the American Statistical Association, 66(336):720–724, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284218.

Proctor:1977:BRB

[Pro77] Charles H. Proctor. Book review: [untitled]. Journal of the American Statistical Association, 72(357):229–230, March 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286947.

Pradhan:1974:EBF

[PS74] Meena Pradhan and Y. S. Sathe. Equivalence between fixed sample size rule and Hoel’s inverse sampling rule for play-the-winner. Journal of the American Statistical Association, 69(346):475–476, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285681.

Pradhan:1975:UES

[PS75] Meena Pradhan and Y. S. Sathe. An unbiased estimator and a sequential test for the correlation coefficient. Journal of the American Statistical Association, 70(349):160–161, March 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285393.

Pradhan:1976:ARC

[PS76a] Meena Pradhan and Y. S. Sathe. Analytical remarks on Canner’s minimax method for finding the better of two binomial populations. Journal of the American Statistical Association, 71(353):239–241, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285777.

Press:1976:MVB

[PS76b] S. J. Press and A. J. Scott. Missing variables in Bayesian regression, II. Journal of the American Statistical Association, 71(354):366–369, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285315.

Proschan:1976:EPM

[PS76c] Frank Proschan and Pasquale Sullo. Estimating the parameters of a multivariate exponential distribution. Journal of the American Statistical Association, 71(354):465–472, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285335.

Puterman:1978:BRB

[Put78] Martin L. Puterman. Book review: [untitled]. Journal of the American Statistical Association, 73(362):446, June 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286704.

Puterman:1979:BRB

[Put79] Martin L. Puterman. Book review: [untitled]. Journal of the American Statistical Association, 74(367):739–740, September 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287025.

Phadke:1974:MCS

[PW74] M. S. Phadke and S. M. Wu. Modeling of continuous stochastic processes from discrete observations with application to sunspots data. Journal of the American Statistical Association, 69(346):325–329, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285651.

Press:1978:CBL

[PW78] S. James Press and Sandra Wilson. Choosing between logistic regression and discriminant analysis. Journal of the American Statistical Association, 73(364):699–705, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286261.

Press:1974:BAS

[PY74] S. James Press and Chung-Fang Yang. A Bayesian approach to second guessing “undecided” respondents. Journal of the American Statistical Association, 69(345):58–67, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285500.

Pyne:1979:SVP

[Pyn79] David A. Pyne. Single-variable Poisson regression: a goodness-of-fit test and the comparison of regression coefficients. Journal of the American Statistical Association, 74(366):489–493, June 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286360.

Q:1974:BRBa

[Q.74a] D. Q. Book review: [untitled]. Journal of the American Statistical Association, 69(346):583–584, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285717.

Q:1974:BRBb

[Q.74b] D. Q. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1053, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286202.

Quandt:1978:EMNa

[QR78a] Richard E. Quandt and James B. Ramsey. Estimating mixtures of normal distributions and switching regressions. Journal of the American Statistical Association, 73(364):730–738, December 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286266.

Quandt:1978:EMNb

[QR78b] Richard E. Quandt and James B. Ramsey. Estimating mixtures of normal distributions and switching regressions: Rejoinder. Journal of the American Statistical Association, 73(364):751–752, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286275.

Quandt:1979:EMN

[QR79] Richard E. Quandt and James B. Ramsey. Estimating mixtures of normal distributions and switching regressions: Rejoinder. Journal of the American Statistical Association, 74(365):56, March 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286721.

Quandt:1972:NAE

[Qua72] Richard E. Quandt. A new approach to estimating switching regressions. Journal of the American Statistical Association, 67(338):306– 310, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284373.

Quandt:1974:SME

[Qua74] Richard E. Quandt. A stochastic model of elections in two-party systems. Journal of the American Statistical Association, 69(346):315– 324, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285650.

Quade:1977:BRB

[Qua77a] Dana Quade. Book review: [untitled]. Journal of the American Statistical Association, 72(357):240, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286966.

Quandt:1977:BRB

[Qua77b] Richard E. Quandt. Book review: [untitled]. Journal of the American Statistical Association, 72(357):239–240, March 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286965.

Quade:1978:BRB

[Qua78] Dana Quade. Book review: [untitled]. Journal of the American Statistical Association, 73(362):444, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286698.

Quade:1979:RAB

[Qua79a] Dana Quade. Regression analysis based on the signs of the residuals. Journal of the American Statistical Association, 74(366):411–417, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286345.

Quade:1979:UWR

[Qua79b] Dana Quade. Using weighted rankings in the analysis of complete blocks with additive block effects. Journal of the American Statistical Association, 74(367):680–683, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286991.

Radlow:1975:AMA

[RA75] Robert Radlow and Edward F. Alf, Jr. An alternate multinomial assessment of the accuracy of the χ^{2} test of goodness of fit. Journal of the American Statistical Association, 70(352):811–813, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285440.

Ryan:1976:NGS

[RA76] Thomas A. Ryan, Jr. and Charles E. Antle. A note on Gupta’s selection procedure. Journal of the American Statistical Association, 71(353):140–142, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285756.

Raghavachari:1971:BRB

[Rag71] M. Raghavachari. Book review: [untitled]. Journal of the American Statistical Association, 66(335):658, September 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283559.

Raghavachari:1975:BRB

[Rag75] M. Raghavachari. Book review: [untitled]. Journal of the American Statistical Association, 70(352):961, December 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285471.

Raghavachari:1978:BRB

[Rag78] M. Raghavachari. Book review: [untitled]. Journal of the American Statistical Association, 73(364):891, December 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286306.

Rahe:1974:TCV

[Rah74] Alton J. Rahe. Tables of critical values for the Pratt matched pair signed rank statistic. Journal of the American Statistical Association, 69(346):368–373, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285660.

Raj:1975:LRR

[Raj75] Baldev Raj. Linear regression with random coefficients: The finite sample and convergence properties. Journal of the American Statistical Association, 70(349):127–137, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285388.

Raj:1977:ENV

[Raj77] Des Raj. On estimating the number of vital events in demographic surveys. Journal of the American Statistical Association, 72(358):377–381, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286802.

Rockette:1974:MLE

[RAK74] Howard Rockette, Charles Antle, and Lawrence A. Klimko. Maximum likelihood estimation with the Weibull model. Journal of the American Statistical Association, 69(345):246–249, March 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285536.

Ramsey:1971:SSP

[Ram71] Fred L. Ramsey. Small sample power functions for nonparametric tests of location in the double exponential family. Journal of the American Statistical Association, 66(333):149–151, March 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284865.

Ramsey:1974:BRB

[Ram74] James B. Ramsey. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1049–1050, December 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286193.

Ramsay:1977:CSS

[Ram77] J. O. Ramsay. A comparative study of several robust estimates of slope, intercept, and scale in linear regression. Journal of the American Statistical Association, 72(359):608–615, September 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286228.

Ramsey:1978:PDBa

[Ram78a] Philip H. Ramsey. Power differences between pairwise multiple comparisons. Journal of the American Statistical Association, 73 (363):479–485, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286584.

Ramsey:1978:PDBb

[Ram78b] Philip H. Ramsey. Power differences between pairwise multiple comparisons: Rejoinder. Journal of the American Statistical Association, 73(363):487, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286586.

Rand:1971:OCE

[Ran71] William M. Rand. Objective criteria for the evaluation of clustering methods. Journal of the American Statistical Association, 66(336):846–850, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284239.

Rand:1973:CBM

[Ran73] William M. Rand. Corrigenda: Back Matter. Journal of the American Statistical Association, 68(341):251, March 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284203.

Rao:1970:EHV

[Rao70] C. Radhakrishna Rao. Estimation of heteroscedastic variances in linear models. Journal of the American Statistical Association, 65(329):161–172, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283583.

Rao:1971:CAM

[Rao71a] M. R. Rao. Cluster analysis and mathematical programming. Journal of the American Statistical Association, 66(335):622–626, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283542.

Rao:1971:SDH

[Rao71b] T. J. Rao. πPS sampling designs and the Horvitz–Thompson estimator. Journal of the American Statistical Association, 66 (336):872–875, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284244.

Rao:1972:EVC

[Rao72] C. Radhakrishna Rao. Estimation of variance and covariance components in linear models. Journal of the American Statistical Association, 67(337):112–115, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284708.

Rao:1973:CSD

[Rao73] T. J. Rao. Corrigenda: πPS Sampling Designs and the Horvitz– Thompson Estimato. Journal of the American Statistical Association, 68(341):251, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284204.

Rao:1975:TPR

[Rao75] Poduri S. R. S. Rao. On the two-phase ratio estimator in finite populations. Journal of the American Statistical Association, 70 (352):839–845, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285445.

Rao:1977:MLA

[Rao77a] J. N. K. Rao. Maximum likelihood approaches to variance component estimation and to related problems: Comment. Journal of the American Statistical Association, 72(358):338–339, June 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286797.

Rao:1977:BRB

[Rao77b] P. S. R. S. Rao. Book review: [untitled]. Journal of the American Statistical Association, 72(360):924, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286491.

Ravindran:1978:BRB

[Rav78] A. Ravindran. Book review: [untitled]. Journal of the American Statistical Association, 73(361):228–229, March 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286572.

Roscoe:1971:IRR

[RB71] John T. Roscoe and Jackson A. Byars. An investigation of the restraints with respect to sample size commonly imposed on the use of the chi-square statistic. Journal of the American Statistical Association, 66(336):755–759, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284224.

Randles:1978:DAB

[RBRH78a] Ronald H. Randles, James D. Broffitt, John S. Ramberg, and Robert V. Hogg. Discriminant analysis based on ranks. Journal of the American Statistical Association, 73(362):379–384, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286669.

Randles:1978:GLQ

[RBRH78b] Ronald H. Randles, James D. Broffitt, John S. Ramberg, and Robert V. Hogg. Generalized linear and quadratic discriminant functions using robust estimates. Journal of the American Statistical Association, 73(363):564–568, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286601.

Robertson:1974:IPE

[RC74] Tim Robertson and J. D. Cryer. An iterative procedure for estimating the mode. Journal of the American Statistical Association, 69 (348):1012–1016, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286181.

Ramsey:1976:SMD

[RC76] James B. Ramsey and Andrew Chesher. Some measures of the “difference” between regression models. Journal of the American Statistical Association, 71(356):972–976, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286871.

Royall:1978:VEF

[RC78] Richard M. Royall and William G. Cumberland. Variance estimation in finite population sampling. Journal of the American Statistical Association, 73(362):351–358, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286664.

Read:1971:PSP

[Rea71] Campbell B. Read. The partial sequential probability ratio test. Journal of the American Statistical Association, 66(335):646–650, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2283549.

Reading:1975:MCP

[Rea75] James C. Reading. A multiple comparison procedure for classifying all pairs out of k means as close or distant. Journal of the American Statistical Association, 70(352):832–838, December 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285444.

Regan:1973:SRA

[Reg73] Opal G. Regan. Statistical reforms accelerated by sixth census error. Journal of the American Statistical Association, 68(343):540–546, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284773.

Reimers:1976:AAR

[Rei76] Cordelia Reimers. Is the average age at retirement changing? Journal of the American Statistical Association, 71(355):552–558, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285582.

Relles:1972:SAG

[Rel72] Daniel A. Relles. A simple algorithm for generating binomial random variables when N is large. Journal of the American Statistical Association, 67(339):612–613, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284449.

Relles:1976:BRBa

[Rel76a] Daniel A. Relles. Book review: [untitled]. Journal of the American Statistical Association, 71(353):252, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285796.

Relles:1976:BRBb

[Rel76b] Daniel A. Relles. Book review: [untitled]. Journal of the American Statistical Association, 71(354):533, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285365.

Revankar:1974:CSF

[Rev74a] Nagesh S. Revankar. Corrigenda: Some Finite Sample Results in the Context of Two Seemingly Unrelated Regression Equations. Journal of the American Statistical Association, 69(346):587, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285721.

Revankar:1974:SFS

[Rev74b] Nagesh S. Revankar. Some finite sample results in the context of two seemingly unrelated regression equations. Journal of the American Statistical Association, 69(345):187–190, March 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285521.

Revankar:1976:URR

[Rev76] Nagesh S. Revankar. Use of restricted residuals in SUR systems: Some finite sample results. Journal of the American Statistical Association, 71(353):183–188, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285765.

Ramsey:1972:MCS

[RG72a] J. Ramsey and R. Gilbert. A Monte Carlo study of some small sample properties of tests for specification error. Journal of the American Statistical Association, 67(337):180–186, March 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284719.

Rao:1972:BOS

[RG72b] J. N. K. Rao and P. D. Ghangurde. Bayesian optimization in sampling finite populations. Journal of the American Statistical Association, 67(338):439–443, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284401.

Randles:1971:CUI

[RH71] Ronald H. Randles and Robert V. Hogg. Certain uncorrelated and independent rank statistics. Journal of the American Statistical Association, 66(335):569–574, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283530.

Royall:1973:REFa

[RH73a] Richard M. Royall and Jay Herson. Robust estimation in finite populations I. Journal of the American Statistical Association, 68 (344):880–889, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284516.

Royall:1973:REFb

[RH73b] Richard M. Royall and Jay Herson. Robust estimation in finite populations II: Stratification on a size variable. Journal of the American Statistical Association, 68(344):890–893, December 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284517.

Rhodes:1975:BRB

[Rho75] George Rhodes. Book review: [untitled]. Journal of the American Statistical Association, 70(351):732, September 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285971.

Rhodes:1977:BRB

[Rho77] George F. Rhodes, Jr. Book review: [untitled]. Journal of the American Statistical Association, 72(359):683, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286243.

Richardson:1972:BRB

[Ric72] J. David Richardson. Book review: [untitled]. Journal of the American Statistical Association, 67(338):490–491, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284413.

Ricardo:1976:BRB

[Ric76] Henry J. Ricardo. Book review: [untitled]. Journal of the American Statistical Association, 71(354):531, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285359.

Rausser:1976:EAW

[RO76] Gordon C. Rausser and Ronald A. Oliveira. An econometric analysis of wilderness area use. Journal of the American Statistical Association, 71(354):276–285, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285298.

Robillard:1972:KDT

[Rob72] Pierre Robillard. Kendall’s S distribution with ties in one ranking. Journal of the American Statistical Association, 67(338):453– 455, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284404.

Robertson:1974:BRB

[Rob74] Tim Robertson. Book review: [untitled]. Journal of the American Statistical Association, 69(346):575, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285700.

Robinson:1975:TAR

[Rob75] J. Robinson. On the test for additivity in a randomized block design. Journal of the American Statistical Association, 70(349):184– 185, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285399.

Robertson:1977:EQE

[Rob77] C. A. Robertson. Estimation of quantiles of exponential distributions with minimum error in predicted distribution functions. Journal of the American Statistical Association, 72(357):162–164, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286930.

Robertson:1978:TAO

[Rob78a] Tim Robertson. Testing for and against an order restriction on multinomial parameters. Journal of the American Statistical Association, 73(361):197–202, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286546.

Robinson:1978:BRB

[Rob78b] Sherman Robinson. Book review: [untitled]. Journal of the American Statistical Association, 73(361):226, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286568.

Robinson:1979:SAE

[Rob79] Peter M. Robinson. Statistical analysis of econometric models: Comment. Journal of the American Statistical Association, 74 (367):646–647, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286984.

Rockwell:1975:IID

[Roc75] Richard C. Rockwell. An investigation of imputation and differential quality of data in the 1970 census. Journal of the American Statistical Association, 70(349):39–42, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285372.

Rodriguez:1976:BRB

[Rod76] Robert N. Rodriguez. Book review: [untitled]. Journal of the American Statistical Association, 71(356):1001–1002, December 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286881.

Rogan:1978:BRB

[Rog78] Walter J. Rogan. Book review: [untitled]. Journal of the American Statistical Association, 73(364):897, December 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286316.

Rohlf:1974:GIJ

[Roh74] F. James Rohlf. Graphs implied by the Jardine–Sibson overlapping clustering methods, B_{k}. Journal of the American Statistical Association, 69(347):705–710, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286005.

Rohr:1976:DTA

[Roh76] R. J. Rohr. On the distribution of two alternative structural t statistics. Journal of the American Statistical Association, 71 (355):731–736, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285611.

Rohatgi:1977:BRB

[Roh77] Vijay K. Rohatgi. Book review: [untitled]. Journal of the American Statistical Association, 72(360):924–925, December 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286492.

Rohatgi:1978:BRB

[Roh78] V. K. Rohatgi. Book review: [untitled]. Journal of the American Statistical Association, 73(361):222, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286560.

Roll:1972:BRB

[Rol72a] Richard Roll. Book review: [untitled]. Journal of the American Statistical Association, 67(337):247, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284745.

Rolph:1972:BRB

[Rol72b] John E. Rolph. Book review: [untitled]. Journal of the American Statistical Association, 67(340):953–954, December 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284670.

Rolph:1976:BRB

[Rol76] John E. Rolph. Book review: [untitled]. Journal of the American Statistical Association, 71(356):1008, December 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286891.

Romberg:1972:CST

[Rom72] Harland F. Romberg. Continuous sequential testing of a Poisson process to minimize the Bayes risk. Journal of the American Statistical Association, 67(340):921–926, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284663.

Rosenberg:1972:ESS

[Ros72] Barr Rosenberg. The estimation of stationary stochastic regression parameters reexamined. Journal of the American Statistical Association, 67(339):650–654, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284459.

Ross:1973:BRB

[Ros73] Sheldon Ross. Book review: [untitled]. Journal of the American Statistical Association, 68(342):497–498, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284118.

Rosenblatt:1974:BRBa

[Ros74a] Judah Rosenblatt. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1052, December 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286198.

Rosenblatt:1974:BRBb

[Ros74b] Judah Rosenblatt. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1052, December 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286199.

Rothman:1971:BRB

[Rot71] E. D. Rothman. Book review: [untitled]. Journal of the American Statistical Association, 66(336):922, December 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284263.

Roth:1978:BRB

[Rot78a] Arthur J. Roth. Book review: [untitled]. Journal of the American Statistical Association, 73(364):893–894, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286309.

Roth:1978:NPS

[Rot78b] Arthur J. Roth. A new procedure for selecting a subset containing the best normal population. Journal of the American Statistical Association, 73(363):613–617, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286610.

Rothenberg:1979:SAE

[Rot79] Thomas J. Rothenberg. Statistical analysis of econometric models: Comment. Journal of the American Statistical Association, 74(367):648–649, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286985.

Royall:1976:LLS

[Roy76] Richard M. Royall. The linear least-squares prediction approach to two-stage sampling. Journal of the American Statistical Association, 71(355):657–664, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285596.

Richardson:1971:DSS

[RR71] David H. Richardson and Robert J. Rohr. Distribution of a structural t-statistic for the case of two included endogenous variables. Journal of the American Statistical Association, 66(334):375–382, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283939.

Reiss:1976:WDF

[RR76] Rolf D. Reiss and Ludger Ruschendorf. On Wilks’ distribution-free confidence intervals for quantile intervals. Journal of the American Statistical Association, 71(356):940–944, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286865.

Relles:1977:SFR

[RR77] Daniel A. Relles and William H. Rogers. Statisticians are fairly robust estimators of location. Journal of the American Statistical Association, 72(357):107–111, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286917.

Rizvi:1972:DFI

[RS72] M. Haseeb Rizvi and K. M. Lal Saxena. Distribution-free interval estimation of the largest α-quantile. Journal of the American Statistical Association, 67(337):196–198, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284722.

Raghunandanan:1973:SPM

[RS73a] K. Raghunandanan and R. Srinivasan. Some product moments useful in sampling theory. Journal of the American Statistical Association, 68(342):409–413, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284087.

Rizvi:1973:SLM

[RS73b] M. Haseeb Rizvi and Herbert Solomon. Selection of largest multiple correlation coefficients: Asymptotic case. Journal of the American Statistical Association, 68(341):184–188, March 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284165.

Rizvi:1974:CSL

[RS74a] M. Haseeb Rizvi and Herbert Solomon. Corrigenda: Selection of Largest Multiple Correlation Coefficients: Asymptotic Case. Journal of the American Statistical Association, 69(345):288, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285575.

Robbins:1974:STI

[RS74b] Herbert Robbins and David O. Siegmund. Sequential tests involving two populations. Journal of the American Statistical Association, 69(345):132–139, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285511.

Ramsey:1976:SFR

[RS76] James B. Ramsey and Peter Schmidt. Some further results on the use of OLS and BLUS residuals in specification error tests. Journal of the American Statistical Association, 71(354):389–390, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285320.

Rasmussen:1979:OAS

[RS79] Shelley L. Rasmussen and Norman Starr. Optimal and adaptive stopping in the search for new species. Journal of the American Statistical Association, 74(367):661–667, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286988.

Rubin:1974:BRB

[Rub74a] Donald B. Rubin. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1050, December 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286194.

Rubin:1974:CEP

[Rub74b] Donald B. Rubin. Characterizing the estimation of parameters in incomplete-data problems. Journal of the American Statistical Association, 69(346):467–474, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285680.

Rubin:1977:FSN

[Rub77] Donald B. Rubin. Formalizing subjective notions about the effect of nonrespondents in sample surveys. Journal of the American Statistical Association, 72(359):538–543, September 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286214.

Rubin:1979:UMM

[Rub79] Donald B. Rubin. Using multivariate matched sampling and regression adjustment to control bias in observational studies. Journal of the American Statistical Association, 74(366):318–328, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286330.

Rao:1977:EVS

[RV77] J. N. K. Rao and K. Vijayan. On estimating the variance in sampling with probability proportional to aggregate size. Journal of the American Statistical Association, 72(359):579–584, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286220.

Richardson:1970:LSG

[RW70] David H. Richardson and De-Min Wu. Least squares and grouping method estimators in the errors in variables model. Journal of the American Statistical Association, 65(330):724–748, June 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284583.

Rippe:1974:FAM

[RW74] Richard D. Rippe and Maurice Wilkinson. Forecasting accuracy of the McGraw–Hill anticipatory data. Journal of the American Statistical Association, 69(348):849–858, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286153.

Robins:1977:MEE

[RW77] Philip K. Robins and Richard W. West. Measurement errors in the estimation of home value. Journal of the American Statistical Association, 72(358):290–294, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286790.

Rogers:1978:TML

[RY78] Gerald S. Rogers and Dennis L. Young. On testing a multivariate linear hypothesis when the covariance matrix and its inverse have the same pattern. Journal of the American Statistical Association, 73(361):203–207, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286547.

Ryan:1978:RMB

[Rya78] Thomas A. Ryan, Jr. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73(361): 90–91, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286524.

Rymes:1971:BRB

[Rym71] T. K. Rymes. Book review: [untitled]. Journal of the American Statistical Association, 66(333):222, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284876.

Ramsey:1971:SET

[RZ71] James B. Ramsey and Paul Zarembka. Specification error tests and alternative functional forms of the aggregate production function. Journal of the American Statistical Association, 66(335):471–477, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2283510.

S:1973:BRB

[S.73] S. M. S. Book review: [untitled]. Journal of the American Statistical Association, 68(342):501, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284125.

S:1974:BRBa

[S.74a] P. K. S. Book review: [untitled]. Journal of the American Statistical Association, 69(345):282, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285564.

S:1974:BRBb

[S.74b] P. K. S. Book review: [untitled]. Journal of the American Statistical Association, 69(346):583, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285714.

S:1974:BRBc

[S.74c] S. M. S. Book review: [untitled]. Journal of the American Statistical Association, 69(347):835, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286040.

S:1975:BRB

[S.75] P. K. S. Book review: [untitled]. Journal of the American Statistical Association, 70(351):736, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285983.

Srivastava:1970:OFF

[SA70] J. N. Srivastava and D. A. Anderson. Optimal fractional factorial plans for main effects orthogonal to two- factor interactions: 2^{m} series. Journal of the American Statistical Association, 65(330):828– 843, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284590.

Sahai:1973:CRV

[SA73] Hardeo Sahai and R. L. Anderson. Confidence regions for variance ratios of random models for balanced data. Journal of the American Statistical Association, 68(344):951–952, December 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284528.

Saaty:1971:BRB

[Saa71] Thomas L. Saaty. Book review: [untitled]. Journal of the American Statistical Association, 66(335):661, September 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283567.

Saboia:1977:AIM

[Sab77] Joao Luiz Maurity Saboia. Autoregressive integrated moving average (ARIMA) models for birth forecasting. Journal of the American Statistical Association, 72(358):264–270, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286787.

Sager:1979:IME

[Sag79] Thomas W. Sager. An iterative method for estimating a multivariate mode and isopleth. Journal of the American Statistical Association, 74(366):329–339, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286331.

Sahai:1976:CEV

[Sah76] Hardeo Sahai. A comparison of estimators of variance components in the balanced three-stage nested random effects model using mean squared error criterion. Journal of the American Statistical Association, 71(354):435–444, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285330.

Samuel:1970:RST

[Sam70] Ester Samuel. Randomized sequential tests. A comparison between curtailed single-sampling plans and sequential probability ratio tests. Journal of the American Statistical Association, 65(329):431–437, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283607.

Sampson:1974:TTR

[Sam74] Allan R. Sampson. A tale of two regressions. Journal of the American Statistical Association, 69(347):682–689, September 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286002.

Samaniego:1976:CCP

[Sam76a] Francisco J. Samaniego. A characterization of convoluted Poisson distributions with applications to estimation. Journal of the American Statistical Association, 71(354):475–479, June 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285337.

Samiuddin:1976:BTH

[Sam76b] M. Samiuddin. Bayesian test of homogeneity of variance. Journal of the American Statistical Association, 71(354):515–517, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285344.

Sampson:1978:SBE

[Sam78] Allan R. Sampson. Simple BAN estimators of correlations for certain multivariate normal models with known variances. Journal of the American Statistical Association, 73(364):859–862, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286294.

Sampson:1979:CSR

[Sam79] Paul D. Sampson. Comment on “Splines and Restricted Least Squares”. Journal of the American Statistical Association, 74 (366):303–305, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286327.

Sands:1973:EDW

[San73] Saul Sands. Estimating data withheld in grouped size distributions. Journal of the American Statistical Association, 68(342):306–311, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284067.

Sanathanan:1974:CPF

[San74a] Lalitha Sanathanan. Critical power function and decision making. Journal of the American Statistical Association, 69(346):398–402, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285668.

Sanathanan:1974:SPL

[San74b] Lalitha Sanathanan. Some properties of the logistic model for dichotomous response. Journal of the American Statistical Association, 69(347):744–749, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286012.

Santini:1974:BRB

[San74c] Christine Santini. Book review: [untitled]. Journal of the American Statistical Association, 69(348):1047–1048, December 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286190.

Sanathanan:1977:EST

[San77a] Lalitha Sanathanan. Estimating the size of a truncated sample. Journal of the American Statistical Association, 72(359):669–672, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286238.

Santner:1977:BRB

[San77b] Thomas J. Santner. Book review: [untitled]. Journal of the American Statistical Association, 72(358):481–482, June 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286823.

Sande:1978:RMB

[San78] Gordon Sande. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73(361):91–92, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286525.

Saperstein:1972:GBP

[Sap72] Bernard Saperstein. The generalized birthday problem. Journal of the American Statistical Association, 67(338):425–428, June 1972.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284398.

Sarndal:1972:BRB

[Sär72] Carl Erik Särndal. Book review: [untitled]. Journal of the American Statistical Association, 67(339):717–718, September 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284483.

Sarndal:1979:BRB

[Sär79] Carl Erik Särndal. Book review: [untitled]. Journal of the American Statistical Association, 74(366):510, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286376.

Saunders:1974:FRV

[Sau74] Sam C. Saunders. A family of random variables closed under reciprocation. Journal of the American Statistical Association, 69(346):533– 539, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285692.

Savage:1970:CWP

[Sav70] Leonard J. Savage. Comments on a weakened principle of conditionality. Journal of the American Statistical Association, 65(329):399–401, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283602.

Savage:1971:EPP

[Sav71] Leonard J. Savage. Elicitation of personal probabilities and expectations. Journal of the American Statistical Association, 66 (336):783–801, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284229.

Saxena:1971:IEL

[Sax71] K. M. Lal Saxena. Interval estimation of the largest variance of K normal populations. Journal of the American Statistical Association, 66(334):408–410, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283946.

Saxena:1976:SSP

[Sax76] K. M. Lal Saxena. A single-sample procedure for the estimation of the largest mean. Journal of the American Statistical Association, 71(353):147–148, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285758.

Steinhorst:1971:DCB

[SB71] R. Kirk Steinhorst and David C. Bowden. Discrimination and confidence bands on percentiles. Journal of the American Statistical Association, 66(336):851–854, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284240.

Sudman:1973:ETM

[SB73] Seymour Sudman and Norman M. Bradburn. Effects of time and memory factors on response in surveys. Journal of the American Statistical Association, 68(344):805–815, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284504.

Shenton:1975:JSK

[SB75] L. R. Shenton and K. O. Bowman. Johnson’s S_{U} and the skewness and kurtosis statistics. Journal of the American Statistical Association, 70(349):220–228, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285407.

Shenton:1977:BMD

[SB77] L. R. Shenton and K. O. Bowman. A bivariate model for the distribution of \sqrt{b_{1}} and b_{2}. Journal of the American Statistical Association, 72(357):206–211, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286939.

Sanathanan:1978:LME

[SB78a] Lalitha Sanathanan and Saul Blumenthal. The logistic model and estimation of latent structure. Journal of the American Statistical Association, 73(364):794–799, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286282.

Shimizu:1978:RRT

[SB78b] Iris M. Shimizu and Gordon Scott Bonham. Randomized response technique in a national survey. Journal of the American Statistical Association, 73(361):35–39, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286512.

Shuster:1979:NMC

[SB79] J. J. Shuster and James M. Boyett. Nonparametric multiple comparison procedures. Journal of the American Statistical Association, 74(366):379–382, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286339.

Scott:1978:FPS

[SBH78] A. J. Scott, K. R. W. Brewer, and E. W. H. Ho. Finite population sampling and robust estimation. Journal of the American Statistical Association, 73(362):359–361, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286665.

Singh:1979:ASB

[SBY79] S. N. Singh, B. N. Bhattacharya, and R. C. Yadava. An adjustment of a selection bias in postpartum amenorrhea from follow-up studies. Journal of the American Statistical Association, 74(368):916–920, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286423.

Samuel-Cahn:1975:RFF

[SC75a] Ester Samuel-Cahn. Remark on a formula by Fisher. Journal of the American Statistical Association, 70(351):720, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285962.

Smith:1975:ULB

[SC75b] Robert Longworth Smith and Robert P. Charrow. Upper and lower bounds for probability of guilt based on circumstantial evidence. Journal of the American Statistical Association, 70(351):555–560, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285932.

Samuel-Cahn:1976:CRF

[SC76] Ester Samuel-Cahn. Corrigenda: Remark on a Formula by Fisher. Journal of the American Statistical Association, 71(353):255, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285806.

Smith:1977:PBC

[SC77] Robert Longworth Smith and Robert P. Charrow. Probabilities based on circumstantial evidence: Comment. Journal of the American Statistical Association, 72(358):318–319, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286795.

Scheffe:1970:PSB

[Sch70a] Henry Scheffé. Practical solutions of the Behrens–Fisher problem. Journal of the American Statistical Association, 65(332):1501–1508, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284332.

Schmitz:1970:BRB

[Sch70b] Andrew Schmitz. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1407–1408, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284312.

Schmidt:1971:BRB

[Sch71] Richard N. Schmidt. Book review: [untitled]. Journal of the American Statistical Association, 66(333):224–225, March 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284881.

Schlesselman:1973:DTT

[Sch73a] James J. Schlesselman. Data transformation in two-way analysis of variance. Journal of the American Statistical Association, 68(342):369–378, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284078.

Schlossmacher:1973:ITA

[Sch73b] E. J. Schlossmacher. An iterative technique for absolute deviations curve fitting. Journal of the American Statistical Association, 68 (344):857–859, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284512.

Schmidt:1973:BRB

[Sch73c] Peter Schmidt. Book review: [untitled]. Journal of the American Statistical Association, 68(341):240, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284182.

Schuster:1973:GFP

[Sch73d] Eugene F. Schuster. On the goodness-of-fit problem for continuous symmetric distributions. Journal of the American Statistical Association, 68(343):713–715, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284804.

Schmidt:1974:MAD

[Sch74a] Peter Schmidt. A modification of the Almon distributed lag. Journal of the American Statistical Association, 69(347):679–681, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286001.

Schuster:1974:CGF

[Sch74b] Eugene F. Schuster. Corrigenda: On the Goodness-of-Fit Problem for Continuous Symmetric Distributions. Journal of the American Statistical Association, 69(345):288, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285576.

Schum:1975:BRB

[Sch75a] David A. Schum. Book review: [untitled]. Journal of the American Statistical Association, 70(352):960–961, December 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285470.

Schwager:1975:BRB

[Sch75b] Steven J. Schwager. Book review: [untitled]. Journal of the American Statistical Association, 70(350):491, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285858.

Schweder:1976:SOM

[Sch76] Tore Schweder. Some “optimal” methods to detect structural shift or outliers in regression. Journal of the American Statistical Association, 71(354):491–501, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285340.

Scheffe:1977:NRMa

[Sch77a] Henry Scheffé. A note on a reformulation of the S-method of multiple comparison. Journal of the American Statistical Association, 72(357):143–144, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286922.

Scheffe:1977:NRMb

[Sch77b] Henry Scheffé. A note on a reformulation of the S-method of multiple comparison: Rejoinder. Journal of the American Statistical Association, 72(357):146, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286924.

Schmidt:1978:BRB

[Sch78a] Richard N. Schmidt. Book review: [untitled]. Journal of the American Statistical Association, 73(362):442–443, June 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286693.

Schucany:1978:RMB

[Sch78b] William R. Schucany. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73(361): 92–93, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286526.

Schwertman:1978:NGG

[Sch78c] Neil C. Schwertman. A note on the Geisser–Greenhouse correction for incomplete data split-plot analysis. Journal of the American Statistical Association, 73(362):393–396, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286671.

Schach:1979:BRB

[Sch79a] Siegfried Schach. Book review: [untitled]. Journal of the American Statistical Association, 74(365):255, March 1979. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286782.

Schmidt:1979:BRB

[Sch79b] Peter Schmidt. Book review: [untitled]. Journal of the American Statistical Association, 74(368):937–938, December 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286445.

Sclove:1973:BRB

[Scl73] Stanley L. Sclove. Book review: [untitled]. Journal of the American Statistical Association, 68(342):499, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284121.

Sclove:1979:OFC

[Scl79] Stanley L. Sclove. The occurrence of fingerprint characteristics as a two-dimensional process. Journal of the American Statistical Association, 74(367):588–595, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286975.

Scott:1978:BPS

[Sco78] Robert C. Scott. The bias problem in smear-and-sweep analysis. Journal of the American Statistical Association, 73(364):714–718, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286263.

Scudo:1970:BRB

[Scu70] Franco Scudo. Book review: [untitled]. Journal of the American Statistical Association, 65(329):471–473, March 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283620.

Sukhatme:1973:NKAa

[SD73a] B. V. Sukhatme and Isidoro P. David. A note on Koop’s approach for finding the bias of the ratio estimate. Journal of the American Statistical Association, 68(342):405–406, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284084.

Sukhatme:1973:NKAb

[SD73b] B. V. Sukhatme and Isidoro P. David. A note on Koop’s approach for finding the bias of the ratio estimate: Rejoinder. Journal of the American Statistical Association, 68(342):407–408, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284086.

Smith:1975:SPF

[SD75] M. H. Smith and J. J. Deely. A secretary problem with finite memory. Journal of the American Statistical Association, 70(350):357– 361, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285822.

Sedransk:1977:SPE

[Sed77] J. Sedransk. Sampling problems in the estimation of the money supply. Journal of the American Statistical Association, 72(359): 516–522, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286209.

Sedransk:1978:BRB

[Sed78] Joseph Sedransk. Book review: [untitled]. Journal of the American Statistical Association, 73(361):230, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286577.

Selvin:1974:MLE

[Sel74] Steve Selvin. Maximum likelihood estimation in the truncated or censored Poisson distribution. Journal of the American Statistical Association, 69(345):234–237, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285533.

Selvin:1975:BRB

[Sel75] Steve Selvin. Book review: [untitled]. Journal of the American Statistical Association, 70(350):488, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285851.

Sen:1971:BRB

[Sen71] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 66(335):659, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283561.

Sen:1972:BRB

[Sen72a] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 67(337):246–247, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284744.

Seneta:1972:PPV

[Sen72b] E. Seneta. Population projection variances and path analysis. Journal of the American Statistical Association, 67(339):617–619, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284451.

Seneta:1973:BRB

[Sen73] E. Seneta. Book review: [untitled]. Journal of the American Statistical Association, 68(342):498, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284119.

Sen:1974:UER

[Sen74] Pranab Kumar Sen. On unbiased estimation for randomized response models. Journal of the American Statistical Association, 69 (348):997–1001, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286178.

Sen:1975:BRB

[Sen75] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 70(352):967, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285487.

Sen:1976:DMG

[Sen76a] A. R. Sen. Developments in migratory game bird surveys. Journal of the American Statistical Association, 71(353):43–48, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285728.

Sen:1976:BRBa

[Sen76b] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 71(355):777, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285639.

Sen:1976:BRBb

[Sen76c] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 71(355):777, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285640.

Sen:1977:BRB

[Sen77] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 72(359):692, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286252.

Sen:1978:BRBa

[Sen78a] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 73(361):221, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286558.

Sen:1978:BRBb

[Sen78b] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 73(363):679, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286626.

Sen:1978:BRBc

[Sen78c] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 73(363):681–682, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286631.

Sen:1979:BRBa

[Sen79a] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 74(367):736–737, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287020.

Sen:1979:BRBb

[Sen79b] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 74(368):935, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286441.

Senter:1979:BRB

[Sen79c] Herman F. Senter. Book review: [untitled]. Journal of the American Statistical Association, 74(368):929, December 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286432.

Sudman:1971:EOC

[SF71] Seymour Sudman and Robert Ferber. Experiments in obtaining consumer expenditures by diary methods. Journal of the American Statistical Association, 66(336):725–735, December 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284219.

Shapiro:1972:AAV

[SF72a] S. S. Shapiro and R. S. Francia. An approximate analysis of variance test for normality. Journal of the American Statistical Association, 67(337):215–216, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284728.

Stephenson:1972:SAE

[SF72b] James A. Stephenson and Helen T. Farr. Seasonal adjustment of economic data by application of the general linear statistical model. Journal of the American Statistical Association, 67(337):37– 45, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284699.

Schucany:1971:BRE

[SGO71] W. R. Schucany, H. L. Gray, and D. B. Owen. On bias reduction in estimation. Journal of the American Statistical Association, 66 (335):524–533, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283519.

Starmer:1974:SRU

[SGS74] C. Frank Starmer, James E. Grizzle, and P. K. Sen. Some reasons for not using the Yates continuity correction on 2 × 2 contingency tables: Comment. Journal of the American Statistical Association, 69(346):376–378, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285662.

Sielken:1973:TLP

[SH73] R. L. Sielken, Jr. and H. O. Hartley. Two linear programming algorithms for unbiased estimation of linear models. Journal of the American Statistical Association, 68(343):639–641, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284792.

Spurrier:1975:DST

[SH75] John D. Spurrier and John E. Hewett. Double sample tests for the mean of a normal population. Journal of the American Statistical Association, 70(350):448–450, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285838.

Spurrier:1976:TSW

[SH76] John D. Spurrier and John E. Hewett. Two-stage Wilcoxon tests of hypotheses. Journal of the American Statistical Association, 71 (356):982–987, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286873.

Searle:1979:DMV

[SH79] S. R. Searle and Harold V. Henderson. Dispersion matrices for variance components models. Journal of the American Statistical Association, 74(366):465–470, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286357.

Shah:1970:LIC

[Sha70a] K. R. Shah. On the loss of information in combined inter- and intrablock estimation. Journal of the American Statistical Association, 65(332):1562–1564, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284337.

Shanbhag:1970:CEG

[Sha70b] D. N. Shanbhag. The characterizations for exponential and geometric distributions. Journal of the American Statistical Association, 65 (331):1256–1259, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284292.

Shapiro:1973:CBR

[Sha73a] S. H. Shapiro. On computing Bayes risks. Journal of the American Statistical Association, 68(342):485–486, June 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284104.

Shaw:1973:PPP

[Sha73b] William H. Shaw. Paradoxes, problems and progress. Journal of the American Statistical Association, 68(341):7–10, March 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284131.

Shaffer:1974:BUP

[Sha74] Juliet Popper Shaffer. Bidirectional unbiased procedures. Journal of the American Statistical Association, 69(346):437–439, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285674.

Shaman:1975:BRB

[Sha75] Paul Shaman. Book review: [untitled]. Journal of the American Statistical Association, 70(349):258, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285425.

Shafer:1976:BRB

[Sha76a] Glenn R. Shafer. Book review: The Emergence of Probability, by Ian Hacking. Journal of the American Statistical Association, 71(354):519–521, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285346.

Sharot:1976:GJF

[Sha76b] Trevor Sharot. The generalized jackknife: Finite samples and subsample sizes. Journal of the American Statistical Association, 71(354):451–454, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285332.

Shaked:1977:FCD

[Sha77a] Moshe Shaked. A family of concepts of dependence for bivariate distributions. Journal of the American Statistical Association, 72 (359):642–650, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286232.

Shapiro:1977:BRB

[Sha77b] Harold T. Shapiro. Book review: [untitled]. Journal of the American Statistical Association, 72(360):928, December 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286497.

Shamseldin:1978:BRB

[Sha78a] Adel Shamseldin. Book review: [untitled]. Journal of the American Statistical Association, 73(363):686, September 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286641.

Shapiro:1978:BRB

[Sha78b] Harold T. Shapiro. Book review: [untitled]. Journal of the American Statistical Association, 73(361):226–227, March 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286569.

Shepp:1972:BRB

[She72] L. A. Shepp. Book review: [untitled]. Journal of the American Statistical Association, 67(340):954–955, December 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284672.

Sheth:1975:BRB

[She75] Jagdish N. Sheth. Book review: [untitled]. Journal of the American Statistical Association, 70(349):250–251, March 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285414.

Shea:1978:SDT

[She78] Gerald Shea. Statistical diagnosis and tests of factor hypotheses. Journal of the American Statistical Association, 73(362):346–350, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286663.

Shenk:1979:BRB

[She79] Debra Shenk. Book review: [untitled]. Journal of the American Statistical Association, 74(367):732, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287010.

Speed:1978:MAL

[SHH78] F. M. Speed, R. R. Hocking, and O. P. Hackney. Methods of analysis of linear models with unbalanced data. Journal of the American Statistical Association, 73(361):105–112, March 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286530.

Shonick:1970:SMO

[Sho70] William Shonick. A stochastic model for occupancy-related random variables in general-acute hospitals. Journal of the American Statistical Association, 65(332):1474–1500, December 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284331.

Shonick:1972:CSM

[Sho72a] William Shonick. Corrigenda: a Stochastic Model for Occupancy-Related Random Variables in General-Acute Hospitals. Journal of the American Statistical Association, 67(337):253, March 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284765.

Shorack:1972:BTE

[Sho72b] Galen R. Shorack. The best test of exponentiality against gamma alternatives. Journal of the American Statistical Association, 67 (337):213–214, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284727.

Shorack:1972:BRB

[Sho72c] Galen R. Shorack. Book review: [untitled]. Journal of the American Statistical Association, 67(339):719–720, September 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284485.

Shonick:1973:CSM

[Sho73] William Shonick. Corrigenda: a Stochastic Model for Occupancy-Related-Random Variances in General-Acute Hospit. Journal of the American Statistical Association, 68(341):251, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284205.

Shore:1975:BRB

[Sho75] Roy E. Shore. Book review: [untitled]. Journal of the American Statistical Association, 70(350):490–491, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285857.

Shoemaker:1976:BRB

[Sho76] Robert W. Shoemaker. Book review: [untitled]. Journal of the American Statistical Association, 71(354):528, June 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285355.

Shumway:1970:ARA

[Shu70] Robert H. Shumway. Applied regression and analysis of variance for stationary time series. Journal of the American Statistical Association, 65(332):1527–1546, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284334.

Sichel:1975:DLW

[Sic75] H. S. Sichel. On a distribution law for word frequencies. Journal of the American Statistical Association, 70(351):542–547, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285930.

Siegmund:1972:BRB

[Sie72] David Siegmund. Book review: [untitled]. Journal of the American Statistical Association, 67(340):956, December 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284675.

Sievers:1978:WRS

[Sie78] Gerald L. Sievers. Weighted rank statistics for simple linear regression. Journal of the American Statistical Association, 73(363): 628–631, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286613.

Sims:1972:BRB

[Sim72a] Christopher Sims. Book review: [untitled]. Journal of the American Statistical Association, 67(337):243, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284738.

Sims:1972:RAP

[Sim72b] Christopher A. Sims. The role of approximate prior restrictions in distributed lag estimation. Journal of the American Statistical Association, 67(337):169–175, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284717.

Simon:1973:AIE

[Sim73] Gary Simon. Additivity of information in exponential family probability laws. Journal of the American Statistical Association, 68 (342):478–482, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284102.

Simon:1974:AAS

[Sim74a] Gary Simon. Alternative analyses for the singly-ordered contingency table. Journal of the American Statistical Association, 69 (348):971–976, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286174.

Sims:1974:SR

[Sim74b] Christopher A. Sims. Seasonality in regression. Journal of the American Statistical Association, 69(347):618–626, September 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285991.

Simon:1975:CAA

[Sim75a] Gary Simon. Corrigenda: Alternative Analyses for the Singly-Ordered Contingency Table. Journal of the American Statistical Association, 70(350):494, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285868.

Sims:1975:NET

[Sim75b] Christopher A. Sims. A note on exact tests for serial correlation. Journal of the American Statistical Association, 70(349):162–165, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285394.

Simons:1976:BRB

[Sim76] Gordon Simons. Book review: [untitled]. Journal of the American Statistical Association, 71(353):242–243, March 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285778.

Simon:1977:MGK

[Sim77] Gary Simon. Multivariate generalization of Kendall’s tau with application to data reduction. Journal of the American Statistical Association, 72(358):367–376, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286801.

Simon:1978:EMA

[Sim78a] Gary A. Simon. Efficacies of measures of association for ordinal contingency tables. Journal of the American Statistical Association, 73(363):545–551, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286598.

Simon:1978:PNA

[Sim78b] Gary A. Simon. Pseudorandom number assignment in statistically designed simulation and distribution sampling experiments: Comment. Journal of the American Statistical Association, 73(363):520–521, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286592;   http://www.jstor.org/stable/
pdfplus/2286592.pdf. See [SM78a].

Simkowitz:1979:BRB

[Sim79a] Michael A. Simkowitz. Book review: [untitled]. Journal of the American Statistical Association, 74(365):252, March 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286777.

Simon:1979:BRB

[Sim79b] Gary A. Simon. Book review: [untitled]. Journal of the American Statistical Association, 74(368):927–928, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286430.

Simpson:1979:FEW

[Sim79c] Joanne Simpson. Field experimentation in weather modification: Comment. Journal of the American Statistical Association, 74(365): 95–97, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286732.

Singh:1971:PMS

[Sin71a] J. Singh. Pooling mean squares. Journal of the American Statistical Association, 66(333):82–85, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284852.

Singh:1971:AOS

[Sin71b] Ravindra Singh. Approximately optimum stratification on the auxiliary variable. Journal of the American Statistical Association, 66 (336):829–833, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284235.

Singpurwalla:1971:PAL

[Sin71c] Nozer D. Singpurwalla. A problem in accelerated life testing. Journal of the American Statistical Association, 66(336):841–845, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284238.

Singh:1976:MRS

[Sin76] C. Singh. Moments of the range of samples from nonnormal populations. Journal of the American Statistical Association, 71 (356):988–991, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286874.

Sirken:1970:HSM

[Sir70] Monroe G. Sirken. Household surveys with multiplicity. Journal of the American Statistical Association, 65(329):257–266, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283590.

Sirken:1972:SSS

[Sir72] Monroe G. Sirken. Stratified sample surveys with multiplicity. Journal of the American Statistical Association, 67(337):224–227, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284732.

Somermeyer:1976:EQV

[SJL76] W. H. Somermeyer, R. Jansen, and A. S. Louter. Estimating quarterly values of annually known variables in quarterly relationships. Journal of the American Statistical Association, 71(355):588–595, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285587.

Stefansky:1973:NDA

[SK73] W. Stefansky and Henry F. Kaiser. Note on discrete approximations. Journal of the American Statistical Association, 68(341):232– 234, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284177.

Sawa:1976:CCR

[SK76] Takamitsu Sawa and Satoru Kanoh. Categorical control in regression. Journal of the American Statistical Association, 71(354):361– 365, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285314.

Skibinsky:1970:CHD

[Ski70] Morris Skibinsky. A characterization of hypergeometric distributions. Journal of the American Statistical Association, 65(330):926–929, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284598.

Sirken:1974:CME

[SL74a] Monroe G. Sirken and Paul S. Levy. Corrigenda: Multiplicity Estimation of Proportions Based on Ratios of Random Variables. Journal of the American Statistical Association, 69(346):587, June 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285722.

Sirken:1974:MEP

[SL74b] Monroe G. Sirken and Paul S. Levy. Multiplicity estimation of proportions based on ratios of random variables. Journal of the American Statistical Association, 69(345):68–73, March 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285501.

Suchindran:1974:EPP

[SL74c] C. M. Suchindran and P. A. Lachenbruch. Estimates of parameters in a probability model for first livebirth interval. Journal of the American Statistical Association, 69(346):507–513, June 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285688.

Schoen:1979:GAE

[SL79] Robert Schoen and Kenneth C. Land. A general algorithm for estimating a Markov-generated increment-decrement life table with applications to marital-status patterns. Journal of the American Statistical Association, 74(368):761–776, December 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286398.

Slater:1973:STE

[Sla73] Paul B. Slater. Spatial and temporal effects in residential sales prices. Journal of the American Statistical Association, 68(343):554–561, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284775.

Slater:1974:DST

[Sla74] Paul B. Slater. Disaggregated spatial-temporal analyses of residential sales prices. Journal of the American Statistical Association, 69(346):358–363, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285658.

Sly:1972:EEN

[Sly72] David F. Sly. Evaluating estimates of net migration and net migration rates based on survival ratios corrected in varying degrees. Journal of the American Statistical Association, 67(338):313–318, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284375.

Swamy:1973:BAE

[SM73] P. A. V. B. Swamy and J. S. Mehta. Bayesian analysis of error components regression models. Journal of the American Statistical Association, 68(343):648–658, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284794.

Swamy:1975:BNB

[SM75] P. A. V. B. Swamy and J. S. Mehta. Bayesian and non-Bayesian analysis of switching regressions and of random coefficient regression models. Journal of the American Statistical Association, 70(351):593–602, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285939.

Schmidt:1977:NAA

[SM77a] Peter Schmidt and W. Robert Mann. A note on the approximation of arbitrary distributed lag structures by a modified Almon lag. Journal of the American Statistical Association, 72(358):442–443, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286815.

Swamy:1977:ELM

[SM77b] P. A. V. B. Swamy and J. S. Mehta. Estimation of linear models with time and cross-sectionally varying coefficients. Journal of the American Statistical Association, 72(360):890–898, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286482.

Schruben:1978:PNAa

[SM78a] Lee W. Schruben and Barry H. Margolin. Pseudorandom number assignment in statistically designed simulation and distribution sampling experiments. Journal of the American Statistical Association, 73(363):504–520, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286590;   http://www.jstor.org/stable/

pdfplus/2286590.pdf. See discussion [Gav78, Kie78a, Mal78, SM78b, Sim78b].

Schruben:1978:PNAb

[SM78b] Lee W. Schruben and Barry H. Margolin. Pseudorandom number assignment in statistically designed simulation and distribution sampling experiments: Rejoinder. Journal of the American Statistical Association, 73(363):524–525, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286595;   http://www.jstor.org/stable/
pdfplus/2286595.pdf. See [SM78a].

Subbaiah:1978:CTT

[SM78c] Perla Subbaiah and Govind S. Mudholkar. A comparison of two tests for the significance of a mean vector. Journal of the American Statistical Association, 73(362):414–418, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286675.

Saniga:1979:PSS

[SM79] Erwin M. Saniga and James A. Miles. Power of some standard goodness-of-fit tests of normality against asymmetric stable alternatives. Journal of the American Statistical Association, 74 (368):861–865, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286412.

Smith:1971:BRBa

[Smi71a] J. E. Keith Smith. Book review: [untitled]. Journal of the American Statistical Association, 66(334):425, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283954.

Smith:1971:BRBb

[Smi71b] V. Kerry Smith. Book review: [untitled]. Journal of the American Statistical Association, 66(336):924, December 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284267.

Smith:1975:MAP

[Smi75a] V. Kerry Smith. Mortality–air pollution relationships: a comment. Journal of the American Statistical Association, 70(350):341–343, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285818.

Smith:1975:SAP

[Smi75b] V. Kerry Smith. A simulation analysis of the power of several tests for detecting heavy-tailed distributions. Journal of the American Statistical Association, 70(351):662–665, September 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285951.

Smith:1976:EPS

[Smi76] V. Kerry Smith. The estimated power of several tests for autocorrelation with non-first- order alternatives. Journal of the American Statistical Association, 71(356):879–883, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286854.

Smith:1977:SSA

[Smi77a] A. F. M. Smith. A simulation study of alternatives to ordinary least squares: Comment. Journal of the American Statistical Association, 72(357):96–97, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286913.

Smith:1977:BRBa

[Smi77b] J. Terry Smith. Book review: [untitled]. Journal of the American Statistical Association, 72(357):232, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286951.

Smith:1977:BRBb

[Smi77c] V. Kerry Smith. Book review: [untitled]. Journal of the American Statistical Association, 72(358):484–485, June 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286829.

Smith:1979:BRBa

[Smi79a] James P. Smith. Book review: [untitled]. Journal of the American Statistical Association, 74(365):250, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286773.

Smith:1979:BRBb

[Smi79b] V. Kerry Smith. Book review: [untitled]. Journal of the American Statistical Association, 74(365):250–251, March 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286774.

Smith:1979:BRBc

[Smi79c] V. Kerry Smith. Book review: [untitled]. Journal of the American Statistical Association, 74(365):251, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286775.

Sheps:1970:TEC

[SMRL70] Mindel C. Sheps, Jane A. Menken, Jeanne Clare Ridley, and Joan W. Lingner. Truncation effect in closed and open birth interval data. Journal of the American Statistical Association, 65(330):678–693, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284579.

Sobel:1971:ABL

[Sob71] Eugene Sobel. Approximate best linear prediction of a certain class of stationary and nonstationary noise-distorted signals. Journal of the American Statistical Association, 66(334):363–372, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283937.

Sokal:1972:BRB

[Sok72] Robert R. Sokal. Book review: [untitled]. Journal of the American Statistical Association, 67(340):960, December 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284681.

Solomon:1972:MEM

[Sol72a] Daniel L. Solomon. Λ minimax estimation of a multivariate location parameter. Journal of the American Statistical Association, 67(339):641–646, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284457.

Solomon:1972:MES

[Sol72b] Daniel L. Solomon. Λ minimax estimation of a scale parameter. Journal of the American Statistical Association, 67(339):647–649, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284458.

Somerville:1970:OSS

[Som70] Paul N. Somerville. Optimum sample size for a problem in choosing the population with the largest mean. Journal of the American Statistical Association, 65(330):763–775, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284586.

Somerville:1974:OSS

[Som74] Paul N. Somerville. Optimum sample size for a problem in choosing the population with the largest mean: Some comments on Somerville’s paper: Rejoinder. Journal of the American Statistical Association, 69(345):270–271, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285543.

Somerville:1975:OSS

[Som75] Paul N. Somerville. Optimum sample size for choosing the population having the smallest variance (large sample size results). Journal of the American Statistical Association, 70(352):852–858, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285447.

Soms:1976:AET

[Som76] Andrew P. Soms. An asymptotic expansion for the tail area of the t-distribution. Journal of the American Statistical Association, 71 (355):728–730, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285610.

Soms:1977:ADF

[Som77] Andrew P. Soms. An algorithm for the discrete Fisher’s permutation test. Journal of the American Statistical Association, 72 (359):662–664, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286236.

Siegel:1976:NEN

[SP76] Jacob S. Siegel and Jeffrey S. Passel. New estimates of the number of centenarians in the United States. Journal of the American Statistical Association, 71(355):559–566, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285583.

Spitzer:1978:MCI

[Spi78] John J. Spitzer. A Monte Carlo investigation of the Box–Cox transformation in small samples. Journal of the American Statistical Association, 73(363):488–495, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286587.

Spitzer:1979:CMC

[Spi79a] John J. Spitzer. Corrigenda: a Monte Carlo Investigation of the Box–Cox Transformation in Small Samples. Journal of the American Statistical Association, 74(367):744, September 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2287028.

Spitzer:1979:SSP

[Spi79b] John J. Spitzer. Small-sample properties of nonlinear least squares and maximum likelihood estimators in the context of autocorrelated errors. Journal of the American Statistical Association, 74(365):41– 47, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286718.

Srikantan:1970:CAB

[Sri70] K. S. Srikantan. Canonical association between nominal measurements. Journal of the American Statistical Association, 65(329):284– 292, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2283592.

Srinath:1971:MSN

[Sri71a] K. P. Srinath. Multiphase sampling in nonresponse problems. Journal of the American Statistical Association, 66(335):583–586, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283533.

Srivastava:1971:GEM

[Sri71b] Surendra K. Srivastava. A generalized estimator for the mean of a finite population using multi-auxiliary information. Journal of the American Statistical Association, 66(334):404–407, June 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283945.

Srivastava:1972:BRB

[Sri72a] J. N. Srivastava. Book review: [untitled]. Journal of the American Statistical Association, 67(339):720–721, September 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284487.

Srivastava:1972:PMS

[Sri72b] S. R. Srivastava. Pooling mean squares in ANOVA model II. Journal of the American Statistical Association, 67(339):676–679, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284466.

Srivastava:1972:DVE

[Sri72c] V. K. Srivastava. Disturbance-variance estimation in simultaneous equations when disturbances are small. Journal of the American Statistical Association, 67(337):164–168, March 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284716.

Srivastava:1973:CDV

[Sri73] V. K. Srivastava. Corrigenda: Disturbance-Variance Estimation in Simultaneous Equations When Disturbances Are Small. Journal of the American Statistical Association, 68(341):251, March 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284206.

Srivastava:1974:TCG

[Sri74] R. C. Srivastava. Two characterizations of the geometric distribution. Journal of the American Statistical Association, 69(345):267– 269, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285541.

Srivastava:1975:LTP

[Sri75] T. N. Srivastava. Life tests with periodic change in failure rate– grouped observations. Journal of the American Statistical Association, 70(350):394–397, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285829.

Scott:1971:BES

[SS71] Alastair Scott and T. M. F. Smith. Bayes estimates for subclasses in stratified sampling. Journal of the American Statistical Association, 66(336):834–836, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284236.

Spjotvoll:1973:EMM

[SS73a] Emil Spjotvoll and Michael R. Stoline. An extension of the T -method of multiple comparison to include the cases with unequal sample sizes. Journal of the American Statistical Association, 68(344):975–978, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284534.

Stekler:1973:FIL

[SS73b] H. O. Stekler and Martin Schepsman. Forecasting with an index of leading series. Journal of the American Statistical Association, 68(342):291–296, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284065.

Scott:1974:ARS

[SS74] A. J. Scott and T. M. F. Smith. Analysis of repeated surveys using time series methods. Journal of the American Statistical Association, 69(347):674–678, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286000.

Schucany:1977:IKT

[SS77a] W. R. Schucany and John P. Sommers. Improvement of kernel type density estimators. Journal of the American Statistical Association, 72(358):420–423, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286810.

Solomon:1977:DSW

[SS77b] Herbert Solomon and Michael A. Stephens. Distribution of a sum of weighted chi-square variables. Journal of the American Statistical Association, 72(360):881–885, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286480.

Singh:1978:SSS

[SS78a] Bahadur Singh and J. Sedransk. Sample size selection in regression analysis when there is nonresponse. Journal of the American Statistical Association, 73(362):362–365, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286666.

Solomon:1978:ADF

[SS78b] Herbert Solomon and Michael A. Stephens. Approximations to density functions using Pearson curves. Journal of the American Statistical Association, 73(361):153–160, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286537.

Sedransk:1979:DAD

[SS79] N. Sedransk and J. Sedransk. Distinguishing among distributions using data from complex sample designs. Journal of the American Statistical Association, 74(368):754–760, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286397.

Subrahmaniam:1975:RST

[SSM75] Kocherlakota Subrahmaniam, Kathleen Subrahmaniam, and J. Y. Messeri. On the robustness of some tests of significance in sampling from a compound normal population. Journal of the American Statistical Association, 70(350):435–438, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285835.

Sukhatme:1975:ASS

[ST75] B. V. Sukhatme and Victor K. T. Tang. Allocation in stratified sampling subsequent to preliminary test of significance. Journal of the American Statistical Association, 70(349):175–179, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285397.

Singh:1976:SED

[ST76a] M. P. Singh and R. Tessier. Some estimators for domain totals. Journal of the American Statistical Association, 71(354):322–325, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285304.

Sobel:1976:ENP

[ST76b] Milton Sobel and Y. L. Tong. Estimation of a normal percentile by grouping. Journal of the American Statistical Association, 71(353):189–192, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285766.

Stark:1975:SEI

[Sta75] A. E. Stark. Some estimators of the integer-valued parameter of a Poisson variate. Journal of the American Statistical Association, 70(351):685–689, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285956.

Stafford:1978:BRB

[Sta78] Frank P. Stafford. Book review: [untitled]. Journal of the American Statistical Association, 73(364):897–898, December 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286317.

Starr:1979:BRB

[Sta79] Ross M. Starr. Book review: [untitled]. Journal of the American Statistical Association, 74(367):737, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287022.

Steffens:1970:PBS

[Ste70a] F. E. Steffens. Power of bivariate Studentized maximum and minimum modulus tests. Journal of the American Statistical Association, 65(332):1639–1644, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284347.

Stevenson:1970:BRB

[Ste70b] Harold W. Stevenson. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1415–1416, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284318.

Sterling:1971:RCE

[Ste71] Theodor D. Sterling. A review of the claim that excess morbidity and disability can be ascribed to smoking. Journal of the American Statistical Association, 66(334):251–257, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283917.

Stephens:1972:MTM

[Ste72] M. A. Stephens. Multisample tests for the von Mises distribution. Journal of the American Statistical Association, 67(338):456–461, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284405.

Sterling:1973:CRC

[Ste73] Theodor D. Sterling. Comment on “review of claim that excess morbidity and disability can be ascribed to smoking”: Rejoinder. Journal of the American Statistical Association, 68(341):87, March 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284145.

Stephens:1974:ESG

[Ste74] M. A. Stephens. EDF statistics for goodness of fit and some comparisons. Journal of the American Statistical Association, 69 (347):730–737, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286009.

Stephenson:1975:BRB

[Ste75] Hal W. Stephenson. Book review: [untitled]. Journal of the American Statistical Association, 70(350):490, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285856.

Stene:1976:OUM

[Ste76a] Jon Stene. On the optimal use of multiauxiliary information. Journal of the American Statistical Association, 71(355):679, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285600.

Steyn:1976:MPN

[Ste76b] H. S. Steyn. On the multivariate Poisson normal distribution. Journal of the American Statistical Association, 71(353):233–236, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285775.

Stewart:1978:OSR

[Ste78a] Theodor J. Stewart. Optimal selection from a random sequence with learning of the underlying distribution. Journal of the American Statistical Association, 73(364):775–780, December 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286279.

Steyn:1978:ACN

[Ste78b] H. S. Steyn. On approximations for the central and noncentral distribution of the generalized variance. Journal of the American Statistical Association, 73(363):670–675, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286622.

Stewart:1979:MUB

[Ste79] Leland Stewart. Multiparameter univariate Bayesian analysis. Journal of the American Statistical Association, 74(367):684–693, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286992.

Stigler:1971:OED

[Sti71] Stephen M. Stigler. Optimal experimental design for polynomial regression. Journal of the American Statistical Association, 66 (334):311–318, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283928.

Stigler:1973:BRB

[Sti73a] Stephen M. Stigler. Book review: [untitled]. Journal of the American Statistical Association, 68(344):1030, December 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284556.

Stigler:1973:SNP

[Sti73b] Stephen M. Stigler. Simon Newcomb, Percy Daniell, and the history of robust estimation 1885–1920. Journal of the American Statistical Association, 68(344):872–879, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284515.

Stigler:1975:BRB

[Sti75] Stephen M. Stigler. Book review: [untitled]. Journal of the American Statistical Association, 70(352):957–959, December 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285467.

Stigler:1976:ESH

[Sti76] Stephen M. Stigler. The effect of sample heterogeneity on linear functions of order statistics, with applications to robust estimation. Journal of the American Statistical Association, 71(356):956–960, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286868.

Stigler:1977:BRB

[Sti77a] Stephen M. Stigler. Book review: [untitled]. Journal of the American Statistical Association, 72(357):239, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286964.

Stigler:1977:ECS

[Sti77b] Stephen M. Stigler. Eight centuries of sampling inspection: The trial of the pyx. Journal of the American Statistical Association, 72(359):493–500, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286206.

Stigler:1977:FOS

[Sti77c] Stephen M. Stigler. Fractional order statistics, with applications. Journal of the American Statistical Association, 72(359):544–550, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286215.

Stidham:1978:BRB

[Sti78a] Shaler Stidham, Jr. Book review: [untitled]. Journal of the American Statistical Association, 73(362):446–447, June 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286705.

Stigler:1978:BRBa

[Sti78b] Stephen M. Stigler. Book review: [untitled]. Journal of the American Statistical Association, 73(361):230, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286576.

Stigler:1978:BRBb

[Sti78c] Stephen M. Stigler. Book review: [untitled]. Journal of the American Statistical Association, 73(362):439, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286684.

Stone:1970:BRB

[Sto70] M. Stone. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1392–1396, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284304.

Stone:1971:CBM

[Sto71] Leroy O. Stone. On the correlation between metropolitan area in- and out-migration by occupation. Journal of the American Statistical Association, 66(336):693–701, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284214.

Stone:1976:SIUa

[Sto76a] Mervyn Stone. Strong inconsistency from uniform priors. Journal of the American Statistical Association, 71(353):114–116, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285744.

Stone:1976:SIUb

[Sto76b] Mervyn Stone. Strong inconsistency from uniform priors: Rejoinder. Journal of the American Statistical Association, 71(353):124– 125, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285752.

Stoline:1978:TSA

[Sto78] Michael R. Stoline. Tables of the Studentized augmented range and applications to problems of multiple comparison. Journal of the American Statistical Association, 73(363):656–660, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286619.

Strand:1971:BRB

[Str71] Otto Neall Strand. Book review: [untitled]. Journal of the American Statistical Association, 66(335):660, September 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283563.

Straf:1972:BRB

[Str72a] Miron L. Straf. Book review: [untitled]. Journal of the American Statistical Association, 67(337):246, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284743.

Stroud:1972:CCMa

[Str72b] T. W. F. Stroud. Comparing conditional means and variances in a regression model with measurement errors of known variances. Journal of the American Statistical Association, 67(338):407–412, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284393.

Stroud:1972:CCMb

[Str72c] T. W. F. Stroud. Comparing conditional means and variances in a regression model with measurement errors of known variances: Rejoinder. Journal of the American Statistical Association, 67(338):413– 414, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284395.

Stroud:1973:CCC

[Str73] T. W. F. Stroud. Corrigenda: Comparing Conditional Means and Variances in a Regression Model with Measurement Errors of Known Variances. Journal of the American Statistical Association, 68(341): 251, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284207.

Stroud:1974:CUE

[Str74] T. W. F. Stroud. Combining unbiased estimates of a parameter known to be positive. Journal of the American Statistical Association, 69(346):502–506, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285687.

Stroud:1975:CCU

[Str75] T. W. F. Stroud. Corrigenda: Combining Unbiased Estimates of a Parameter Known to be Positive. Journal of the American Statistical Association, 70(350):494, June 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285869.

Strahan:1976:BRB

[Str76] Robert F. Strahan. Book review: [untitled]. Journal of the American Statistical Association, 71(353):251, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285793.

Strawderman:1978:MAG

[Str78] William E. Strawderman. Minimax adaptive generalized ridge regression estimators. Journal of the American Statistical Association, 73(363):623–627, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286612.

Stuart:1971:DOS

[Stu71] Alan Stuart. Does an observed sequence of numbers follow a simple rule? (another look at Bode’s law): Comment. Journal of the American Statistical Association, 66(335):564–565, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283525.

Shumway:1974:LDF

[SU74a] R. H. Shumway and A. N. Unger. Linear discriminant functions for stationary time series. Journal of the American Statistical Association, 69(348):948–956, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286169.

Singh:1974:ESU

[SU74b] Balvir Singh and Aman Ullah. Estimation of seemingly unrelated regressions with random coefficients. Journal of the American Statistical Association, 69(345):191–195, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285522.

Subrahmanya:1972:TEQ

[Sub72] M. T. Subrahmanya. Theory of equitable quality level and errorareas under the OC curves of acceptance sampling plans. Journal of the American Statistical Association, 67(337):234–240, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284735.

Suchindran:1974:BRB

[Suc74] C. M. Suchindran. Book review: [untitled]. Journal of the American Statistical Association, 69(346):582–583, June 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285712.

Sudman:1970:MUP

[Sud70] Seymour Sudman. The multiple uses of primary sampling areas of national probability samples. Journal of the American Statistical Association, 65(329):61–70, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283575.

Sudman:1972:SVR

[Sud72] Seymour Sudman. On sampling of very rare human populations. Journal of the American Statistical Association, 67(338):335–339, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284378.

Sudman:1978:OCD

[Sud78] Seymour Sudman. Optimum cluster designs within a primary unit using combined telephone screening and face-to-face interviewing. Journal of the American Statistical Association, 73(362):300–304, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286656.

Sun:1979:BRB

[Sun79] T. C. Sun. Book review: [untitled]. Journal of the American Statistical Association, 74(366):503–504, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286366.

Sutherland:1978:BRB

[Sut78] Michael Sutherland. Book review: [untitled]. Journal of the American Statistical Association, 73(361):215–217, March 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286551.

Susarla:1976:NBE

[SV76] V. Susarla and J. Van Ryzin. Nonparametric Bayesian estimation of survival curves from incomplete observations. Journal of the American Statistical Association, 71(356):897–902, December 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286858.

Sobel:1971:PWR

[SW71] Milton Sobel and George H. Weiss. Play-the-winner rule and inverse sampling in selecting the better of two binomial populations. Journal of the American Statistical Association, 66(335):545–551, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283521.

Schmidt:1973:ALT

[SW73] Peter Schmidt and Roger N. Waud. The Almon lag technique and the monetary versus fiscal policy debate. Journal of the American Statistical Association, 68(341):11–19, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284132.

Serfling:1976:ATS

[SW76] R. J. Serfling and D. D. Wackerly. Asymptotic theory of sequential fixed-width confidence interval procedures. Journal of the American Statistical Association, 71(356):949–955, December 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286867.

Serfling:1977:CAT

[SW77] R. J. Serfling and D. D. Wackerly. Corrigenda: Asymptotic Theory of Sequential Fixed-Width Confidence Interval Procedures. Journal of the American Statistical Association, 72(358):488, June 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286838.

Shapiro:1978:BSP

[SW78a] C. P. Shapiro and Robert L. Wardrop. The Bayes sequential procedure for estimating the arrival rate of a Poisson process. Journal of the American Statistical Association, 73(363):597–601, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286607.

Skillings:1978:DFT

[SW78b] John H. Skillings and Douglas A. Wolfe. Distribution-free tests for ordered alternatives in a randomized block design. Journal of the American Statistical Association, 73(362):427–431, June 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286678.

Summers:1978:EHI

[SW78c] Anita A. Summers and Barbara L. Wolfe. Estimating household income from location. Journal of the American Statistical Association, 73(362):288–292, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286654.

Sinha:1979:UIT

[SW79] Bimal Kumar Sinha and H. S. Wieand. Union-intersection test for the mean vector when the covariance matrix is totally reducible. Journal of the American Statistical Association, 74(366):340–343, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286332.

Switzer:1973:BRB

[Swi73] Paul Switzer. Book review: [untitled]. Journal of the American Statistical Association, 68(341):243, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284187.

Switzer:1976:BRB

[Swi76] Paul Switzer. Book review: [untitled]. Journal of the American Statistical Association, 71(355):777, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285641.

Switzer:1978:BRB

[Swi78] Paul Switzer. Book review: [untitled]. Journal of the American Statistical Association, 73(363):684, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286636.

Swindel:1972:ASS

[SY72] Benee F. Swindel and David O. Yandle. Allocation in stratified sampling as a game. Journal of the American Statistical Association, 67(339):684–686, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284468.

Sykes:1974:BRB

[Syk74] Zenas M. Sykes. Book review: [untitled]. Journal of the American Statistical Association, 69(345):276, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285547.

Solomon:1970:ODS

[SZ70] H. Solomon and S. Zacks. Optimal design of sampling from finite populations: a critical review and indication of new research areas. Journal of the American Statistical Association, 65(330):653–677, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284578.

Srivastava:1973:MCC

[SZ73] J. N. Srivastava and M. K. Zaatar. A Monte Carlo comparison of four estimators of the dispersion matrix of a bivariate normal population, using incomplete data. Journal of the American Statistical Association, 68(341):180–183, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284164.

Shaked:1979:NAA

[SZB79] Moshe Shaked, William J. Zimmer, and Clifford A. Ball. A nonparametric approach to accelerated life testing. Journal of the American Statistical Association, 74(367):694–699, September 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286993.

Taubman:1978:STL

[TA78] Stephen B. Taubman and Sidney Addelman. Some two-level multipleresponse factorial plans. Journal of the American Statistical Association, 73(363):607–612, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286609.

Taeuber:1976:BRB

[Tae76] Richard C. Taeuber. Book review: [untitled]. Journal of the American Statistical Association, 71(354):524–526, June 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285350.

Taeuber:1977:BRB

[Tae77] Richard C. Taeuber. Book review: [untitled]. Journal of the American Statistical Association, 72(359):689–690, September 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286248.

Takacs:1971:BRB

[Tak71a] Lajos Takacs. Book review: [untitled]. Journal of the American Statistical Association, 66(334):429, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283963.

Takeuchi:1971:UAE

[Tak71b] Kei Takeuchi. A uniformly asymptotically efficient estimator of a location parameter. Journal of the American Statistical Association, 66(334):292–301, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283926.

Tamura:1972:BRB

[Tam72] H. Tamura. Book review: [untitled]. Journal of the American Statistical Association, 67(339):715, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284479.

Tamhane:1979:CPM

[Tam79] Ajit C. Tamhane. A comparison of procedures for multiple comparisons of means with unequal variances. Journal of the American Statistical Association, 74(366):471–480, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286358.

Tanis:1978:BRB

[Tan78] Elliot A. Tanis. Book review: [untitled]. Journal of the American Statistical Association, 73(361):219, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286554.

Tarter:1979:NSD

[Tar79a] Michael E. Tarter. Nonparametric statistical data modeling: Comment. Journal of the American Statistical Association, 74(365):128, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286739.

Tarter:1979:TML

[Tar79b] Michael E. Tarter. Trigonometric maximum likelihood estimation and application to the analysis of incomplete survival information. Journal of the American Statistical Association, 74(365):132–139, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286742.

Tatsuoka:1971:BRB

[Tat71] Maurice M. Tatsuoka. Book review: [untitled]. Journal of the American Statistical Association, 66(335):651–652, September 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283550.

Taylor:1974:BRBa

[Tay74a] Lester D. Taylor. Book review: [untitled]. Journal of the American Statistical Association, 69(347):830, September 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286030.

Taylor:1974:BRBb

[Tay74b] Lester D. Taylor. Book review: [untitled]. Journal of the American Statistical Association, 69(347):832, September 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286034.

Tate:1970:NCT

[TB70] Merle W. Tate and Sara M. Brown. Note on the Cochran Q test. Journal of the American Statistical Association, 65(329):155–160, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2283582.

Tietjen:1972:TUM

[TB72] G. L. Tietjen and R. J. Beckman. Tables for use of the maximum F -ratio in multiple comparison procedur. Journal of the American Statistical Association, 67(339):581–583, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284442.

Thomas:1977:ATM

[TB77a] Marlin U. Thomas and Donald R. Barr. An approximate test of Markov chain lumpability. Journal of the American Statistical Association, 72(357):175–179, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286933.

Turner:1977:SCB

[TB77b] David L. Turner and David C. Bowden. Simultaneous confidence bands for percentile lines in the general linear model. Journal of the American Statistical Association, 72(360):886–889, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286481.

Turner:1979:SCB

[TB79] David L. Turner and David C. Bowden. Sharp confidence bands for percentile lines and tolerance bands for the simple linear model. Journal of the American Statistical Association, 74(368):885–888, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286418.

Turnbull:1974:SAH

[TBH74] Bruce W. Turnbull, Bryon Wm. Brown, Jr., and Marie HU. Survivorship analysis of heart transplant data. Journal of the American Statistical Association, 69(345):74–80, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285502.

Tan:1972:SCM

[TC72] W. Y. Tan and W. C. Chang. Some comparisons of the method of moments and the method of maximum likelihood in estimating parameters of a mixture of two normal densities. Journal of the American Statistical Association, 67(339):702–708, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284472.

Tsai:1975:SSB

[TDL75] W. S. Tsai, B. S. Duran, and T. O. Lewis. Small-sample behavior of some multisample nonparametric tests for scale. Journal of the American Statistical Association, 70(352):791–796, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285438.

Teichmoeller:1971:NDS

[Tei71] John Teichmoeller. A note on the distribution of stock price changes. Journal of the American Statistical Association, 66(334):282–284, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283922.

Tenenbein:1970:DSS

[Ten70] Aaron Tenenbein. A double sampling scheme for estimating from binomial data with misclassifications. Journal of the American Statistical Association, 65(331):1350–1361, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284301.

Theil:1973:RPC

[TF73] Henri Theil and Yoram Friedman. Regional per capita incomes and income inequalities: Point estimates and their standard errors. Journal of the American Statistical Association, 68(343):531–539, September 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284772.

Taubman:1971:AWC

[TG71a] Paul Taubman and Peter Gottschalk. The average workweek of capital in manufacturing. Journal of the American Statistical Association, 66(335):448–455, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283507.

Thomas:1971:SSP

[TG71b] Donald G. Thomas and John J. Gart. Small sample performance of some estimators of the truncated binomial distribution. Journal of the American Statistical Association, 66(333):169–177, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284868.

Tarone:1975:NUR

[TG75a] Robert E. Tarone and Gary Gruenhage. A note on the uniqueness of roots of the likelihood equations for vector-valued parameters. Journal of the American Statistical Association, 70(352):903–904, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285455.

Thomas:1975:CIE

[TG75b] David R. Thomas and Gary L. Grunkemeier. Confidence interval estimation of survival probabilities for censored data. Journal of the American Statistical Association, 70(352):865–871, December 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285449.

Thomas:1977:TEC

[TG77] Donald G. Thomas and John J. Gart. A table of exact confidence limits for differences and ratios of two proportions and their odds ratios. Journal of the American Statistical Association, 72(357):73– 76, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286908.

Tarone:1979:CNU

[TG79] Robert E. Tarone and Gary Gruenhage. Corrigenda: a Note on the Uniqueness of Roots of the Likelihood Equations for Vector-Valued Parameters. Journal of the American Statistical Association, 74(367):744, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287030.

Tate:1973:ITG

[TH73] Merle W. Tate and Leon A. Hye. Inaccuracy of the χ^{2} test of goodness of fit when expected frequenies are small. Journal of the American Statistical Association, 68(344):836–841, December 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284509.

Thall:1979:BRB

[Tha79a] Peter F. Thall. Book review: [untitled]. Journal of the American Statistical Association, 74(365):245–246, March 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286764.

Thall:1979:HSR

[Tha79b] Peter F. Thall. Huber-sense robust M-estimation of a scale parameter, with applications to the exponential distribution. Journal of the American Statistical Association, 74(365):147–152, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286744.

Theil:1970:CLR

[The70] Henri Theil. The cube law revisited. Journal of the American Statistical Association, 65(331):1213–1219, September 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284287.

Theil:1974:TRR

[The74] Henri Theil. A theory of rational random behavior. Journal of the American Statistical Association, 69(346):310–314, June 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285649.

Thisted:1977:BRB

[Thi77a] Ronald A. Thisted. Book review: [untitled]. Journal of the American Statistical Association, 72(359):692–693, September 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286254.

Thisted:1977:SSA

[Thi77b] Ronald A. Thisted. A simulation study of alternatives to ordinary least squares: Comment. Journal of the American Statistical Association, 72(357):102–103, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286915.

Thoni:1970:CTE

[Tho70] Hanspeter Thoni. Corrigenda: a Table for Estimating the Mean of a Lognormal Distribution. Journal of the American Statistical Association, 65(330):1011–1012, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284623.

Thornton:1973:BRB

[Tho73a] Judith A. Thornton. Book review: [untitled]. Journal of the American Statistical Association, 68(342):496–497, June 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284116.

Thorp:1973:NSA

[Tho73b] Edward O. Thorp. Nonrandom shuffling with applications to the game of Faro. Journal of the American Statistical Association, 68 (344):842–847, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284510.

Thompson:1975:GSF

[Tho75] Mark Thompson. On any given Sunday: Fair competitor orderings with maximum likelihood methods. Journal of the American Statistical Association, 70(351):536–541, September 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285929.

Thomas:1976:RML

[Tho76] D. L. Thomas. Reciprocal moments of linear combinations of exponential variates. Journal of the American Statistical Association, 71(354):506–512, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285342.

Thomsen:1977:ESW

[Tho77] Ib Thomsen. On the effect of stratification when two stratifying variables are used. Journal of the American Statistical Association, 72(357):149–153, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286926.

Thompson:1979:BMG

[Tho79] Rory O. R. Y. Thompson. Bias and monotonicity for goodnessof-fit tests. Journal of the American Statistical Association, 74 (368):875–876, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286415.

Toothaker:1978:OSS

[THP78] Larry E. Toothaker, Jana L. Hicks, and James M. Price. Optimum subsample sizes for the Bartlett–Kendall homogeneity of variance test. Journal of the American Statistical Association, 73(361):53–57, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286515.

Thursby:1979:ASE

[Thu79] Jerry G. Thursby. Alternative specification error tests: a comparative study. Journal of the American Statistical Association, 74(365):222– 225, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286754.

Tiahrt:1975:BRB

[Tia75] Kenneth J. Tiahrt. Book review: [untitled]. Journal of the American Statistical Association, 70(351):732–733, September 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285972.

Tihansky:1972:PBN

[Tih72] Dennis P. Tihansky. Properties of the bivariate normal cumulative distribution. Journal of the American Statistical Association, 67(340):903–905, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284658.

Tiku:1971:PFT

[Tik71] M. L. Tiku. Power function of the F -test under non-normal situations. Journal of the American Statistical Association, 66 (336):913–916, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284254.

Tiku:1972:MTP

[Tik72] M. L. Tiku. More tables of the power of the F -test. Journal of the American Statistical Association, 67(339):709–710, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284473.

Tin:1970:CCS

[Tin70] Myint Tin. Corrigenda: Comparison of Some Ratio Estimators. Journal of the American Statistical Association, 65(330):1012, June 1970.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284624.

Tiao:1970:UOE

[TL70] G. C. Tiao and David R. Lund. The use of OLUMV estimators in inference robustness studies of the location parameter of a class of symmetric distributions. Journal of the American Statistical Association, 65(329):370–386, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283599.

Tobler:1979:SPIa

[Tob79a] Waldo R. Tobler. Smooth pycnophylactic interpolation for geographical regions. Journal of the American Statistical Association, 74 (367):519–530, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286968.

Tobler:1979:SPIb

[Tob79b] Waldo R. Tobler. Smooth pycnophylactic interpolation for geographical regions: Rejoinder. Journal of the American Statistical Association, 74(367):535–536, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286970.

Tong:1970:SPI

[Ton70] Yung Liang Tong. Some probability inequalities of multivariate normal and multivariate t. Journal of the American Statistical Association, 65(331):1243–1247, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284290.

Thomas:1979:NSG

[TP79] David R. Thomas and Donald A. Pierce. Neyman’s smooth goodnessof-fit test when the hypothesis is composite. Journal of the American Statistical Association, 74(366):441–445, June 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286351.

Troskie:1975:BRB

[Tro75] C. G. Troskie. Book review: [untitled]. Journal of the American Statistical Association, 70(349):251, March 1975. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285416.

Trumbo:1970:BRB

[Tru70] Bruce E. Trumbo. Book review: [untitled]. Journal of the American Statistical Association, 65(330):1007–1008, June 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284617.

Trumbo:1977:BRB

[Tru77] Bruce Edward Trumbo. Book review: [untitled]. Journal of the American Statistical Association, 72(360):926–927, December 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286494.

Trussell:1978:BRBa

[Tru78a] T. James Trussell. Book review: [untitled]. Journal of the American Statistical Association, 73(362):443, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286694.

Trussell:1978:BRBb

[Tru78b] T. James Trussell. Book review: [untitled]. Journal of the American Statistical Association, 73(362):444, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286699.

Tarter:1975:IAB

[TS75] Michael Tarter and Abraham Silvers. Implementation and applications of bivariate Gaussian mixture decomposition. Journal of the American Statistical Association, 70(349):47–55, March 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285374.

Thursby:1977:SPT

[TS77] Jerry G. Thursby and Peter Schmidt. Some properties of tests for specification error in a linear regression model. Journal of the American Statistical Association, 72(359):635–641, September 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286231.

Tsay:1976:SCO

[Tsa76] Jia-Yeong Tsay. On the sequential construction of D-optimal designs. Journal of the American Statistical Association, 71(355):671–674, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285598.

Tsutakawa:1972:DEB

[Tsu72] R. K. Tsutakawa. Design of experiment for bioassay. Journal of the American Statistical Association, 67(339):584–590, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284443.

Terrell:1971:ELS

[TT71] R. D. Terrell and N. E. Tuckwell. The efficiency of least squares in estimating a stable seasonal pattern. Journal of the American Statistical Association, 66(334):354–362, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283936.

Tin:1972:EDM

[TT72] Myint Tin and Than Toe. Estimation for domains in multistage sampling. Journal of the American Statistical Association, 67(340):913–916, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284661.

Tuckwell:1978:BRB

[Tuc78] Henry C. Tuckwell. Book review: [untitled]. Journal of the American Statistical Association, 73(364):896–897, December 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286315.

Tufte:1970:BRB

[Tuf70] Edward R. Tufte. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1414–1415, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284317.

Tufte:1979:BRB

[Tuf79] Edward R. Tufte. Book review: [untitled]. Journal of the American Statistical Association, 74(368):935, December 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286442.

Tukey:1979:CEP

[Tuk79a] John W. Tukey. Comments on Emanuel Parzen, Nonparametric statistical data modeling. Journal of the American Statistical Association, 74(365):121–122, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286735. See [Par79c].

Tukey:1979:MSR

[Tuk79b] John W. Tukey. Methodology, and the statistician’s responsibility for BOTH accuracy AND relevance. Journal of the American Statistical Association, 74(368):786–793, December 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286400. An expanded version of ASnum-

ber 121.

Turnovsky:1970:EEF

[Tur70] Stephen J. Turnovsky. Empirical evidence on the formation of price expectations. Journal of the American Statistical Association, 65 (332):1441–1454, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284328.

Turnbull:1974:NES

[Tur74] Bruce W. Turnbull. Nonparametric estimation of a survivorship function with doubly censored data. Journal of the American Statistical Association, 69(345):169–173, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285518.

Tan:1977:RTA

[TW77] W. Y. Tan and S. P. Wong. On the Roy–Tiku approximation to the distribution of sample variances from nonnormal universes. Journal of the American Statistical Association, 72(360):875–880, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286479.

Tsutakawa:1974:PTA

[TY74] R. K. Tsutakawa and S. L. Yang. Permutation tests applied to antibiotic drug resistance. Journal of the American Statistical Association, 69(345):87–92, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285504.

Uhler:1970:SMD

[UB70] Russell S. Uhler and Paul G. Bradley. A stochastic model for determining the economic prospects of petroleum exploration over large regions. Journal of the American Statistical Association, 65(330):623– 630, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284573.

Upton:1976:MMT

[Upt76] G. J. G. Upton. More multisample tests for the von Mises distribution. Journal of the American Statistical Association, 71 (355):675–678, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285599.

Urquhart:1975:BRB

[Urq75] N. Scott Urquhart. Book review: [untitled]. Journal of the American Statistical Association, 70(351):734, September 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285976.

Uthoff:1970:OTP

[Uth70] Vincent A. Uthoff. An optimum test property of two well-known statistics. Journal of the American Statistical Association, 65 (332):1597–1600, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284342.

VanNess:1972:BRU

[Van72] John W. Van Ness. Book review: [untitled]. Journal of the American Statistical Association, 67(338):497–498, June 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284425.

VanRyzin:1973:BRU

[Van73] John Van Ryzin. Book review: [untitled]. Journal of the American Statistical Association, 68(342):499–500, June 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284122.

Vannman:1976:EBO

[Van76] Kerstin Vannman. Estimators based on order statistics from a Pareto distribution. Journal of the American Statistical Association, 71 (355):704–708, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285605.

Nostrand:1978:BRB

[Van78] R. Craig Van Nostrand. Book review: [untitled]. Journal of the American Statistical Association, 73(364):889, December 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286303.

Veale:1975:IEE

[Vea75] James R. Veale. Improved estimation of expected life when one identified spurious observation may be present. Journal of the American Statistical Association, 70(350):398–401, June 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285830.

Venezia:1978:DTA

[Ven78] Itzhak Venezia. A decision-theoretic approach to the aggregation problem at the pre-data-collection stage. Journal of the American Statistical Association, 73(363):552–558, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286599.

Visscher:1978:OES

[VG78] William M. Visscher and Aaron S. Goldman. Optimization of earnings in stochastic industries, with applications to casinos. Journal of the American Statistical Association, 73(363):499–503, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286589.

Vijayan:1975:EVU

[Vij75] K. Vijayan. On estimating the variance in unequal probability sampling. Journal of the American Statistical Association, 70 (351):713–716, September 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285960.

Villegas:1976:SIU

[Vil76] C. Villegas. Strong inconsistency from uniform priors: Comment. Journal of the American Statistical Association, 71(353):124, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285751.

Villegas:1977:ISI

[Vil77a] C. Villegas. Inner statistical inference. Journal of the American Statistical Association, 72(358):453–458, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286818.

Villegas:1977:RI

[Vil77b] C. Villegas. On the representation of ignorance. Journal of the American Statistical Association, 72(359):651–654, September 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286233.

Vinod:1976:BRB

[Vin76a] H. D. Vinod. Book review: [untitled]. Journal of the American Statistical Association, 71(353):244–245, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285781.

Vinod:1976:ANR

[Vin76b] Hrishikesh D. Vinod. Application of new ridge regression methods to a study of Bell system scale economies. Journal of the American Statistical Association, 71(356):835–841, December 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286847.

Vinod:1976:EAE

[Vin76c] Hrishikesh D. Vinod. Effects of ARMA errors on the significance tests for regression coefficients. Journal of the American Statistical Association, 71(356):929–933, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286863.

Vitaliano:1979:RMT

[Vit79] Peter Paul Vitaliano. The relative merits of two statistics for testing A hypothesis about the equality of three correlated proportions. Journal of the American Statistical Association, 74(365):232–237, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286756.

Wackerly:1977:ASA

[Wac77] Dennis D. Wackerly. An alternative sequential approach to the problem of selecting the best of k populations. Journal of the American Statistical Association, 72(358):464–468, June 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286820.

Wachter:1979:BRB

[Wac79] Kenneth W. Wachter. Book review: [untitled]. Journal of the American Statistical Association, 74(366):507–508, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286373.

Wagner:1970:MLE

[Wag70] S. S. Wagner. The maximum-likelihood estimate for contingency tables with zero diagonal. Journal of the American Statistical Association, 65(331):1362–1383, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284302.

Wahi:1971:BRB

[Wah71] Pran N. Wahi. Book review: [untitled]. Journal of the American Statistical Association, 66(333):226–227, March 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284885.

Waksberg:1978:SMR

[Wak78] Joseph Waksberg. Sampling methods for random digit dialing. Journal of the American Statistical Association, 73(361):40–46, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286513.

Walsh:1970:SSA

[Wal70] John E. Walsh. Sample sizes for approximate independence of largest and smallest order statistics. Journal of the American Statistical Association, 65(330):860–863, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284593.

Wallis:1974:SAR

[Wal74] Kenneth F. Wallis. Seasonal adjustment and relations between variables. Journal of the American Statistical Association, 69(345):18–31, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285495.

Walker:1979:AEU

[Wal79a] Joseph J. Walker. An asymptotic expansion for unbalanced quadratic forms in normal variables. Journal of the American Statistical Association, 74(366):389–392, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286341.

Walter:1979:BRB

[Wal79b] S. D. Walter. Book review: [untitled]. Journal of the American Statistical Association, 74(368):930–931, December 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286435.

Wampler:1970:RAS

[Wam70] Roy H. Wampler. A report on the accuracy of some widely used least squares computer programs. Journal of the American Statistical Association, 65(330):549–565, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284566.

Wang:1971:PTE

[Wan71] Ying Y. Wang. Probabilities of the Type I errors of the Welch tests for the Behrens–Fisher problem. Journal of the American Statistical Association, 66(335):605–608, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283538.

Warner:1971:LRR

[War71] Stanley L. Warner. The linear randomized response model. Journal of the American Statistical Association, 66(336):884–888, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2284247.

Ware:1972:FSL

[War72] James H. Ware. The fitting of straight lines when both variables are subject to error and the ranks of the means are known. Journal of the American Statistical Association, 67(340):891–897, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284656.

Ware:1974:BRB

[War74] James H. Ware. Book review: [untitled]. Journal of the American Statistical Association, 69(345):281, March 1974. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285557.

Warner:1975:ASU

[War75] Stanley L. Warner. Advocate scoring for unbiased information. Journal of the American Statistical Association, 70(349):15–22, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285369.

Warner:1977:BRB

[War77] Dennis L. Warner. Book review: [untitled]. Journal of the American Statistical Association, 72(360):928, December 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286496.

Warner:1978:PIU

[War78a] Kenneth E. Warner. Possible increases in the underreporting of cigarette consumption. Journal of the American Statistical Association, 73(362):314–318, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286658.

Warrack:1978:BRB

[War78b] Giles Warrack. Book review: [untitled]. Journal of the American Statistical Association, 73(364):889, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286304.

Ware:1979:BRB

[War79] James H. Ware. Book review: [untitled]. Journal of the American Statistical Association, 74(367):735, September 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2287017.

Wasserman:1979:BRB

[Was79] Stanley Wasserman. Book review: [untitled]. Journal of the American Statistical Association, 74(367):727–728, September 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287006.

Watson:1975:BRB

[Wat75] G. S. Watson. Book review: [untitled]. Journal of the American Statistical Association, 70(351):730, September 1975. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285967.

Watson:1977:BRB

[Wat77] Geoffrey S. Watson. Book review: [untitled]. Journal of the American Statistical Association, 72(359):687–689, September 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286247.

Watson:1978:BRB

[Wat78] Geoffrey S. Watson. Book review: [untitled]. Journal of the American Statistical Association, 73(363):685–686, September 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286639.

Waternaux:1979:BRB

[Wat79] Christine Waternaux. Book review: [untitled]. Journal of the American Statistical Association, 74(365):244–245, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286762.

Weed:1971:SOS

[WB71] Harrison D. Weed, Jr. and Ralph A. Bradley. Sequential one-sample grouped signed rank tests for symmetry: Basic procedures. Journal of the American Statistical Association, 66(334):321–326, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283930.

White:1970:TPS

[WC70] R. E. White and H. A. Cowles. A test procedure for the simulated plant-record method of life analysis. Journal of the American Statistical Association, 65(331):1204–1212, September 1970. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284286.

Woodruff:1976:CMA

[WC76] Ralph S. Woodruff and Beverley D. Causey. Computerized method for approximating the variance of a complicated estimate. Journal of the American Statistical Association, 71(354):315–321, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285303.

Waller:1972:CBR

[WD72] Ray A. Waller and David B. Duncan. Corrigenda: a Bayes Rule for the Symmetric Multiple Comparisons Problem. Journal of the American Statistical Association, 67(337):253–255, March 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284766.

Watson:1976:EES

[WD76] William D. Watson, Jr. and Paul B. Downing. Enforcement of environmental standards and the central limit theorem. Journal of the American Statistical Association, 71(355):567–574, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285584.

Wei:1978:RPW

[WD78] L. J. Wei and S. Durham. The randomized play-the-winner rule in medical trials. Journal of the American Statistical Association, 73(364):840–843, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286290.

Wyckoff:1979:IPS

[WE79] James Wyckoff and Max Engelhardt. Inferential procedures on the shape parameter of a gamma distribution from censored data. Journal of the American Statistical Association, 74(368):866–871, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286413.

Weber:1971:ARP

[Web71] Donald C. Weber. Accident rate potential: An application of multiple regression analysis of a Poisson process. Journal of the American Statistical Association, 66(334):285–288, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283923.

Weber:1976:BRB

[Web76] Warren E. Weber. Book review: [untitled]. Journal of the American Statistical Association, 71(356):1003–1004, December 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286883.

Weber:1978:BRB

[Web78a] Warren E. Weber. Book review: [untitled]. Journal of the American Statistical Association, 73(361):227–228, March 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286570.

Wright:1978:ITP

[WEB78b] F. T. Wright, Max Engelhardt, and Lee J. Bain. Inferences for the two-parameter exponential distribution under Type I censored sampling. Journal of the American Statistical Association, 73(363):650–655, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286618.

Wegman:1974:BRB

[Weg74] Edward J. Wegman. Book review: [untitled]. Journal of the American Statistical Association, 69(345):280, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285555.

Weitzman:1970:IMC

[Wei70] R. A. Weitzman. Ideal multiple-choice items. Journal of the American Statistical Association, 65(329):71–89, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283576.

Weiss:1971:BRB

[Wei71a] Irving Weiss. Book review: [untitled]. Journal of the American Statistical Association, 66(335):659–660, September 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283562.

Weiss:1971:APM

[Wei71b] Lionel Weiss. Asymptotic properties of maximum likelihood estimators in some nonstandard cases. Journal of the American Statistical Association, 66(334):345–350, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283934.

Weisberg:1972:ULP

[Wei72] Herbert Weisberg. Upper and lower probability inferences from ordered multinomial data. Journal of the American Statistical Association, 67(340):884–890, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284655.

Weiss:1973:APM

[Wei73] Lionel Weiss. Asymptotic properties of maximum likelihood estimators in some nonstandard cases, II. Journal of the American Statistical Association, 68(342):428–430, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284091.

Weisberg:1974:BRB

[Wei74] Sanford Weisberg. Book review: [untitled]. Journal of the American Statistical Association, 69(346):581, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285710.

Weiss:1975:ADL

[Wei75] Lionel Weiss. The asymptotic distribution of the likelihood ratio in some nonstandard cases. Journal of the American Statistical Association, 70(349):204–208, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285404.

Weisberg:1976:BRBa

[Wei76a] Herbert I. Weisberg. Book review: [untitled]. Journal of the American Statistical Association, 71(354):522–523, June 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285348.

Weisberg:1976:BRBc

[Wei76b] Herbert I. Weisberg. Book review: [untitled]. Journal of the American Statistical Association, 71(356):1006, December 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286886.

Weisberg:1976:BRBb

[Wei76c] Sanford Weisberg. Book review: [untitled]. Journal of the American Statistical Association, 71(355):768, September 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285620.

Wei:1977:CDS

[Wei77a] Lee-Jen Wei. A class of designs for sequential clinical trials. Journal of the American Statistical Association, 72(358):382–386, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2286803.

Weisberg:1977:BRB

[Wei77b] Sanford Weisberg. Book review: [untitled]. Journal of the American Statistical Association, 72(360):925–926, December 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286493.

Weiss:1977:ESA

[Wei77c] Randall D. Weiss. Elasticities of substitution among capital and occupations in U.S. manufacturing. Journal of the American Statistical Association, 72(360):764–771, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286457.

Wei:1978:AUM

[Wei78a] L. J. Wei. An application of an urn model to the design of sequential controlled clinical trials. Journal of the American Statistical Association, 73(363):559–563, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286600.

Weisberg:1978:BRB

[Wei78b] Sanford Weisberg. Book review: [untitled]. Journal of the American Statistical Association, 73(361):223–224, March 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286563.

Weiss:1978:MKS

[Wei78c] Marc S. Weiss. Modification of the Kolmogorov–Smirnov statistic for use with correlated data. Journal of the American Statistical Association, 73(364):872–875, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286297.

Weissman:1978:EPL

[Wei78d] Ishay Weissman. Estimation of parameters and larger quantiles based on the k largest observations. Journal of the American Statistical Association, 73(364):812–815, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286285.

Welsch:1973:BRB

[Wel73] R. E. Welsch. Book review: [untitled]. Journal of the American Statistical Association, 68(341):243–244, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284189.

Welsch:1977:SMC

[Wel77] Roy E. Welsch. Stepwise multiple comparison procedures. Journal of the American Statistical Association, 72(359):566–575, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286218.

Welsch:1979:NSD

[Wel79] Roy E. Welsch. Nonparametric statistical data modeling: Comment. Journal of the American Statistical Association, 74(365):123–124, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286736.

West:1975:BEK

[Wes75] Sandra A. West. Bias in the estimator of Kendall’s rank correlation when extreme pairs are removed from a sample. Journal of the American Statistical Association, 70(350):439–442, June 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285836.

West:1977:ULP

[Wes77] Sandra A. West. Upper and lower probability inferences pertaining to Poisson processes. Journal of the American Statistical Association, 72(358):448–452, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286817.

Westley:1979:DUR

[Wes79] Glenn D. Westley. The demand for urban rail transportation. Journal of the American Statistical Association, 74(367):576–587, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286974.

Wetherill:1978:BRB

[Wet78] G. Barrie Wetherill. Book review: [untitled]. Journal of the American Statistical Association, 73(363):679, September 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286627.

Winkler:1979:WRR

[WF79] Robert L. Winkler and Leroy A. Franklin. Warner’s randomized response model: a Bayesian approach. Journal of the American Statistical Association, 74(365):207–214, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286752.

Williams:1971:SMC

[WG71] W. H. Williams and M. L. Goodman. A simple method for the construction of empirical confidence limits for economic forecasts. Journal of the American Statistical Association, 66(336):752–754, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284223.

Williams:1972:ACT

[WG72] O. Dale Williams and James E. Grizzle. Analysis of contingency tables having ordered response categories. Journal of the American Statistical Association, 67(337):55–63, March 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284701.

Wheeler:1973:BRB

[Whe73] Alan C. Wheeler. Book review: [untitled]. Journal of the American Statistical Association, 68(341):243, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284188.

White:1970:TNP

[Whi70] John S. White. Tables of normal percentile points. Journal of the American Statistical Association, 65(330):635–638, June 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284575.

White:1976:BRB

[Whi76] Colin White. Book review: [untitled]. Journal of the American Statistical Association, 71(355):776, September 1976. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285636.

White:1979:BRB

[Whi79a] Kenneth J. White. Book review: [untitled]. Journal of the American Statistical Association, 74(365):248, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286769.

Whitt:1979:NIS

[Whi79b] Ward Whitt. A note on the influence of the sample on the posterior distribution. Journal of the American Statistical Association, 74(366):424–426, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286347.

Wichern:1979:BRBa

[Wic79a] Dean W. Wichern. Book review: [untitled]. Journal of the American Statistical Association, 74(365):246–247, March 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286765.

Wichern:1979:BRBb

[Wic79b] Dean W. Wichern. Book review: [untitled]. Journal of the American Statistical Association, 74(367):733–734, September 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287014.

Wiggins:1970:BRB

[Wig70] Nancy Wiggins. Book review: [untitled]. Journal of the American Statistical Association, 65(332):1683–1685, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284354.

Wigton:1977:BRB

[Wig77] William H. Wigton. Book review: [untitled]. Journal of the American Statistical Association, 72(357):236, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286957.

Wilcken:1971:CDB

[Wil71] C. L. Wilcken. Comment on “Discriminatory Bias in Rates Charged by the Canadian Automobile Insurance Industry”. Journal of the American Statistical Association, 66(334):289–290, June 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283924.

Wilson:1973:CRC

[Wil73] Ronald W. Wilson. Comment on “review of claim that excess morbidity and disability can be ascribed to smoking”. Journal of the American Statistical Association, 68(341):85–87, March 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284144.

Williford:1974:BRB

[Wil74] William O. Williford. Book review: [untitled]. Journal of the American Statistical Association, 69(346):581, June 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285709.

Williams:1976:BRB

[Wil76] O. Dale Williams. Book review: [untitled]. Journal of the American Statistical Association, 71(353):251, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285794.

Williams:1978:CAF

[Wil78] J. S. Williams. Canonical analysis: a factor-analytic method of comparing finite discrete distribution functions. Journal of the American Statistical Association, 73(364):781–786, December 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286280.

Winkler:1971:PPS

[Win71] Robert L. Winkler. Probabilistic prediction: Some experimental results. Journal of the American Statistical Association, 66 (336):675–685, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284212.

Winkler:1972:DTA

[Win72a] Robert L. Winkler. A decision-theoretic approach to interval estimation. Journal of the American Statistical Association, 67(337):187– 191, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284720.

Winkler:1972:SPP

[Win72b] Robert L. Winkler. Some probability paradoxes in choice from among random alternatives: Comment. Journal of the American Statistical Association, 67(338):376–378, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284386.

Winkler:1974:BRB

[Win74a] Othmar W. Winkler. Book review: [untitled]. Journal of the American Statistical Association, 69(346):576–577, June 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285703.

Winterbottom:1974:LCL

[Win74b] Alan Winterbottom. Lower confidence limits for series system reliability from binomial subsystem data. Journal of the American Statistical Association, 69(347):782–788, September 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286018.

Winkler:1975:BRBa

[Win75a] Robert L. Winkler. Book review: [untitled]. Journal of the American Statistical Association, 70(352):965, December 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285481.

Winkler:1975:BRBb

[Win75b] Robert L. Winkler. Book review: [untitled]. Journal of the American Statistical Association, 70(352):965, December 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285482.

Winkler:1975:BRBc

[Win75c] Robert L. Winkler. Book review: [untitled]. Journal of the American Statistical Association, 70(352):965–966, December 1975. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285483.

Winkler:1975:CPA

[Win75d] Robert L. Winkler. Cognitive processes and the assessment of subjective probability distributions: Comment. Journal of the American Statistical Association, 70(350):290–291, June 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285809.

Winkler:1976:BRB

[Win76a] Robert L. Winkler. Book review: [untitled]. Journal of the American Statistical Association, 71(354):527–528, June 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285354.

Winter:1976:BRB

[Win76b] Sidney G. Winter. Book review: [untitled]. Journal of the American Statistical Association, 71(355):770, September 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285624.

Winkler:1977:BRB

[Win77] Robert L. Winkler. Book review: [untitled]. Journal of the American Statistical Association, 72(358):482, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286824.

Winkler:1978:BRBa

[Win78a] Robert L. Winkler. Book review: [untitled]. Journal of the American Statistical Association, 73(361):222–223, March 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286561.

Winkler:1978:BRBb

[Win78b] Robert L. Winkler. Book review: [untitled]. Journal of the American Statistical Association, 73(364):894, December 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286310.

Winston:1978:BRB

[Win78c] Wayne Winston. Book review: [untitled]. Journal of the American Statistical Association, 73(361):225, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286566.

Winkler:1979:BRB

[Win79] Robert L. Winkler. Book review: [untitled]. Journal of the American Statistical Association, 74(368):922–923, December 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286426.

Wiorkowski:1975:BRB

[Wio75] John J. Wiorkowski. Book review: [untitled]. Journal of the American Statistical Association, 70(349):255–256, March 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285420.

Wipf:1971:BRB

[Wip71] Larry J. Wipf. Book review: [untitled]. Journal of the American Statistical Association, 66(335):658–659, September 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2283560.

Wisniewski:1972:PTH

[Wis72] T. K. M. Wisniewski. Power of tests of homogeneity of a binomial series. Journal of the American Statistical Association, 67 (339):680–683, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284467.

Wittes:1974:AMC

[Wit74] Janet T. Wittes. Applications of a multinomial capture-recapture model to epidemiological data. Journal of the American Statistical Association, 69(345):93–97, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285505.

Waikar:1971:TSE

[WK71] Vasant B. Waikar and S. K. Katti. On a two-stage estimate of the mean. Journal of the American Statistical Association, 66(333):75– 81, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284851.

White:1977:PCA

[WL77] Robert F. White and Thomas M. Lewinson. Probabilistic clustering for attributes of mixed type with biopharmaceutical applications. Journal of the American Statistical Association, 72(358):271–277, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2286788.

Williams:1970:SBP

[WM70] W. H. Williams and C. L. Mallows. Systematic biases in panel surveys due to differential nonresponse. Journal of the American Statistical Association, 65(331):1338–1349, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284300.

Wiorkowski:1971:PCS

[WM71] John J. Wiorkowski and William R. McLeod. Prediction and control of the size of an input-output system. Journal of the American Statistical Association, 66(336):712–719, December 1971. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284217.

Wecker:1977:BAP

[WM77] William E. Wecker and Thomas E. Morton. Bounds on absorption probabilities for the m-dimensional random walk. Journal of the American Statistical Association, 72(357):154–156, March 1977. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286927.

Wolf:1973:TCV

[WN73] Edward H. Wolf and Joseph I. Naus. Tables of critical values for a k-sample Kolmogorov–Smirnov test statistic. Journal of the American Statistical Association, 68(344):994–997, December 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284538.

Wallenstein:1974:PSL

[WN74a] Sylvan R. Wallenstein and Joseph I. Naus. Probabilities for the size of largest clusters and smallest intervals. Journal of the American Statistical Association, 69(347):690–697, September 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286003.

Wolf:1974:CTC

[WN74b] E. H. Wolf and J. I. Naus. Corrigenda: Tables of Critical Values for a k-Sample Kolmogorov–Smirnov Test Statistic. Journal of the American Statistical Association, 69(345):288, March 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285577.

Wang:1977:BDA

[WNIO77] Ming-Mei Wang, Melvin R. Novick, Gerald L. Isaacs, and Dan Ozenne. A Bayesian data analysis system for the evaluation of social programs. Journal of the American Statistical Association, 72(360):711–722, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286451.

Wolfe:1973:SGR

[Wol73] Douglas A. Wolfe. Some general results about uncorrelated statistics. Journal of the American Statistical Association, 68(344):1013–1018, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284542.

Wolfe:1974:CPW

[Wol74a] Douglas A. Wolfe. A characterization of population weightedsymmetry and related results. Journal of the American Statistical Association, 69(347):819–822, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286025.

Wolfowitz:1974:BRBa

[Wol74b] J. Wolfowitz. Book review: [untitled]. Journal of the American Statistical Association, 69(345):278–279, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2285551.

Wolfowitz:1974:BRBb

[Wol74c] J. Wolfowitz. Book review: [untitled]. Journal of the American Statistical Association, 69(345):279, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285553.

Wolfe:1977:CPS

[Wol77a] Douglas A. Wolfe. On a class of partially sequential two-sample test procedures. Journal of the American Statistical Association, 72(357):202–205, March 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286938.

Wolfowitz:1977:CCS

[Wol77b] J. Wolfowitz. Conditional confidence statements and confidence estimators: Comment. Journal of the American Statistical Association, 72(360):820–822, December 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286467.

Wolf:1979:BRB

[Wol79a] Charles Wolf, Jr. Book review: [untitled]. Journal of the American Statistical Association, 74(365):249–250, March 1979. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286772.

Wolter:1979:CEF

[Wol79b] Kirk M. Wolter. Composite estimation in finite populations. Journal of the American Statistical Association, 74(367):604–613, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286977.

Woodruff:1971:SMA

[Woo71] Ralph S. Woodruff. A simple method for approximating the variance of a complicated estimate. Journal of the American Statistical Association, 66(334):411–414, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283947.

Woodward:1976:APT

[Woo76] W. A. Woodward. Approximation of Pearson Type IV tail probabilities. Journal of the American Statistical Association, 71(354):513– 514, June 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2285343.

Woodroofe:1979:OAB

[Woo79] Michael Woodroofe. A one-armed bandit problem with a concomitant variable. Journal of the American Statistical Association, 74 (368):799–806, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286402.

Worm:1974:SAF

[Wor74] George H. Worm. Sensitivity analysis of the fixed point vector of transition matrices. Journal of the American Statistical Association, 69(348):961–962, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286171.

Worsley:1979:LRT

[Wor79] K. J. Worsley. On the likelihood ratio test for a shift in location of normal populations. Journal of the American Statistical Association, 74(366):365–367, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286336.

Wright:1978:ESI

[Wri78] F. T. Wright. Estimating strictly increasing regression functions. Journal of the American Statistical Association, 73(363):636–639, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286615.

Walker:1978:DLC

[WS78] Glenn A. Walker and John G. Saw. The distribution of linear combinations of t-variables. Journal of the American Statistical Association, 73(364):876–878, December 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286298.

Wiley:1973:SCC

[WSB73] David E. Wiley, William H. Schmidt, and William J. Bramble. Studies of a class of covariance structure models. Journal of the American Statistical Association, 68(342):317–323, June 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284069.

Wilson:1971:TSC

[WT71] P. David Wilson and James Tonascia. Tables for shortest confidence intervals on the standard deviation and variance ratio from normal distributions. Journal of the American Statistical Association, 66(336):909–912, December 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284253.

Wainer:1975:MSM

[WT75] Howard Wainer and David Thissen. Multivariate semi-metric smoothing in multiple prediction. Journal of the American Statistical Association, 70(351):568–573, September 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285934.

Wolde-Tsadik:1976:GSC

[WT76] Girma Wolde-Tsadik. A generalization of an SPRT for the correlation coefficient. Journal of the American Statistical Association, 71(355):709–710, September 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2285606.

Warren:1974:EVA

[WWF74] Richard D. Warren, Joan Keller White, and Wayne A. Fuller. An errors-in-variables analysis of managerial role performance. Journal of the American Statistical Association, 69(348):886–893, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286158.

Wynn:1977:MPS

[Wyn77] Henry P. Wynn. Minimax purposive survey sampling design. Journal of the American Statistical Association, 72(359):655–657, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286234.

Wyzga:1978:EAP

[Wyz78] Ronald E. Wyzga. The effect of air pollution upon mortality: a consideration of distributed lag models. Journal of the American Statistical Association, 73(363):463–472, September 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286581.

Yackel:1974:BRB

[Yac74] James W. Yackel. Book review: [untitled]. Journal of the American Statistical Association, 69(347):831, September 1974. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286033.

Yakowitz:1976:SSH

[Yak76] Sidney J. Yakowitz. Small-sample hypothesis tests of Markov order, with application to simulated and hydrologic chains. Journal of the American Statistical Association, 71(353):132–136, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285754.

Yarnold:1970:MEG

[Yar70] James K. Yarnold. The minimum expectation in χ^{2} goodness of fit tests and the accuracy of approximations for the null distribution. Journal of the American Statistical Association, 65(330):864– 886, June 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284594.

Yancey:1972:SFS

[YBJ72] T. A. Yancey, M. E. Bock, and G. G. Judge. Some finite sample results for Theil’s mixed regression estimator. Journal of the American Statistical Association, 67(337):176–179, March 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284718.

Yeats:1973:EPA

[Yea73] A. J. Yeats. An evaluation of the predictive ability of the FRB sensitive price index. Journal of the American Statistical Association, 68(344):782–787, December 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284500.

Yoshihara:1971:PAP

[YFS71] Kunio Yoshihara, Kenichi Furuya, and Takao Suzuki. The problem of accounting for productivity change in the construction price index. Journal of the American Statistical Association, 66(333):33–41, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284843.

Ylvisaker:1977:TR

[Ylv77] Donald Ylvisaker. Test resistance. Journal of the American Statistical Association, 72(359):551–556, September 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286216.

Young:1970:BRB

[You70] Allan H. Young. Book review: [untitled]. Journal of the American Statistical Association, 65(331):1411–1413, September 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284314.

Young:1972:MDA

[You72a] J. C. Young. The moments of a distribution after repeated selection with error. Journal of the American Statistical Association, 67(337):206–210, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284725.

Young:1972:STS

[You72b] Kan Hau Young. A synthesis of time-series and cross-section analyses: Demand for air transportation service. Journal of the American Statistical Association, 67(339):560–566, September 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284438.

Young:1973:BRB

[You73] Kan Hua Young. Book review: [untitled]. Journal of the American Statistical Association, 68(342):498–499, June 1973. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284120.

Young:1976:ICM

[You76a] Dennis L. Young. Inference concerning the mean vector when the covariance matrix is totally reducible. Journal of the American Statistical Association, 71(355):696–699, September 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285603.

Young:1976:VSF

[You76b] J. C. Young. Varietal screening from finite normal populations. Journal of the American Statistical Association, 71(353):87–92, March 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2285736.

Yalovsky:1972:BRB

[YS72] Morty A. Yalovsky and George P. H. Styan. Book review: [untitled]. Journal of the American Statistical Association, 67(340):957–958, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284676.

Yang:1977:ETT

[YS77] Grace Yang and Mete Sirvanci. Estimation of a time-truncated exponential parameter used in life testing. Journal of the American Statistical Association, 72(358):444–447, June 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2286816.

Yu:1975:BPS

[Yu75] George C. S. Yu. Bounds on the power for some Renyi-type tests. Journal of the American Statistical Association, 70(349):233–237, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2285409.

Z:1972:BRB

[Z.72] S. Z. Book review: [untitled]. Journal of the American Statistical Association, 67(340):963–964, December 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2284687.

Zabell:1976:BRB

[Zab76] Sandy Zabell. Book review: [untitled]. Journal of the American Statistical Association, 71(356):996–998, December 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286877.

Zabell:1978:BRB

[Zab78] Sandy Zabell. Book review: [untitled]. Journal of the American Statistical Association, 73(361):220–221, March 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286557.

Zacks:1970:UMA

[Zac70] S. Zacks. Uniformly most accurate upper tolerance limits for monotone likelihood ratio families of discrete distributions. Journal of the American Statistical Association, 65(329):307–316, March 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283594.

Zacks:1973:SDF

[Zac73] S. Zacks. Sequential design for a fixed width interval estimation of the common mean of two normal distributions. I. The case of one variance known. Journal of the American Statistical Association, 68(342):422– 427, June 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284090.

Zahn:1977:BRB

[Zah77] Douglas A. Zahn. Book review: [untitled]. Journal of the American Statistical Association, 72(357):233–235, March 1977. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2286955.

Zarembka:1970:BRB

[Zar70] Paul Zarembka. Book review: [untitled]. Journal of the American Statistical Association, 65(332):1691–1693, December 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284359.

Zar:1972:STS

[Zar72a] Jerrold H. Zar. Significance testing of the Spearman rank correlation coefficient. Journal of the American Statistical Association, 67(339):578–580, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2284441.

Zarnowitz:1972:BRB

[Zar72b] Victor Zarnowitz. Book review: [untitled]. Journal of the American Statistical Association, 67(338):495–496, June 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2284421.

Zeckhauser:1971:COI

[Zec71] Richard Zeckhauser. Combining overlapping information. Journal of the American Statistical Association, 66(333):91–92, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284854.

Zellner:1971:BNB

[Zel71] Arnold Zellner. Bayesian and non-Bayesian analysis of the log-normal distribution and log-normal regression. Journal of the American Statistical Association, 66(334):327–330, June 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283931.

Zellner:1975:BAR

[Zel75] Arnold Zellner. Bayesian analysis of regression error terms. Journal of the American Statistical Association, 70(349):138–144, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2285389.

Zellner:1976:BNB

[Zel76a] Arnold Zellner. Bayesian and non-Bayesian analysis of the regression model with multivariate Student-t error terms. Journal of the American Statistical Association, 71(354):400–405, June 1976. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285322. See corrigenda [Zel76b].

Zellner:1976:CBN

[Zel76b] Arnold Zellner. Corrigenda: “Bayesian and Non-Bayesian Analysis of the Regression Model with Multivariate Student-t Error Terms”. Journal of the American Statistical Association, 71(356):1010, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286897.

See [Zel76a].

Zellner:1979:SAEa

[Zel79a] Arnold Zellner. Statistical analysis of econometric models. Journal of the American Statistical Association, 74(367):628–643, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286981.

Zellner:1979:SAEb

[Zel79b] Arnold Zellner. Statistical analysis of econometric models: Rejoinder. Journal of the American Statistical Association, 74(367):650–651, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2286986.

Zerbe:1979:RAC

[Zer79] Gary O. Zerbe. Randomization analysis of the completely randomized design extended to growth and response curves. Journal of the American Statistical Association, 74(365):215–221, March 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2286753.

Zacks:1973:DAS

[ZF73] S. Zacks and J. Fennell. Distribution of adjusted stock levels under statistical adaptive control procedures of inventory systems. Journal of the American Statistical Association, 68(341):88–91, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2284146.

Zeigler:1972:DSP

[ZG72] R. K. Zeigler and Aaron Goldman. A double sampling plan for comparing two variances. Journal of the American Statistical Association, 67(339):698–701, September 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284471.

Zinger:1972:NMC

[Zin72] Alexis Zinger. A note on multiple-choice items. Journal of the American Statistical Association, 67(338):340–341, June 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284379.

Zacks:1971:MSE

[ZM71] Shelley Zacks and Roy C. Milton. Mean square errors of the best unbiased and maximum likelihood estimators of tail probabilities in normal distributions. Journal of the American Statistical Association, 66(335):590–593, September 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283535.

Zellner:1979:MEL

[ZP79] Arnold Zellner and Soo-Bin Park. Minimum expected loss (MELO) estimators for functions of parameters and structural coefficients of econometric models. Journal of the American Statistical Association, 74(365):185–193, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286749.

Zahn:1971:ECT

[ZR71] Douglas A. Zahn and Gail C. Roberts. Exact χ^{2} criterion tables with cell expectations one: An application to Coleman’s measure of consensus. Journal of the American Statistical Association, 66(333):145– 148, March 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2284864.

See corrigenda [ZR74].

Zahn:1974:CEC

[ZR74] Douglas A. Zahn and Gail C. Roberts. Corrigenda: Exact χ^{2} Criterion Tables with Cell Expectations One: An Application to Coleman’s Measure of Consensus. Journal of the American Statistical Association, 69(345):288, March 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2285578. See [ZR71].

Zyskind:1973:BRB

[Zys73] George Zyskind. Book review: [untitled]. Journal of the American Statistical Association, 68(344):1021–1022, December 1973. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2284545.

Zyskind:1975:BRB

[Zys75] George Zyskind. Book review: [untitled]. Journal of the American Statistical Association, 70(349):252–254, March 1975. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2285418.

Zacks:1972:ESF

[ZZ72] S. Zacks and W. J. Zimmer. Estimators of severity factors in a multiplicative Poisson model. Journal of the American Statistical Association, 67(337):192–195, March 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2284721.